A Bibliography of Publications in ACM SIGACT News

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

28 August 2024 Version 1.33

Title word cross-reference

#1 [Mat14a]. #53 [Hes03]. #P [HV95].

{0, 1} [Vio09]. $139.95 [Cha09]. 2 [Hem06a, SS06]. $28.00 [Jou94]. 3 [Sto73a, Sto73b]. $30.00 [Gas22f]. $33.95 [Gas05a]. $37.50 [Sta93]. $39.50 [Gas94, O’R92a]. $39.95 [Mak93, Pan93]. 3x + 1 [Sch97]. $46.25 [Vla94]. $49.50 [Hen93]. $49.95 [Pur93, Rie93, Sis93, Zwa93]. 4 × 4 [Shy78a]. $51.48 [Pap05b]. $54.95 [Oli04b]. $59.99 [Gre22h]. $60.00 [Ada11]. $69.00 [Mik93]. $72.00 [Oli04c]. $79.95 [Bea07]. $99.99 [Tra22]. _{2} [FdV23]. _{∞} [Lit92]. A = B [Kre00, PWZ96]. e [Bla06, Mao94]. i [GKP96]. i, n [GKP96]. k [GG86]. λ [Di 95, Lea97]. N [Nis78, GKP96]. n^{2} [BBG94]. O(ln n) [Qia87]. O(log log n) [AM75]. O(n^{log_{2} 7} log n) [Boo77]. p [Gal74, LP92]. P = NP [Lip10]. P^{3} [BCG^{+}22]. π [Bla97, Hem03a, Mil99, Puc00, SW01]. Σ∗ [Swe87]. \sqrt{−1} [Nah98].

-Aha [Gas09h]. -Calculus [Di 95, Hem03a, Lea97, Mil99, SW01, Puc00].

-colorability [Sto73a, Sto73b]. -complete [Gal74]. -hard [BBG94]. -shovelers [LP92]. -stage [Hem06a, SS06].

//cstheory.stackexchange.com [CEG^{+}10]. /Heidelberg [Gri10].

0 [Kri08, Law06, Lou05, Mos05, Oli06, Pap10, Pri10, Vio04, Wan05, dW07]. 0-201-54886-0 [Vla94]. 0-262-07143-6 [Sta93]. 0-262-08306-X [Wil04]. 0-262-19475-9 [Wil04]. 0-262-52258-6 [Mos05]. 0-387-00152-2 [Law06]. 0-387-55597-8 [Mik93]. 0-387-97664-7 [Sis93]. 0-444-82841-9 [Oli06]. 0-521-39115-6 [Lou05]. 0-521-39196-2 [JK92]. 0-521-40044-9 [Pur93]. 0-521-40347-2 [Rie93]. 0-521-40436-3 [Pan93]. 0-521-40446-0 [O’R92a]. 0-521-42036-9 [Mak93]. 0-521-43111-5 [Zwa93]. 0-521-80177-X [Wan05]. 0-521-80816-2 [Pri10]. 0-8176-3593-9 [Jou94]. 0-8176-3625-0 [Hen93]. 0-8176-3719-2 [Gas94]. 0-8176-4127-0 [dW07]. 0-8186-7200-5 [Mos05]. 0-8218-2121-0 [Kri08]. 0-8218-2161-X [dW07]. 00PM [Gas06e]. ’07 [Bor07b].

1 [Ful10, Knu11, Spr06, Was08a, vS16b, Che79c, LaT11, Ros23]. 1-58488-421-5 [Spr06]. 1-58488-518-1 [Was08a]. 1-58488-551-3 [Ful10]. 100th [Pud06]. 113pp [Hen93]. 14th [Kan14]. 16271 [FRRS17, Wel17b]. 179pp [JK92]. 19 [Hem20c]. 1975 [Eil76]. 1976 [DL79, Her77]. 1977 [Che79a, Sub78]. 1978 [Che79c, Che80]. 1979 [Che82, Mai80]. 1983 [Smo21]. 1994 [Puc05c, Zeu94]. 1996 [LB99, Rod96]. 1997 [AP98, Lon97, PD98, RW99, Sim04, To06, Wan97, Gol00a]. 1998 [Cha05, Lon98]. 1999 [Gas06e, Lee03b, Pet04, Rie99b].

2 [Law06, Pri10, vS16c]. 2000 [Bla06, Irw03a, Irw08, Kel08e, Mar03a, Wid03]. 2000-Year [Gas22e]. 2001 [Dav06a, Ess05, Hes03, Irw03b, Lee03a, Mas03, Náv04, Pap05a, Wei03]. 2002 [Dav05, Kin05, Kin06, McN04, Tay04]. 2003 [Bor07a, Dan06, Gas05b, Gle05, Kre07, Nat07, Nat10, Rau08, Sha10a, Zha04, Zha05]. 2004 [Dav06b, Dav08b, Gas04a, Kat06b, Les06, Raj04e]. 2005 [Aar07, Afr05, Ben06, Dav08c, Fah05, Gas10g, Kat06a, Kat07, Rog08, Sch06, Sta07, Tal09, Ter08, Wer10]. 2006 [Dol07, GKKK06, Jan08a, Kye09, Lee10, Raz07, Upa10]. 2007 [Col10]. 2008 [Azi10a, Azi10b, Cou09a, DGW08, Jan10a, Mei10, Neu10, Wil10a, Wil10c, Wil10b]. 2009 [Apo10, Cal10, Das09, Gas10e, Jan10b, Mos10, Sri10]. 2010 [Bar10, Bon10c, Gas10f, KM10, Van10]. 2011 [HM11, Hel11, KPS11]. 2012 [Arf13, Göö12, Sen12]. 2013 [Ano13, BS13, Fan13, Tim13]. 2014 [Den14, Gro14, PT14, Rad14]. 2015 [AG15, Bur15, Cal15, Tse15]. 2016 [Sch16, Su16, Tan16, Tse17, Wel17a, Wel17c]. 2017 [Ass17, Dav17, FM17, Gee17, Kon17, Kuz18]. 2018 [Are18, BD18, BK18, Dhu19, Mil18b, Sch18, Ten18, Wel19a]. 2019 [BDCDL19, BFU19, Koc19]. 2020 [BBDD20, AF20, AK20]. 2021 [Pic21].

2022 [ABLO23, Ali23a]. 21st [For21, Gas20e, Ham90]. 226pp [Gas94]. 22nd [BBAGP09]. 233pp [Mak93]. 267pp [Pur93]. 26th [Agu12]. 273pp [Zwa93]. 2nd [Bón13b, HMU01a, Lab13, Mas00a, Mas01a, dV16, Zeu89].

3 [Ful10, Vio04]. 3-540-66783-0 [dW07, Vio04]. 372pp [Rie93]. 3rd [CK18, Tra22].

4 [Rog08]. 419pp [Sta93]. 427pp [Sis93]. 4a [Rog13]. 4th [Bón11b, Bón11c, FK89, Pro90].

5 [Mos05, Spr06].

6 [Kos74a, Lou05, Mur10]. 60 [GMS74]. 60th [Wel15b]. 691 [Sta89]. 6th [Par00c].

717pp [Vla94].

88/89 [Hem15d].

9 [Lev10, Oli06]. ’94 [TT95]. ’96 [Hul96]. 978 [Lev10, Mur10, Pap10]. 978-0-521-73833-0 [Pap10]. 978-0-521-87282-9 [Lev10]. 978-3-540-49243-6 [Mur10]. ’98 [Par98b].

= [CT23, Gas10f, Gas12f, Gas19h, Yli04].

A.K [Con03]. Aanderaa [LS74]. Aaron [Fur11, Gas16c]. Aaronson [Gas07h, Gre13]. Aarts [Mar09a]. Abadi [Pos99]. Abdennadher [Oli04a]. abelian [Vio19]. Abelson [Gas09a, Gas21c]. ability [DLM07]. Abiteboul [Nag21b]. Abstract [Bau17, McN04, Par02b, Akm92, Bat93, Coh93b, CH06, Cul82, Di 93, Ead93, EKP79, GW97, Kan93, Kau93, Lio93, Mil72a, O’H98, Pac93, Sug93, Tam93, Tol93, dFR93, vLL76]. abstract/concrete [Cul82]. Abstractions [Jac06, Lee10]. Abstracts [Sta76a, SIG83a, SIG83b, SIG83c, SIG84a, SIG84b, SIG84c, Cha73a, SIG82b]. academe [Cor83]. Academic [Bor07a, Kre07]. acceptance [Her00b]. accepting [Pet94]. Access [Aar07, Wil06, Lew14, SvL74]. Accountability [Ros02a]. Aceto [Zwa93]. achievements [Lou91]. Ackermann [McB92, PM80]. ACM [ACM12, Kos74a, Chr04a, Kei08a, Kei08b, Kei08c, Kei09a, Kei09b, Kei09c, Puc05a, Puc06a, Raj01a, Raj01b, Raj01c, Raj02b, Raj02c, Raj02d, Raj02e, Raj03a, Raj03b, Raj03c, Raj04a, Raj04b, Raj05a, Raj05b, Raj06a, Raj06b, SP07, Vit00]. Action [Ace92, Mos92, Rie93, Zwa93]. Active [Jen01, McC01]. Activities [Gur94, SP07, Ros95c]. activity [GP89]. acyclic [Kos97]. ad [Raj02a]. Adam [Bar12]. Adams [Gas10e, Bón12]. Add [Kre01]. Addendum [McB87a, SS04a]. Addison [Che82, Gas09a, Mai80, Puc03, Rog08, Vla94, Fah04]. Addison-Wesley [Che82, Mai80, Rog08, Vla94, Fah04]. additive [Tre09, Tew14]. address [FKR08, Via08b]. Addressing [Mah19]. Adiabatic [CCD15]. Adjeroh [Neu10]. Adleman [Her83]. advance [SIG82b]. Advanced [AW98, AW04, Bra08, Jan10a]. Advances [AG18, Ali21a, Ali23c, Ass23, CKN21, Gas13e, Wel18c, DBHHT94, Din13, Hem99]. Adventures [Con09, Gas11a, Por12, HV95, Hro09]. advertisement [Joh96b]. Advice [BFK^{+}16]. Adware [Ayc11, Wer13]. Affairs [Gas07c]. affine [Bha13]. affine-invariant [Bha13]. Afoot [Kri08, Meh00]. afraid [Par98a]. After [ALL08, Gas09a, Gas21c, Ing83]. again [Ros05b]. Against [LST22]. Agarwal [O’R92a]. Age [Bla12, Nag21b]. agenda [BCvR09]. Agents [Mil09, Pap10]. Aggregability [KS06b]. aggregation [Chr14]. Agreement [Ali21a, CKN21]. Aha [Gas09h, Gar06, Gas09h]. Aho [Ben73a, Weg74]. Aigner [Bón11c, Boo10, Nag16]. al [Irw03a]. Alan [All98, Dav11, McN04, Sha13, Wil10a, Fen21a, Hem14a, Pet08]. Albert [Cha09, Gas10a, Gle05, Lou18b, Ste11b]. Albert-Làaszlò [Lou18b]. Alberto [Gas15i, Irw03b]. Aldous [Mas00b]. Alecos [Gas10b]. Alekhnovich [Raz07]. Alessandro [Sri10]. Alevras [Oli04b]. Alex [Gas06e, Gas14h, Har11]. Alexander [Gas09c, Gas16f, Jam98, Kri08, Wil04, Zha05, Gas15d]. Alexandru [Kel16a]. Alfred [FF04, Les06]. Algebra [BDK01, Cal12, DFM07, FM08, Gas12a, MMW06, Sho05, Sho09, TM11, Tse91, Vin09, Xie12, Che06, Gri10, HWR09, Ses09, Tay02, Gre16f, Kat06a, Mik02, Nag19]. Algebraic [Bar09, BCS97, Coh93a, Coh95, Coh96, EKP79, FM15, Gin75, Her00b, Kel08e, LST22, Mos12, Sti93, Sti09, TM11, Vin09, Vol21, BHHR99, Bur72, Eil76, Fit11, Kop11, Nag17a, To06]. Algebras [Ace92, GPP83, Zwa93]. ALGO [vS12b]. Algol [GMS74]. Algorithm [Che21, GT02, Kar11, Kel16a, Kel09a, Knu77b, Lei07, McN99, Sav74, Ski98, Ski08, Tra22, Buh96, HN14, Her83, Hua03, Kei10b, Kos97, KT98, Mof83, Mor97, Qia87, Ram94, Weg74, Ang76, Nar14, Náv04]. algorithmes [Lew78]. Algorithmic [AG18, APR16, Azi10a, BS08, BV07, Chi86, Chr05a, Col10, CHK^{+}09, DH10, Edw08, FM09, FRRS17, Hro09, IP05, Jou09, KL10, LO05, NRTV07, Puc05b, Sha19, SLB09, dV16, VGS05, Wel16a, Wel17b, Wel18c, All72, AY12, Bón09a, Gas13d, Gas15c, Law06, Lev10, Por12, Ste12, Teu13, Upa21]. Algorithmics [Gas13f, Hro03, Kel16b, Mar15a, Mas03]. Algorithms [Aga91, Amb17, ASvS23a, ASvS23b, AL07, Ata99, AB09b, AB10, Baa78, Baa88, BV00, BFM^{+}15, BS79, BP05, BS98, Bor09a, BFK^{+}16, BO06, CH15, Cha09, CS91, CHL07, CL17, CK18, Das92, DPV08, Dea09, DO07, Dey07, Die04, DLO05, DP09, Ead93, Edm08, FB10, FGT17, FS19, Gas14i, Gho07, GT97, GT01, GT04, GT06, GT10, Gre16f, Gre19g, GK81, GK82, GK90, GK07, Gus97a, HT03, Hen09, Her02a, Hoc97b, HSvS19, HSvS20, HSvS21, Hua14, Jan10b, Jen01, JS03, Jun99, Jun05, Jun08, Kat07, Khu01, Khu02, Khu12, Kin15, Knu97, Knu00, Kor93, Kor02, KV08, KV12, KR03, Kye09, Lab12, Lee03b, Lei92, Lev07, Lev12, Mas00a, Mas01a, Mav09, McC01, MS08, MB05, MU05, Mur13, O’R92a, OZ06, Pit00, RZ09, RS91b]. Algorithms [Ros93c, Sch04, Sha92, She08, She10, Sot10, Sta07, Sta92, Sto02, SS06, Tam93, Tau06, Tel00, Val09, dV18, Wei95e, Wel15a, Wel15c, Zeu89, Zeu92, dB^{+}97, dB^{+}00, dB^{+}08, vS15, vS16a, vS16b, vS16c, vS17a, vS17b, vS18a, vS18b, vS21, vS22, Ala03, Amb04, Azi14, BY95, BS84, yC08, Che13, Chr04e, Chr04f, Chr05d, CKM06, Chr08b, Chr09a, Chr09b, Chr10a, Chr10b, Chr11, DKW05, DBHHT94, Fot11, Fra85, Gal75, GL80, Goo93, HN14, HWR09, Hem06a, HW12, HES72, Iva02b, Jaw05, Jon72, Knu11, KV00, KV02, Kor06, KS99, KO11, KM10, LP10, Mat97, MM99, Pas99, Pru98, Puc05h, RS88, RS91a, Ros03a, Ros04d, Ros04e, Ros05c, SB14, Ski99, Sub95, Sud00, Vis91, Wel14b, WS11, vS12a, vS12b, vS13, vS14a, vS14b, Hoc97a]. Algorithms [Kel09b, MR95, Par95a, Thu90, Wel17c, Cha12, Con03, Dav06b, EA10, Fah04, Fah05, FM11, For04, Gre17g, Lab13, Lam13, Mar03a, McN01b, Mik11, Nag21b, Nag21a, Pop12, Sga09, Sri10, Ste11a, Wil04, Xie12, dW07, Apo14a, Cad13, Kin14, Ada11, Ben98, Elr11, Gal15, Gas99a, Gus97b, Iva02a, Jam98, Kel09b, Kin15, Mav02, Pap12b, Sch06, Wid03]. alikes [BA05]. Alive [Ros96c, Aaz18]. Allan [You99]. Allouche [Ber04]. Almost [Bra95]. Alon [Upa10]. alphabet [AM75]. Alt [Kin15]. alternate [Ben73a]. Alternative [Puc07a]. alternatives [Mor83]. Always [Bra95, Cou11, Pol09]. Amar [Neu10]. America [dOG10]. American [Dav05, FM09, Fri11, Kat05, Kob98, Kri08, McN99, Ogi01, Sch09, Wil10d, dW07, GL09]. Amin [To06]. Amir [Cad21, Gas15d, Sis93]. amortized [HR00, IMP11]. Amos [Dav08c, Dav09]. AMS [Fri10, Gas10e, Gol00a, Mas00b]. Amy [Mat14a, Nag13a]. analog [Kug76a]. Analyses [Gas14i]. Analysis [Baa78, Baa88, BV00, BEY05, Dey07, DP09, Fah04, Gas16d, GT02, GK81, GK82, GK90, GK07, Har11, Ios07, Jac06, JK92, Jen97, Knu97, Knu00, Lei07, Lev07, Lev12, Mav09, McC01, McN01b, Mik93, Mik11, MU05, Náv04, Puc03, Rie97, Sri10, Tse91, Wel17c, BS84, Dav06a, Dwy90, HR00, Hut01, Kel16a, MR97, Ram94, Sti08, Zeu89, Kat07, Kel09a, Lee10, McN99, Sot10, Sta07, You99, Apo16, Jen01]. Analytic [BMMR20, FS09, Sha10a, Shp03, Bón10b, Bón14]. analyze [Lew78]. Ananay [Fah04]. Anany [Gas13d]. Ancestral [Kel16b]. Anchor [Ree01]. Ancient [Ben06, Mol05]. Anderson [Sga10]. Andreas [Bor09a, FB14, Koz14]. Andrew [Dav11]. Andrews [LRV10]. Anecdotes [Gas06e]. Angular [Kau93]. Anil [Dek99a, MT04, TA95]. Animated [Ros95b]. Animating [Sta92]. Animation [Ros94b, Vin09, Ros95c, TM11]. Anita [Les06]. Ankur [Upa21]. Anna [Gre17f]. annealing [Has85a]. Annie [Gas10b]. anniversary [Pud06]. Annotated [Pet08, Wil10c]. Announcement [GKP88, EKP79, Sta89]. Announcements [Lat95a, Lat95b, Lat95c]. Annual [Ali19, Ali20d, Kei10d, Wan97, Wel14d, Wel15d, Wel16d, Wel18d, All00, FK89, Kei12d, Vit00, Wel13]. Answer [SF89]. answered [CEG^{+}10]. answers [Kat14a]. Anthology [She15]. Anthony [Kre07, Kye09, Sch09]. Antiquity [Gas22e]. Antoine [AY12]. Antoniou [Bor09a]. Anup [Cad21]. Aperiodic [FGH18]. Apocryphia [Gas06e]. Apostolos [Gas10b]. Applegate [Spr09]. Application [Ham76, BBAGP09, Sug93]. Applications [AH06, AL07, Bat93, Ber04, BC^{+}91, Bón13b, BM07, BC09, DK02, DK07, Fur11, KN01, LV93, LV97, Li08, LV08, Men04, MM07, NS93, NS97, Nie02, OZ06, Pro90, Sha10a, Shp03, TSS03, Ter11, Weg00, dB^{+}97, dB^{+}00, dB^{+}08, AS03, BCRV03, Esc93, Gue83, KL91, KM10, Mil72b, Rav90b, Sud00, Tur84a, Tur84b, Ven02, Bón10a, Bor09a, Cad21, Cha09, Dek99a, Gas97a, Gas16g, Joh10, Kri99, Mas00a, MT04, Mur13, Sga10, Sha12a, Spr08, TA95, Yan11b, Zha05, Azi14]. Applied [BSS11, HWR09, Joh10, Lot05, Rob84, RT05, Gop08, Bón11a, Dav08a, Kat13b, Pap11, Xie12, Yan20]. Applying [Bar12, Elb09, Wil08]. appreciation [Hem14a]. Approach [AB09a, Bón10a, Bra77, BC09, BV07, DFM07, Dun86b, HK05, Mar98, Mar08a, McC01, Nag20, Sch00, Tse91, Vel94, BSS11, Gas02a, MB05, Sav98, Apo10, Coh08, Abr11, Azi11, Che79b, Col10, Jen01, Pet04, Smi09, dV16, Wid03, Wil10d, Cad14, Koz18]. approaches [vdBN04]. APPROX [vS12b]. Approximate [BT20, Con95]. Approximately [Bru13]. Approximating [FRT04, O’R04a]. Approximation [Cha12, Hem06a, Hoc97a, Hoc97b, Hro03, Mas03, RW19, SS06, Chr04c, CK98, Elk04, GL80, WS11, BS98]. Apress [Ter08]. April [Gol00a, PD98]. Apt [Aro98]. Arbitrary [Che79c, VV78, Sta87, XL93]. architecture [BBAGP09]. Architectures [Das92, Lei92, Eya13]. Arcs [Ros77]. Area [Di 93, FFMS94]. Argumentation [Bau17]. arguments [CP13]. Ariel [Nag17b, Zim13]. arising [Gas13c]. Arithmetic [BZ11, Con03, Edw08, FM09, Kay18, Kor93, Kor02, All97, Lib99, Yan12]. Arnaud [Pop12]. Arora [Apo10]. Arrangements [Aga91, O’R92a]. Arrays [ABM08, Das92, Lei92, Neu10]. Art [BQ03, Gas05b, Gas10b, Knu11, Rog08, Rog13, Rog18, DM22, Gre17f]. Arthur [Gas05b, Gre16g, Gre17e, Mar03a, Sha12b, dW07]. articles [Gas14d, Abr13]. Artificial [Mos05, FC83]. ask [Par98a]. asking [Pap95]. aspect [Vak13]. Aspects [Mar09a, MAK07, Upa21, BS09, Geo09]. assertions [BMO77, Bur72]. Assessing [Ros99c]. assignment [Mit01b]. assignments [Mit01a]. Associates [Jan10b]. Association [dOG10]. Asymptopia [Gas16b]. Asymptotics [Sha04, Bea07]. Asynchronous [Ray05, Chr05b, Rey05]. Atallah [Pap12b, Gas99a]. Attacks [Ira12, Kei11c, Yan07]. Attiya [Her00a]. attribute [Eng89]. atypical [HW12]. Auctions [Nag20]. August [Wan97]. Ausiello [Kin15]. Authentication [Dwo95]. Author [Gas22d, Gas22e, Hac09]. Automata [AH06, DKV09, Eil76, Gal13, Gas94, Gas13f, Gin75, GTW02, Hav74, HU79, HMU01a, HMU01b, HMU07, KN01, Kin12, LR76, Mai80, MP21, Rod96, Sak11, Sch75, Sha09, Str94, Wil10b, But71, Cul82, Gop08, Kar00, OS14, Pet94, Rod99, MT04, Cad12, Her77, MT04, Sga10]. Automata-Theoretic [Eil76, Gin75]. automata/ [Cul99]. Automated [BK12, Lov78, Che79a]. Automath [Che80, van79]. Automatic [AS03, AT17, KVY12, Wel17a, Lio93, Ber04]. automating [CW72]. automaton [Hem08a, Lei80, Lit03, Mar08b, Paz70, Ree71, Sta84a, Sta84b]. available [GL02, SIG83b]. Average [BMMR20, PV21, Dwy90]. Average-case [PV21, Dwy90]. Avi [Gre21e]. award [Rod09, Ano13, Rou05].

Awarded [Ten18, Pel13]. awareness [San84]. axiomatic [All72]. Axioms [Lon71, FKW94, Li04]. Aycock [San10, Wer13].

B [Cal13, Cha12, Coh07, Kat14b, Kin15, Kum13, LaT11, Mas01b, Mat14b]. Baader [Nar00]. Baase [Zeu89]. bachelor [Tho93]. back [FFGH08]. Backlash [Ros95d]. backlog [Rav90a, Wei95a, Wei95b, Wei96a]. Backwards [Ros00c]. Bad [Mar20]. Badly [Ree71]. Baeten [Cal12]. bags [GLMW96]. Bake [Gre17e]. Bakhadyr [MT04]. balanced [Woo78]. Ballard [Mos05]. Barabàasi [Lou18b]. Barak [Apo10]. Barbara [Rod09]. Barber [Gal14]. Bard [Fit11]. Barnes [Bla06]. Barrel [FPE04]. Barriers [Gas15c]. Barrow [Azi10b]. Barrow-Green [Azi10b]. Barry [Bón11a, Pap11]. Barwise [Dek98]. Baseball [Alb03, Gle05]. Based [Amo99, LB99, OR99a, Ogi01, RW99, WG00, HR00, Jon72, Jor83, Kho05, Khu05b, KGO10, Par00c, Kat13a]. Basel [For99]. Basic [Che21, Elr11, Gas09b, Gol01, Hin97, Jen97, Kat05, Kre01, MS08, Mik93, Puc03, TS96, TS00, Avi01, FB14, Pos98]. Basics [Dwy03, Wan05, Gol10a]. Basin [Dav11, Kat13b]. Basis [Lov78, MW93, Che79a]. Basten [Cal12]. Bauer [Azn13, Kil12]. Baum [OR99a]. Baumgart [Gre18e]. Baur [Mor85]. Baur-Strassen [Mor85]. Bayesian [Gal14]. BDA [Rad14]. Be [Gas20f, AF14, Har95, Sei01a, Weg86]. Beautiful [Hem14a]. Beaver [Aar20, Dun86b, Wal82, YDX97, Jul92]. Beck [Gas12e]. Beckmann [Bla06]. Bedroom [Gas09h, Hay08]. Beedle [Jan10b]. Beginner [Gre19e, Rau08, dV18, Wal03, Wal12]. Beginners [Gre16e]. Behind [Alp18, Mos10]. Behrends [Gas09h, Nag16]. being [Gas07a, Slo07]. Beissinger [Ayu12]. Bela [Ist12, Bón09c, Cen00, Kri99]. Belazzougui [Kel16a]. Bell [Neu10]. Bella [Sta10]. Ben [Gas13c, Vol21]. Bender [Mos05]. Benjamin [Gas05b, Gre16g, Gre17e, Kin05, Kin06, Sha12b, She15]. Bent [Ada09a, Gas10e]. Benthem [Che80]. Berg [Mas00a]. Berkeley [Sim84a, Sim84b]. Berlekamp [Nag17a, She13]. Berlin [Dav06b, Dav08c, Dav09, For99, Gri10, Kye09, Vio04]. Berlinski [Cen15]. Berman [Sch06]. Bernhard [Wil04]. Best [Gas10e, Nah04, Rou05, CK98]. Bets [Ski01, Gas02b]. better [Bra85, DO13, YDX97]. between [KLMY21, Mil72a, Ven13b, vL74]. Beware [Ros95d]. Beyond [GW18, Hua14, Raz16, SS02, HV95, HS12, Wil04]. BFTW [KR09]. Bibliography [Did93, Joh90]. Bierbrauer [Spr06]. Big [Knu76a, VM85, Pap95]. bigger [Mat14b]. Biggs [Low12]. Bijective [Loe11, Bón13a]. Bin [vS15]. Binary [BV07, Joh10, Weg00, Bur78, Coo83, Kan88, Col10]. bio [RB11, Azn13]. bio-ontologies [RB11, Azn13]. Biography [Bla06]. Bioinformatics [Dwy03, Gas02a, Wan05, Azn14, Pap13a]. Bioinspired [Cha09, OZ06, Her13]. Biological [Kel16a]. Biologically [BO06, Kye09]. Biologists [Pap13a]. Biology [FC^{+}99, Gus97a, Lab12, Val09, JKL99, FC^{+}99, Ben98, Neh99, Gus97b]. Biomolecular [Hem99]. Bioprocesses [CHK^{+}09, Ste12]. bipartite [BM08, Kri75]. birds [Gal78]. Birkäuser [Sha10a]. Birkhäuser [For99, Gas94, Gas00a, Hen93, Jam98, Jou94, Lea97, Mar03a, MT04, Pap05a, Puc05c, Rau08, dW07, EA10]. BIRS [Wel17c]. Birthday [Wel15b]. Biscuits [Sha12b]. Bishop [Cho12]. bisimulation [Sti99]. Bit [GW18, Boo77]. Bits [ALL08, Gas09a, Gas21c]. Bixby [Spr09]. Black [Gre21f]. Blackburn [Hes03]. Blanchet [Bón09a]. Blanchet-Sadri [Bón09a]. Blanchette [Lew13]. Blanton [Pap12b]. Blatner [Bla06]. block [Kat14b]. blocks [BCRV03, BCRV04]. blog [Gas07h, Gas15h]. Blogs [Gas07h]. Bloomsbury [Gas10b]. Blown [ALL08, Gas09a, Gas21c]. blues [Ros04b]. Blum [Gre07, McN01a, XL93]. Boaz [Apo10]. Bollobás [Ist12, Bón09c, Cen00, Kri99]. Bóna [Gas15b]. Bonato [Sch09]. Bonnie [Pin10]. Book [Aar07, Abr11, Ada11, All98, Amo99, Apo11, Apo12, Aro98, Avi01, Ayu12, Azi10a, Azi10b, Azi11, Bar12, Bea07, Bel11, Ben06, Ben98, Ber04, BS98, Bón09b, Bón09a, Bón09c, Bón10b, Bón10a, Bón11a, Bón11b, Bón11c, Bón12, Boo10, Bor08, Bor07a, Bor09a, Bra95, Cad11, Cal10, Cen00, Cen03, Cen11, Cha09, Cha05, Che10, Che77, Che79b, Che79a, Che79c, Che80, Che82, Coh07, Col10, Con03, Cou09b, Cou09a, Cou10, Cou11, Dan06, Das92, Dav02, Dav05, Dav06a, Dav06b, Dav08b, Dav08c, Dav08a, Dav11, DL79, Dea09, Dek97, Dek98, Dek99b, Dek99a, Dye11, Eil76, EA09, EA10, Elr11, Ess05, Fab11, Fah04, Fah05, Fal07, FM08, FM09, FM11, Fit11, For99, For04, Fri10, Fri11, Ful10, Fur11, Gas94, Gas97a, Gas97b, Gas98a, Gas98b, Gas98c, Gas98d]. Book [Gas99a, Gas99b, Gas99c, GK99, Gas99e, Gas00a, Gas00b, Gas00c, Gas00d, Gas00e, Gas01b, Gas01a, Gas02b, Gas04a, Gas05b, Gas05a, GKKK06, Gas07c, Gas07b, Gas07d, Gas09a, Gas09b, Gas09c, Gas10b, Gas10e, Gas10d, Gas10a, Gas10c, Gas10f, Gas10g, Gas11a, Gas11b, Gas12a, Gas14c, Gas15a, Gle05, Gol00a, Gol00b, Gre07, Gre15a, Gre15b, Gre15c, Gre16a, Gre16b, Gre16c, Gre16d, Gre17a, Gre17b, Gre17c, Gre17d, Gre18a, Gre18b, Gre18c, Gre18d, Gre19d, Gre19a, Gre19b, Gre19c, Gre20a, Gre20b, Gre20c, Gre20d, Gre21d, Gre21a, Gre21b, Gre21c, Gre22b, Gre22c, Gre22d, Gre22e, Gri10, dOG10, Gup98, Har10, Har11, Hem03a, Hen09, Hen93, Her00a, Her77, Hes03, Ira12, Irw03b, Irw03a, Irw08, Ist00, Iva02a, JK92, Jam98, Jan08a, Jan08b, Jan10b, Jan10a, Jen01, Joh10]. Book [Jou94, Kar11, Kat06a, Kat06b, Kat07, Kat12, Kel08e, Kel09a, Khu98a, Kin05, Kin06, Kin11, Kob98, Kop11, Kre00, Kre07, Kri99, Kri00, Kri08, Kye09, LaT11, Law06, Lea97, Lea00, Lee03a, Lee03b, Lee10, Les06, Lev10, Lew78, Lou05, Low12, Lur99, Mai80, Mak99, Mak93, Mar03a, Mar09a, Mas00b, Mas00a, Mas01a, Mas01b, Mas02, Mas03, Mav02, Mav09, McN04, McN99, McN01a, Mei10, Mel11, Mie11, Mik02, Mik93, Mik11, MT04, Mou10, Mur10, Nar00, Nat07, Nat10, Náv04, Neh99, Neu10, O’R92a, OR99a, Ogi01, Oli04a, Oli04c, Oli04b, Oli06, Pan93, Pap11, Pap12a, Pap05b, Pap05a, Pap10, Par11, Pet04, Pin10, Pos98, Pos99, Pri10, Pru00, Puc00, Puc01b, Puc01a, PC03, Puc05c, Pur93, Pux97, Rau08, Ree01, Ree77, Rey11]. Book [Rie93, Rie10, Rog08, Roj12, Rot07, San10, Sar03, Sch09, Sch06, Sch08b, Sga09, Sga10, Sha10a, Sha12a, Sim04, Sin99, Sis93, Smi00, Smi09, Soi09, Sot10, Spr03, Spr06, Spr08, Spr09, Sri06, Sri10, Sta07, Sta10, Sta11, Sta93, Ste11b, Ste11a, Sub78, Sza12, Tal09, TA95, Tas99, TBG95, Tay04, Ter08, Ter11, TM11, To06, Tra23a, Tra23b, Tra23c, Tra23d, Tra24, Upa10, Vio04, Vla94, Wan05, Was08a, Wei03, Wer10, Wid03, Wil04, Wil10a, Wil10c, Wil10b, Wil10d, Yan11a, Yan11b, You99, Zeu89, Zeu92, Zha04, Zha05, Zwa93, yea00a, Gas99d, Gas01c, Gas01d, Gas01e, Gas02c, Gas02d, Gas02e, Gas02f, Gas03a, Gas03b, Gas03c, Gas03d, Gas04b, Gas04c, Gas04d, Gas05c, Gas05d, Gas05e, Gas06a, Gas06b, Gas06c, Gas06d, Gas07e, Gas07f, Gas07g, Gas08c]. book [Gas08d, Gas08a, Gas09d, Gas09e, Gas09f, Gas09g, Gas10h, Gas10i, Gas10j, Gas10k, Gas11c, Gas11d, Gas11e, Gas11f, Gas12b, Gas12c, Gas12d, Gas13b, Gas13a, Gas14a, Gas14b, Hac09, Kat05, Puc07a, AZ98, AZ99, AZ01, AZ04, AZ10, Bón11c]. Books [Bla06, GKKK06, Gas09b, Gas00f, Ree01]. Boolean [Apo16, Gre22f, Yel78, Ang76, ASU86, BCRV03, BCRV04, Boo77, CK02, CH10, HHH98, Mul69, RST15, Shy78b, Gas13e, Kam13, Tay04, Azi14]. Boosting [Kin14]. Börger [Sim04]. Borodin [You99]. Borut [Gal21]. Boston [For99, Jam98, MT04]. both [Gre81]. bound [Che74, CK89, Jul92, Kan88, Wil11, dJvL75]. boundary [Mil72a]. Bounded [GM99, GHOP22, KA89b, GK80, KA89a, LS78, Mac95, Pet94, For99]. Bounds [Ali23c, Ass23, Kay18, LST22, Sha10a, Shp03, HHR97, Vio09, Vio17, Wel14c]. Boxer [Wid03]. Boxes [Ber00, She13]. Boyce [TF82]. BPP [CT23]. Brabazon [Kye09]. Bradfield [Hen93]. Bradley [Fur11]. Brady [Che79b]. Branching [Weg00, Joh10]. Brandt [Nag17b]. Brass [Gas07d, Jan10a]. Bravely [Gol06]. Braverman [Cal10]. Brazil [Agu12]. Breadth [Ros93c]. Breadth-First [Ros93c]. Break [Ayu12, BP06]. Breaking [Bar15, DKW05]. Bremen [Paz13]. Brent [Yan12]. Brewer [GL02]. Brewing [Ros96b]. Brian [Abr13, Gas09h, Gas15f]. Brief [ABLO23, Che24, TZ16, Grä08, Joh96a]. Brigitte [Nar14]. Brook [Ski99]. Browder [Gas13c]. Brown [Azi10a, Sha12b]. Brualdi [Bón10a]. Bruce [Gas16c]. Bruno [Cad14]. Buchmann [Col10, Lee03a]. Buettcher [Rub13]. buffer [Eng18, Gol10b]. Building [Zho09]. Bundy [Dav11]. Burdens [Bla12, Lew13]. Burdman [Les06]. Burger [Gas10e]. Bürgisser [To06]. Buried [Cou11, Pol09]. burning [Dun86a]. Burrows [Neu10, ABM08]. Burton [Che82]. Business [Sta78a, FK89]. Busy [Aar20, Wal82, YDX97, Dun86b, Jul92]. Bytes [Ban15, Ban15]. byways [Nag13a]. Byzantine [Ali21a, CKN21, Kei10c, KS10, KR09, Vuk10].

C [Apo12, Cal12, Che10, Dav05, Jan10b, Kin05, Kin06, Kin15, Mar15a, Pap12a, Pap05a, Ter08, Wan12, HS01, O’R94d, RZ09, Wei95e, WK01, Wel05, Dek99b]. Cabinet [Gas09h, Ste09]. Cai [Gre22f]. Calc [Ada09a, Gas10e]. Calculated [Ski01, Gas02b]. Calculating [Sin72]. calculations [McB85, McB87a]. Calculi [Lea00, Han94]. Calculus [BM07, Di 95, Fur11, Hem03a, Lea97, McG72a, Mil99, SW01, CR74, Fis72, Puc00]. call [Kan14]. Cambridge [Ada11, Apo10, Avi01, Azi10a, Ben98, Ber04, Bón10b, Bón10a, Coh07, Cou09b, Cou09a, Dav06a, Dav11, DL79, Dea09, Dek98, Dek99b, Gas10a, Gup98, Hes03, Irw08, JK92, Jan10a, Kat05, Kat06a, Kat07, Les06, Lev10, Lou05, Mak93, Mou10, Nar00, O’R92a, Pan93, Pap10, Pos98, Pri10, Puc03, Pur93, Rie93, Rie10, Rot07, Sga09, Sga10, Smi09, Sot10, Sri10, Sta07, Wan05, Wil04, Wil10b, Zwa93]. Can [BT08, Con09, Gas11a, Gas20f, Sei01a, Vis91, AF14, Cas13]. canonical [Fas86]. Cantone [Irw03b]. Capocelli [Gal92]. capturing [Hof00]. Cardelli [Pos99]. Carl [Mat14a, Sha12a, Sta93]. Carlo [Gas00a, Maz14, Mou10]. Carmit [Kat12]. Carol [Xie12]. Carsten [Her13]. Casanova [Pop12]. Case [RS91b, Wil06, BHHR99, Dwy90, HW12, KW81, PV21, RS91a, DLO09]. casual [Wil11]. Category [Tse91, GMS74]. Catherine [Dav05, Mar15a, Mou10]. causes [McG72b]. CBS [Gas06e]. CCS [Ali21d, BB21]. ceiling [Ang91]. Celebration [Wel15b]. cellular [Hem08a, Mar08b]. census [HG00]. Center [FS19, Oli04c, Wel19b, KW07]. Central [Bón12, Ada74, Ada09b]. Centre [Che80]. Centrum [Che79c]. Century [For21, Fre01, Gas07b, Gas20e, Had15, Kar00, Gas15f]. certain [Gab75]. Chabane [Mie11]. Chaitin [Tas99]. Challenge [Kre01, GJM02, Raj00, RW94, Dav05]. Challenges [CCD15, Con09, DM22, FS15, Gas11a, Har15, Mah19, GJM02, Lou91, SR03, Vio17]. Challenging [Mit01a]. Chance [AN09, Gas10a, Now02, Gas10a]. Change [Ken05, Fri10]. changed [She14]. changing [Che07]. Chapman [Ben06, Cha09, Che10, Che79b, Ful10, Kat05, Kel09a, Spr06, Was08a]. characterization [LM73, LL75]. characterizations [Vol99]. Charles [Bla06, Cen14, Hen93, Mat14b, Wil10c]. Charms [Ros23]. Chartier [Ban15]. Chartrand [Fab11, Gre16g]. Chases [Cen09]. Chattarjee [Kat13a]. Chazal [BF17]. ’Cheat [She97, Sei96]. cheatlist [Lev96]. Chechik [Ano13]. Checking [van79, LP07, Che80]. Chemoinformatics [FB10, Ste11a]. Chen [Gre22f]. Cheng [Sil12, Gre17e]. Cheng-Kuan [Sil12]. Chi [Li14]. Chiang [Kat14a]. Chicago [Nat07, Nat10, Ros04b]. Child [Ber00, She13]. China [Wan97]. Choice [Hes92, Nag17b, Nag18b, Pan93]. Chris [Lea00]. Christian [Gas11a]. Christof [Mou14]. Christophe [Gal15]. Christos [Gas10b, Kel09b]. Chromatic [CZ09, GHOP22, Fab11]. Chung [Par11, Lur99]. Church [BA05, Ros75]. Chvátal [Spr09, Ros23]. cipher [Kat14b]. Ciphers [Wag03, Xie12, HWR09, Kat05]. Circle [Moh93]. Circles [BM96, Pet02, Dek98, Gas10c]. Circuit [Gas94, Kay18, Str94, Gol77, Lad75, Lit03, Tur84a, Tur84b, Wil11]. Circuits [LST22, RST15, dRGR22, All97, BGH07, Dew79, Wat11]. circularity [Eng89]. Circumstances [Smo21]. Ciril [Koz14]. clarity [GH00, HG00]. Clark [Fri10, Ree77]. Clarke [Rub13]. class [Ang76, Cal83, FKR08, Gas13c, Via08b]. Classes [Che82, HH24, Lew79, Boo84a, Boo84b, Gue83, HV95, Hof00, Raj03d, Vol99]. Classic [Gas21d, Maz14]. Classical [BGG97, Bra94a, BT20, HS01, KSV02, Sim04, TS99, Zom06, Bra94b, dW07, Pap05a]. Classifiers [Her02a, Wil04]. classroom [Abr13]. Clausal [Dav02, Wol93]. Clausen [To06]. claw [BHT97]. claw-free [BHT97]. Clever [Nah04]. clique [Koz78]. close [Hil07]. closure [Shy78b, Yel78]. Clote [Tay04]. cloth [O’R92a]. Cloud [BBAGP09, BCvR09, CKS09]. clouds [Kei09b]. Clustering [Azn14]. CMO [Wel17c]. Co [Bla06, Che79a, Her77, Cul99, FdV23]. co-finite [Cul99]. Coalgebras [Kur06]. COCOON’97 [Wan97]. Codd [TF82]. Code [Dwy03, Ree01, Sin99, Ten87, Kho05, Khu05b, Wan05, Mar20]. Codebreakers [Gas01b, Kah67a, Kah67b, Kah74, Kah96]. Codes [BP06, Big08, KS16, Mol05, Sal07, Sti93, Sti09, Xie12, DO13, Gur04, HWR09, Nas13, O’R04d, Jor83, Ayu12, Kop11, Ben06, Low12, Nag19]. Coding [Bie05, HHL00, Nag17a, Sal05a, Spr06, Cas13, Kei13a, Wie83, Irw03a, Tal09, Wer10]. Cognition [Gas22f]. Cohen [Kel08e, Was08a, Grä08]. Coin [Blu83]. Colin [Gal13, Gas10e, Wei03]. collapse [HHH98, You92, Cen14]. collapses [HHH98]. Collection [BR99, Fre01, Gar06, Gas11g, Gas07b, SP07]. Collective [Dan06, Mou03]. College [Gas16h]. colorability [Sto73a, Sto73b]. Colorful [Gas09c, Soi09]. Coloring [Gas09c, Gas16a, Soi09, HK84a, HK84b]. Colorings [Gas16a]. colors [Khu05a, Raj05c]. Coloured [Jen97, Mik93]. COLT [Zeu94]. Colton [Neh99]. Column [Aar05, AK90, ÁDFT15, Ali19, Ali20a, Ali20b, Ali20c, Ali20d, Ali21d, Ali21a, Ali21b, Ali21c, Ali22, Ali23a, Ali23c, Ali23b, All23, AH91, Amb17, ASvS23a, ASvS23b, Are18, AdW17, BKM18, Bau17, Bra94a, BMMR20, BT20, CCD15, Car15, Cha20, CT23, DJ14, Dum16, Dum18, DJ19, Dwo95, ED17, FGT17, FS15, Gas97b, Gas98a, Gas98b, Gas98c, Gas98d, Gas99c, GK99, Gas99e, Gas00b, Gas00c, Gas00d, Gas00e, Gas14c, Gas15a, Gas17a, Gas17b, Gas17c, Gas18a, Gas18b, Gas18c, Gas18d, GHMP19, Gas19h, Gas19d, Gas19a, Gas19b, Gas19c, GB20, Gas20a, Gas20b, Gas20c, GM21, Gas21b, Gas22a, Gas22c, Gas23b, Gas23c, Gas23d, Gas24, Gee17, Gha23, Gol06, Gre15a, Gre15b, Gre15c, Gre16a, Gre16b, Gre16c, Gre16d, Gre17a, Gre17b, Gre17c]. Column [Gre17d, Gre18a, Gre18b, Gre18c, Gre18d, Gre19d, Gre19a, Gre19b, Gre19c, Gre20a, Gre20b, Gre20c, Gre21d, Gre21a, Gre21b, Gre21c, Gre22b, Gre22c, Gre22d, Gre22e, Gro14, HH24, Hem14e, Hem15a, Hem15b, Hem15c, Hem15d, Hem16a, Hem16b, Hem16c, Hem17a, Hem17b, Hem17c, HS18, Hem18c, Hem18a, Hem18b, Hem19d, Hem19a, Hem19b, Hem19c, Hem20a, Hem20b, Hem20c, Hem20d, Hem21d, Hem21a, Hem21b, Hem21c, Hem22a, Hem22b, Hem22d, Hem23a, Hem23b, HSvS19, HSvS20, HSvS21, Hua14, KA89b, Kay18, Kei08b, Kel14d, Kel15a, Kel15b, Kel15c, Kel15d, Kel16c, Kel16d, Kel16e, Kel16f, Kel17a, Kel17b, Kel17c, Kel17d, Kel18a, Kel18b, Kel18c, Kel18d, Kel19d, Kel19a, Kel19b, Kel19c, Kel20a, Kel20b, Kel20c, Kel20d, Kel21d, Kel21a, Kel21b]. Column [Kel21c, Kel22a, Kel22b, Kel23a, Kel23b, Kel23c, Kel23d, Kel24, KL16, KLMY21, Koc19, KS16, LST22, Lin92b, MP21, Ngo23, OW17, O’R94c, PV21, Pic21, Pig15, Raz16, RST15, Sha17, SS06, Tan16, Tao20, Tra23a, Tra23b, Tra23c, Tra23d, Tra24, Vio19, Vol21, Vol23a, Vol23b, Vol24, Wei95c, Wei95d, Wei96b, Wel14d, Wel15a, Wel15b, Wel15c, Wel15d, Wel16a, Wel16b, Wel16c, Wel16d, Wel17a, Wel17b, Wel17c, Wel17d, Wel18a, Wel18b, Wel18c, Wel18d, Wel19a, Wel19b, Wel19c, Zhu20, dRGR22, vS15, vS16a, vS16b, vS16c, vS17a, vS17b, vS18a, vS18b, vS21, vS22, yea00a, AAV13, BCG^{+}22, Ber92, Bha13, Bra90a, Bra89b, Bra89a, Bra90b, BI08, yC08, Che11b, Chr04a, Chr04b, Chr04e, Chr04f, Chr05c, Chr05d, CKM06, Chr08a, Chr08b, Chr09a, Chr09b, Chr10a, Chr10b, Chr11, CP13]. column [DO99, Dum12, DT12, DJ13a, DJ13b, DT14, DLM07, Dvi09, Dwo89, Dwo90, Dwo96, Gas99d, Gas01c, Gas01d, Gas01e, Gas02c, Gas02d, Gas02e, Gas02f, Gas03a, Gas03b, Gas03c, Gas03d, Gas04b, Gas04c, Gas04d, Gas05c, Gas05d, Gas05e, Gas06a, Gas06b, Gas06c, Gas06d, Gas07e, Gas07f, Gas07g, Gas08c, Gas08d, Gas08a, Gas09d, Gas09e, Gas09f, Gas09g, Gas10h, Gas10i, Gas10j, Gas10k, Gas11c, Gas11d, Gas11e, Gas11f, Gas12b, Gas12c, Gas12d, Gas13b, Gas13a, Gas14a, Gas14b, Gas22b, Goo93, Gur04, Hag99, Hag02, Hem94a, Hem97b, Hem01b, Hem01c, Hem01d, Hem02a, Hem02b, Hem02c, Hem03b, Hem03c, Hem03d, Hem04, Hem05b, Hem05c, Hem05a, Hem06c, Hem06d, Hem06e, Hem06b, Hem07c, Hem07d, Hem08b, Hem08c, Hem08d]. column [Hem09a, Hem09b, Hem09c, Hem09d, Hem10b, Hem10c, Hem10d, Hem11a, Hem11b, Hem11c, Hem11d, Hem12a, Hem12b, Hem12c, Hem13c, Hem13a, Hem13b, Hem14b, Hem14c, Hem14d, Hem22c, Kei08c, Kei08d, Kei09a, Kei09b, Kei09c, Kei09d, Kei10a, Kei10d, Kei11a, Kei11b, Kei11c, Kei11d, Kei12a, Kei12b, Kei12c, Kei12d, Kei13a, Kei13b, Kei13c, Kel05, Kel06a, Kel06b, Kel06c, Kel06d, Kel07a, Kel07b, Kel07c, Kel07d, Kel08a, Kel08b, Kel08c, Kel08d, Kel09c, Kel09d, Kel09e, Kel09f, Kel10a, Kel10b, Kel11a, Kel11b, Kel11c, Kel11d, Kel12a, Kel12b, Kel12c, Kel12d, Kel13d, Kel13a, Kel13b, Kel13c, Kel14a, Kel14b, Kel14c, Kel22c, Kho05, Khu01, Khu02, Khu05b, Khu06, Khu12, KM10, KR03, Lam14, Len08, Mat97, Mos12, O’R88, O’R89b, O’R89c]. column [O’R89a, O’R90, O’R93, O’R94b, O’R95b, O’R95c, O’R96a, O’R97b, O’R97c, O’R98, O’R99b, O’R00a, O’R00b, O’R00c, O’R01, O’R02, O’R03, O’R04b, O’R04c, O’R04d, O’R06a, O’R06b, O’R06c, O’R07, O’R08a, O’R08b, O’R12, Par89b, Par90a, Par90b, Par98b, Puc05a, Puc05f, Puc05g, Puc06a, Puc06c, Puc07a, Puc10, Raj01a, Raj01b, Raj01c, Raj02c, Raj02d, Raj02e, Raj03a, Raj03b, Raj03c, Raj04a, Raj04b, Raj04d, Raj05a, Raj05b, Raj06a, Raj06b, Raj06c, Raj06d, Rie99a, Ros03a, Ros04d, Ros04e, Ros05c, Sin98, SP07, Su09, Sud11, Tre09, Vad07, Van10, Van06, Ven13a, Via05, Via08a, Wat11, Wei93, Wel13, Wel14a, Wel14b, Wel14c, Wil11, dB06, vS12a, vS12b, vS13, vS14a, vS14b, Bra95, Gas12f, Hag00, Hem95, Hem96, Hem10a, HW12]. Column [Kei08a, Kei10b, Kei10c, O’R96b, O’R97a, Raj02b, Raj05d, Vio09]. Column1 [Gre20d]. Columns [Hem23b]. combinatoires [Knu76b, Lew78]. Combinatorial [Bec08, BQ03, Bón10a, BC09, CH15, Cor01, Gas05b, Gas16d, Gre19g, Her13, Hro03, Knu97, KV00, KV02, Kor06, KV08, KV12, KS99, Mas03, McN99, Sti08, Upa16, Val09, Wel15a, GL80, Li14, Knu11, Bor08, Gas12e, Iva02a, Kum13, Lab12]. Combinatorics [BS08, BC^{+}91, Bón08a, Bón09a, Bón13b, BMMR20, Dav08a, FS09, GGL95, Har11, HHM00, HHM08, Ios07, Kri99, KRY09, Lot05, Mar98, Maz14, Maz10, Rob84, RT05, SD08, Vio19, Wan97, Bón14, Cha08, Loe11, Tre09, Bón10b, Bón11a, Bón13a, Gas99b, Gas15b, Ham13, Lee03b, Mie11, Pap11, FB14, Tew14, Abr11, Bón09b, Cad11, Mou10]. Combinatory [Sta84a, Sta84b]. combining [AAH11]. Come [Ros96c]. comic [Gas10b]. command [OB83]. comment [Ben73b]. Comments [Gre81, Lei91, Ric73]. Common [Gol06, Kei10a]. Communicating [Cal12, Dav06a, Mil99, Mil09, Pap10, GL84a, GL84b, Hut01, Puc00]. Communication [Ali21a, Big08, CKN21, Gas13f, Gup02, HH24, Low12, Vio19, dRGR22, FKW94, Hac09, KLMY21, San84, TS99, Cal13, Cad21, Lib17]. Communications [Sal05a, AF14, Tal09, Wer10]. community [Gol92, Gas22f, Gas22d]. Commutative [Rup91]. Compact [AKR83, Koz18]. Companion [Azi10b, GBGL08, Kel09a, Lei07, Cen02, HO01, Kat14b, Gas14e]. Company [Sub78]. Comparative [Kat05]. Comparison [vdBN04]. compendium [HW94]. competition [Gol12b]. Competitive [BFM^{+}15, BEY05, Wel15c, You99]. Competitiveness [Chr07, CKM06, IMP11]. compiler [Lon72]. complementary [CC78]. complete [Aar05, Fra85, Gal74, GJ78, Gol77, Kri77, Lad75, McB91, Raj05d, Sch81, Sto73a, Sto73b]. completeness [Eps87, Gol10a, Nag13b]. Completions [Dun00]. Complex [AK90, VR07, Fer05, Kin11]. Complexity [Apo10, AB09a, Bau17, Ben23, BCSS98, BMMR20, BCS97, Cad21, Che11b, DF99, DH10, DK00, FFG96, FS19, FS15, Gas97a, Gas07h, Gas13f, Gas13e, Gas16a, GHOP22, GL80, Gol08b, Gre21f, Gre22g, GW18, HH24, Hem94a, Hem95, Hem96, HS97, Hem01a, Hem10a, HW12, Hem14e, Hem15a, Hem15b, Hem15c, Hem15d, Hem16a, Hem16b, Hem16c, Hem17a, Hem17b, Hem17c, Hem18c, Hem18a, Hem18b, Hem19d, Hem19a, Hem19b, Hem19c, Hem20a, Hem20b, Hem20c, Hem20d, Hem21d, Hem21a, Hem21b, Hem21c, Hem22a, Hem22b, Hem22d, Hem23a, Hem23b, LV97, LV08, OS14, Oli06, PV21, Raz16, RSZ21, RST15, Rot05, Sha19, Shp03, Str94, TW06, To06, Vio19, Vol23a, Vol23b, Vol24, Wei15, Wel17c, Wil18, Zim04]. complexity [Afs82, AFG^{+}96, ASU86, BA04, BCRV03, Boo84a, Boo84b, BI08, Cen02, Cha71, Cha73b, Che06, Eng89, FKW94, GC86, Gol08a, Gol10a, Gra82, GLMW96, Gup02, HW94, Hav71, Hem97b, Hem98, HS98, HO01, Hem01b, Hem01c, Hem01d, Hem02a, Hem02b, Hem02d, Hem02c, Hem03b, Hem03c, Hem03d, Hem04, Hem05b, Hem05c, Hem06c, Hem06d, Hem06e, Hem08b, Hem08c, Hem09a, Hem09b, Hem09c, Hem09d, Hem10b, Hem11a, Hem11b, Hem11c, Hem11d, Hem12a, Hem12b, Hem12c, Hem13c, Hem13a, Hem13b, Hem14b, Hem14c, Hem14d, Hem22c, Hof00, Kei08d, Knu77a, Knu84b, Knu84a, KGO10, KL05, Kul13, LV93, O’R06c, PGW^{+}96, Reg93, San84, Sha02, Sha13, TS99, Via05, Vol99, Wil11, YZ91, Bir91, Gas23a]. Complexity [Hem94b, Lon97, Lon98, Cen03, Chi15, Gas94, Ist00, Mil13, Teu13, All98, Fal07, Gre22f, McN01a, Rot07]. complexity-theoretic [BA04]. composition [Can06, Puc06b]. compound [Sug93]. Comprehensive [MMW06, LaT11]. Compression [ABM08, Neu10, Sal07, Wei15, Dav93, Hau12, Nas13]. Computability [Ali21b, BY09, Cen02, Gal21, Gas13f, Gas16d, Gas16g, dOG12, Hem20c, HRR12, Pet08, Pip97, Ray21, Esc93, Fer09, Kan05, Kei12c, Gup98, Cal10, Sha13]. Computable [McN04, Par02b, LM73, Mil72b, Nou80]. Computation [Amo05, Ata99, AB09b, AB10, Bal97, BCSS98, BEY05, BM07, CC76, CK02, Fur11, Gas99a, GM10, GW18, dOG12, HD09, Her13, HU79, HMU01b, HMU07, Hua98, Irw08, KSV02, Koz06, KM10, Pap12b, Ree77, Sav74, Sim00, Tay04, You99, All97, CPV04, Coo12, Cul82, FH06, Far19, FM06, Fer09, Goo93, Gop08, Hac09, HK84a, HK84b, Iva02b, KS10, KLMY21, Lee89, Lee90a, Lee90b, Lev91, Mac95, Par86, Raj04c, Raj06e, XL93, LP98, Sip96, dW07, Apo11, Apo12, Azi16, Dav08c, Dav09, Gas20f, Gre16h, Gre21e, HMU01a, Mai80, McN01a, Mos05, Sga12]. Computational [ÁDFT15, AO98, ABCC06, AB09a, BKM18, Ben98, Car15, Cen03, Cha71, Chr04b, Coh93a, Coh95, Coh96, DO99, DK00, Dum12, DT12, DJ13a, DJ13b, DT14, DJ14, Dum16, Dum18, DJ19, ED17, FM15, Gas07h, GFW17, GG21, Gol08a, Gol08b, Gus97a, Gus97b, Kat06a, KV95, Kel08e, KL16, Lab12, Mar03b, MO01, Nag17b, Nag18b, O’R88, O’R89b, O’R89c, O’R89a, O’R90, O’R92b, O’R93, O’R94a, O’R94b, O’R94c, O’R94d, O’R95a, O’R95b, O’R95c, O’R96a, O’R96b, O’R97a, O’R97b, O’R97c, O’R98, O’R99b, O’R00a, O’R00b, O’R00c, O’R01, O’R02, O’R03, O’R04b, O’R04c, O’R06a, O’R06b, O’R07, O’R08a, O’R08b, O’R12, PW01, PW09, Raj04d, Raj06c, Raj06d, Rie99a, Sha17, Sho05, Sho09, Sof03, Spr09, Val09, Ven13a, Zhu20, Zim04, dB^{+}97, dB^{+}00]. Computational [dB^{+}08, AFG^{+}96, Cha73b, FM14, Gol10a, Gra82, JKL99, PGW^{+}96, Pap13b, Pat72, Sha02, Vio17, Apo10, Hem94b, Apo10, Bor07a, Dek99b, Ess05, Kre07, Mas00a, Oli06]. Computations [Gre22g, BFK^{+}79, CM72, Pud06]. Compute [Mor85]. Computed [Gas20f, AF14]. Computer [Ayc06, Baa78, Baa88, BV00, Ben98, Bón13b, Bra77, BZ11, Cha73a, Che79b, Coo04, Cou09a, CH10, Dav11, DFM07, FM08, Gas13f, Gas20d, Gas21d, GKP88, Gus97a, Gus97b, HH76, Jan08a, Kam13, Khu98a, Kil14b, Knu96, Knu11, Kor93, Kor02, Kri00, LaT11, Lea00, Mac20, MW93, Mar09a, MMW06, Mer07, Nag19, Nie02, Par88, Par00b, Pos99, Pro90, Puc05c, Rie10, Rog08, Rog13, Rog18, Ros93a, Ros93b, Ros93c, Sal05a, Sal05b, Sch89, SP98, Sch08a, Tal09, Ter11, TM11, Vin09, Vla94, Wer10, Yan12, YM08, AJK^{+}97, ABL^{+}95, Cul82, Dew86, HR02, Han94, Hes03, Joh84, Kil14a, KW07, Lan82, Lee89, Lee90a, Lee90b, Li12, Man82a, Par89a, Par94, Par02a, Ros00d, Sei96, Sei01a, SY87, Sta76a, Tay05, Tay02, Tho93, Toy93, Tre09, Man82b, Ros97b, Sim84a, Sim84b]. Computer [Zeu89, Con03, San10]. Computers [Amo99, LB99, OR99a, Ogi01, RW99, Ski01, WG00, Kei08b, Par00c, Yli05]. Computing [Agu12, AF20, AK20, Ali19, Ali20a, Ali20b, Ali20c, Ali20d, Ali21d, Ali21a, Ali21b, Ali21c, Ali22, Ali23a, Ali23c, Ali23b, Ano13, AW98, AW04, Bra94a, BT20, CCD15, COP01, Che24, CC76, Cou09a, Dwo95, FGK^{+}02, Gol00b, HS01, Har15, Hir01, Irw03b, Kat07, Kei08a, Kei08b, KS01, Kuz18, Kuz19, Lev97, Mar03a, MU05, Pap05a, Pit00, Puc01a, Raj02b, RR20, Sga13, Sri18, Sta07, Upa16, Wan97, Wel14d, Wel15a, Wel15b, Wel15c, Wel15d, Wel16a, Wel16b, Wel16c, Wel16d, Wel17a, Wel17b, Wel17c, Wel17d, Wel18a, Wel18b, Wel18c, Wel18d, Wel19a, Wel19b, Wel19c, Yan02, YM08, Zal97a, Zom06, dW07, AAH11, BBAGP09, Bau10, BCvR09, BR98, Bor09b, Bra94b, BT08, Cas02, CBS07, Chr04a, Cle08, CFF^{+}97, DP08, Dwo89, Dwo90, Dwo96, Fic97, Geo09]. computing [GW97, Gre13, Hem99, Her83, HL08, HRR12, KS06a, Kei08c, Kei08e, Kei09a, Kei09b, Kei09c, Kei09d, Kei10a, Kei10d, Kei11a, Kei11b, Kei11c, Kei11d, Kei12a, Kei12b, Kei12c, Kei12d, Kei13a, Kei13b, Kei13c, Lev96, Lon71, Lou91, LP92, Pel13, Puc05a, Puc06a, Raj00, Raj01a, Raj01b, Raj01c, Raj02c, Raj02d, Raj02e, Raj03a, Raj03b, Raj03c, Raj04a, Raj04b, Raj05a, Raj05b, Raj06a, Raj06b, SP07, Slo02a, Slo02b, SIG83c, Wel13, Wel14a, Wel14b, Wel14c, Zal97b, Zal99, ACM12, Ano13, CP74, Kei10b, Kei10c, Lat95a, Lat95b, Lat95c, Raj01d, Ric00, Zal97c, dW07, Her00a, Kil12, Ree77, Ste11b, Vio04, Amo20, Lou89, Lou90, Lou95a, Lou95b, Lou95c, Lou96a, Lou96b, Lou97, Sar03, Sei97a, Sei97b, Sei97c, Sei98a, Sei98b, Sei98c, Sei98d, Sei99a, Sei99b]. Computing [Sei00a, Sei00b, Sei00c, Sei01b, Sei01c, Sei02a, Sei02b, Sei02c]. comun [Hem02d]. Concentration [DP09, Sri10]. concept [FKR08, Via08b]. Concepts [AB09b, Jen97, Mik93, Kra07]. Conceptual [Apo10, Gol08b, Gol08a]. Concerning [GKKK06, BMO77, Fla05, Yli04]. Concise [Hau12, Par08]. Concrete [GKP89, Cul82, GKP88]. concurrency [Kei09a, Lam09, MM13]. Concurrent [AF20, Ali20b, AO91, AO97, AdBO09, Aro98, Hen09, Kuz18, MP92, Old91, RR20, RG06, Sch00, Sis93, Tau06, Wel18a, Kei12b, Pet04, Upa10]. condition [vL76]. Conditions [Gas06e, Kas05]. Condliffe [Gas16h]. Condon [Ste12]. Conference [AF20, LRV10, Wan97, Dol07, Sim85, Sta76a, Cha73a]. Conferences [Ali23b, FdV23]. confidentiality [Puc04b]. Configurations [Gre20e]. Confinement [VS98]. Conflict [GH09, Wil13]. congestion [CK04, Puc04a]. Congruences [Coh08]. Conitzer [Nag17b]. conjecture [AAV13, CT23, GL02, LS74]. Conjectures [Mos19, Hem94a]. Conjugate [Wie83]. Connect [Lou18a]. connection [vL74]. Connections [Cou15]. Connectivity [Gol00a, PD98, GI91]. conquer [BHS80, Kao97]. Consensus [Ali20a, BCN20, TV16, Wel16c, KR01]. consequences [GH00]. Consistency [Ali22, AMN22, KSRS14, Wel14a]. consistent [GL02]. Consortium [Ham90]. Constable [Che80]. Constance [Ban14a, Bel11]. Constanda [Gas11a]. Constant [HH24, LST22]. Constant-Cost [HH24]. Constant-Depth [LST22]. Constraint [FA03, Oli04a, BCRV04, Kan94, Lib99]. construct [Pys76]. Constructing [Cha20, Lei80]. construction [Eng89, Mos12, Ten87]. Constructions [HL10, Sti08, Kat12]. Constructive [Cou10, Edw05, Swe87]. constructors [Nou80]. constructs [GS80]. consumption [Lon71]. contest [SR03].

Context [Fas83, Sch18, Spe93, Col93, Ell70, Fas86, GS80, Hut70]. Context-Free [Fas83, Spe93, Col93, Fas86, Hut70]. Continuations [Thi99, dRCR94]. Contour [McG72a]. contracts [Eve83]. Contributions [ABL^{+}95, Hem14a]. Control [Lib17, MM13, Raj02a]. controversies [Gol92]. Conundrum [Ros95d]. Convergence [Has85a]. conversion [Kug76a]. converting [Gal75]. convex [Dwy90]. convincing [Bra91]. Convolution [Smo21]. Cook [Mil13, Spr09, Azi12a]. Cooperation [GH09, Bau10, Wil13]. Coordination [CK04, Puc04a]. Copernicus [GKKK06]. Cormack [Rub13]. Cornuéjols [Bor08]. correct [Jon72, Sha10b, Bru13]. corrected [Rob70]. Correcting [KS16, She97, Gur04, O’R04d]. Correction [Kos74a, McG72b]. Corrections [CR74]. Correctness [Bel07, Lon72, Sta10, Ash72, HES72, Weg74]. correlation [Vio09]. Correspondence [Irw08, Sim00]. Corrigendum [Knu84b]. Cosmo [Lea97]. Cost [Ali23b, FdV23, HH24, KR01]. Costs [GW18]. Count [BQ03, Con09, Gas05b, Gas11a, All97]. countability [Swe87]. counter [Pet94]. counterexample [Eng89]. Counting [Gas23a, Gre22f, BCG^{+}22, Che11b, DLM07, HV95]. Courcelle [Cad14]. Course [Aig07, Bon08b, Boo10, Cha98, Coh93a, Coh95, Coh96, Cou11, GG21, Kel08e, Pol09, Ros93c, Sch09, Sch06, Sha09, Sta89, Wil10b, Sav98, SP07, Gre17h]. courses [Toy93, Ros93d]. cover [Khu02]. Covering [Cor01, Bor08]. COVID [Hem20c]. COVID-19 [Hem20c]. cow [Chr04d, Ros04c]. Cowell [Ree77]. Cozzens [Led16]. CRA [Ber92]. Cracker [FPE04]. cracking [FPE04]. Crama [Azi14, Kam13]. Crampton [Nat07, Nat10]. CRC [Ben06, Cha09, Ful10, Gas04a, Kat05, Kel09a, Spr03, Spr06, Spr08, Was08a]. Created [Gas21d]. Creative [Wal63, Wal09, Mit01a, Dye11]. Creators [Soi09, Gas09c]. creatures [Agu04, Khu04]. crisis [Lew14]. Cristopher [Azi16, Gre16h]. Criteria [Lev95]. Criterion [Ali22, AMN22]. Critique [Gol92, Har86]. Crochemore [Ada11, Gal15, Mar15b]. Crossing [GHOP22]. crossovers [Dew79]. Crusade [Bra85]. Cryptanalysis [Bar09, Hin10, Jou09, Wag03, BHT97, Kat05, AY12, Fit11, Roj12]. Cryptanalytic [Yan07, HKR83, Ira12]. cryptarithms [Eps87]. Crypto [Hom17]. Cryptoclub [Ayu12, BP06]. Cryptocomplexity [Rot05, Fal07]. Cryptographic [Bla12, Dif82, Shp03, Can06, Puc06b, Sha10a]. Cryptography [Bar12, Big08, Bra94a, Buc01, Buc04, CFA^{+}06, DK02, DK07, Elb09, Gol01, Gre19e, HHL00, HPS08, Irw03a, Kat05, KL08, Lee03b, Mol03, Nie02, Puc03, Rot07, Sin99, Sti95, Sti02, Sti06, TW06, Was08b, WK01, Wel05, BB89, BHHR99, Bra94b, BC96, Gol99, Kra07, KGO10, Mou14, Wie87, Kat05, Cou12, Ful10, Gas13f, Lee03a, Low12, Mei10, Mur10, Ree01, Ter11, Was08a, Kat05, Spr03, Ter08]. Cryptological [Lew00, Gas05a]. Cryptology [Bra90a, Bra89b, Bra89a, Bra90b, Bra94a, Bra95, Fal07, Nag19, Rot05]. cryptosystem [Bra83, Jor83]. CS&E [Bay95]. CSE [Sta89]. CSE/OR [Sta89]. CSLI [Oli04c]. CSNET [Lan82]. CSP [Pet04, Sch00]. cstheory.stackexchange.com [CEG^{+}10]. cube [Gar08, Gas09h]. Cucker [McN01a]. Cult [Gas15i]. Culture [Bla12]. Cultures [Sha92]. Cunial [Kel16a]. Cunningham [Gre17h]. Curiosities [Ste09, Gas09h]. Current [Ros95b]. Currently [Gas06e]. Curriculum [Ros93b]. Curry [Irw08, Sim00]. Curve [CFA^{+}06, Dey07, Was08a, Sot10]. Curves [Gas10d, Gre19e, Kob08, Nag19, Was08b, Che10]. cut [Sub95]. Cuzzolin [Gal22]. CV [COJ78]. Cvetković [Bón10a]. Cyber [Cho12]. Cyberspace [Cra03a, Cra03b, Nat07, Nat10]. cycle [Dun86a]. cycles [ER85].

D [Apo11, Azi12b, Azn14, Cao13, Cen11, Che79a, Che82, FM11, Kin15, Mai80, Mat14a, Nag17b, Oli04c, Puc01a, Rau08, Ree77, Rie93, Spr08]. D.R [Irw03a]. Dagstuhl [FRRS17, KSRS14, Wel14a, Wel17b]. Dahan [Nag13a]. Dahan-Dalmedico [Nag13a]. Dale [Bel14]. Dalmedico [Nag13a]. Dan [Ben98, Kel16b, Kum12, Mie11]. Dana [Mos05]. Dangerous [Bla06, Fas04, Gas15d]. Daniel [Abr11, Azi11, FM08, Gre17h, Gre22h, JH20, Lee10, Mik11, Nag15b, Smi09]. Danny [Rou05]. Danupon [Ano13]. Darel [Xie12]. Dasgupta [Kel09b]. Data [ABM08, Bar12, Bra08, EA10, Elb09, Elr11, FGK^{+}02, FS19, GT97, GT01, GT04, GT06, GT10, Hau12, Jan10a, Koz18, Mah19, MS08, Mie11, Mil17a, Mil17b, Mil18a, Neu10, RZ09, Ros93c, Sal03, Sal05a, Sal07, SD08, Sto02, Tal09, Wei98, Wel19b, Wer10, EKP79, GN07, HR00, KM10, LP92, Nas13, Raj07b, Suc98, GJM02, Wei95e, Mil17b, Dav05, Jan10b, Mas01a, Pap05b]. Database [AK90, Are18, Chr08a, Gee17, Gro14, KA89b, Kan94, Koc19, Len08, Ngo23, Pap95, Par98b, Pic21, Su09, Tan16, Tao20, Van10, Van06, Via08a, dB06, KA89a, Yan96]. Databases [AK90, GMN87, Har86, Lib99, Suc08]. d’autre [Lew78]. d’autres [Knu76b]. Davey [Coh07]. David [Amo99, Bla06, Cad11, Cen15, Cha12, Che80, Cou15, Dav05, Dav11, Gal14, Gas01b, GKKK06, Gas22e, Gre20e, Hau12, Jan08a, Jan10b, Kat13b, Mas00b, Mas01b, Nag21a, Nas13, Ogi01, Pap05b, Rie10, Tal09, Wil13]. Davis [Kil14a, Kil14b]. dawn [BB89]. day [BBAGP09]. Daylight [Gas15c, Gas15c]. Deadlock [GL84a, GL84b]. Death [yea00b]. Debate [Gas10e, Gas10e]. Deborah [Kob98]. decades [HS12]. Decidability [BJK^{+}16, Sti99, Wel16b, Mil72a]. Deciding [KA89a, KA89b]. Decision [BGG97, BM07, COP01, Che82, Fur11, Irw03b, Joh10, Pat72, Sim04, Weg00, BCG^{+}22, KLMY21, Kul13]. Decisions [Saa01, Pri10]. Declarative [COP01, Irw03b]. Decoding [KS16, Sud00]. decomposable [Sch82]. Decomposition [Aga91, O’R92a, TF82, Mor76]. Deconstructing [Ros03b]. Deductive [MW93, Vla94]. Deep [Cou11, Pol09]. defense [Par86]. defenses [Yu11]. Defining [Sch18]. Definition [Cha70, FM05, Ash72, Ben73b, Lee72, Rob70]. definitional [HV95]. definitions [All72, CM72, KGO10]. Degree [BT20, Cud69]. degrees [Ros98c, Tho93]. Dekker [Irw03a]. Delayed [JM07]. Delfs [Mur10]. Demaine [FM11, Apo11]. demands [Che07]. Democritus [Gre13]. Denis [Gas16d]. Density [Gre19f]. denumerable [Fer05]. department [Che77]. dependence [Lon71]. Depth [LST22, BGH07]. derandomization [Sha10b]. Derivation [All72, Sim00, Irw08]. Derivatives [Mor85, dJvL75]. description [Bur72]. Descriptional [BMMR20]. Descriptions [TZ16]. Design [Baa78, Baa88, BV00, Che21, Dav06a, Di 95, Fah04, Gol00a, GT02, JK92, Kar11, Lev07, Lev12, LR97, Mav09, Nag20, Náv04, PD98, Ski98, Ski08, Tra22, Tse91, Zeu89, Afs82, Apo14a, Che07, Hut01, Kei10b, Pru98, WS11, Kel16a, Lea97, Azn14, Cha12, Mav09, Pap12a]. Designing [Ali21c, Har21]. designs [FC83, Sti08]. Deszö [Ist12]. Details [All23]. detection [GL84a, GL84b]. detectors [Chr05b, Ray05, Rey05]. determining [Kos97]. Deterministic [Gab10, Zim13]. Devadas [Nag23]. Devadoss [FM14]. Devdatt [Sri10]. Developing [DLO09]. Development [LR76, SS07, Jon72, Li12, Her77]. developments [Gas13c, Ven13a]. Dexter [Apo12]. Dey [Sot10]. Diagonalization [HS18]. diagram [AS91, Cud69]. Diagrams [Joh10, Weg00, Zhu20]. Dialogue [GKKK06, Fla05]. Dichotomies [Gre22f, Che11b]. dictionaries [GJM02]. did [Hil07]. Dieter [Dav11, Fah05, Lam13]. Dietzfelbinger [Kat06b, Kin15]. Difference [EA09, Ela96, Ela99, Ela05]. Difficulties [Mah19]. Diffie [Jan08b, Mak99]. Digital [ALL08, Gas09a, Gas21c, Kug76a, Che13, Mou15]. digraphs [Sug93]. Dijkstra [Ano13, Mos10, Sch18, Alp18, SS04a, SS04b]. Dilemmas [Pin10, GS08]. DIMACS [AP98, Dav05, Gol00a, GJM02, LB99, PD98, RW99, WG00, Fei89, Fei90, Ven02]. dimensions [vS15]. Dimitris [Oli04b]. D’Informatique [GP89]. Ding [Cen03]. Ding-Zhu [Cen03]. Diophantine [Gre22a]. direct [Gal74]. Directed [TV16, Wel16c]. direction [San84]. directions [CFF^{+}97]. director [Gas07a, Slo07]. DISC [Agu12, AG15, BK18, Bon10c, Bor07b, DGW08, Dol07, FdV23, FM17, HM11, PT14, Su16, Tie09, Tim13, Göö12]. DISC’00 [Hig01]. DISC’01 [Her02b]. DISC’09 [BBAGP09]. disclosure [JM07]. Discovered [GW18, Nah04]. Discovery [Lak76, DL79]. Discrete [AP98, BMP05, Eri10, Gas07d, Gas17d, Gas20d, Gre20e, Knu03b, Mas00b, Oli04c, Rau08, RFR97, Ros23, Sza12, Ven02, Wal03, Wal12, Xie12, Abr13, Apo14b, HWR09, Her83, FM14, Kob98]. diseases [KL10]. Disjoint [GHSW14]. disjointness [CP10]. disjunctive [BBF^{+}73]. disposition [DH79a]. dispositions [DH79b]. Dissertation [Ano13]. distance [Gas14h, Yuv76]. distances [Ven13b]. Distributed [ACM12, AAH11, AF20, AK20, Ali19, Ali20a, Ali20b, Ali20c, Ali20d, Ali21d, Ali21a, Ali21b, Ali21c, Ali22, AMN22, Ali23a, Ali23c, Ali23b, AT17, Ano13, AKGN18, AW98, AW04, APR16, Bar15, Cas02, CH15, Chr04c, Cle08, DP08, Dwo89, Dwo90, Dwo95, Dwo96, Elk04, Gav01, Har15, HL08, Kei08a, Kei08b, Kei09d, Kei10a, Kei10b, Kei10c, Kei10d, Kei11a, Kei11b, Kei11c, Kei11d, Kei12a, Kei12b, Kei12c, Kei12d, Kei13a, Kei13b, Kei13c, Kuz19, Mav02, Raj01d, Ray05, Ray21, Tel00, Wel13, Wel14a, Wel14b, Wel14c, Wel14d, Wel15a, Wel15b, Wel15c, Wel15d, Wel16a, Wel16b, Wel16c, Wel16d, Wel17a, Wel17b, Wel17c, Wel17d, Wel18a, Wel18b, Wel18c, Wel18d, Wel19a, Wel19b, Wel19c, BS09, BP05, BK12, BT08, CBS07, Chr04a, Chr05b, Chr06b, DO13, FKR08]. distributed [Geo09, HRR12, Kei08c, Kei08e, Kei09a, Kei09b, Kei09c, KSRS14, Pel13, Puc05a, Puc06a, Raj00, Raj01a, Raj01b, Raj01c, Raj02c, Raj02d, Raj02e, Raj03a, Raj03b, Raj03c, Raj04a, Raj04b, Raj05a, Raj05b, Raj06a, Raj06b, Rey05, SP07, Ver06, Via08b, Zho09, Agu12, Raj02b, Sch06, Her00a, Upa16, dV16]. Distributing [Küp13, Kei13b]. distribution [BB87]. distributions [Fra85, Ven13b]. Diversions [Gar08, Hay08, Gas09h]. Divide [BHS80, Kao97]. Divide-and-conquer [BHS80, Kao97]. Division [Dan06, Gas00f, Mou03]. divisions [dJvL75]. Djamal [Kel16a]. Djeraba [Mie11]. DLBA [Iba91]. Dmitry [Upa16]. DNA [Dav08c, Dav09, WG00, Amo05, KL10, LB99, Par00c, RW99, WG00, Amo99, OR99a, Ogi01]. DNA-based [Par00c]. do [Cas13, Dos87a, Joh96b]. Doctoral [Ano13]. Documents [Bla12]. does [Gur86]. dogmas [CBS07]. Domain [Gre22f]. domains [Hes02, KGO10]. Domenico [Irw03b]. Dominic [Rot07]. Dominich [Gri10]. Donald [Apo14b, Apo14a, Gas14e, Gas14g, Gas15c, Khu98a, Lew78, McN99, McN01b, Mik11, Neu10, Pri10, Rog08, Rog13, Rog18]. Donna [Gas10b]. Dorit [BS98]. Doron [Kre00]. Dorst [FM08]. Dots [Ber00, She13]. Dots-and-Boxes [Ber00]. doubling [CKM06]. Douglas [Spr03]. Dovetailing [Pys76, Ham76]. Dowek [Cad13, Nag21b]. Downey [Chi15, Ist00, Teu13]. downward [HHH98]. Doxiadis [Gas10b]. dozen [Jaw05, Puc05h]. Dragos [Bón10a]. Drawing [Sug93, TT95, Bat93, Coh93b, Ead93]. drawings [Kan93, Kau93, Pac93, Tam93, dFR93]. Dreben [Che82]. Dress [FB14]. driven [Kan05, OS14]. Droste [Cad12]. drug [Azn14]. DSN’08 [Kli08]. Dual [Hua14, Chr07]. Duality [Fra84a, Fra84b]. Duan [Mel11]. Dubhashi [Sri10]. Dude [Con09, Gas11a]. Dudley [dOG10, Yan11a]. due [FdV23]. Dunlap [Cad16]. Durrett [Bón11b]. duties [Gol09]. Duval [Mor03]. Dwyer [Wan05]. Dynalab [Ros94b]. Dynamic [APR16, Che80, Coh93b, Har79, HKT00, HKT01, KO11, Li08, Wel16a, BS09, Kei11a, Pru98, vdBN04, Puc01a, Yan11b]. Dynamics [Ali20a, BCN20]. dystrophy [KL10].

E.-R [Pur93]. Earley [Jon72]. earth [Ven13b]. easy [Gre81, HHH98]. easy-hard [HHH98]. EATCS [Mar09a]. Economics [Ros97b]. ed [Mas00a, Mas01a, Tra22, Zeu89]. Edelsbrunner [GG21]. Edgar [Gas15c]. edge [GI91, HK84a, HK84b]. edge-coloring [HK84a, HK84b]. Edited [All98, BS98, Gas15f, Gas20c, Gas21d, Gre18e, Kob98, Nag17b, Nag18b, Abr13, Amo99, Cad12, Cha09, Gas99a, Gas99b, Gas04a, Gas07b, Gas10a, Gas13c, Kri99, Lab13, Lev10, Mas00b, Nar14, Neh99, OR99a, Pap13a, Pin10, Puc01b, Sga12, Sol12, Ste12, Wan12, Maz14]. editing [HS98]. Edition [Dek99b, Dek99a, Gal21, Gas97a, Gas16c, Nag17a, dV16, Avi01, Bón09b, Bón09c, Bón11b, Bón11c, Bón13b, Coh08, EA09, Fah05, Ham13, Irw03a, Kar11, Kei09d, Lab13, Nag14, Smi00, Spr03, Gre19f, HMU01a]. editor [dOG10, Par00a]. Editors [Joh96a]. Edmonds [Sga09]. Eds [Nag16]. Edsger [Sch18, Alp18]. Education [Che11a, Che13, Rog08, Ros92a, Ros92b, Ros92c, Ros93b, Ros93c, Ros93d, Ros94b, Ros94a, Ros95b, Ros95c, Ros95d, Ros95a, Ros96b, Ros96a, Ros96c, Ros96d, Ros97a, Ros97b, Ros97c, Ros97d, Ros98a, Ros98b, Ros99c, Ros99a, Ros99b, Ros00a, Ros00b, Ros01a, Ros01b, Ros03c, Ros03d, Ros05a, Ros05b, Lew06, Ros00d, Gas07c, Gas16h]. Edward [Gas10e, Gas15e, Gre07, Mos05]. Edwards [Cou10, FM09]. effect [Mof83]. Efficient [HL10, Kat12]. efficiently [Fra84a, Fra84b]. efforts [Ros00d]. Egon [Sim04]. Ehrhard [Gas09h, Nag16]. El-Yaniv [You99]. Elastic [Ali22, AMN22]. Elaydi [EA09]. Elbirt [Bar12]. Elche [BS09]. Eleanor [Sga13]. Elections [HK05, Pri10, Saa01, Wil10d]. Electronic [Bla12]. elegant [Sei01a]. Elementary [Cou11, Dud09, Led16, Pol09, Yan11a]. Elements [LP98, Sak11, Nis78, Cen15, Amo20, Kin12]. Eleusis [Gas09h, Gar08]. Eli [Bla06, Kat07, Sta07]. Elimination [Ali20c, Suo20, Chi86]. Ellen [Gas16h]. Elliott [Spr08]. Elliptic [CFA^{+}06, Gre19e, Was08a, Was08b, Che10]. Elsevier [Oli06]. elusive [Lam14]. Elwyn [Nag17a, She13]. embedding [Tur84a, Tur84b]. Emerging [AJK^{+}97, Zom06]. Emile [Mar09a]. Emissions [FdV23]. Empire [Vuk10]. Enablers [FS19]. encodings [Gal74]. Encryption [DL98, DL07, Led16, Jan08b, Kat13a]. encryptors [Mor83]. End [Bra94a, Bra94b, Sta87]. Endriss [Nag17b, Nag18b]. Engelfriet [Cad14]. Engineering [AL07, BL06, Bor09a, CH10, Bat93, Gop08, Har95, Mor97, Ros97b, Gre07, Kam13]. Engineers [Mar20]. Engines [BCC10, Gas22f]. England [Ben98]. enhance [Vis91]. Enhanced [Ros98b]. Entertaining [Gas23a]. Enumeration [Aig07, KS99, Boo10]. Enumerative [Cha08, Ham13]. Environmental [Ali23b, FdV23]. Epic [DPP09]. epistemic [LP07]. Eppstein [Gre20e]. equality [Li04]. Equation [Cou09b, Mor03, Mor05, Hen83]. Equational [Cal12, Puc05e]. Equations [Ela96, Ela99, Ela05, Gas12a, Gre22a, Ses09, She14, EA09]. equilibria [CPV04, Raj04c]. Equivalence [Ali21d, BB21, Ber72, Sti99]. equivalent [Koz78, Ten87]. era [BB89, Kel16a]. erasing [LS78]. Erasures [GW18]. Erdos [Gas14h, Par96, Ros23, CEG98, Par11]. Eric [OR99a, Ogi01]. Erich [Sim04]. Erickson [Sza12]. Erik [Apo11, Gre22h]. Ernest [Mak93]. Ernst [Aro98, Kel09a]. Ernst-Rüdiger [Aro98]. Errata [Lev97]. Error [KS16, She97, Din08, Gur04, McG72b, O’R04d]. Error-Correcting [KS16, Gur04, O’R04d]. escapes [Cen09]. Essays [Edw05, Fre01, Gas22f, Gas22d, PSTM00, Puc01b, Gas07b, Cou10]. Essential [Gas15c, Gas20d]. Essentials [FA03, HHL00, Irw03a, Oli04a]. estimating [YDX97]. estimation [FL83]. Etude [SS04b, SS04a]. Euclid [Cen15]. Eugene [Ric94]. Eugenia [Gre17e]. Eugenio [Irw03b]. Eulogy [Raz07]. Europe [Gur94]. Eva [Lev10]. evaluate [dJvL75]. Evaluating [BCC10, vL78]. Evangelos [Tay04]. Even [Lab13, Lab13]. Events [Paz70]. Everyday [Fis08, Gas09b]. Everything [Par98a]. Everywhere [Nag16]. Evidence [Bla12]. Evolution [Ree01, Sin99]. Exact [Lam13]. Exam [Ada09a, Gas10e]. example [Jon72, Ros95b]. Examples [Dur10, GT02, Bón11b, Náv04]. Excellence [Lew06, Gas07c]. exciting [Raj00]. exclusion [Ala03, Bur78]. exercises [JP00]. exhaustive [HKR83]. expanded [Irw03a]. expander [Gur04, O’R04d]. expansion [Yeh12]. Experimental [Amo05, BB89, Dav08c, Dav09, Mar15a, FC83, Hem99, MM99, Mit01b]. Experiments [Ali21c, Dew86, Har21, Hac09, Joh96b]. expert [FL83]. explain [KL10]. Explaining [Pri10, Saa01]. Explanation [Kre01]. Explicit [Kel16b]. Exploring [Mac20]. Explosion [ALL08, Gas09a, Gas21c]. Exponential [Lam13, McB85, McB87a, McB91]. expression [Lei80, OB83]. Expressions [Ali21d, BB21]. expressive [GLMW96]. extended [Akm92, Coo76, FL83, GW97, Joh96b, Mil72a]. extension [Che74]. Extensions [AP01, Oli04b, Pad10]. Extracting [Rag91]. Extraction [Gab10, Zim13, Dvi09]. Extractors [Cha20]. Extrapolation [FGH18]. Extremal [Pac93, Bón13b]. Ezra [Sha12b].

F [Das92, OR99a, Ree77, Dos87b]. F-gsm [Dos87b]. Fabio [Gal22, Kel16a]. faces [Gha23]. Facilities [Gol00a, PD98]. facility [Fot11, Lio93]. factor [GPP83, Gas19f]. factoring [Buh96, Gas16e]. Factorization [Cou12, FS15, Ven13a]. Fagin [Dek97]. failure [Chr05b, Ray05, Rey05]. Fair [Gas00f, Mou03, Dan06]. Falling [Gas15c]. Famous [Pet09, Ala03, Rey11]. Fan [Lur99, Par11]. Fanzhang [Yan11b]. far [vS14b]. Farach [Neh99]. Farach-Colton [Neh99]. Fascicle [Rog18]. Fascicles [Rog08]. Fascinating [Gas11b, Gre16g, dK09]. Fast [Gas19e, Koz87, Mor85]. Faulon [Ste11a]. Fault [Eya13, TV16, Wel16c, CBS07, KR01, Kei11b, KR09]. Fault-tolerant [TV16, Wel16c, KR01]. faults [KR01]. feasibility [GL02, Her83]. Feasible [For04, HT03, Gre22g]. Federal [SY87]. Fedor [JH20, Lam13]. Feferman [Les06]. Felix [Gas13c, Nag17b]. Fellows [Chi15, Ist00]. females [Tho93]. fences [KVY12]. Fernandez [Cad13, dOG12]. Fernando [Kin11]. Few [Ali21b, Ray21]. fewer [Tho93]. Fibonacci [Cad16]. fiction [DP08, Kas05, Gas06e]. Fields [Che79c, Kop11, MM07, Sha12a, Sti93, Sti09, VV78, AAH11, Bha13, Gol12b, Nag15b]. fifteen [Dif82]. Fifth [CH14, Dav05, LRV10, GJM02]. Fifty [Ros98a]. fighting [Vak13]. Figures [Tom89, Tom90, Mat14b]. file [Gup02]. Financial [BO06, Kye09]. find [CK98, HR00, Ram94]. Finding [GKP96]. Fine [Wil18]. Fine-Grained [Wil18]. Finite [Ali21d, BB21, Coh08, McB91, MP21, MM07, Nag15b, RS91b, Str94, Bha13, But71, Cul99, Fra84a, Fra84b, GL84a, GL84b, Lei80, Puc05d, Yli05, Gas94, Sha12a]. First [Gas15f, Gas15g, Gre17h, Har79, Rod96, Ros93c, DLM07, FKR08, HS12, Sav98, Via08b, Che80, Kuz18]. First-Order [Har79, DLM07]. Fisher [Gas09b]. fit [KGO10]. Five [Ali21c, Beh08, Gas15g, Har21, HN14, Gas09h]. Five-Minute [Beh08]. Fixed [FGH18, BI08, Kei08d]. Fixed-Parameter [FGH18, BI08, Kei08d]. Flajolet [Bón10b]. Flannery [GKKK06]. Flipping [Blu83]. Florescu [Gas16b]. Flow [Nag21a]. Flows [Cao13]. Floyd [Knu03a]. FOCS [Joh90]. focus [KM10]. fog [Chr04d, Ros04c]. Fogel [Mas01b]. Fold [FM13]. Folding [DO07, FM11]. folk [Har80]. following [Ros93c]. Fomin [JH20, Lam13].

Fontijne [FM08]. Fool [Ali21c, Har21]. Forbidden [Gre20e]. Ford [Cao13]. forecast [Dif82]. Forgot [Gas07c, Lew06]. forgotten [vL74]. form [BBF^{+}73, Dos87a, Par94, TF82]. Formal [Bel07, Cha98, Dav06b, Eil76, Gas94, Gin75, Jon72, KGO10, McN04, Par02b, Sch04, Sha09, Str94, Wil10b, Cud69, Sta10]. formalistic [Li12]. Formalizing [AKGN18, Wel18b]. Forms [BV07, Col10, Sav74, McB91]. Formulas [Lew79, Old91, Che82, Che82, Pur93]. Fortnow [Gas07h, DW13]. Forum [Che11a, Ros92a, Ros92b, Ros93b, Ros93c, Ros93d, Ros94b, Ros94a, Ros97a, Ros98a, Che13, Ros92c, Ros96a, Ros96d, Ros97d, Ros98b, Ros99c, Ros99a, Ros00b, Ros01a, Ros01b, Ros03c, Ros03d, Ros05a, Ros05b, Ros95b, Ros95c, Ros95d, Ros96b, Ros96c, Ros97b, Ros97c, Ros99b, Ros00a]. Foundation [GKP88, Mil13]. Foundations [APR16, FRRS17, Gal21, Gol01, GT02, Kil12, Mah19, MW93, Náv04, Nie02, Puc01a, Sal05b, SLB09, Vla94, Wel16a, Wel17b, Att08, Bau10, Chr05a, Kin14, LO05, Man82a, Man82b, Kat05, Azi10a, Jan08a, Puc03, Ter11]. Four [Gol06, Khu05a, Raj05c, Car77, Gas14d, Ang76]. four-gadget [Car77]. fourth [GM12]. Framework [JK92, Tse91]. Frameworks [Bau17]. Francine [Bón09a]. Francisco [Mur13]. Francois [Lew13]. Frank [Her13]. Franklin [Jan10b]. Franz [Gre18e, Nar00]. Franzblau [Kob98]. Fred [Bón11a, Kob98, Neh99, Pap11, Xie12]. Frédéric [BF17, Ful12]. Fredriksson [Gas07b]. Free [Fas83, Gas06e, Had15, BHT97, Col93, Ell70, Fas86, GS80, Hut70, Spe93]. French [Kin12]. Frenkel [Gas15e]. Freund [Kin14]. Frey [Was08a]. Fridays [Gas06e]. Friedrich [Kil12]. Frontier [Aar20, DP08]. Frontiers [Gas13e]. Fruhwirth [Oli04a]. Fulkerson [Cao13]. fun [Gas14g]. Function [Gas13e, Kop11, Mor85, Sti93, Sti09, Hem08a, Jul92, Kat78, Mar08b, McB92, YDX97]. functional [Esc93, FL83]. Functions [CK02, Gas19e, Tay04, Thi99, Azi14, BHHR99, BHT97, CM72, Dun84, Kos74a, Kos74b, LM73, Lon71, Mil72b, Mul69, VM85, She91a, She91b, Apo16]. fundamental [ER85]. Fundamentals [AW98, AW04, Lev96, Lev97, Chi15]. Funded [Ros94a]. funding [SY87]. further [Ric73]. Fused [Kre01]. Future [Gas21d, AFG^{+}96, PGW^{+}96]. Fuzzy [Li08, Yan11b, Par91].

G [Che79c, Chi15, Gas15c, Her77, Kob98, Mak93, Pri10, Tas99, Was08a]. Gabizon [Zim13]. Gabriel [Lab12]. gadget [Car77]. Gadi [Hen09]. Galore [Gas15g]. Gambling [Ski01]. Game [Azi10a, Ber00, Fis08, Gas09b, Geo09, Kei11b, Kri08, Lev10, Meh00, Men04, NRTV07, She13, SLB09, Spr08, Str93, AAH11, Cor83, Spr08, Aaz18]. Game-Theoretic [Azi10a, SLB09, Geo09]. Games [AN09, Gas10a, Gas12e, Gas23a, GTW02, HD09, Now02, TM11, Vin09, Bec08, CK04, Puc04a, MT04, Gas10a, Gas14g, Apo11, Cou15, Gas10a, Nag18a]. Gammerman [Law06]. Gardner [Gas09h, Gas15f, BR99, Gar08, Gas11g]. Garibaldi [Gas14h]. Garrett [Mil17b]. Garrity [Gas10e]. Gary [Fab11, Gre16g, Nag15b, Sha12a]. Gasarch [For99, Gre22h, Gas20c]. Gathen [Hom17]. Gay [Sga12]. Gebert [Cou14]. Gems [Kul13, SP98, Kri00]. genealogy [Coo04, Joh84]. General [AB09b, BHS80, Grä78a, Grä78b, Grä98, Jou94, Sny91, Hen83, Sud11, Tur84a, Tur84b, Coh08]. generalization [HK84a, HK84b]. generalizations [AS03, Ber04]. generalized [But71, McB87b, KB79a]. Generals [Kei10c]. generate [FC83]. Generating [ER85]. generation [KS99, Kei10c]. generic [Dew86]. Genome [Kel16a]. Genome-Scale [Kel16a]. Genomic [Dwy03, Wan05]. Gentle [Gol00b, KS01, Sga13]. gentler [Dwy90]. Genus [GHOP22]. Geometric [DO07, DFM07, Vin09, Zhu20, Yuv76, FM08, FM11, TM11]. Geometry [ÁDFT15, BKM18, BMP05, Car15, Dek99b, DFM07, DJ14, Dum16, Dum18, DJ19, ED17, Gal22, GG21, KL16, Mas00a, MMW06, O’R94c, O’R94d, PW01, PW09, Sha17, Zhu20, dB^{+}97, dB^{+}00, dB^{+}08, AO98, Che13, Chr04b, DO99, Dum12, DT12, DJ13a, DJ13b, DT14, MO01, O’R88, O’R89b, O’R89c, O’R89a, O’R90, O’R92b, O’R93, O’R94a, O’R94b, O’R95a, O’R95b, O’R95c, O’R96a, O’R97b, O’R97c, O’R98, O’R99b, O’R00a, O’R00b, O’R00c, O’R01, O’R02, O’R03, O’R04b, O’R04c, O’R06a, O’R06b, O’R07, O’R08a, O’R08b, O’R12, Raj04d, Raj06c, Raj06d, Rie99a, Ven13a, O’R96b, O’R97a, Ess05, FM15, Gas07d, Gre20e, Cou14, FM14]. George [Coh08, Nag15a, Vak13]. Georgia [For99]. Gerard [LaT11, Bor08]. Gessel [Maz14]. get [Hil07]. Ghosh [Dea09, dV16]. Giampaolo [Sta10]. Gian [Maz14, Mou10]. Gian-Carlo [Maz14, Mou10]. Gifford [Ogi01]. Gilbert [Nag15a]. Gillen [Gre16e]. Gilles [Cad13, Nag21b]. Gillman [Wil13]. Gina [Gas00a]. Gina-Carlo [Gas00a]. Ginsberg [Gas19f]. Ginsburg [Eil76]. Giorgio [Kin15]. given [Lei80]. glass [Ang91]. Glisse [BF17]. global [Hem08a, Mar08b, Sub95]. Go [Far19]. Gödel [Gas10f, Lip10, Cha71, Pud06, Sha94, Hil07]. Gödels [Pux97]. Going [Ros00c]. Golan [Pap13b]. Gold [Pin10]. Golden [Cad16, DW13, Ves21, Hem22d]. Goldfarb [Che82]. Goldreich [Apo10, Kat05, Lee03b, Puc03, Nag13b, Upa20]. Goldwasser [Dav05]. Gollman [Cho12]. Gone [Ros97d]. Gonzalo [Koz18]. Good [Che20]. Goodrich [Mas01a, Náv04]. Google [AAC^{+}08, Gas08b]. Gopalan [Bel14]. Goppa [Jor83]. Gotcha [Gar06, Gas09h]. Gowers [Azi10b]. GRA [Pro90]. GRA-GRA [Pro90]. gracefully [Gab75]. Grädel [MT04, Sim04]. Graduate [Bay95, Gas97a, Ros98c]. Graham [Gas99b, Par11]. Grained [Wil18]. Grammars [Pro90, Ros75, But71, Cud69, Eng89, Fas86, Gal13]. Grammatical [Gal13]. Graph [Ali23c, Ass23, Bol98, Bón09b, Bon08b, Cen00, CZ09, Chu97, CC78, Fab11, Gas04a, Gre16g, Gre19g, GY04, HHM00, HHM08, HL09, Lur99, Mar08a, Pro90, Ros75, Bat93, Coh93b, ER85, GJ78, Kan93, Koz78, Kri77, Pac93, Tur84a, Tur84b, TT95, Sch09, Azi11, Cad14, Lab13, Sil12]. graphical [OB83]. Graphs [Bol85, Bol01, CEG98, Gas23a, Jun99, Jun05, Jun08, Kel16b, MMW06, Par11, Ros73a, Ros73b, CK89, CC78, GHOP22, Gur04, HK84a, HK84b, Kri75, O’R04d, Ros77, Tol93, Bón09c, Rau18, Fah05]. Grätzer [Coh08]. Great [Gas07c, Gas10e, Pet09, Rey11, Sri14, Lew06].

Green [Azi10b, Fri11]. Greene [Mik11]. Gregory [Fit11]. Grid [Gas16a, Bor09b, Cas02, Kan93]. Gröbner [Cou09b, Mor05]. Grolemund [Mil17b]. Gross [Gas04a]. Große [Dav08b]. Große-Rhode [Dav08b]. Grötschel [Gas99b]. Group [Hay08, Gas09h]. Grundlagen [Che80, van79]. GSM [Dos87a, Dos87b]. Guerino [LaT11]. Guerre [DCV82, FFGH08]. Guest [Aar05, AAV13, All23, AdW17, BCG^{+}22, Bau17, Bha13, BMMR20, BT20, CCD15, Cha20, Che11b, CT23, CP13, Dvi09, FGT17, Gas12f, GHMP19, Gas19h, Gha23, Gol06, Gur04, HH24, Kay18, Kho05, Khu05b, KLMY21, KS16, Lam14, LST22, Lin92b, MP21, Mos12, OW17, O’R04d, PV21, Pig15, Raj05d, Raz16, Sud11, Tre09, Vio09, Vio19, Vol21, Wat11, Wil11, dRGR22, BI08, yC08, DLM07, Kei08d, Vad07]. Guidance [Dav11, BHI05]. Guide [Ben06, Dud09, Gas20f, Lea00, Mar15a, Mol05, Par88, Par00b, Rau08, Wal03, Wal12, Yan11a, Han94, Par89a, Gre19e, dV18]. Guided [Cad11, Maz10, Pet08]. Gunter [Sta93, Bón11c]. Gurevich [Sim04]. Gusfield [Ben98, Kel16b]. Guy [Lab13].

H [Avi01, Coh07, Dav05, Dye11, Ess05, FM09, JK92, Kin15, Lew14, Mar09a, Mik11, Mos05, Mou10, Pan93, Tew14, Was08a]. Håstad [Ten18]. Hadley [Mil17b]. Hadlock [Cen14]. Hal [Gas09a, Gas21c]. Half [Kar00, BBAGP09]. HALG [CK18, CL17]. Hall [Ben06, Cha09, Ful10, Gol00b, Hen09, Kat05, Kel09a, Kel09b, Spr06, Was08a, Wid03, Che10, Che79b]. Hall-CRC [Ful10]. Hall/CRC [Ben06, Kat05, Kel09a, Spr06, Was08a]. Halpern [Dek97, Zha04]. halting [Bur87, Lap91, Wil86, Yli05]. Hamiltonian [Lit03]. Hammer [Azi14, Kam13]. Hancart [Gal15]. Hanchart [Ada11]. Handbook [Ata99, AB09b, AB10, BvBW07, Cha09, CFA^{+}06, DKV09, FB10, GGL95, Gre19g, GY04, OZ06, Zom06, Gas99a, Pap12b, Cad12, Gas99b, Gas04a, Ist12, Nag15b, Nag17b, Ste11b, Ste11a, Was08a]. Hands [HK05, Wil10d, BSS11]. Hands-On [HK05, Wil10d, BSS11]. Hankerson [Irw03a]. Hankin [Lea00]. Hanoi [Koz14, Gas09h, GC86, KB79a, Sta87]. Hans [Fal12b, Gas17d, Mur10, Pap05a]. Happiness [ALL08, Gas09a, Gas21c]. Hard [FGK^{+}02, Hoc97a, Hoc97b, Hro03, Mas03, BS98, BBG94, Dem09, GH00, Gre81, HHH98, Hua03, Knu74a, KS06b, Kri75]. Hardcover [Ada11, Tra22, Pap05b, Bón10b, Bón10a, Boo10, Cou10, dOG10, Pin10]. Harder [For21]. hardest [GS80]. Hardware [Har15]. Hardy [Pap05a, Xie12]. Harel [Che80, Puc01a, Ste12]. Harlan [Amo99]. Harmful [CBS07]. harmony [vS12a]. Harold [Cou10, Irw08]. Harris [Bón09b]. Harry [Che82, Gas07c, Gas09a, Gas16h, Gas20d, Gas21c, Gas21d]. Hartmanis [Gre22g, Hem22d]. Harvey [Amo99]. hash [BHT97]. Hayes [Gas09h]. Hazay [Kat12]. heapsort [KW81]. Hearn [Apo11]. Heart [Gas15e]. heaven [Ven13b]. Heidelberg [Dav06b, Dav08c, Kye09, Dav09, Gri10]. height [Woo78]. height-balanced [Woo78]. Heiko [Cad12]. Heinz [Bau10]. Heisenberg [Lap91]. Helmut [Mur10]. help [BT08]. Hemaspaandra [All98, For04]. Henle [Gas15f]. Henning [Kop11]. Henri [Kel08e, Pop12]. Herbert [GG21, Kre00].

Herbrich [Wil04]. Herlihy [Upa16, Wel15b]. Hervé [Dan06]. Hesselink [Pan93]. Heterogeneous [Dav08b, GR04]. heuristic [HW12]. Heuristics [Hro03, MF02, Mas01b, Mas03]. Hexaexagons [Gas09h]. Hidden [Gas15e, FL83, Nag18a]. hiding [Bra90b]. hierarchical [Coo83]. Hierarchy [RST15, HHH98, Sch81, You92]. High [KS16, Kel16a]. High-Rate [KS16]. high-throughput [Kel16a]. Higher [CH06, Edw08, Gas16h, Bel14, FM09]. Higher-order [CH06, Bel14]. Highlights [CL17, CK18]. highways [Nag13a]. Higuera [Gal13]. Hilbert [Gas13c, Gas21a, Li12, Mat93, Gas13c, Pru00]. Hill [Kel09b]. Hindley [Pos98]. Hinek [Roj12]. Hinz [Koz14]. Hirschfeldt [Gas16d, Teu13]. Hirst [Bón09b]. Hirvensalo [Vio04, dW07]. Historic [Pet08]. historical [Mat14b]. History [Bla06, Gas12a, Kap99, Sch18, Abr13, Kei10a, Ses09, Joh96a, Har10, Nag13a]. hoc [Raj02a]. Hochbaum [BS98]. Hodge [Wil10d]. Hoffstein [Mei10]. hole [Coo76]. Holes [Gre21f]. Holland [Che79a, Eil76, Her77, Sub78]. Holographic [yC08]. home [Hem98, Ros97b]. Homer [Sha13]. honor [Gas13c]. Honour [PSTM00, Puc01b]. Hooker [Nag14]. Hopcroft [Mai80]. Hopkins [Abr13, Gas15f]. Horizon [Ros93d]. Hosam [Sta11]. hot [Bra90a]. house [BHHR99]. Housman [Wil13]. Howard [Gas94, Irw08, Sim00]. Hromkovic [Mas03, Mav09, Gas13f, Por12]. Hsing [Sil12]. Hsu [Sil12]. http [CEG^{+}10]. Huber [FB14]. Huffman [Ten87]. Huffman-equivalent [Ten87]. hulls [Dwy90]. Hunker [Cho12]. Huth [Dav06a]. Hüttel [Fal12b]. Hutter [Dav11]. Hydra [AF20, AF20]. hyperbolic [Hem08a, Mar08b]. Hypercubes [Lei92, Das92]. Hyperelliptic [CFA^{+}06, Was08a]. hypothesis [Vél78].

Ian [Gas09h, Gre15d, Sga12, She14, Sri14, Ste13]. ID [KGO10]. ID-based [KGO10]. idea [Sug93, Bla06]. ideas [HN14, Gas21d]. idempotent [KL05]. Identity [Kat13a]. Identity-based [Kat13a]. IEEE [Mos05]. IFIP’94 [Kud95]. Igor [Sha10a]. II [Cou09b, LB99, She91b, AFG^{+}96, BCRV04, CR03, GH00, HS97, Mor05, RS91a, All98, OR99a]. III [RW99, Amo99, Gas10a]. Imaginary [Bla06, Nah98]. Immunocomputing [TSS03, Zha05]. Impact [Ros94a]. Implementation [Dav05, Mil72b, EKP79, GJM02, Hut01, Vis91, Dav06a]. Implementing [AKGN18, BCC10, Rod96, Wel18b]. implications [Cha73b]. implicit [FFB89]. Import [Mil17b]. importance [Smi99]. Impossibility [Gas22e]. Impossible [Blu83, DW13]. Improved [GG86, dJvL75]. improvement [Fra84a, Fra84b]. Imre [Azi10b, DL79, Gas01a]. In-place [Pas99]. in-theory.blogspot.com [Gas07h]. inapproximability [Kho05, Khu05b, Lam14]. incompleteness [Cha71]. Incorporating [Ros93b]. Increasing [OW17]. Incredible [Gre15d]. incremental [Fot11]. Independence [HH76, FMOP70]. independent [Cal83, FKW94, Li04]. Indiscrete [RP97, Gas00a]. induction [McB87b, Ros96a]. Inductive [MNV72]. Industrial [Joh10]. Inevitable [Dud08, dOG10]. Inference [Gal13, Fas86, KVY12]. Inferring [SSC14]. Infinite [Cen15, GTW02, MT04, FKW94, Puc05d, Wal82, Yli05]. infinitely [Agu04, Khu04]. Infinitesimal [Gas15d]. Infinity [Bel11, Gas19e, Rei06, Sri14, Ban14a]. Information [Bay95, Big08, BCC10, Cha98, Di 95, Kat13b, Lea97, Low12, Mou16, Oli04c, Wei15, BSS11, Bra90b, Cha73b, Gri10, KL10, Raj03d, Wat11, Fen21b, Rub13]. information-theoretic [Cha73b]. Infrastructure [Fic97]. Ingo [Joh10]. Initial [Tse91]. injective [Ing83]. injectivity [Hem08a, Mar08b]. innovation [Kei13b, Pel13]. Innovative [Ste11b, Zom06]. input [Fra85, OS14]. input-driven [OS14]. Insider [Cho12]. Insight [Gar06, Gas09h]. insights [AAH11]. Inspecting [dRCR94]. Inspired [BO06, Kye09, Ste11b, Zom06]. Institute [Gas15g, WG00, All00]. instructions [SSC14]. instrumental [Gol12a]. Integer [Cou12, LS06, Mel11, Wol98, ASU86, Cha05]. integers [GK80, Mat94, Gas16c]. Integrated [Nag14]. Integrating [Zom06]. Integration [Dav08b, GR04]. intellectual [Gol12a]. intelligence [FC83, Mos05]. Interaction [PSTM00, Puc01b]. Interactive [PV21, Wei15, Bra90a, CP13]. intercloud [Vuk10]. Interconnection [Sil12, HL09]. Interdisciplinary [Mah19]. interested [Ski99]. Interesting [Rei06, Yli04, Bel11]. intermittent [BMO77]. International [Agu12, LRV10, Pro90, Rod96, Ros00d, Wan97, Par00c, Sta78a]. Internet [Náv04, LO05, Chr05a, GT02, KM10]. interpretations [Gue83]. Intersection [Aga91, Bir91, YZ91, O’R92a]. Interval [FGK^{+}02, Has85b]. intervals [DR12]. interview [Rou05]. Intractability [Sta89]. Introducing [Men04, Spr08]. Introduction [Baa78, Baa88, BV00, Bal97, Bie05, Buc01, Buc04, Che24, Chr09a, CC76, COJ78, DP90, DP02, DK02, DK07, Edw08, EA09, EA10, Ela96, Ela99, Ela05, Fal07, Fal12b, FM09, Gas97a, Gas12a, Gas13f, Gas19e, Gre19f, dOG12, Hau12, Hem07a, Hem07b, Hem15b, Hem15c, Hem15d, Hem16b, HPS08, HU79, HMU01b, HMU07, Hro03, Hüt10, JORS99, Kat06a, KL08, Kei07, Khu07, KS01, Knu97, Lei92, Lev07, Lev12, LV93, LV97, LV08, Low12, Mar03a, Mar09b, Mas03, Mav09, McN99, Mei10, Mer07, Par02b, Par08, Pit00, Puc07b, Puc07c, Raj07a, Ree77, RV10, RB11, Rot05, Sch03, Sho05, Sho09, Sip96, Sto02, TW06, Tel00, Zeu89, dW07, Big08, Che13, Chr05b, Fer09, Hem08b, Kei08e, Lew78, Mat94, Ray05, Rey05, Ses09, Wat11, Ros23, Gol00b]. Introduction [Kat05, Led16, Rie10, Rot07, Sga13, Azn13, Coh07, Das92, Fah04, Fal12a, Ful10, Ful12, HMU01a, Lee03a, Mai80, Mav02, McN04, Mos05, Mur10, Pap13b, Sch08b, Spr06, Upa20, Wil10a]. Introductory [Ros93b, Ros93c, Ros00c]. invariant [Bha13]. invisible [Ang91]. Invitation [GN07, Raj07b]. IOS [Gas07b]. Iosevich [Gas14h, Har11]. Ira [Maz14]. Ireland [Dav11]. ISBN [Gas94, Hen93, JK92, Jou94, Law06, Lev10, Mak93, Mik93, Mos05, Mur10, O’R92a, Oli06, Pan93, Pap10, Pri10, Pur93, Rie93, Sis93, Spr06, Sta93, Vio04, Vla94, Wil04, Zwa93, dW07]. Island [Sch09]. isomorphism [CC78, Koz78, RW94]. Isomorphisms [Di 95, Lea97]. Israel [Con03]. Issue [Gas20c, CP74]. issues [Cas02, Fic97]. István [Bor07a].

Itanium [Kre01]. Iteration [Che21]. Iterative [GPP83, Li14]. Ivars [Ban14b, Gas10c]. I’ve [Bra95]. ix [Zwa93].

J [Azi12a, Bar12, Bón09b, Bón12, Cal12, Cen11, Che79b, Che80, Ess05, FM11, FM13, Gas05b, Gas10f, Gre16f, Jen01, Kat14b, Kum13, Led16, Mar09a, Pap12b, Pos98, Puc01a, Smi09, Tas99, TM11, Wil04, Cen09]. J. [Grä08]. Jackson [Lee10]. Jacob [Gre22h]. Jacobo [Gas14i]. Jacobus [FB14]. Jacques [Gas12a, Kin12]. Jain [Smi00]. James [EA10, Gas15g, Mas00b, Sga10]. Jan [Gas19e, Mar09a, Mou14]. Janet [Ayu12]. Jason [Roj12]. Java [Pap05a, Ros96b, GT97, GT01, GT04, GT06, GT10, HS01, Wei98, Mas01a]. JavaHot [Ros96b]. Je [Sga09]. Jean [Gas11b, Lew13]. Jean-Francois [Lew13]. Jean-Marie [Gas11b]. Jean-Paul [Ber04]. Jeanne [Fri11, Nag13a]. Jeffrey [Ber04, Jen01, Mai80, Mei10, Wil10b]. Jeffry [Bón09b]. Jennifer [Gas05b]. Jensen [Mik93]. Jérôme [Nag17b]. Jeremy [Nat07, Nat10]. Jerome [Sch06]. Jesús [Mur13]. Jewels [Mar15b]. JFLAP [Rod99]. Jill [Mei10]. Jim [Gle05]. Jin [Gre22f]. Jin-Yi [Gre22f]. Joachim [Hom17]. Job [FM05]. Jody [LaT11]. Joel [Gas16b]. Johan [Ten18]. Johannes [Col10, Lee03a]. John [Aar07, Azn14, Bea07, Bón09b, Cen03, Fen21b, Gas20f, Jan10b, Mai80, Mil17a, Nag14, Náv04, Rot07, San10, Wer13]. Johnson [Dav05]. Johnsonbaugh [Kel09b]. Joint [Apo10, Bla06, Coh08, Gas07h, Gas09h, Gas11g, Gas13c, Gas14d, Gas23a, Gre17e, Gre18e, Kel09b, Lou18a, Mos05, Puc03, Vak13, dW07, Gas06e, Grä08]. Jon [Dek98]. Jonathan [Ful10, Mou15, Wil10d]. Joost [Cad14]. Jörg [Fal07, Mur13]. Joseph [Dek97, Dek99b, FM14, Kob98, Mei10, Mou10, Zha04]. Joswig [FM15]. Journal [CP74, Rav90a, Wei95a, Wei95b, Wei96a, Par89b, Par90a, Par90b]. Journals [Hem97a]. Journeys [Gas10d, Kob08]. Joux [AY12]. Joy [Bla06, Bla97, Gas16e]. joys [Gas07a, Slo07]. Jozsef [Gas12e]. Jr [Cao13, Kat05]. Jr. [Cen11]. Juba [Cal13]. Judith [Cen11]. Judy [Fri11]. Juergen [Spr06]. Jukna [Bón13b, Gas13e]. Julia [Cal10, Gas14h, BY09]. Julian [Hen93]. Jumps [Thi99]. June [AP98, Azi10b, LB99, LRV10, RW99, WG00]. Jungnickel [Fah05]. Juraj [Gas13f, Mas03, Mav09, Por12]. Jürgen [Cou14, Gas17d]. Juris [Gre22g, Hem22d, Koz22, Sim22]. Jutting [Che80].

K&CK [Mos10]. Kahn [Gas01b]. Kalman [Kum12]. Kanellakis [Smo21]. Kaplan [Bla06]. Kasman [Gas06e]. Katharina [FB14]. Katz [Ful10, Gas19e, Mou15]. Kauffman [Das92]. Kaufmann [FM08]. Ken [Gas09a, Gas14f, Gas21c]. Kenneth [Gre16f, Sch06]. Kenschaft [Fri10]. Kepes [Gre17f]. Ker [Cen03, DW19]. Ker-I [Cen03, DW19]. Kernel [Her02a, Wil04]. kernelization [GN07, Raj07b, JH20]. Kernels [SS02, Wil04]. Key [Cou12, Kat05, Mol03, BB87, Hen83, Jor83]. Khoussainov [MT04]. Killing [Gal78]. Kinber [Gol00b]. Kind [Mas02, Wol02]. Kinder [Dwy90]. kinesthetic [SP07]. King [Cen15]. Kitaev [Rau18, dW07]. Klaus [Dav06b]. Klavzar [Koz14]. Klima [Wil10d]. Kluwer [Bor07a, Kre07]. knapsacks [AKR83, Ing83]. Knebl [Mur10]. Know [Ali21b, Ray21, Par98a]. Knowledge [FHMV95, FHMV03, Por12, RG06, Hro09, Kei10a, Dek97, Upa10]. Knudsen [Kat14b]. Knuth [Gas15c, Kan14, Lew78, McN99, Oli04c, Rog08, Ten18, Apo14b, Apo14a, Gas14e, Gas14g, Gas15c, Hua98, Khu98a, McN01b, Mik11, Pag70, Rog13, Rog18, Gas14e]. Ko [Cen03, DW19]. Koblitz [Gas10d]. Kok [Ste12]. Kolmogorov [Gas97a, KL05, LV93, LV97, LV08, RSZ21, Sha19]. Koninck [Gas11b]. Koolen [FB14]. Koren [Con03]. Korst [Mar09a]. Korte [Kum13]. Kozen [Puc01a, Apo12]. Kozlov [Upa16]. Kozma [Ist12]. Kranakis [Tay04]. Kranz [Gas06e]. Kratsch [Lam13]. Kreveld [Mas00a]. Kronecker [Mor03]. Krusemeyer [Nag15a]. Krzysztof [Aro98]. Kuan [Sil12]. Kuich [Cad12]. Kumar [Dea09]. Kung [Mou10]. Kurt [Elr11, Hil07, Mik93].

L [All98, Azi14, Bón09b, Cao13, Che80, FM14, Gas99b, Kam13, Kil12, Lib17, Low12, Mar09a, Nag15b, Ree77, Ric94, Sch06, Sha12a, Sha13, Wid03, Xie12]. Làaszlò [Lou18b]. Lab [Ros93c]. Laboratoire [GP89]. laboratories [Ros93a]. laboratory [CW72, SIG83c]. LaDuke [Fri11]. Lagemann [Gas16h]. Lakatos [DL79, Gas01a]. Lambda [Fis72, Han94, McG72a, Lea00]. Lance [DW13, Gas07h]. land [Agu04, Khu04]. Landau [Che80, Mak99, Jan08b, Mor13, van79]. Landman [Gas16c]. Landweber [OR99a]. Lane [All98, For04]. Lang [Nag17b]. Language [Di 95, Fas04, Jac06, Lea97, Lee10, Oli04c, Pig15, PSTM00, Puc01b, AM75, Ash72, Cad14, GS80, HR02, Pys76]. language-theoretic [Cad14]. Languages [AK90, BMMR20, Fas83, Gin75, Gun92, Her77, HU79, HMU01a, HMU01b, HMU07, LR76, Mai80, McN04, Par02b, Par08, Pie02, Sha09, Sta93, Wil10a, Wil10b, All72, Col93, Cul99, Dos87a, Dos87b, Ell70, Gre81, GLMW96, Hof00, Hut70, Jaf78, Lib99, OB83, Pet94, Rup91, Sei74, Spe93, SW82, VS98, Weg72, YZ91, Ros00c, Bir91, Eil76, Kin05, Kin06]. Langville [Mat14a]. Lap [Li14]. Large [Ist12, Nah04, Kei09c, Kei13c, KM10, Zho09]. Large-Scale [Ist12, Kei13c, KM10, Zho09]. largest [GKP96]. Larry [Raj04e]. Lars [Kat14b]. Larson [Nag15a]. Lasse [Gre16e]. last [Joh90]. László [Azi12b]. Lattice [Coh08, Grä78a, Grä78b, Grä98, BCRV03, Dos87a]. Lattices [Coh07, DP90, DP02]. Lau [Li14]. Laura [Gas16b, OR99a]. Laurence [Cha05]. Law [Bla12, Gre18e]. Lawler [Ric94]. Lawrence [Che10, Dek98]. layout [Lio93, Tur84a, Tur84b]. lazy [Esc93]. Leader [Azi10b]. leads [KL05]. Learn [JORS99, Nag23, Tay05, Smi00]. Learning [AdW17, Fas83, Gal13, Her02a, JORS99, Jen01, Law06, McC01, Ros99c, SS02, VGS05, Gas02a, Ros01c, SP07, KV95, Wil04, Upa21, Gal14, Wil04]. Least [Nah04]. Lecroq [Ada11, Gal15]. Lecture [Che24, Ros93c]. Lectures [Gre21f]. Ledeen [Gas09a, Gas21c]. Ledger [AKGN18, Wel18b]. Leen [For04]. Legacy [CEG98, Par11]. Legrand [Pop12]. Leibniz [Kil14a, Kil14b].

Leighton [Das92]. Leiss [Kel09a]. lemma [Col93, Jaf78]. Lemmermeyer [Gre18e]. Len [Gas09b]. length [Kan88, Sal07, Nas13]. lengths [CR74]. Lens [PV21]. Leo [FM08]. Leonard [Gas10g, Gre21f]. Leslie [Bru13]. Less [Ali21a, CKN21, Gas22f]. lessons [Ski99]. LessWrong [Gas22d, Gas22d]. letter [Lip10, Par00a, Gas10f]. leurs [Knu76b, Lew78]. Level [Dav11, BHI05]. Levitin [Fah04, Gas13d, Gas13d]. Lewand [Gas05a]. Lewin [Rou05]. Lewis [Che82, Gas07c, Gas09a, Gas20d, Gas21c, Gas16h, Gas21d]. Leyton [Azi10a]. Leyton-Brown [Azi10a]. Li [Gas97a, Mel11, Yan11b]. Liberty [ALL08, Gas09a, Gas21c]. Library [FM09]. LICS [FK89]. Life [ALL08, Ada74, Ada09b, Bón12, Cha70, FF04, Fis08, Gas09a, Gas09c, Gas21c, Gre17f, Les06, Soi09, Kat14a, Gas09b, Ste13]. light [Gol92]. Lih [Sil12]. Lih-Hsing [Sil12]. like [Weg86]. Limit [Ada74, Ada09b, Bón12]. Limited [GLM96]. Limits [Cha98, Tas99, Coo12]. Lin [Sil12]. Lindell [Ful10, Kat12]. Lindenmayer [Her77]. Lindner [Pap12a]. Line [DL98, DL07, Ess05, FM05, Jan08b, PW01, PW09, vS16b, vS16c, BM08, Che74, DLO05, Eng18, Gal75, Kau93, Mak99]. Linear [AP01, Gre16f, Hua14, Lin92a, Lin92b, LO94, MMW06, Nag20, Pad10, Sav74, Bur78, CH93, Gol95, Reg93, Sud11, Oli04b]. Linguistic [Zam77, Sub78]. Linial [Ano13]. link [HHH98]. Linkages [DO07]. linking [RGL10]. Linton [Sol12]. Lipton [Gas10f, Gas14f, Gre16f]. Liskov [Rod09]. Lisp [Lon72]. List [DM22, Sud00]. LITP [GP89]. Liveness [Sch18]. Lives [Lou18a]. LLL [Nar14]. Local [Bar15, KS16, Mar09a, MAK07, IMP11, Rob70, Sha02]. Locality [Ali20c, Suo20]. Location [Gol00a, PD98, Fot11]. Loehr [Bón13a]. log [Gol77, Lad75, LM73, LL75]. log-space [LM73]. logarithms [Her83]. Logic [AH91, BdRV01, BvBW07, COJ78, DL79, Dew79, FF04, GMN87, Gas94, Gur94, HR02, Har79, Har86, HKT00, Jac06, KA89b, Lak76, Lee10, Li08, Lin92b, MP92, MMW06, McN04, NS93, NS97, Nie02, Par02b, Puc01a, Sch89, Sch08a, Sis93, Str94, Ter11, Yan11b, Cad14, Che06, Chr05c, DLM07, Gop08, HKT01, KA89a, Lin92a, Mil72b, Pal90, Par91, Puc05d, Puc05f, Puc05g, Puc06c, Puc07a, Puc10, Che80, Hes03, Les06, McN04, Bel14, Dek99a, Puc05c, TA95]. Logical [Azi10a, Che79a, Kra07, Lov78, MW93, Puc05e, SLB09, Car98, Mil13, PC03]. Logicomix [DPP09, Gas10b]. Logics [DH79a, GTW02, MT04, Puc07a, Kur06, LP07]. Lokshtanov [JH20]. Long [Kho05, Khu05b]. Longest [OW17]. look [BA05, Chr06b, Ver06]. look-alikes [BA05]. looking [Chr04d, Ros04c]. Loops [GMS74, Ros96a]. Lorea [Mur13]. Loren [Nag15a]. lost [Gas10f, Lip10]. Lototsky [Gre07]. Lotus [Dwo95]. Louridas [dV18]. Lovász [Azi12b, Gas99b]. Love [Gas15e]. Loveland [Che79a]. Lower [Ali23c, Ass23, Kay18, LST22, Sha10a, Shp03, Che74, CK89, HHR97, Kan88, Vio17, Wel14c]. Lowness [Hem93]. Lozin [Rau18]. LR [Ben73a]. Lu [Bor09a]. Luca [Gas07h, Pos99, Zwa93].

M [Bón09b, Bón11c, Cal12, Che79b, Che79c, Cou10, Dav08c, Dav09, FM09, Gas99b, Gas16c, Ist00, Jan10b, Kin15, Koz14, Mar09a, Mou16, dW07]. MAA [Gas05b, Gas05a, Gas06e, Gas10e, Gas10c, Gle05, Pin10, Spr08]. Maarten [Hes03]. Macaulay [Cou09b, Mor05]. MacCormick [Gas20f]. MacCuish [Azn14, Azn14]. Machine [BA95, Pig15, Reg93, Upa21, Gal14, Gas02a, Kan05, LMZ10, McB92, Nir86, McG72a, Pet08]. Machines [CC76, MMW06, McN04, Par02b, Par08, Pux97, Ree77, SS02, Sha94, Wil04, Ben74, Chi86, Gal74, GL84a, GL84b, Her00b, Kei10a, Par86, SvL74, Wil86, Sta78a, Wil10a]. Mackie [Sga12]. made [BM08, DO13]. Magic [Gas10c, Gas23a, Gre17e, Pet02, Hro09, Por12]. Mahajan [Vak13]. Mahmoud [Sta11]. Make [Ayu12, BP06, Nah04]. Makes [Bel11, Rei06]. Makespan [DM22]. Mäkinen [Kel16a]. Making [All97, HN14, Ros96c]. Malware [Ayc06, San10]. Man [Alp18, Gas19f]. management [Gol10b, IP05, Puc05b]. Manes [Mak93]. Manfred [Cad12, Oli04b]. manipulation [But71]. Mann [FM08]. Manna [Sis93, Vla94, Fur11]. Mannucci [Cou09a]. Manual [Ski98, Ski08, SR03, Tra22, Kar11]. Many [Nah04, Agu04, Khu04]. Maor [Bla06]. Map [Gas22d, SSC14]. Mapping [Cra03a, Cra03b, Nat07, Nat10]. mappings [Dos87a, Dos87b]. Maps [Has85b]. Marc [BF17, Mas00a]. Marcel [Irw03a]. Marcus [Kel09b, Abr11, Azi11]. Marek [Fal12a]. Margaret [Led16]. Maria [Gas13d]. Mariages [Knu76b, Lew78]. Maribel [Cad13, dOG12]. Marie [Gas11b]. Marina [Pap12b]. Marius [Oli06]. Mark [Bón14, Cal10, Gas15g, Mas00a, Mou16, Nag15a]. market [CPV04, Raj04c, Nag20]. Marko [Kre00]. Maros [Bor07a]. Marriage [Knu97, McN99]. Martin [Bón11c, Boo10, Dav08b, FFGH08, For99, Gas09h, Gas11g, Kat06b, Kil14a, Kil14b, Nag16, Neh99, Pos99, Sza12, BR99, DCV82, Gar08, Gas15f]. Martinez [Gas15i]. Martins [Bla06]. Mary [Ree01, Sin99]. Mass [Wil04]. Massachusetts [WG00]. Mastering [Ali20b, RR20]. Masters [Ali21b, Ray21]. Matching [ABM08, FGT17, Lab12, LP86, LP09, Val09, vS16b, vS16c, BM08, Gal75, GG86, LO94, Neu10, Azi12b, Nag20]. Math [Ban15, Gas15i, Gas15e, Mac20, Gas23a, Gre17e, Ban15]. Mathemagician [BR99, Gas11g]. Mathematical [Cha70, Che80, Dav05, Dav11, DL79, Dey07, FC^{+}99, FM09, Gar08, Gas06e, Gas09c, Gas09h, Gas10g, Gas15h, Gas16d, Gas22e, dOG10, Hay08, Hes02, HPS08, Kat05, Knu97, Kob98, Kri08, Lak76, McN99, Mei10, Nag15a, Pet02, Ros23, SS07, Sch09, SD08, Ski01, Soi09, Sot10, Sri14, Ste09, Wap05, Wil10d, dW07, BHI05, Gas15d, Kas05, Vak13, Ven02, Gas06e, Gre22h, Mos05, Ban14b, Gas10c, Gas13c, Mie11, Neh99]. Mathematically [Ada09a, Gas10e]. Mathematician [Kob08, Gas10d]. Mathematicians [Gas06e, Nah04, Pet09, Gre17f, Rey11]. Mathematics [Ayu12, BM96, Beh08, BP06, BL06, Cha98, CR03, Con09, Cou15, CH10, Dud08, Edw05, Eri10, Fal12a, Fri11, Gas09c, Gas10e, Gas20d, Gas23a, GBGL08, GKP88, GKP89, GL09, GK81, GK82, GK90, GK07, dOG10, HK05, Joh10, Kam13, Ken05, Knu03b, Kob98, LaT11, Led16, Lew00, MMW06, Nie02, Ogi01, Pin10, RFR97, Ros98d, She15, Soi09, Ste13, Ter11, Wal63, Wal09, Wal03, Wal12, Wil10d, Abr13, All00, Coo12, Coo04, Fas04, GS08, Mar09b, Nag13a, Nag18a, Tay05, Gas09h, Vak13, Azi10b, Cou10, Dye11, Fri10, Gas05a, Gas11a, Gas15g, Har10, Oli04c, Rau08, Sza12, Tas99, Apo14b, Gas15g, Gas23a, Gre07, Gre21e, Mik11, Nag16]. mathématique [Lew78]. Mathematisch [Che79c]. MathML [Ros05b]. maths [Koz14]. Matiyasevich [Pru00]. Matrices [Che79c, VV78, Shy78a]. Matrix [Bón10a, BC09, Ang76, ASU86, Boo77, Khu12, Shy78b, Ven13a, Yel78]. matroid [Din13]. Matt [Gas19f]. Matter [FRRS17, Wel17b, Ros97c]. Matthew [Gas19e, Kat14b]. Maurice [Upa16, Wel15b]. maximal [Dwy90]. Maxime [Gal15, Mar15b]. maxitive [KL05]. may [KL10]. Mazur [Cad11]. Mazzola [LaT11]. McConnell [Jen01]. McGeoch [Dav05, Mar15a]. McGraw [Kel09b]. McGraw-Hill [Kel09b]. Mealy [Chi86]. mean [Gur86, Yuv76]. Measure [DP09, Sri10]. Measures [DLO05, Hav71, KL05]. mechanics [BT08]. mechanisms [CK04, Her00b, Puc04a]. medical [Ven02]. meet [Kei08b]. meeting [All00, FK89, Par00c, Lie90]. meets [AAH11, Zho09]. Mehlhorn [Elr11]. Mehlmann [Kri08]. Meirav [JH20]. membership [DBHHT94]. Memoriam [DW19, Ric94, Knu03a]. memories [Att11]. Memory [Sco15, AH83, Att08, Gra09, GM12, HS12, Kei12a, KVY12, Mof83, RGL10, SSC14, SB14, Wel14b, CH14, KR12]. Mendelson [Spr08]. Mental [Ros04a]. mentoring [Cun95]. Mergesort [Par98a]. merging [Nis78, San85]. merit [Tom89, Tom90]. Mermin [Rie10]. Mertens [Azi16, Gre16h]. Mesopotamian [Gas12a, Ses09]. Meta [Dav11, BHI05]. Meta-Level [Dav11, BHI05]. Metamathematics [Sha94, Gas19e, Pux97]. metatheory [Pap95]. Method [BHS80, Mar03b, Vem04, Kan93, Li12, Bor07a, Sri06, Vak13]. Methodologies [Che21]. Methodology [Dav05, GJM02]. Methods [CH10, Dav06b, FM15, Jen97, Kam13, Mik93, Mos05, Sch04, Ash72, Li14, MNV72, Nag14]. metrics [FRT04, O’R04a]. Metz [Gre22h]. Meulen [Che79c]. Meyer [Mat14a]. Michael [Azi12b, Bón09b, Cal10, Che80, Chi15, Dav05, Dav06a, FM15, Gas10a, Gas15f, Kat07, Kat13b, Kye09, Mas01a, Náv04, Neh99, Sta07, Ter08, To06, Raz07]. Michalewicz [Mas01b]. Michele [Wan12]. Michiels [Mar09a]. Microprogrammed [SvL74]. Microprogramming [Jon73]. Microsurveys [AP98, Mas00b]. Midwest [HK84a, HK84b, Ham90]. Mihai [Tho13]. Mika [Vio04, dW07]. Mikhail [Gas99a, Pap12b]. Miklós [Gas15b, Ist12]. Miller [Bel14, Wid03, Led16]. Milmeister [LaT11]. Milner [Pap10, PSTM00, Puc00, Puc01b]. Milutinović [Koz14]. Ming [Gas97a]. minimal [Dew79]. Minimax [Did93]. Minimization [DM22]. minimum [Sub95]. Mining [Mie11, SD08]. Minorities [Fri10, Ken05]. minus [Nah98]. Minute [Beh08, Gas09h]. Minutes [FK89]. Miodrag [Rey11]. Mirco [Cou09a]. Miscellany [Dud08, dOG10]. Miseries [Shi19, Wel19c]. Misha [Raz07]. mismatches [GG86]. Mitzenmacher [Kat07, Sta07]. Mladen [Hac09]. MMDS [KM10]. Mobile [Hem03a, Mil99, Puc00, SW01]. Modal [BdRV01, BvBW07, Sti01, Hes03, Wei03]. Model [McG72a, Mil17b, BK12, LP07, LP10, OB83]. Modeling [Ski01]. Modelling [BO06, Puc03, Kye09]. Models [CK02, CH10, Fer09, GH09, Hem20c, Kam13, KLMY21, Mah08, Tse91, Zom06, Ber72, Chr06b, Goo93, HN14, HR02, KO11, Mat97, Pat72, Reg93, Ros04a, SB14, Sin72, Ver06, Wel14b, DLO09, Kei10b, JK92, Sta11, Tay04, dOG12, Wil13]. moderated [Gas10e]. Moderately [Gol06]. Modern [AH06, Apo10, AB09a, Ben06, Bol98, BZ11, Ful10, Gol99, KL08, Mas01b, MF02, Mol05, Sga10, Tay02, Gas15d, Gri10, Cen00, Gre19e, Lee03b, Mil18a, Yan12]. Modification [Weg74]. Modularity [FKR08, Via08b]. module [FL83]. modules [Abr13, BF17]. Modulo [DR12, Ing83]. Mohit [Li14]. Moitra [Upa21]. Molasses [Ros01c]. Molecular [FC^{+}99, Neh99, JKL99]. Mollin [Ben06, Kat05]. moment [GH00, HG00]. monadic [Cad14]. Monographs [Mar09a, Pos99]. monotone [Gol77]. Monotonicity [Ves21]. monster [Chr04d, Ros04c, yea00b]. Montréal [Lew78]. Moore [Azi16, Gre16h]. Mora [Cou09b]. Morgan [Das92, FM08]. Morris [Hua98]. Morsels [Gre22h]. Moser [Gas07d]. Moses [Dek97]. Moshe [Dek97]. Moss [Dek98]. Mosses [Rie93]. Mossinghoff [Bón09b]. most [McG72b]. Motion [Mil09, Pap10]. motivated [Hav74]. Moulin [Dan06]. Moulton [FB14]. Mount [Mil17a]. Moving [Ven13b]. Mueller [Knu77b]. Muffin [GHMP19, Gre22h]. Mukherjee [Neu10]. Mullen [Nag15b, Sha12a]. Multi [Ali23c, Ass23]. Multi-Pass [Ali23c, Ass23]. Multiagent [SLB09, Azi10a]. multicore [HL08]. multihead [Ben74]. Multiple [Kao97, DH79b, vS15]. Multiple-size [Kao97]. multiplication [Ang76, ASU86, Boo77, Che74, Ing83, Khu12, Shy78a]. multiplications [dJvL75]. Multiply [Kre01]. Multiply-Add [Kre01]. Mummert [Sha12a]. Mung [Kat14a]. MuseumsForum [Bau10]. Mutual [Bur78, Ala03]. myotonic [KL10]. Myth [Kri08, Meh00]. myths [Ala03, Koz14, Gas15i].

N [Azn13, Mat14a, Pux97, Rie10, Sha19, dW07, Gur86]. Nadathur [Bel14]. NAE3SAT [Mor88]. Nahin [Bla06, Cen09]. Nanongkai [Ano13]. Nati [Ano13]. Natural [Bal97, Kap99, Mos05, Vol21, Esc93]. Nature [Azi16, Gre16h, Ste11b, Zom06]. Nature-Inspired [Ste11b, Zom06]. Navarro [Koz18]. NC [Ten87]. Neal [Gas10d]. Near [Dav05, GJM02]. Nearly [RS91b, Zeu92, RS88, RS91a]. necessary [Jaf78, dJvL75]. necessity [HKR83]. Needed [Att08]. needles [Dvi09]. Neighbor [Dav05, GJM02]. Nerode [Dek99a, MT04, TA95]. Net [BDK01, Mik02]. Nets [Jen97, Mik93, Old91, Cie88, Kos97, KT98, Pur93]. network [Afs82, Che07, PD98, Gol00a, Lou18b, Nag21a]. networked [DO13, FM06, Kat14a]. Networks [APR16, Fah05, FS19, Gav01, Jun99, Jun05, Jun08, Lib17, Lou18a, TV16, VR07, Wan12, Wel16a, Wel16c, Wel19b, GK09, HN14, HL09, Kei10b, Kei11a, Kei11c, KO11, OW09, Raj02a, Yu11, Cao13, Ist12, Kel16b, Kin11, Sil12]. Neumann [Her13]. Neural [Cie88]. News [AH91, ASvS23a, ASvS23b, Hem95, Hem96, Hem10a, HW12, Hem16b, Hem16c, Hem18c, Hem18a, Hem18b, Hem19d, Hem19a, Hem19b, Hem19c, Hem20a, Hem20b, Hem20c, Hem20d, Hem22b, Hem22d, Hem23a, Hem23b, HSvS19, HSvS20, Kei08a, Kei08b, Kos74a, Vol23a, Vol23b, Vol24, vS16b, vS16c, vS18a, vS18b, vS22, Bra90a, Che85, Chr04a, Chr04e, Chr04f, Chr05c, Chr05d, CKM06, Chr08b, Chr09a, Chr09b, Chr10a, Chr10b, Chr11, Hem01b, Hem01d, Hem02a, Hem02b, Hem02d, Hem02c, Hem03b, Hem03d, Hem04, Hem05b, Hem05c, Hem06c, Hem06d, Hem06e, Hem08b, Hem08c, Hem09a, Hem09b, Hem09c, Hem09d, Hem10b, Hem11a, Hem11b, Hem11c, Hem11d, Hem12a, Hem12b, Hem12c, Hem13c, Hem13a]. news [Hem14b, Hem14c, Hem14d, Hem22c, Kei08c, Kei09a, Kei09b, Kei09c, KM10, O’R06c, Puc05a, Puc05f, Puc05g, Puc06a, Puc06c, Puc07a, Puc10, Raj01a, Raj01c, Raj02b, Raj02c, Raj02d, Raj02e, Raj03a, Raj03c, Raj04a, Raj04b, Raj05a, Raj05b, Raj06a, Raj06b, Ros03a, Ros04d, Ros04e, Ros05c, SP07, Via05, vS12a, vS12b, vS14a, vS14b, Hem97b, Hem01c, Hem03c, Hem13b, Hem14e, Hem15a, Hem15b, Hem15c, Hem15d, Hem17a, Hua14, Joh96a, Raj01b, Raj03b, vS13, vS15, vS17b, Amb17, Hem16a, Hem17b, Hem17c, Hem21d, Hem21a, Hem21b, Hem21c, Hem22a, HSvS21, vS16a, vS17a, vS21]. Newsflash [FFG96]. newsletter [Kat78]. Next [Kei10c]. Nguyen [Nar14, Mil13]. Nicholas [Bón13a, Gup98]. Nievergelt [Ter11]. Nik [Sol12]. Nina [Mil17a]. nineties [All97]. Nipkow [Nar00]. Nisan [Lev10]. Nixdorf [Bau10]. No [AN09, Gas10a, Now02, KR01, Kos74a]. Noam [Lev10]. Noble [Bla06]. Nobody [Gre22h]. nodes [Ros77]. nomenclature [Ram89]. nominations [Kan14]. Non [BM96, Ell70, Vio19, BFK^{+}79, Coo83, LS78, Gre18e]. Non-abelian [Vio19]. Non-context-free [Ell70]. non-erasing [LS78]. non-hierarchical [Coo83]. non-oblivious [BFK^{+}79]. Non-Residues [Gre18e]. Non-Wellfounded [BM96]. Nondeterminism [Ric21, Chi86, GLM96]. nondeterministic [Gal74, LL75, Ree71]. Nonlinear [LS06, Mel11]. nonnegative [Ven13a]. Norah [Azn14]. Normal [TF82]. Norman [Low12]. normed [Sti99]. North [Che79a, Eil76, Her77, Sub78]. North-Holland [Che79a]. Noson [Cou09a]. not-ready-for-prime-time [Hem94a]. notation [Bra85, DR12]. notations [BA91, HV95]. Note [TV16, Wel16c, BBF^{+}73, BMO77, Buh96, Eng89, FL83, FMOP70, GJ78, GK80, Kan05, Kri77, KW81, Lei91, Mof83, Nou80, Ree71, Sch81, Sei74, Shy78a, Shy78b, Yel78]. Notes [Yel78, Dwo95]. Nothing [Bla06, Kap99, Ric73]. Notices [Mar08b, Hem08a]. notions [BA91]. novel [Kos97]. Nowakowski [Gas10a]. NP [BS98, DW13, GJ78, Gas10f, Kri77, Nag13b, Aar05, Dem09, Eps87, FGK^{+}02, Fra85, Gas12f, Gas19h, Gha23, GH00, GHSW14, Gol10a, Hem93, HHR97, Hoc97a, Hoc97b, Hua03, Knu74a, KS06b, Kri75, Raj05d, Sch82, Sto73a, Yli04, Nag13b]. NP-complete [GJ78, Kri77, Aar05, Fra85, Raj05d, Sto73a]. NP-completeness [Eps87, Gol10a, Nag13b]. NP-decomposable [Sch82]. NP-hard [BS98, FGK^{+}02, Hoc97a, Hoc97b, Dem09, Hua03, Knu74a, KS06b, Kri75]. NP-P [Hem93]. NP-Pairs [GHSW14]. NSF [Che85, Gas07a, Lat95a, Lat95b, Lat95c, Lee89, Lee90a, Lee90b, Ric00, Ros94a, Slo02a, Slo02b, Slo07, Zal97a, Zal97b, Zal97c, Zal99]. NSF-Funded [Ros94a]. null [Ben73b]. Numb3rs [Gas06e]. Number [Coh93a, Coh95, Coh96, Cou11, Dud09, Edw08, FM09, GKKK06, Gas10e, GHOP22, Gra82, GW18, Har11, Ios07, Kat05, Kat06a, Kel08e, Mao94, Pol09, Sha10a, Sha12b, Sho05, Sho09, Shp03, Wag03, Was08b, Yan02, Yan11a, Yan20, Fla05, Gab75, KL10, dJvL75, Bla06, Gre19f, Sar03]. Numbers [Bel11, Gas10c, MMW06, Pet02, Rei06, dK09, Esc93, Fer05, She15, Cad16, Gas11b, Gre15d]. numerical [HES72].

O [Kat05, Mar03a, Puc03, Zha05, dW07, Gur86]. obituary [Tho13]. Object [AK90, DFM07]. Object-Oriented [DFM07]. Objects [AC96, AKGN18, Wel18b, Zhu20, O’H98, Pos99]. oblivious [BFK^{+}79]. obtaining [Tho93]. October [Agu12]. Oded [Apo10, Lee03b, Nag13b, Upa20]. O’Donnell [Che80, Apo16]. offer [Cun95]. offine [Chr08b]. offline [Chr06a]. Olariu [Cha09, Wan12]. old [Mac95]. Olderog [Aro98, Pur93]. omega [VM85, Knu76a]. Omicron [Knu76a]. Omodeo [Irw03b]. On-Line [FM05, BM08, DLO05, Che74, Gal75]. One [BHHR99, Che24, Gas15h, MW93, Mik93, Pig15, Woo78, AM75, Gal78, IZ70, Nah98, Pet94, Wal82]. one-counter [Pet94]. One-sided [Woo78]. one-symbol [IZ70]. One-Tape [Pig15]. One-Volume [MW93]. One-way [BHHR99, Wal82]. O’Neill [Kye09]. Online [Amb17, ASvS23a, ASvS23b, BEY05, BFK^{+}16, Chr14, DM22, Fot11, HSvS19, HSvS20, HSvS21, Hua14, vS15, vS16a, vS16b, vS16c, vS17a, vS17b, vS18a, vS18b, vS21, vS22, Chr04e, Chr04f, Chr05d, CKM06, Chr08b, Chr09a, Chr09b, Chr10a, Chr10b, Chr11, FFMS94, IMP11, Jaw05, Lew14, Puc05h, Ros03a, Ros04d, Ros04e, Ros05c, vS12a, vS12b, vS13, vS14a, vS14b, Hua14, You99]. only [AM75, Boo77]. ontologies [RB11, Azn13]. Open [Bor09b, DM22, FH06, FGH18, For21, Gas16a, Gas17a, Gas17b, GFW17, Gas17c, Gas18a, Gas18b, Gas18c, Gas18d, Gas19d, Gas19a, Gas19b, Gas19c, GB20, Gas20a, Gas20b, Gas20c, GM21, Gas21b, Gas22a, Gas22b, Gas22c, Gas23b, Gas23c, Gas23d, Gas24, Gav01, Khu89, Khu90, Khu95, Khu96a, Khu96b, Khu97, Khu98b, Raj06e, RSZ21, Wil18, Wil06, FKW94, Gre22a, IZ70, JKL99, Pag70, Tho13]. Open-source [Bor09b]. operating [Gal74]. Operation [Kre01]. Operational [Fal12b, Hüt10, Weg72]. Operations [Che79c, VV78, Boo77]. operator [FMOP70]. operators [Coo83]. Opportunities [Har15, AJK^{+}97, Cun95, Lou91]. opportunity [Jon73]. Optimal [RS91b, Zeu92, Ang76, RS88, RS91a]. optimally [Bra83]. Optimization [AP01, Ali22, AMN22, AL07, Bor08, Bor09a, Cor01, Gre19g, Hro03, Iva02a, KV08, KV12, Lib17, Mas03, Oli04b, Pad10, SS02, Wil04, Fra84a, Fra84b, Hem06a, KV00, KV02, Kor06, SS06, Her13, Kum13, Li14, Nag14]. Optimized [Gas07h]. optimizing [FC83]. Oracle [For21, LS78]. Oracles [RST15, Bra90b]. Orchard [Nag15a]. Order [Che80, DP90, DP02, FGH18, Har79, Bel14, Cad14, CH06, DLM07, Coh07]. Ordered [Sch03, Nis78, Sch08b]. Orders [Mie11, SD08, dFR93]. organizer [Sug93]. Oriented [Abr11, Azi11, DFM07, Mar98, Mar08a]. Origami [DO07, Gar08, Gas09h]. Origins [Bau10, Kil12]. O’Rourke [Dek99b, FM11, FM13, FM14]. orthogonal [Tam93]. orthogonalization [Koz87]. OSDI [Cle08]. Osherson [Smi00]. Oswald [Gre18e]. Otfried [Mas00a]. Other [Ada09a, Gas09h, Gas10e, Hay08, Knu97, McN99, Pin10, Fas04, GS80, GS08, HV95, Moh93, Ros02b]. Oudot [BF17]. Our [Ali23b, FdV23, Gol09, Lou18a, Ram89]. outer [Toy93]. outline [Sta89]. overcome [Kei11c]. Overmars [Mas00a]. Overview [Ali20a, BCN20, Gav01, DO13, Khu12, Suc98]. Oxford [Bla06].

P [Cha12, DW13, Gas10f, Ham13, Mar09a, McN04, Mou10, Nag21a, Puc03, Sri10, Wil10a, Zha05, CT23, Die04, Gas12f, Gas19h, Gol10a, Gol77, Hem93, Lad75, Mor88, Yli04, Nag13b]. P=NP [Gas15c]. Paar [Mou14]. PACELC [Wel18a, Gol18]. Pach [Gas07d]. Packet [Ves21, Gol10b]. Packing [Bor08, Cor01, Moh93, vS15]. Padberg [Oli04b]. paged [Mof83]. Pages [Gas15h, Ada11, Bea07, Bón10b, Bón10a, Boo10, Cha09, Cou10, Fri10, Gas10b, Gas10e, Gas22f, Gre22h, dOG10, Jou94, Mik93, Mou10, Nar00, Oli04b, Pan93, Pin10, Tra22, Wil04, Fal07, Gas07c, Gas10d, Gas10c, Har10, Joh10, Rie10, Rot07, Sga10, Wil04, Wil10d]. Paging [DLO09]. Pairs [GHSW14, Ros77]. Palash [Kat13a]. palindrome [Gal75]. Panario [Nag15b]. Pancakes [FPE04]. Panconesi [Sri10]. Pankaj [O’R92a]. panorama [BCG^{+}22]. Panos [dV18]. Papadatos [Gas10b]. Papadimitriou [Gas10b, Kel09b]. Paper [Ali21d, Fis08, Gas09b, Par88, Par00b, Pet08, Rou05, MM99]. paperback [Oli04c]. Papers [FC^{+}99, Gas21d, Khu98a, Knu96, Knu00, Knu03b, Maz14, McN01b, Oli04c, Rag91, Apo14b, Apo14a, Gas14e, Gas14g, GJM02, Jaw05, Puc05h]. Paradigm [Cou09b, Kay18, Mor05, Ros93c]. Paradigms [Mav09]. Paradox [Meh00, Wap05, Gas10g, Kri08]. Paradoxes [Gas09h]. Parallel [Amo20, CS91, Das92, FGT17, Goo93, Iva02b, Lei92, Mat97, Par86, Sch06, Thu90, Wid03, BP05, CP13, HHR97, HG00, Koz87, Mil72a, MB05, Mor76, Qia87, Vis91, vL74, vL76, Pop12]. parallel-rewriting [vL74]. Parameter [FGH18, BI08, Kei08d]. Parameterized [BJK^{+}16, Chi15, DF99, JH20, Wel16b, HW94, Ist00]. Paris [GP89]. Parkes [McN04, Wil10a]. Part [Puc06b, She91a, She91b, AFG^{+}96, BCRV03, BCRV04, Can06, CR03, Gol93, Gol94, PGW^{+}96, You92, Knu11, RS91b, vS16b, vS16c]. Partial [BS08, Bón09a, Mie11, Raj03d, SD08, dFR93, CM72]. Partially [Dun00]. parties [Kei13b, Küp13]. Parting [All23]. partition [GL02]. partition-tolerant [GL02]. partly [Ing83]. Party [HL10, Kat12]. Pass [Ali23c, Ass23]. password [JM07]. path [Kan88]. pathogenic [KL10]. paths [Ram97]. Patrascu [Tho13]. Patricia [Fri10]. Patrick [Hes03, Kat13b]. Pattern [ABM08, Lab12, Neu10, Val09, LO94]. Patterns [LRV10, Sol12]. Paul [Ber04, Bla06, Cen11, Cou11, Gas97a, Ros23, Sch06, Yan12, Cen09, Par96]. Pavel [Pap13a]. Pavicic [Hac09]. Paxos [Ros03b]. PCP [AAV13, Mos19]. PCPs [Din08, Kho05, Khu05b, Mos12]. Pea [Gas10g, Wap05]. Pearls [Eri10, Sza12]. Pearson [Hen09, Kel09b, Rog08]. Pearson/Prentice [Hen09, Kel09b]. Pearson/Prentice-Hall [Kel09b]. Peeper [AS91]. Peiffer [Nag13a]. Pelzl [Mou14]. Pemantle [Bón14]. People [Gas14f]. Perception [Ros99c]. Perfect [FGT17, GH00, HG00]. Performance [Ali21c, DLO05, Har21, Mof83]. Periodic [Lie75, Mat94]. Perl [Lab12, Wan05, Dwy03, Val09]. Permutation [LRV10, Afs82, Sol12]. permutations [Bón08a]. persepctive [Chr06a]. Persistence [BF17, Zhu20]. Perspective [BB21, Gol08b, Zim04, Chr08b, GW97, Gol08a, Apo10, Lib17, Oli06]. Perspectives [Ali21d, Yan96, Cou14]. Pessimal [BS84]. Pessimization [Dun87]. Peter [Azi14, Azn13, Elr11, Jan10a, Kam13, Rie93, Tay04, To06]. Peters [Gas10g, Con03]. Petersburg [Kuz19]. Peterson [Gas10c, Ban14b]. Petković [Rey11]. Petkovsek [Kre00]. Petr [Bla06, Koz14]. Petreschi [Kin15]. Petri [Mik93, BDK01, Jen97, Kos97, KT98, Mik02]. Petzold [Wil10c]. Pevzner [Pap13a]. Ph.D. [Pas99]. PhD [GL09]. PhD’s [Fri11]. Phenomena [BM96]. Philip [Spr08]. Philippe [Bón10b]. philosophical [Cha73b]. Philosophy [Mor03, GS08, Pin10]. Phong [Nar14]. phrase [But71]. Phuong [Mil13]. Phylogenetic [Kel16b, FB14]. physical [Aar05, Raj05d]. Physics [BL06, Gre07]. Pi [Bla06, Bla06]. Picture [Coh08]. piece [Gre22h]. Pied [BR99, Gas11g]. Pierce [Kin05, Kin06]. pigeon [Coo76]. Ping [Fab11, Gre16g]. Pioneering [GL09, Fri11]. Pipher [Mei10]. Pippenger [Gup98]. Pittenger [Mar03a, dW07]. PKIs [KGO10]. PL [COJ78]. PL/CV [COJ78]. place [Pas99]. Planar [Aga91, Mor88, O’R92a, Sto73a, Sto73b, CK89, Gol77, Kan93, Tol93]. planarity [dFR93]. Plane [Gho07, Dew79, Hem08a, Mar08b, Moh93, Dea09]. Plans [Ves21]. Play [Ber00, She13]. Playing [BCRV03, BCRV04]. pleasant [Agu04, Khu04]. Pless [Ayu12]. Plotkin [Puc01b]. Plummer [Azi12b]. Pnueli [Sis93]. PODC [ABLO23, Ali23a, BBDD20, Bar10, BD18, BDCDL19, BS13, Cas08, CL09, Dav17, Den14, FdV23, Hel11, Sch16, Tse15, Sen12]. PODS [Afr05, Are18, Ben12, Cal15, Fan13, Gee17, Got00, Gro14, Hul96, Koc19, Len08, Pic21, Raj05e, Su09, Tan16, Van10, Van06, dB06, Par98b]. PODS’95 [Abi95]. PODS’97 [Özs97]. Poetry [Rag91]. Point [Bar15, Mac20]. Pointer [BA95]. Polak [Sga13]. policies [Gol10b]. Policriti [Irw03b]. Political [Cra03a, Cra03b, Nat07, Nat10]. Politics [DL98, DL07, Jan08b]. Poll [Gas19h, Gas12f]. Pollack [Cou11]. Polya [Vak13, Mah08, Sta11]. Polyhedra [DO07]. Polyhedral [FM15]. polylog [DLM07]. Polymorphism [O’H98]. Polynomia [Kum12]. Polynomial [Cou09b, Die04, FS15, Kat06b, Mor03, Mor05, RST15, Gal74, Has85b, HHH98, Her00b, Hua03, Sch81, Sim76, Sto73b, You92, dJvL75, vL78]. polynomial-time [Hua03, Sch81, Sim76, You92]. polynomially [AKR83]. polynomials [FPE04, Vio09]. POPL [Rie99b, Weg86]. Popular [She15].

Population [AG18, Wel18c]. Portugal [Gre17g]. posed [Lan10]. positions [Sta87]. Possible [Fri10, Ken05, Nah04]. Post [BCRV03]. postage [Sha02]. Postscript [Knu74a]. potential [HR00]. potential-based [HR00]. Pottmann [Ess05]. Power [Kin15, Lou18a, GLMW96, IP05, Puc05b, vS12a, Nag18a]. pp [O’R92a]. Practical [AL07, Gas20f, Jen97, Mik93, Koz18, Nir86, Bor09a, Mil17a]. Practice [AK20, Kuz18, Kuz19, MP21, Har95, Hua03, Kei09c, KR09, OW09, RGL10, Sti95, Sti02, Sti06, Wil08, Zho09, Spr03, Wan12]. practitioners [Mou14]. Pratt [Hua98]. pre [Fri11, GL09]. pre-1940 [Fri11, GL09]. predecessor [FFB89]. Predicate [Hes92, Man92, Mak93]. Predicate-Transformation [Hes92]. prefix [Ben73a, Ten87]. prefixes [Sei74]. preliminary [CR74, Gra82, Joh84, KR01, SY87]. Prentice [Gol00b, Hen09, Wid03]. Prentice-Hall [Gol00b, Kel09b]. preorder [CH93]. Preprocessing [JH20]. Present [Par88, Par00b, MM99]. Presentation [GP89]. preserved [SW72]. preserving [FMOP70, Kos74a, Kos74b]. Presidential [Gas10e]. Press [Aar07, Apo10, Azi10a, Azi10b, Ben98, Bla06, Bón10b, Bón10a, Cou09b, Cou09a, Dan06, Dav06a, DL79, Dea09, Dek98, Dek99b, Gas10a, Gas22f, Gup98, Har10, Hes03, Irw08, JK92, Jan10a, Kat05, Kat06a, Kat07, Kin05, Kin06, Lee10, Les06, Lev10, Lou05, Mak99, Mak93, Mos05, Mou10, Nar00, Nat07, Nat10, O’R92a, Pan93, Pap10, Pos98, Pri10, Puc03, Pur93, Rie93, Rie10, Rot07, Sga09, Sga10, Smi09, Spr08, Spr09, Sri10, Sta07, Sta93, Wan05, Wil04, Wil10b, Zha04, Zwa93, Avi01, Ber04, Coh07, Gas07b, Sot10, Spr03, Spr08]. Presses [Lew78]. pretty [Har95]. Priestley [Coh07]. Primal [Hua14, Chr07]. primal-dual [Chr07]. Primality [Die04, Vél78, Cou12, Gre16e, Kat06b]. prime [Hem94a, Mor76]. Primer [OW17, You92]. Primes [Die04, Gre19f]. Primitive [FFB89, Sha92]. princess [Chr04d, Ros04c]. Princeton [Azi10b, Bla06, Har10, Spr09, GBGL08]. Principle [Wil06, Coo76, Lap91, Aar07]. Principles [ACM12, Ano13, DK02, DK07, Lou18a, Raj00, TSS03, Zha05, Kei08e, Gas16d]. Prinz [Gre22h]. priority [GJM02]. Privacy [DL98, DL07, Pap05b, Sal03, Lew14, Jan08b, Mak99]. Prize [Alp18, Ano13, Pel13, Kei13b, Mos10, Sch18, Ten18, Kan14]. Prizes [All23]. Probabilistic [BC^{+}91, Kat07, Lee03b, MU05, Sta07, Suc08, Ell70, Fra85, Gol99, Mac95, Paz70, Kri99]. probabilities [KL05]. Probability [AP98, Dur10, Gas09h, MU05, Sri18, Mas00b, Bón11b, Kat07, Sta07]. Probably [Bru13]. Problem [Abr11, ABCC06, Azi11, Ben23, BGG97, Che82, Gas14i, GHMP19, Mar98, Mar08a, Mar20, Mat93, Ros73a, Ros73b, Spr09, Yli04, Ben73b, Bur87, Dem09, Din13, Dun86a, Eng89, Eng18, FKW94, GC86, GN07, Hav74, Khu02, Koz78, Kri75, Lad75, Lap91, Lit03, LP92, Pag70, Qia87, Raj07b, Ram97, San85, Sch97, Sha02, Sta87, Sub95, Tur84a, Tur84b, Wei98, Wil86, Yli05, Dun86b, Gas21a, Wei95e, Pru00, Sim04, Gas14h]. problèmes [Knu76b, Lew78]. Problems [AP01, Ali21d, BB21, Blu83, BMP05, CEG98, DM22, FM05, FGH18, Gas07d, Gas16a, Gas17a, Gas17b, GFW17, Gas17c, Gas18a, Gas18b, Gas18c, Gas18d, Gas19d, Gas19a, Gas19b, Gas19c, GB20, Gas20a, Gas20b, Gas20c, GM21, Gas21b, Gas22a, Gas22c, Gas22e, Gas23b, Gas23c, Gas23d, Gas24, Gav01, Gre22f, Hoc97b, Hro03, Jam98, Knu97, Mac20, Mas03, McN99, Nag15a, Oli04b, Par95a, RSZ21, SS07, She08, She10, Wil18, Aar05, BCG^{+}22, BBG94, BCRV04, BI08, Cal83, Che11b, Chr14, Con95, Cud69, Fra84a, Fra84b, Fra85, Gal74, GJ78, Gas13c, Gas14f, Gas22b, GL80, Gol77, Hem06a, Hua03, IP05, JKL99, Kei08d, Khu89, Khu90, Khu97, Khu98b, Knu74a, Kri77, Pac93, Pas99, Pat72, Puc05b, Raj05d, SS06, Tho13, Ven02, WM87, Hoc97a, Khu95]. Problems [Khu96a, Khu96b, BS98, Par11, Sri14, Gas13c]. Probst [Cho12]. Procaccia [Nag17b]. Procedures [BM07, COP01, Fur11, Irw03b]. Proceedings [Gol00a, LRV10, ACM12, Agu12]. Process [Ace92, Zwa93, Cal12]. Processes [Ali21d, BB21, Old91, SW01, Sti01, Mit01b, Sti99, Cal12, Wei03, Hem03a]. Processing [Zhu20, Iva02b, Zam77, Sub78]. processor [AH83]. Professor [Ste09, Gas09h]. Program [Ash72, Nag23, CP74, GL72, Gas07a, Ham76, Mil72a, Slo02a, Slo02b, Slo07, SvL74, XL93, Zal97b, Zal99, Lat95a, Lat95b, Lat95c, Ric00, Sim84a, Sim84b, Zal97a, Zal97c]. Programmable [FRRS17, Wel17b]. programmation [GP89]. Programmer [Kel09a, Lei07]. Programming [Bra77, COP01, Che79b, Che79c, Che80, COJ78, FA03, Gun92, Hof00, Irw03b, Jam98, Kin05, Kin06, Knu11, LS06, MW93, Nag20, Pie02, Rie97, Rog08, Rog18, Ros00c, She08, She10, SR03, Sta93, Tau06, VV78, All72, BBAGP09, CW72, Esc93, Gol95, Kan94, MR97, Pru98, VS98, Weg72, Wol98, vdBN04, Cha05, Hen09, Mel11, Oli04a, Rog13, Vla94, Bel14, Nag23]. Programs [AO91, AO97, AdBO09, CC76, Hes92, Hua14, Joh10, KA89b, Weg00, BK12, Bur72, KA89a, Kei12b, MNV72, Sin72, Aro98, Pan93, Ree77]. progress [Joh90, Khu12]. Project [Che20, Coo04, Ros94a]. Projection [Sri06, Vem04, Mos12]. projective [Cou14]. Projects [Ros94a, Abr13]. Prömel [Gas17d]. Proof [Avi01, BQ03, Bla12, Coh08, GHSW14, GS08, Gol93, Gol94, Jou94, Mil13, PSTM00, Raz16, Sha94, Sny91, TS96, TS00, BA04, Ben73a, Coo76, DLM07, McB87b, Pal90, Sei01a, Swe87, Weg74, Gas05b, Lew13, Pux97, Pin10, Puc01b]. Proof-by-Picture [Coh08]. Proofs [AZ98, AZ99, AZ01, AZ04, AZ10, BQ03, Lak76, Lee03b, Vol21, dRGR22, CR74, Gol99, Ros96a, Gas14f, Bón11c, Cad13, DL79, Gas01a, Gas05b]. Properties [Bra92, Eil76, Gin75, Hen93, Ros73a, Ros73b, Sti01, SW72, Wei03, BHHR99, Bha13, Lie75, MNV72, Sin72, Sud11, VS98, Gre22g]. Property [Upa20]. proposal [Knu74b, Ven99]. proposed [DR12]. Propositional [GHSW14, CR74]. Propositions [DH79b]. Propp [Mas00b]. proteomics [Pap13b]. Protocol [Blu83, Bra91, Eve83]. Protocols [AG18, AT17, Bel07, HL10, Kat12, Wel17a, Wel18c, Bra90a, Can06, Puc06b, Puc03, Sta10]. prototype [BB89]. Provable [Gre22g]. Prove [Smi09, Vel94]. Providence [Sch09]. Proving [Che79a, Gol18, Lov78, Yeh12, MNV72, Wel18a]. Pruim [Kri00]. Pseudorandomness [Lee03b, Shp03, Gol99, Vad07, Sha10a]. PSPACE [Con95]. Pub [Das92]. Public [Cou12, Gas07c, Gas16h, Kat05, Mol03, BB87, Hen83, Jor83, Ros02a]. Public-Key [Kat05, Mol03]. Publication [Oli04c]. Published [Ada11, Bón10a, Dav09, Gas10g, Mav09, Apo10, Bón10b, Cou10, Fri10, Gas10b, Gas10e, Gas10d, Gas10a, Gas10c, Mou10]. Publisher [Cal10, Ful10, Mur10, Nat10, Wil10c, Wil10b]. Publishers [Bor07a, Che80, FM08, Kre07]. Publishing [Che79a, Fre01, Gas07b, Her77, Sub78]. Pulse [Ros99c]. pumping [Col93, Jaf78, SW82]. Purely [Zhu20]. Purpose [Gas16h]. Pursuit [Azi12a, Coo12, She14]. pushdown [OS14, Sti99]. Puzzled [Nag23]. Puzzler [BR99, Gas11g]. Puzzles [Apo11, HD09, PV21, Pet09, Rey11, FPE04, Nag23, Gas13d]. PWS [BS98]. Pythagoras [Gas15i]. Python [Jan10b, RZ09].

Q [Nar14, Ing83, Gas19g]. Quadratic [BV07, Col10, Gre18e, Gre18e]. Quantificational [Che82, Lew79]. quantifier [LL75]. quantifiers [Gra82]. Quantitative [Oli06, Zim04]. Quantum [Amb04, AdW17, BB87, Bra94a, Bra94b, BHT97, BT20, CCD15, Che24, Fen21b, GM10, Hac09, HS01, Hir01, KSV02, Mar03a, MP21, Mer07, OW17, Pap05a, Pit00, Ree01, Sga12, Sin99, Wie87, YM08, dW07, AAV13, BB89, BGH07, BC96, BT08, DKW05, DP08, Gha23, Kei08b, TS99, Wat11, Wel14c, dW07, Gas19g, Cou09a, Gre13, Gre16f, Gre17g, Mou16, Rie10, Sga13, Vio04]. Quarter [Had15]. Quarter-Century [Had15]. quasilinear [RW94]. Queen [Ree01, Sin99]. Queries [For99, GM99, Pal90, Pap95]. Query [AK90, GLMW96, Lib99]. querying [Kan94]. Quest [Gas22e, Wei15]. question [Gas10f, Lip10]. Questions [CEG^{+}10, For21, GW18, FH06, IZ70, Kat14a, Raj06e]. queues [GJM02]. Quinn [Gas05b].

R [Aro98, Bea07, Cad11, Cao13, Cen14, Che82, Chi15, Dav06a, Fur11, Gas99b, Gre19e, Ist00, Kat05, Kat14b, Kin15, Li14, Lib17, Lur99, Pur93, Spr03, Wid03, Val09, Lab12, Mil17a, Mil17b]. Rachel [Gas20d]. radix [Lee72]. Raised [GW18]. Raising [HHR97]. Rajsbaum [Upa16]. Ralf [Wil04]. Rambau [Mur13]. Ramsey [Gas19e, Gas17d, Gas16c]. Ran [You99]. Randall [Kri00]. Random [Bol85, Bol01, Gab10, Ist12, Kob08, Law06, RST15, Sri06, Vem04, VGS05, Zim13, Mit01b, SvL74, Bón09c, Gas10d]. Randomization [Gas13f, Hro03, Mas03]. Randomized [Die04, DP09, Kat07, Mav09, MU05, MR95, Sha92, Sri10, Sta07, Gol95]. Randomness [DH10, Gas15h, Teu13, Dvi09, Sha19]. range [GK80]. rank [Mat14b]. ranking [Mat14a]. ranks [Ang91]. Rao [Cad21]. rarely [Par02a]. Rat [Lit92]. Rate [KS16]. rating [Mat14a]. Ratio [Cad16, Ves21]. Ravi [Li14]. Ravikumar [Lei91]. reachability [Kos97, KT98]. Reactive [Dav06b, MP92, Sch04, Sis93]. Read [All23]. ready [Hem94a]. Real [BCSS98, McN01a, Pet04, Sch00, Fer05, Gal75, dV18]. Real-Time [Pet04, Gal75]. Real-World [dV18]. realistic [HN14]. Reality [Kas05, Aar05, Raj05d, Gas15e, Gas06e]. Really [BQ03, Gas05b, Mat14b]. Realms [Kum12]. Reasoning [Cha98, FHMV95, FHMV03, Hal03, BHI05, Gal14, Dav11, Dek97, Zha04]. Rebecca [Gre16e]. Recently [GW18]. Recipe [Cha20]. Reciprocity [Gre18e]. recogniser [Jon72]. Recognition [Pig15, Gal75]. Recognize [Ros73a, Ros73b]. Recombination [Kel16b]. ReCombinatorics [Kel16b]. Reconfigurable [FS19, Wel19b]. Reconfiguring [Kei10a, LMZ10]. Reconstruction [Dey07, Sot10]. record [CH06, DKW05]. record-breaking [DKW05]. Recreational [Gas23a]. Recurrences [BHS80, Kao97]. Recursion [BA04, Che21, For99, GM99, Hes92, KA89b, Pan93, FFB89, GL72, KA89a, Pru98, SW72]. Recursion-Theoretic [BA04]. Recursive [CM72, Kat78, Lon71]. Redondo [Kin11]. reducibilities [Sim76]. reducibility [LS78]. Reducing [GI91]. reduction [Dew86, GN07, Raj07b]. reductions [GH00, You92]. Redux [Gas06e]. Reed [Jan10b]. reentry [Sim84a, Sim84b]. referee [Weg86]. refereeing [Smi99]. referees [Par89a, Par94]. Refinement [Ace92, Zwa93]. Refinements [Gas21a]. Reflections [Ros98a]. Reflects [Gas22d]. Refutations [DL79, Lak76, Gas01a]. Regan [Gas14f, Gre16f]. Regular [BMMR20, Dos87a, Dos87b, GPP83, Jaf78, Lei80, Rup91, Sei74, SW82, YZ91, Bir91]. Regularity [Kos74a, Kos74b, FMOP70, vL76]. regularity-preserving [FMOP70]. Regularization [SS02, Wil04]. Reid [Ban14a, Bel11]. Reimann [Gas19e]. Reingold [Zeu92]. reinvented [BB87]. Reischuk [Kin15]. rejected [vLL76]. Related [FC^{+}99, Kum12, BFK^{+}79, Cal83, Fic97, GJM02]. Relation [Knu97, McN99, TF82]. relations [Knu76b, KGO10, Lew78]. Relationship [Old91]. Relativizations [Boo84a, Boo84b]. relativized [Bra83]. Reliability [GK09, OS85]. reliable [Zho09]. remain [Ang91]. Remembering [Sim22]. Remembrances [Fen21a]. Reminiscences [Koz22]. removal [SW72]. Rempe [Gre16e]. Rempe-Gillen [Gre16e]. Renaissance [Ses09, Gas12a]. Renato [Gal92, Gre17g]. rendezvous [Che06]. Reniers [Cal12]. reordering [Eng18]. repair [BK12]. repairability [DO13]. repeated [LO94]. repeats [KL10]. repetition [CP13]. replication [Kei13c]. Report [Abi95, Afr05, All00, Are18, Ben12, Cal15, Che20, CS91, Fan13, Fei89, Fei90, Gee17, Got00, Gro14, Ham90, Hig01, Hul96, Kel14d, Kel15a, Kel15b, Kel15c, Kel15d, Kel16c, Kel16d, Kel16e, Kel16f, Kel17a, Kel17b, Kel17c, Kel17d, Kel18a, Kel18b, Kel18c, Kel18d, Kel19d, Kel19a, Kel19b, Kel19c, Kel20a, Kel20b, Kel20c, Kel20d, Kel21d, Kel21a, Kel21b, Kel21c, Kel22a, Kel22b, Kel23a, Kel23b, Kel23c, Kel23d, Kel24, Koc19, Kud95, Lie90, Lon97, Lon98, Özs97, Par00c, Pic21, Pro90, Raj05e, Rie99b, Rod96, TT95, Tan16, Thu90, Wan97, Wei95c, Wei95d, Wei96b, Zeu94, BS09, FFMS94, Gra82, Hag99, Hag02, Hem05a, Hem06b, Hem07c, Hem07d, Hem08d, Hem10c, Hem10d, Joh84, Joh90, Kel05, Kel06a, Kel06b, Kel06c, Kel06d, Kel07a, Kel07b, Kel07c]. report [Kel07d, Kel08a, Kel08b, Kel08c, Kel08d, Kel09c, Kel09d, Kel09e, Kel09f, Kel10a, Kel10b, Kel11a, Kel11b, Kel11c, Kel11d, Kel12a, Kel12b, Kel12c, Kel12d, Kel13d, Kel13a, Kel13b, Kel13c, Kel14a, Kel14b, Kel14c, Kel22c, Khu06, Lee89, Lee90a, Lee90b, Len08, Mor97, Par98b, Rav90a, SY87, Sin98, Slo02a, Slo02b, Sta74a, Su09, Van10, Van06, Vit00, Wei93, Zal97b, Zal99, dB06, Hag00, Lub96, Ric00, Wei95a, Wei95b, Wei96a, Zal97a, Zal97c]. reports [FFMS94, KB79b, New76, Sta73a, Sta73b, Sta73c, Sta74b, Sta74c, Sta74d, Sta75a, Sta75b, Sta76b, Sta76c, Sta76d, Sta77a, Sta77b, Sta77c, Sta78b, Sta79a, Sta79b, SIG80, SIG81a, SIG81b, SIG82a, SIG82c, SIG83b]. repository [Ski99, Ven99]. representable [Paz70]. representation [KL05, Moh93, Nir86, XL93]. representations [Tol93]. Representing [WM87]. Reprint [Lou89, Lou90, Lou96a]. Reprints [Lou95a, Lou95b, Lou95c, Lou96b, Lou97, Sei97a, Sei97b, Sei97c, Sei98a, Sei98b, Sei98c, Sei98d, Sei99a, Sei99b, Sei00a, Sei00b, Sei00c, Sei01b, Sei01c, Sei02a, Sei02b, Sei02c]. Required [Ros73a, Ros73b]. requirements [Di 93]. requiring [AM75]. Research [BMP05, DM22, Gas23a, Hem08a, Mar08b, SS07, SIG82b, Wil06, AAC^{+}08, All00, BCvR09, Cas02, CFF^{+}97, Fic97, Gas08b, Kei08e, Lee89, Lee90a, Lee90b, Ros03e, SIG83c, Ros95d, Cen11, Gas07d]. researchers [Lan82]. resetting [Iba91]. Residues [Gre18e, Gre18e]. resolution [Coo76, Kau93]. Resource [BFM^{+}15, Wel15c, Lon71]. Resource-Competitive [BFM^{+}15, Wel15c]. Resources [Gas00f, Hem94b, Abr13]. respect [GH00]. response [GMN87, Grä08]. responsibility [OS85]. restated [Col93]. Restricted [GHOP22]. result [Mor76]. Results [Ali21b, DM22, HH76, Ray21, Che74, CK98, Fra85, HW94, Kho05, Khu05b, KGO10, Ram97, Sim76, YDX97, Yli04]. Retour [FFGH08, DCV82]. Retrieval [BCC10, Di 95, Lea97, Rub13, Gri10]. Retrospective [All98, HS97, Pud06, Ros96d]. Reuben [Kin12]. Reverse [Gas16d, vL78]. Review [Aar07, Aaz18, Abr11, Abr13, Ada11, AF20, AK20, Ali19, Ali20d, All98, Amo20, Amo99, AY12, Apo11, Apo12, Apo14b, Apo14a, Apo16, Aro98, Ass17, Avi01, AG15, Ayu12, Azi10a, Azi10b, Azi11, Azi12a, Azi12b, Azi14, Azi16, Azn13, Azn14, BBDD20, Ban14b, Ban14a, Ban15, Bar12, Bea07, Bel11, Bel14, BF17, BD18, BDCDL19, Ben06, Ben98, Ber04, BK18, Bla06, Bón09b, Bón09a, Bón09c, Bón10b, Bón10a, Bón11a, Bón11b, Bón11c, Bón12, Bón13a, Bón13b, Bón14, Bon10c, Boo10, Bor08, Bor07a, Bor09a, Bor07b, BFU19, Bru13, Bur15, Cad11, Cad12, Cad13, Cad14, Cad16, Cad21, Cal10, Cal12, Cal13, Cao13, Cen00, Cen03, Cen11, Cen14, Cen15, Cha09, Cha12, Cha05, Che10, Che79b, Che79a, Che79c, Che82, Chi15, Cho12, Coh07, Col10, Con03, Cou09b, Cou09a]. Review [Cou10, Cou11, Cou12, Cou14, Cou15, DGW08, Dan06, Das92, Das09, Dav02, Dav05, Dav06a, Dav06b, Dav08b, Dav08c, Dav08a, Dav09, Dav11, Dav17, DL79, Dea09, Dek97, Dek98, Dek99a, Den14, Dhu19, DW13, DM22, Dye11, Eil76, EA09, EA10, Elr11, Ess05, Fab11, Fah04, Fah05, Fal07, Fal12a, Fal12b, FM08, FM09, FM11, FM13, FM14, FM15, Fen21b, FM17, Fit11, FB14, For04, Fri10, Fri11, Ful10, Ful12, Fur11, Gal21, Gal22, Gal13, Gal14, Gal15, Gas94, Gas97a, Gas97b, Gas98a, Gas98b, Gas98c, Gas98d, Gas99a, Gas99b, Gas99c, GK99, Gas99e, Gas00b, Gas00c, Gas00d, Gas00e, Gas01b, Gas01a, Gas02b, Gas04a, Gas05b, Gas05a, GKKK06, Gas07c, Gas07b, Gas07d, Gas09a, Gas09b, Gas09c, Gas10b, Gas10e, Gas10d, Gas10a, Gas10c, Gas10f, Gas10g]. Review [Gas11a, Gas11b, Gas12a, Gas12e, Gas13d, Gas13f, Gas13e, Gas14c, Gas14e, Gas14h, Gas14f, Gas14g, Gas14i, Gas15a, Gas15c, Gas15i, Gas15d, Gas15e, Gas15f, Gas15g, Gas15h, Gas15b, Gas16b, Gas16e, Gas16c, Gas16f, Gas16d, Gas16g, Gas16h, Gas17d, Gas19f, Gas19e, Gas19g, Gas20d, Gas20e, Gas20f, Gas21c, Gas21d, Gas22f, Gas22d, Gas22e, Gas23a, GG21, Gle05, Gol00a, Gol00b, Göö12, Gop08, Gre07, Gre13, Gre15a, Gre15b, Gre15c, Gre15d, Gre16a, Gre16b, Gre16c, Gre16d, Gre16g, Gre16h, Gre16e, Gre16f, Gre17a, Gre17b, Gre17c, Gre17d, Gre17e, Gre17f, Gre17g, Gre17h, Gre18a, Gre18b, Gre18c, Gre18d, Gre18e, Gre19d, Gre19a, Gre19b, Gre19c, Gre19g, Gre19e, Gre19f, Gre20a, Gre20b, Gre20c, Gre20d, Gre20e, Gre21d, Gre21a, Gre21b, Gre21c, Gre21e, Gre21f, Gre22b, Gre22c, Gre22d, Gre22e, Gre22f, Gre22g]. Review [Gre22h, Gri10, dOG10, dOG12, Gup98, Guz08, Ham13, Har10, Har11, Hau12, Hem03a, Hen09, Hen93, Her00a, Her77, Her13, Hes03, Hom17, Ira12, Irw03b, Irw03a, Irw08, Ist00, Ist12, Iva02a, JH20, Jam98, Jan08a, Jan08b, Jan10b, Jan10a, Jen01, Joh10, Jou94, Kam13, Kar11, Kat06a, Kat06b, Kat07, Kat12, Kat13b, Kat13a, Kat14b, Kat14a, Kel16a, Kel16b, Kel08e, Kel09a, Khu98a, Kil12, Kil14a, Kil14b, Kin05, Kin06, Kin11, Kin12, Kin14, Kin15, Kli08, Kob98, Kon17, Kop11, Koz14, Koz18, Kre00, Kre07, Kri99, Kri00, Kri08, Kum12, Kum13, Kuz19, Kye09, LaT11, Lab12, Lab13, Lam13, Law06, Lea97, Lea00, Led16, Lee03a, Lee03b, Lee10, Les06, Lev10, Lew78, Lew13, Lew14, Li14, Lib17, Lou05, Lou18a, Lou18b, Low12, Lur99, Mac20, Mai80]. Review [Mak99, Mak93, Mar03a, Mar15a, Mar15b, Mar20, Mar09a, Mas00b, Mas00a, Mas01a, Mas01b, Mas02, Mas03, Mat14b, Mat14a, Mav02, Mav09, Maz14, McN04, McN99, McN01a, Mei10, Mel11, Mie11, Mik02, Mik93, Mik11, Mil13, Mil17a, Mil17b, Mil18a, Mil18b, Mor13, MT04, Mou14, Mou15, Mou16, Mou10, Mur10, Mur13, Nag13b, Nag13a, Nag14, Nag15b, Nag15a, Nag16, Nag17a, Nag17b, Nag18a, Nag18b, Nag19, Nag20, Nag21b, Nag21a, Nag23, Nar14, Nas13, Nat07, Nat10, Náv04, Neh99, Neu10, O’R92a, OR99a, Ogi01, Oli04a, Oli04c, Oli04b, Oli06, Pan93, Pap11, Pap12a, Pap13a, Pap13b, Pap05b, Pap05a, Pap10, Pap12b, Par11, PT14, Pet04, Pin10, Pop12, Por12, Pos98, Pos99, Pri10, Pru00, Puc00, Puc01b, Puc01a, PC03, Puc05c]. Review [Pur93, Pux97, Rad14, Rau08, Rau18, Ree01, Ree77, Rey11, Rie93, Rie10, Rog08, Rog13, Rog18, Roj12, Ros23, Rot07, Rub13, San10, Sar03, Sch09, Sch06, Sch08b, Sch16, Sen12, Sga09, Sga10, Sga12, Sga13, Sha19, Sha10a, Sha12a, Sha12b, Sha13, She13, She14, She15, Sil12, Sim04, Sis93, Smi00, Smi09, Sol12, Sot10, Spr03, Spr06, Spr09, Sri06, Sri10, Sri14, Sta07, Sta10, Sta11, Sta93, Ste11b, Ste11a, Ste12, Ste13, Su16, Sub78, Sza12, Tal09, Tas99, Tay04, Ter08, Ter11, TM11, Teu13, Tew14, To06, Tra22, Tra23a, Tra23b, Tra23c, Tra23d, Tra24, Tse15, Tse17, Upa10, Upa16, Upa20, Upa21, dV16, dV18, Vio04, Vla94, Wan05, Wan12, Was08a, Wei03, Wel14d, Wel15d, Wel16d, Wel17a, Wel17c, Wel18d, Wel19a, Wer10, Wer13, Wid03]. Review [Wil04, Wil10a, Wil10c, Wil10b, Wil10d, Wil13, Xie12, Yan11a, Yan11b, Yan12, Yan20, You99, Zeu89, Zeu92, Zha04, Zha05, Zim13, Zwa93, yea00a, ASvS23a, ASvS23b, Apo10, Arf13, Bar10, BS98, BS13, Cas08, CL09, Che77, Coh08, Dek99b, Dol07, For99, Gas99d, Gas00a, Gas01c, Gas01d, Gas01e, Gas02c, Gas02d, Gas02e, Gas02f, Gas03a, Gas03b, Gas03c, Gas03d, Gas04b, Gas04c, Gas04d, Gas05c, Gas05d, Gas05e, Gas06a, Gas06b, Gas06c, Gas06d, Gas06e, Gas07e, Gas07f, Gas07g, Gas07h, Gas08c, Gas08d, Gas08a, Gas09d, Gas09e, Gas09f, Gas09g, Gas09h, Gas10h, Gas10i, Gas10j, Gas10k, Gas11c, Gas11d, Gas11e, Gas11f, Gas11g, Gas12b, Gas12c, Gas12d, Gas13b, Gas13a, Gas13c, Gas14a, Gas14b, Grä08, HM11, Hel11, HSvS19, HSvS20, HSvS21, JK92, Kat05, Kei08c]. review [Kei10d, Kei11d, Kei12d, Kei13b, Kel09b, KSRS14, McN01b, Mos05, Nar00, Par89b, Par90a, Par90b, Paz13, Puc03, Puc07a, Tie09, Tim13, Vak13, Wel13, Wel14a, Wel14c, dW07, vS17a, vS17b, vS18b]. Reviews [Che80, Gas00f, Spr08, TA95, TBG95, Gas14d, Lou89, Lou90, Lou95a, Lou95b, Lou95c, Lou96a, Lou96b, Lou97, Sei97a, Sei97b, Sei97c, Sei98a, Sei98b, Sei98c, Sei98d, Sei99a, Sei99b, Sei00a, Sei00b, Sei00c, Sei01b, Sei01c, Sei02a, Sei02b, Sei02c]. Revised [Nag17a, Irw03a]. revisited [Eng18, Kul13, Ros95d]. revolution [HL08]. Rewriting [BN98, KL91, Lou05, Nar00, BKdV03, vL74, vL76]. Rex [Wan05]. Rhode [Dav08b, Sch09]. Rice [Cal83]. rich [Har95]. Richard [Ben06, Bón10a, Cad16, Dek99a, Gas10a, Gas10f, Gas14f, Gre16f, Ham13, Kel09b, Lew14, TA95, Vla94, Wil10d, Yan12]. Richeson [Gas22e]. Richman [Xie12]. Richter [Cou14]. Richter-Gebert [Cou14]. Rick [Bón11b, Wil13]. Rieffel [Sga13]. Riemann [Vél78]. Rijke [Hes03]. ring [XL93]. Riot [Ada09a, Gas10e]. Rippling [BHI05, Dav11]. risks [Lan10]. Road [Kil14b, Kil14a]. Robert [Apo11, Bla06, Bón10b, Che80, Ful12, Gas05a, Gas16g, Ist12, Kin14, Lew14, Knu03a, Pop12]. Roberto [Lea97, Mas01a, Náv04]. Roberts [Bón11a, Kob98, Neh99, Pap11]. Robertson [Gas16c]. Robic [Gal21]. Robin [Bón14, Pap10, Puc00, Puc01b, PSTM00]. Robinson [Azn13]. Robshaw [Kat14b]. robustly [Att11]. robustness [Che07]. Rock [Fis08, Gas09b]. Rodger [Pap12a]. Rodney [Chi15]. Roger [Pin10, Pos98, Pel13]. Ron [Pap13a, Par11]. Ronald [Dek97]. Root [GKKK06, Fla05, Nah98]. Roots [GS80, SS07, Gre81, Cen11]. Rosen [Upa10]. Rosenberg [LS74]. Rosenstein [Kob98]. Rossella [Kin15]. Rosser [Ros75]. Rota [Gas00a, Mou10, KRY09, Maz14]. rotating [Dvi09]. Rothe [Fal07]. Roughgarden [Lev10]. Round [Ali20c, Suo20]. Rounds [Shi19, Wel19c]. Routing [Che07, Gav01, Raj02a]. Royer [Smi00]. Rozenberg [Her77].

RSA [Roj12, Hin10, Kat05, Mol03, Yan07, Ira12]. RTA’91 [KL91]. Rubin [Amo99]. Rüdiger [Aro98]. Rudolph [Gas19g]. rule [Kan05]. Rules [Hes92, Hem22d]. running [Gas06e, Par98a, Sim85]. Ruskuc [Sol12]. Russian [Ang76]. Ryan [Apo16, Puc03]. Rytter [Mar15b].

S [Avi01, BS98, Bón11a, Che79c, Che80, Cou09a, Dav05, Dye11, Kar11, Kat05, Kob98, Kre00, Mou10, Neh99, Oli04a, Pap11, Puc03, Rey11, Tra22]. Saari [Pri10]. Saber [EA09]. Sadri [Bón09a]. safety [Pal90]. Sakarovitch [Kin12]. Saket [JH20]. Sakoda [Rav90b]. salary [Kla88]. Salesman [ABCC06, Spr09, Coo12, Azi12a]. Salesmanship [Ros97c]. Sally [Cen11]. Salomaa [Ste12]. Salomon [Jan08a, Pap05b, Tal09, Wer10, Hau12, Nas13]. Salvador [Agu12]. same [FFGH08]. Samson [Knu77b]. Samuel [Gas16e]. Sanders [Elr11]. Sandi [Koz14]. Sandor [Gri10]. Sanjeev [Apo10]. Sanjit [Kat13a]. Sanjoy [Kel09b, Vak13]. Sankt [Kuz19]. sans [Pru98]. Santos [Mur13]. Santosh [Sri06]. Sarkar [Kat13a]. SAT [DKW05]. satanic [HV95]. satisfaction [BCRV04]. Satisfiability [Gas14i, Mar09b, Fal12a, Rog18]. Satyan [FM14]. Saurabh [JH20]. Saving [Ben74]. Scalable [KS10, KS06a]. Scale [Ist12, Mos19, Bor09b, Kei13c, KM10, Zho09, Kel16a]. scenes [Mos10]. Schaefer [Kel09b]. Schaller [Kat13b]. Schapire [Kin14]. Scheduling [DM22, FM05, RV10, Ves21, IMP11, Ful12]. Scheideler [Kin15]. schemas [CP74]. schemata [Fis72, Mil72a]. scheme [TF82]. schemes [GL72]. Schläpfer [Kat13b]. Schneider [Dav06b, Pet04, Puc03]. Scholar [Gas16f]. Scholarship [Wil06]. Scholkopf [Wil04]. Schöning [Gas14i, Kri00, Puc05c]. School [AK20, Kuz18, Kuz19, Wel18a, Hom17]. Schools [RFR97, Kob98]. Schwarzkopf [Mas00a]. Schwichtenberg [Avi01]. Science [Ben98, Bra77, Cha73a, Che79b, CH10, Dav11, DFM07, Fre01, Gas13f, Gus97a, Gus97b, HH76, Kam13, Knu96, Mac20, Mah19, Mar09a, Mer07, Mil17a, Mil17b, Nie02, Par88, Par00b, Rie10, Ros93b, Ros93c, Ros94a, SP98, Sim84a, Sim84b, Ter11, Wol02, AJK^{+}97, ABL^{+}95, Ang91, Coo04, DP08, Gol12a, HR02, Hes03, Joh84, KW07, Lan82, Li12, Man82a, Mat14a, Par89a, Par94, Ros93a, Ros00d, Sei96, SY87, Sta76a, Tay05, Tho93, Toy93, Tre09, Man82b, Pos99, Pro90, Ros97b, Bón13b, FM08, Gas20d, Gas21d, Khu98a, Kri00, Lou18b, GKP88, Mas02, Mil18a]. scientic [Gol12b]. Scientific [Gas07b, Gre22h, GW97, GP89, KM10, Gre07]. scientist [Par02a]. Scientists [LaT11, MMW06, Sch89, Sch08a, YM08, Gol09, Han94, Cou09a, Lea00, Puc05c]. Scissors [Fis08, Gas09b]. Scotland [LRV10]. Scots [Ree01, Sin99]. Scott [Gas07h, Gre13, Mil72b]. Scribner [Gas01b]. Search [BCC10, DW13, DPP09, Gas16f, Gre17g, MAK07, Amb04, CH93, HKR83, KS99, Mat97, Woo78, Mar09a]. Searches [Dav05, GJM02]. Searching [RS91b, RS88, RS91a, Zeu92]. Sebastian [Azn13]. Second [Cou11, Dek99b, Dek99a, Gal21, Gas97a, Gas16c, Pol09, Sha09, Wil10b, Avi01, Bón09b, Bón09c, Cad14, Coh08, Fah05, Gas12f, Gre19f, Ham13, Irw03a, Kar11, Nag14, Smi00, Spr03, CS91, Kuz19]. second-order [Cad14]. Secrecy [Ben06, Mol05, Ree01, Sin99]. Secret [Ayu12, BP06, Gas01b, Kah67a, Kah67b, Kah96]. secretary [Din13]. Secure [HL10, Hut01, Kat12, Bra83, Dav06a]. Security [Bel07, Can06, Dwo95, Elb09, Mor13, Puc03, Puc06b, Sal03, Sal05b, Sta10, BSS11, BHHR99, GK09, Lan10, Bar12, Cho12, Jan08a, Kat13b, Pap05b, Lew14]. Sedgewick [Bón10b]. Seife [Bla06]. Selected [Knu96, Knu00, Knu03b, Vio17, Apo14b, Apo14a, Gas14g, Khu98a, McN01b, Oli04c]. Self [Her02b, CC78, FC83]. self-complementary [CC78]. self-optimizing [FC83]. Self-stabilization [Her02b]. Selman [All98, Hem14a, Sha13]. Seltzer [Gas21c]. Semantic [Cal13, GM10, GR04, Dav08b, Sga12]. Semantics [Gun92, Hes92, Hüt10, Man92, Mos92, Tse91, But71, Car98, CP74, Weg72, Fal12b, Mak93, Rie93, PC03, Sta93]. Semi [DBHHT94, For04, HT03]. Semi-Feasible [For04, HT03]. Semi-membership [DBHHT94]. semifeasible [FH06, Raj06e]. semilattices [Ber86]. Seminar [FRRS17, Wel17b, Fei89, Fei90, KSRS14, Wel14a]. semistructured [Suc98]. Senger [Gas14h]. Sense [San84]. Sensitivity [Fra85]. Sentences [Fas83]. Sept [BBAGP09]. September [BS09]. sequel [Tom90]. Sequence [Ros93c, Dun84, Fla05, Kel16a, GKKK06]. Sequences [Ben98, Ber04, Gus97a, Gus97b, AS03]. sequencing [Kel16a]. Sequential [Ali20b, AO91, AO97, AdBO09, Aro98, RR20, Sch06, Wid03, BP05, MB05, Par86]. Serge [Nag21b]. Sergey [Gre07, Rau18]. Sergio [Upa16]. serial [Vis91]. Series [Gas97a, Mar09a, Pin10, Pos99, Sof03, Wil10d, Fei89, Fei90, Kre07]. Seriously [Sim00, Irw08]. service [FFMS94]. service-technical [FFMS94]. services [GL02]. Sesiano [Gas12a]. Set [COP01, Mie11, SD08, CP10, ER85, FKW94, Gre17h, Irw03b]. SETH [RW19]. Sets [BS79, BY09, COP01, MMW06, Sch03, Sch08b, Akm92, CK89, Fer05, GH00, Mat94, Sch81, Sch82, You92, Cal10, Irw03b]. setting [CH06]. several [Bón14]. Seymour [Eil76]. Shackell [Bea07]. Shafer [Law06]. Shallit [Ber04, Wil10b]. Shamir [Pap13a]. Shanghai [Wan97]. Shankar [Pux97]. shaped [Gas15d]. shared [Bur78]. Sharing [Att11]. Sharma [Smi00]. Sheet [She97, Sei96]. Shemanske [Gre19e]. Shen [Jam98, Sha19, Zeu92, dW07]. Sheng [Bor09a]. shift [Jul92, YDX97]. Shimon [Lab13]. Shiri [Ano13]. Shmoys [Cha12]. Shoham [Azi10a]. Shokrollahi [To06]. Shor [Buh96]. Shore [Dek99a, TA95]. Short [Gas06e, GG21, TZ16, Buh96, Chr05b, Coo76, Kas05, Mof83, Ray05, Rey05]. Shortened [OW17]. Shortest [Ben23, Ram97]. Shots [All23]. Should [Ali21b, Ray21, Tay05]. Shoup [Kat06a]. shovelers [LP92]. show [Gas06e]. Shparlinski [Sha10a]. Shtetl [Gas07h]. Shtetl-Optimized [Gas07h]. Shub [McN01a, XL93]. SIAM [CP74]. SICACT [Lub96]. sided [Woo78]. SIGACT [Kei08a, Kos74a, Raj01b, Raj03b, AH91, ASvS23a, ASvS23b, Chr04a, Chr04e, Chr04f, Chr05c, Chr05d, CKM06, Chr08b, Chr09a, Chr09b, Chr10a, Chr10b, Chr11, Hem95, Hem96, Hem97b, Hem01b, Hem01c, Hem01d, Hem02a, Hem02b, Hem02d, Hem02c, Hem03b, Hem03c, Hem03d, Hem04, Hem05b, Hem05c, Hem06c, Hem06d, Hem06e, Hem08b, Hem08c, Hem09a, Hem09b, Hem09c, Hem09d, Hem10a, Hem10b, Hem11a, Hem11b, Hem11c, Hem11d, Hem12a, Hem12b, Hem12c, HW12, Hem13c, Hem13a, Hem13b, Hem14b, Hem14c, Hem14d, Hem14e, Hem15a, Hem15b, Hem15c, Hem15d, Hem16b, Hem16c, Hem17a, Hem18c, Hem18a, Hem18b, Hem19d, Hem19a, Hem19b, Hem19c, Hem20a, Hem20b, Hem20c, Hem20d]. SIGACT [Hem22b, Hem22c, Hem22d, Hem23a, Hem23b, HSvS19, HSvS20, Hua14, Joh96a, Jon73, Kei08b, Kei08c, Kei09a, Kei09b, Kei09c, Kla88, KM10, O’R06c, Puc05a, Puc05f, Puc05g, Puc06a, Puc06c, Puc07a, Puc10, Raj01a, Raj01c, Raj02b, Raj02c, Raj02d, Raj02e, Raj03a, Raj03c, Raj04a, Raj04b, Raj05a, Raj05b, Raj06a, Raj06b, Ros03a, Ros04d, Ros04e, Ros05c, SP07, Via05, Vit00, Vol23a, Vol23b, Vol24, vS12a, vS12b, vS13, vS14a, vS14b, vS15, vS16b, vS16c, vS17b, vS18a, vS18b, vS22, Amb17, Hem16a, Hem17b, Hem17c, Hem21d, Hem21a, Hem21b, Hem21c, Hem22a, HSvS21, vS16a, vS17a, vS21]. sight [Sch97]. sigma [KL05]. sigma-maxitive [KL05]. Signatures [Mou15]. signing [Eve83]. Silva [BF17]. Silverman [Mei10]. similar [KL10]. Simmons [Irw08]. Simon [Ree01, Sga12]. Simons [Pin10]. Simovici [Mie11]. Simple [Hin97, Pos98, Ros23, BM08, Dem09, DLM07, Yli04]. Simpler [Dun86b, Eng89, Ram94]. Simplex [Bor07a, Mar03b]. simplexity [BS84]. simplified [GJ78, Kri77]. simulated [Has85a]. simulation [Ben74, Bur72]. Simulations [AW98, AW04, HS01, Pap05a]. since [Gre13]. Singh [Ree01, Li14]. single [Ram97]. single-source [Ram97]. SINR [LP10]. Sipser [Rav90b]. SIROCCO [Arf13, BFU19, Bur15, Das09, Kei13b, Kon17, KPS11, Mil18b, Tse17, Wel17a]. Six [Lou18a, Cen14]. Sixth [Dav05, GJM02]. Size [Che79c, VV78, CK89, Kao97]. Skiena [Mat14b, Tra22, Kar11]. skip [Hua98]. Skormin [Zha05]. Slicing [Gas16d]. Sliding [Mos19]. Sloan [Lew14]. Smale [McN01a, XL93]. Small [BGH07, Nah04, Din08, GKP96, Gre22h]. smallest [Gre22a]. Smart [Mar20]. Smith [Gol00b]. Smola [Wil04]. Smolyak [Gre22h]. Smullyan [IZ70]. Snyder [Jou94]. Soare [Gas16g]. Social [Kin11, Nag17b, Nag18b, VR07, Kei11c, Yu11]. Society [Dav05, FM09, Joh10, Kat05, Kob98, Kri08, McN99, Sch09, Wil10d, dW07, Ogi01]. Sofo [Kre07]. Softcover [Gre22h, Gas10b, Gas10e, Gas05a, Fri10, Mou10]. Software [Dav08b, GR04, Jac06, Mar20, Bat93, Lee10]. Soifer [Gas09c, Gas16f]. Sokolova [Zha05]. Solomon [Les06]. solution [Hen83, Pag70, Sch97]. Solutions [AP01, She08, She10, Con95, Jam98, Nag15a, Oli04b]. Solvable [Che82, AKR83, Ing83]. Solve [Gas22e, Mas01b, MF02, Vak13]. solved [Fra84a, Fra84b]. solver [Gas13c]. Solving [BHS80, Blu83, Gas12a, Mor03, Mor05, Nag23, Hav74, Hua03, Ses09, Wei98, Wei95e, Cou09b]. Soma [Gas09h, Gar08]. ’Some [GJ78, Ala03, BA91, Cha73a, Cha73b, JKL99, Kri77, Lei91, Rav90b, Ric73, Sim85, Sud11, Wil18, YDX97, Yli04, DBHHT94, Gal74, Gol92, Gue83, Har95]. Sommersby [FFGH08]. Song [Cou12, Ira12]. songs [Knu77a, Knu84b, Knu84a]. Sons [Cha05, Náv04, Cen03]. Sophisticated [Ber00, She13]. sorting [BFK^{+}79, GK80, Lee72, Pas99]. SOSP [Cle08]. Soul [Gas07c, Lew06]. soundness [Din08]. Source [Cha20, Bor09b, Ram97]. Sources [Cen14, Gab10, Hut70, Zim13]. SPAA [Ass17, Dhu19, Wel19a]. SPAA’08 [Guz08]. Space [Cen15, Mac95, Mil09, Pap10, Ros02b, AM75, BFK^{+}79, Gol77, Has85b, Lad75, LM73, LL75, LS78]. Space-bounded [Mac95]. Spain [BS09]. Spanner [MM13]. sparse [GH00, You92]. Speaker [Par88, Par00b]. Special [AB10, CP74, FC^{+}99]. specific [Sin72]. Specification [Sis93]. Specifications [GR04, Dav08b]. Specifying [Puc04b]. Spectral [Chu97, Lur99]. Spectrum [Pin10]. speculate [KS06a]. speculation [Kug76a]. speech [Mos10]. speed [Rob70]. speed-up [Rob70]. speedup [Par86]. Spencer [Gas16b]. Spillner [FB14]. Splaying [CH93]. Splendors [Shi19, Wel19c]. Spring [Pap05b, Sta89]. Spring-Verlag [Pap05b]. Springer [All98, Aro98, Bea07, Boo10, Bor09a, Cal10, Che80, Col10, Cou10, Dav06b, Dav08b, Dav08c, Dav09, Dek99a, Ess05, Fah05, Fal07, Gas97a, Gas10d, Gas10f, Gri10, Irw03b, Jan08a, Kat06b, Kel08e, Kri00, Kye09, Law06, Lee03a, Lee03b, Mar09a, Mas03, Mav09, McN04, Mei10, Mik93, Mur10, Neu10, Oli04b, Pos99, Sim04, Sis93, Tal09, Tas99, Tay04, To06, Tra22, Upa10, Vio04, Wei03, Wer10, Wil10a, Zha05, dW07, For04, Gas07d, McN01a, MT04, Oli04a, Sta10]. Springer-Verlag [Aro98, Che80, Dav06b, Dav08b, Dav08c, Dav09, Fah05, Gri10, Irw03b, Jan08a, Kat06b, Kri00, Kye09, McN04, Mei10, Mik93, Oli04b, Pos99, Sim04, Sis93, Tay04, Upa10, Vio04, Zha05, Gas07d, McN01a, MT04, Oli04a, Sta10]. SPTCC [Kuz18]. SPTDC [AK20, Kuz19]. spurious [Chi86]. Spyware [Ayc11, Wer13]. Square [GKKK06, Fla05, Nah98]. squeezed [Ros03e]. SRC [Ros95c]. Srikant [Lib17]. Srini [Nag23]. St [Bla06, Gas15g, LRV10]. Stability [BF17, Ros04f]. stabilization [Her02b]. Stable [Knu97, McN99]. stables [Knu76b, Lew78]. stage [Hem06a, SS06]. stamp [Sha02]. Stanley [Ham13]. start [Lev91, Sta87]. Stasys [Bón13b, Gas13e]. State [Ali21d, BB21, DM22, Gas16f, GL84a, GL84b, Kei10a, LMZ10, YZ91, Bir91]. State-of-the-art [DM22]. States [Gas10e, OW17]. Statistics [Alb03, Gle05, OW17]. status [Ber92]. Steeb [Pap05a]. Stephan [Azi16, Cha09, Gre16h, Wan12]. Stephen [FM08, Mil13]. steps [Yeh12]. Steve [BF17, Gre18e, Pet04, Sol12]. Steven [Gas06e, Gas14h, Kar11, Led16, Mat14b, Sha13, Tra22]. Stewart [Gas09h, Gre15d, She14, Sri14, Ste13, Ste09]. Stichtenoth [Kop11]. Stinson [Spr03]. Stirling [Puc01b, Wei03]. STOC [Joh90, Lev95]. STOC/FOCS [Joh90]. Stochastic [Hut70, Lib17, Hem06a, SS06]. Stockmeyer [Raj04e]. stone [Gal78]. Stony [Ski99]. storage [Cas13, DO13, Kei13a]. stored [SvL74]. Storer [EA10]. Stories [Ada09a, Con09, Fri10, Gas06e, Gas11a, Ken05, Mac95, Gas10e]. Story [Bla06, Gas01b, Kah67a, Kah67b, Kah96, Mao94, Nah98, CP10]. Straffin [Spr08]. straight [CH06, Kau93]. straight-line [Kau93]. Strassen [Mor85]. Strategic [CFF^{+}97]. Strategy [Str93, Spr08]. Straubing [Gas94].

Streaming [Ali23c, Ass23]. Street [Vak13]. Street-fighting [Vak13]. string [Gal75, GG86, WM87]. string-matching [Gal75]. Stringology [Mar15b]. Strings [Ben98, CHL07, Gus97a, Gus97b, Lie75, Ada11, Gal15]. stroll [Agu04, Khu04]. Structural [Fal12b, Hüt10]. Structure [BF17, HH24, LP10, Mur13, Rie97, But71, Cad14, HR00, McB85, McB87a, MR97, Gas15h]. Structured [Fas83, JK92, Smi09, Tse91, Vel94]. Structures [Bra08, CH15, Dav05, EA10, Elr11, GT97, GT01, GT04, GT06, GT10, Gun92, Jan10b, Koz18, Mas01a, MS08, RZ09, Ros93c, Sta93, Sto02, Sub78, Wel15a, GJM02, Hem14a, Wei98, Zam77, Wei95e, Gas17d, Jan10a]. struggle [Gol12b]. Student [FM09, Ros99c, Rou05]. Students [Ali21b, Che24, Par88, Par00b, Ray21, Mit01a, Mou14, Tay05]. Study [ABCC06, Oli04c, CW72, DLO09, Spr09]. studying [CT23]. Subexponential [She91a, She91b, Gol95, Her83]. Subir [Dea09]. Subjects [Gas23a]. sublinear [KR03]. Subsequences [OW17]. subset [Lon72, Qia87]. substrings [LO94]. Subtle [Cou15]. Subtractive [Mor83]. suffice [Khu05a, Raj05c]. sufficient [Jaf78]. Suffix [ABM08, Neu10]. suitable [Mat97]. Suitening [Ram89]. Sukumar [dV16]. sum [Qia87]. Summary [ABLO23, Ali23a]. Summation [Kre07, Sof03]. Summer [AK20, Kuz18, Kuz19, Wel18a, All00]. Sun [Gas10g, Wap05, Mel11]. Superpolylogarithmic [She91a, She91b]. superpowers [PM80]. Support [FC^{+}99, Neh99, SS02, Wil04]. Surface [Dey07, Sot10]. surfaces [Moh93]. Surfing [Par95b]. Surreal [Gas10c, Pet02]. Surveillance [Lan10, Mor13]. Survey [AdW17, BFK^{+}16, DLO05, FS19, Gue83, Wel19b, Chr04c, Elk04, GL80, Gol95, Gol10b, HW12, Kla88, Raj02a, Yu11]. Surviving [Sta89]. Susan [Jan08b, Mak99, Mor13]. Susskind [Gre21f]. Svetlana [Zha05]. swell [Ang91]. swimming [Chr04d, Ros04c]. switches [Gol10b]. Sybil [Kei11c, Yu11]. Syllabus [Ros93c]. symbol [AM75, IZ70]. Symbolic [LP07, Sha04, Bea07]. symmetric [Gab75]. Symmetry [Bar15]. symposia [BR98]. Symposium [ACM12, Agu12, HK84a, HK84b]. Synchronization [Had15, Tau06, Gup02, Hen09]. Syntax [But71, CH06]. Syntax-semantics [But71]. Synthesis [AT17, Wel17a]. synthesizing [Kei12b]. System [Che79c, TM11, VV78, Vin09, van79, Hen83, Ros94b, Che80]. Systems [Azi10a, Bra92, Cou09b, Dav06a, Dav06b, GHSW14, Gol93, Gol94, JORS99, MP92, Mil99, Mor03, Mor05, Pet04, Puc00, Ray05, Sch00, Sch04, SLB09, Sis93, dV16, BS09, BKdV03, But71, Cas13, Chr05b, Chr06b, Coo83, DO13, DP08, FL83, FKR08, Har95, Hut01, IZ70, Kan05, Kei08e, Kei09c, KSRS14, Mor76, Rey05, Ver06, Via08b, Wel14a, Zho09, vL74, vL76, Hen93, Lou05, Smi00]. Szemerédi [Gre17f].

T [Cal12, Das92, FB14, Gas05b, JK92, Mas01a, MT04, Náv04, Oli04a, Sha12b]. Table [Cul99]. Table-automata [Cul99]. Table-automata/ [Cul99]. tables [Hua98]. Tac [Gas12e, Bec08]. tag [vL74]. tag-systems [vL74]. tailor [DO13]. tailor-made [DO13]. Take [Hem98]. Take-home [Hem98]. takes [CH93]. Taking [Irw08, Ros99c, Sim00]. Talbot [Rot07]. Tale [Nah98, Bla06]. tales [HS98, Gas22e]. talks [Par02a]. Tamal [Sot10]. Tamassia [Náv04, Mas01a]. Tanton [Gas15g]. Tao [Tew14, Gas15h]. Tape [Pig15, Wal82]. tapes [Ben74, LS78]. Tarakanov [Zha05]. Tardos [Lev10]. Tarski [Les06, FF04]. tasking [Ber72]. Taubenfeld [Hen09]. tautologies [BBF^{+}73]. Taxonomy [Gol93, Gol94]. TCS [Gol92]. Teaching [Alb03, BY95, Fek09, Hem94b, Hem20c, Lam09, Ric21, Rod99, Ros03e, Abr13, Far19, Kei09a, Man82a, Man82b, Ros95d, Gle05]. Team [HS18]. Technical [Hag99, Hag00, Hag02, Hem05a, Hem06b, Hem07c, Hem07d, Hem08d, Hem10c, Hem10d, Kel05, Kel06a, Kel06b, Kel06c, Kel06d, Kel07a, Kel07b, Kel07c, Kel07d, Kel08a, Kel08b, Kel08c, Kel08d, Kel09c, Kel09d, Kel09e, Kel09f, Kel10a, Kel10b, Kel11a, Kel11b, Kel11c, Kel11d, Kel12a, Kel12b, Kel12c, Kel12d, Kel13d, Kel13a, Kel13b, Kel13c, Kel14a, Kel14b, Kel14c, Kel14d, Kel15a, Kel15b, Kel15c, Kel15d, Kel16c, Kel16d, Kel16e, Kel16f, Kel17a, Kel17b, Kel17c, Kel17d, Kel18a, Kel18b, Kel18c, Kel18d, Kel19d, Kel19a, Kel19b, Kel19c, Kel20a, Kel20b, Kel20c, Kel20d, Kel21d, Kel21a, Kel21b, Kel21c, Kel22a, Kel22b, Kel22c, Kel23a, Kel23b, Kel23c, Kel23d, Kel24, Khu06, Rag91, Sin98, Sta74d, Sta75b, SIG80, SIG81a, SIG81b, SIG82c, Wei93, Wei95c, Wei95d, Wei96b, FFMS94, KB79b]. technical [New76, Sta73a, Sta73b, Sta73c, Sta74b, Sta74a, Sta74c, Sta75a, Sta76b, Sta76c, Sta76d, Sta77a, Sta77b, Sta77c, Sta78b, Sta79a, Sta79b, SIG82a, SIG83b]. technique [FC83, HHH98, HG00, Lee72, Rav90b]. Techniques [AB09b, AB10, Bor07a, GM10, Gun92, HL10, Kat12, Kre07, Mar03b, Sga12, Sof03, KL91, Sta93]. Technologies [Zom06, Bor09b, Lan10]. Technology [Mor05, Sch06, Dif82, WG00, Cou09b]. telegram [Tuc66]. Telephone [Blu83]. Temporal [Bra92, Hen93, MP92, Sis93, Sti01, Wei03, LP07, Puc05d]. temporal-epistemic [LP07]. Tenth [Gas21a, Mat93, Pru00]. Teo [Cou09b]. Terence [Gas15h, Tew14]. Terese [Lou05]. Term [BN98, BKdV03, Lou05, Nar00]. terminological [Knu74b]. Terms [Old91, Pur93, Jul92, YDX97]. Territory [Gas22d]. Terry [Gas19g]. Tesman [Bón11a, Pap11]. test [Par91]. Testing [Ber86, Cou12, Die04, Gre16e, Kat06b, KS16, Sud11, Bha13, Ham76, Upa20]. Tests [Vél78]. Text [Gas97a, Dav93, Gas10b]. Textbook [GKP88, Mou14]. Textbooks [Ros05d, Ros95b, Ros98b]. th [GKP96]. Their [Old91, Pro90, CM72, Fas83, Gas13c, Gue83, Kur06]. Theme [Gol06, SS04a, SS04b, Sud11]. Theobald [FM15]. Theorem [BA04, Che79a, Lov78, Mor85, Ros75, Ben73a, Cal83, Cha71, KL05, SW82, Ada74, Ada09b, Bón12]. theorems [CP13, Gre81, Har80, Kug76b, Gas20e]. Theoretic [Azi10a, Eil76, Gas16d, Gin75, Kat05, SLB09, Wag03, BA04, Cad14, Cha73b, Geo09, KW07]. Theoretical [Amo05, BS09, Dav11, Kre01, Kri00, Mar09a, MAK07, Par88, Par00b, SP98, Sei96, AJK^{+}97, ABL^{+}95, Hem99, Hes03, Joh84, Lan82, Man82a, Man82b, Par89a, Par94, Tre09, Dav08c, Dav09, Gas13f, Mar09a]. Theories [Cal12, Pip97, Ros01c, Gup98]. théorique [GP89]. Theorists [Gas07h].

Theory [Aaz18, AC96, AK90, AAC^{+}08, AK20, All98, AH06, Are18, AdW17, Ata99, AB09b, AB10, Azi11, Ber04, Bie05, Bol98, Bón10a, Bón11b, Bra77, BC09, BCS97, COP01, CH14, Cha98, CZ09, Che79b, Chu97, CC76, Coh93a, Coh95, Coh96, Cul82, Dav02, DK00, Dud09, Dur10, Edw08, Fal07, FFG96, Fen21b, Fis08, FS15, GM99, Gas99a, Gas08b, Gas09b, Gas09h, Gas16c, Gas16g, Gas17d, Gas19e, Gas20f, Gee17, GW97, Grä78a, Grä78b, Grä98, Gre17h, Gre19g, Gre19f, Gro14, GY04, Ham90, HHL00, HHM00, HHM08, Hay08, Hem95, Hem96, HS97, Hem03a, HT03, Hem10a, HW12, Hem14e, Hem15a, Hem15b, Hem15c, Hem15d, Hem16a, Hem16b, Hem16c, Hem17a, Hem17b, Hem17c, Hem18c, Hem18a, Hem18b, Hem19d, Hem19a]. Theory [Hem19b, Hem19c, Hem20a, Hem20b, Hem20c, Hem20d, Hem21d, Hem21a, Hem21b, Hem21c, Hem22a, Hem22b, Hem22d, Hem23a, Hem23b, Her02a, Hin97, HU79, HMU01a, HMU01b, HMU07, Ios07, Irw03b, Irw03a, Iva02a, JH20, JORS99, Joh10, Jou94, KA89b, Kat06a, KN01, KS01, Koc19, KV08, KV12, Koz06, Kri08, KR12, Kuz18, Kuz19, Lev91, LR97, Lou91, Mai80, Mar08a, Meh00, Men04, MP21, Mie11, MT04, Ngo23, NRTV07, Pap12b, Pic21, Pol09, Pos99, Ree77, Ros96c, RST15, Rot05, Sak11, SW01, Sar03, Sga10, Sha09, Sha10a, Sho05, Sho09, Shp03, Sil12, SD08, Sny91, Spr03, Spr08, Str93, Tan16, Tao20, TS96, TS00, Tse91, Vol23a, Vol23b, Vol24, Wan12, Was08b, Weg00, Wil04, Wol93, Yan02, AAH11, AFG^{+}96]. theory [AS03, Ang91, Azi14, Bec08, BR98, BCRV03, Cas13, Cen02, Chr08a, CEG^{+}10, CFF^{+}97, Esc93, FH06, Far19, FM06, Fer09, Fic97, FKW94, Gas15d, GMS74, Gop08, Gra09, GM12, Hac09, HK84a, HK84b, HN14, Har95, Hav74, Hem94a, Hem97b, HS98, Hem01a, HO01, Hem01b, Hem01c, Hem01d, Hem02a, Hem02b, Hem02d, Hem02c, Hem03b, Hem03c, Hem03d, Hem04, Hem05b, Hem05c, Hem06c, Hem06d, Hem06e, Hem08b, Hem08c, Hem09a, Hem09b, Hem09c, Hem09d, Hem10b, Hem11a, Hem11b, Hem11c, Hem11d, Hem12a, Hem12b, Hem12c, Hem13c, Hem13a, Hem13b, Hem14b, Hem14c, Hem14d, Hem22c, Hes02, HL09, IZ70, Iva02b, Kar00, Kat78, Kei09c, Kei11b, Kei12a, KV00, KV02, Kor06, KL10, KR09]. theory [Len08, Lit03, LP86, LP09, O’R06c, OW09, PGW^{+}96, Par98b, Raj06e, RGL10, Rod99, Ros04f, Sav98, Sch75, Slo02a, Slo02b, Sti95, Sti02, Sti06, Su09, Vad07, Van10, Van06, Ven99, Via05, Via08a, Wil08, XL93, Yan96, Zal97b, Zal99, Zho09, dB06, KV95, Lat95a, Lat95b, Lat95c, LP98, Ric00, Ros93b, Sip96, Zal97a, Zal97c, Gas07h, Avi01, Azi12b, Bón09b, Cen00, Coh08, Cou11, Fab11, FM09, For99, Gal21, Gas04a, Gas12e, Gre16g, dOG12, Har11, Kel08e, Kin12, Lev10, Lur99, Mou16, Nag17a, Pap12a, Pos98, Sha12b, Spr06, To06, Wil10b, Yan11a, Sha13, Apo12, Cen03, For04, Gol00b, Yan20]. theory.blogspot.com [Gas07h]. there [KR01]. Thermodynamic [GW18]. thesis [BA05, Par86, Pas99]. Theta [Knu76a]. Thierry [Gal15]. Things [Nah04, Par02a, Fas04, Ros02b]. Think [Edm08, Sga09]. Thinking [Ali20b, RR20]. Third [Gas19h, KR12, EA09, Kei13b, Küp13, Wan97, AK20]. Thirty [Hem23b]. Thomas [Gas10e, Gre19e, MT04, Kin12]. Thomson [Sch06]. Thorsten [FM15]. Those [dK09, Gas11b]. Thoughts [All23, RP97, Har95, Sim85, Gas00a]. threading [Fek09]. Threats [Cho12]. Three [Ali21d, BB21, IZ70, Jaw05, Old91, Puc05h, Nis78, Yeh12, Gas00f, Gas07h, Gre21f]. throughput [Kel16a]. Tic [Gas12e, Bec08]. Tic-Tac-Toe [Gas12e, Bec08]. Ticket [DW13]. Tidy [Mil17b]. Tim [Ban15, Lev10]. Time [AH91, AH83, Die04, GC86, Hem20c, Pet04, Ros73a, Ros73b, Sch00, BFK^{+}79, CH93, Gal74, Gal75, Gol95, Has85b, Hem94a, Her00b, Hua03, KR03, Lon71, LP92, Par98a, Reg93, Sch81, Sim76, You92, Kat06b]. Time-memory-processor [AH83]. time-space [BFK^{+}79]. time-varying [LP92]. Times [Ada74, Ada09b, Bón12, Gas12a, Mol05, Ses09, Ben06]. Timothy [Azi10b, Neu10]. Tiuryn [Puc01a]. Tobias [Nar00]. Today [Sco15]. Toe [Gas12e, Bec08]. Tofte [Puc01b]. tolerance [Kei11b, KR09]. tolerant [CBS07, Eya13, GL02, KR01, TV16, Wel16c]. Tomescu [Kel16a]. tool [Lan82]. Toolbox [MS08, Elr11]. Tools [Gol01, Mie11, SD08, HS98, Kat05, Puc03]. Top [Har11, Ios07, Ang91]. Topics [AB10, AW98, AW04]. topological [San84]. Topology [GFW17, Raj02a, GG21, Upa16]. Torán [Gas14i]. Torenvliet [For04]. TORIX [VV78, Che79c]. Tour [Maz10, Pet08, Wil11, Cad11]. Tower [Gas09h, Koz14, Sta87]. Towers [KB79a, GC86]. Tracts [Dav11, Hes03]. tradeoff [BFK^{+}79]. tradeoffs [AH83]. traffic [Che07]. training [SR03]. transaction [Eya13, Kei13c]. Transactional [CH14, HS12, KR12, RGL10, Sco15, SB14, Wel14b, Att08, Gra09, GM12, Kei12a]. Transactions [FFGH08]. transfinite [RS91a]. Transform [ABM08, Mil17b, Neu10]. Transformation [Hes92]. transformations [WM87]. Transformer [Mak93, Man92]. Transitions [Hüt10, Fal12b]. transitive [Shy78b, Yel78]. translated [Gre18e]. Translating [GL72]. translation [Cad13]. Translator [Kin12]. Travel [FdV23]. Traveling [ABCC06, Azi12a, Spr09, Coo12]. Travelling [Chr06b, Ver06]. Treasure [Hem94b]. Treasurer [Lub96]. tree [CH93, FRT04, Kul13, McB85, McB87a, O’R04a]. Trees [Ben98, Das92, Did93, Fal12b, Gus97a, Gus97b, Hüt10, Kug75, Lei92, Ead93, Gab75, Kan88, KLMY21, McB91, Woo78]. Treks [Gas10c, Pet02, Ban14b]. Trends [Har15, Gol92, Nag18b]. Trevisan [Gas07h]. Triangulations [Mur13]. Tribute [BR99, Gas11g]. Trifonov [Rou05]. Troelstra [Avi01]. trust [KGO10, Ros02a]. trusted [Kei13b, Küp13]. Trusting [CKS09]. Truth [DPP09, Gas16d, DH79a, DH79b]. trying [Ros05b]. Tse [JK92]. TSP [Lam14]. TTPs [KGO10]. Turing [BA05, Ben74, Gal74, Kan05, McB92, Nir86, Pet08, Pig15, Rod09, You92, Kil14a, Kil14b, Wil10c, Gas16g]. Turning [Ros98a, Wil86]. tutorial [BBAGP09, Can06, IMP11, Puc06b, Yu11, HRR12]. TV [Gas06e]. Twenty [Gas15f]. Twenty-First [Gas15f]. Two [Cha20, Che21, Gar06, HL10, Hem22d, JP00, Kat12, Pet94, Sim76, Sub95, AAH11, Gal78, Gas99b, Gol92, HS12, vdBN04]. Two-Party [HL10, Kat12]. Two-Source [Cha20]. Two-way [Pet94]. Type [Car98, Hin97, Pos98, KL10, PC03]. Type-logical [Car98, PC03]. Types [Di 95, Lea97, Pie02, Wol93, EKP79, McB91, O’H98, Dav02, Kin05, Kin06]. typical [HW12]. typical-case [HW12]. Typically [Sha10b].

Typically-correct [Sha10b].

U.C. [Sim84a, Sim84b]. Ulle [Nag17b, Nag18b]. Ullman [Mai80, Ben73a, Weg74]. Ulrich [Col10]. ultimately [Mat94]. Umesh [Kel09b]. un-merging [San85]. Unauthorized [Lew14]. Unbounded [Hes92, Pan93, RS91b, Zeu92, RS88, RS91a]. uncertain [Che07]. Uncertainty [Hal03, Lap91, Gal22, Zha04]. uncomputable [Dun84]. Undaunted [Akm92]. undecidability [Mil72a, Pal90]. undecidable [Cud69, Gup02, Hem08a, Mar08b]. undelivered [Mos10]. Undergraduate [Che24]. Understand [Ali20c, Suo20]. Understanding [Ali21c, Elb09, Har21, Bar12, Mou14]. Underwood [dOG10, Yan11a]. Unexpected [Saa01, Pri10]. Unification [Sny91, Jou94]. Unified [Wid03, MB05, Vad07]. Uniform [Vol99]. Unifying [JK92, Tse91]. uninteresting [Kug76b]. union [HR00]. union-find [HR00]. United [Gas10e]. Universal [Kil14b, Mul69, CK89, Gra82, Kil14a, Cal13]. universe [Nou80]. Université [Lew78]. University [Apo10, Avi01, Azi10a, Azi10b, Ben98, Ber04, Bla06, Coh07, Cou09b, Cou09a, Dav06a, DL79, Dea09, Dek98, Dek99b, Gas07c, Gup98, Har10, Hes03, Hil07, Irw08, JK92, Jan10a, Kat05, Kat06a, Kat07, Les06, Lev10, LRV10, Lou05, Mak93, Nar00, Nat07, Nat10, O’R92a, Pan93, Pap10, Pos98, Pri10, Puc03, Pur93, Rie93, Rot07, Sga09, Sga10, Smi09, Sot10, Spr09, Sri10, Sta07, Wan05, Wil10b, Zwa93, Lew06]. unknown [She14]. Unlocking [Nag18a]. Unplugged [Kin15]. Unsolvable [Lew79, Dun87, Che82]. Unsolved [CEG98, Par11]. Upfal [Kat07, Sta07]. upon [Mof83]. Urn [Mah08, Sta11]. Uroz [Koz14]. USA [Sch09]. Use [Jen97, Mik93]. used [Sch75]. useful [HN14]. Using [Alb03, Ali20c, Ayu12, BP06, Far19, Gle05, Jan10b, Kei11c, Suo20, Tri88, Tse91, Val09, Boo77, Bur78, Coo76, HHH98, Her83, Lab12, RZ09, Wei98]. Uspensky [Sha19]. Uwe [Gas14i, Kri00, Puc05c].

V [Fit11, Gre07, JH20, Kat05, Lam13, Lev10, Sha19, WG00, Ogi01]. Vadim [Rau18]. Vague [Dav93]. Valiant [Bru13]. validation [Lee72]. Valiente [Lab12]. Vallée [Nar14]. Valuable [All23]. valuations [Sin72]. value [Gol77, Lad75]. values [DH79b, FL83, Gol12a]. Variable [Che79c, Sal07, VV78, Nas13]. Variable-length [Sal07, Nas13]. variables [Bur78, FL83, Bón14]. Variance [FGK^{+}02]. variant [Jor83]. Variants [Gas21a, Hin10, Pig15, Roj12]. Variation [Mor85]. Various [Gas23a]. varying [LP92]. Vasek [Ros23]. Vassilevska [Khu12]. Vatter [Sol12]. Vazirani [Kel09b, Lev10]. Vector [Ben23, BS79, SS02, Wil04]. Vectors [Che79c, VV78, Dwy90]. Vega [Kin11]. Vega-Redondo [Kin11]. Vehicular [OW09, Wan12]. Veldhorst [Che79c]. Veli [Kel16a]. Velleman [Smi09]. Vempala [Sri06]. Venema [Hes03]. Vera [Ayu12]. Verdi [Dek97]. Vereshchagin [Sha19]. Verification [AO91, AO97, AdBO09, Bau17, BJK^{+}16, BM07, Sch04, Wel16b, Bur72, Puc05e, Fur11, Aro98, Dav06b]. Verifier [COJ78]. Verifying [Bra92, Hen93]. Verlag [All98, Aro98, Bea07, Bor09a, Che80, Dav06b, Dav08b, Dav08c, Dav09, Dek99a, Ess05, Fah05, Gri10, Irw03b, Jan08a, Kat06b, Kri00, Kye09, Lee03a, Lee03b, McN04, Mei10, Mik93, Oli04b, Pap05b, Pap05a, Pos99, Sim04, Sis93, Tas99, Tay04, Upa10, Vio04, Wei03, Zha05, Gas07d, McN01a, MT04, Oli04a, Sta10]. Version [MW93, OW17, BCG^{+}22, CR74, FKW94, KR01]. versions [DKW05]. versus [TS99, VM85]. vertex [GI91, Khu02]. Vertical [SS07]. vi [Jou94]. via [CKM06, Chr07, Gre16f, Gre19f, Kho05, Khu05b, Yu11]. viable [Ben73a]. Vicious [BM96, Dek98]. Victor [Fal12a, Kat06a, Zha05]. video [Nag18a]. View [Bar15, Har11, Ios07]. Views [Old91]. vii [Mik93]. viii [Hen93, Sga10]. Vijay [Lev10]. Viking [Bla06]. Vin [BF17]. Vince [TM11]. Vincent [FB14, Nag17b, Sol12]. Vingron [Neh99]. Virginia [Khu12]. Viruses [Ayc06, San10]. Visibility [Gho07, Tol93, Dea09]. Visions [Sri14]. Visualize [Mil17b]. Vitanyi [Gas97a]. Vivien [Ful12]. Vladimir [Rou05]. Vöcking [Kin15]. Vogler [Cad12]. Vol [FM09, Gas23a, Che79c, Kat05, Lee03b]. Volk [Vol21]. Vollmer [Col10, Kin15]. Volume [Gar06, Gas20e, Gre22f, Knu11, MW93, Mik93, Rog08, Ham13, Rog13, Ven02, Wil10d]. volumes [Gas99b]. Voronoi [AS91]. Voting [HK05, Wil10d]. Vovk [Law06]. VS [Gas10e, Che21, GPP83, Har95, Ros95d, RW19]. Vu [Tew14]. Vyalyi [dW07]. Vygen [Kum13].

W [Alp18, Che79a, Fal12a, Gre16f, Gre17h, MT04, Nat07, Nat10, Rau08, Sch18, Xie12, Knu03a]. Waerden [Gas16f]. Wagner [Kin15]. Wagstaff [Kat05, Gas16e]. Wait [Had15]. Wait-Free [Had15]. waiting [Bur78]. Waldecker [Gre16e]. Waldinger [Vla94]. Walk [Gas15b]. Walker [Bla06, Xie12]. Walks [Gre17g]. Wall [Dye11]. Wallis [Rau08]. Wallner [Ess05]. Want [KS06a]. Wanted [Bra95, Par98a]. wants [Gre22h]. Wapner [Gas10g]. Ward [Mat14b]. Wardhaugh [She15]. Warner [Lew14]. Warren [Che82]. wars [Ros02b]. Warshall [Weg74]. Washington [Che10]. Waterman [Neh99]. Watrous [Fen21b]. Wattenhofer [Pel13]. Way [KRY09, BHHR99, Pet94, Wal82, Mou10]. Wayne [Jou94]. Ways [Ali21c, Har21, Nah04, CT23]. Weak [Gab10, Hav71, Zim13, GH00]. Wealth [She15]. Web [Sch09, Bor09b, KGO10, Bon08b, CR03, GL02, Par95b, Ros95a, Ros96b, Ros98b, Ros98c, Ros98d, Ven99]. web-scale [Bor09b]. weblog.fortnow.com [Gas07h]. Wegener [Joh10]. Weighted [Cad12, DKV09]. Weigle [Wan12]. Weissmann [LaT11]. Welch [Her00a]. Welfare [Mou03, Dan06]. Wellfounded [BM96]. Wells [Cou15]. Welschenbach [Ter08]. Welsh [Rot07]. Wendy [Gas21c]. were [Par98a]. Werner [Cad12]. Wesley [Che82, Gas09a, Mai80, Rog08, Vla94, Fah04, Puc03]. Wheeler [Neu10, ABM08]. Where [KR09, Fek09, Mat14b, Ros97d]. Which [Gas10e, BCG^{+}22, Paz70]. While [Nag23]. Whitfield [Jan08b, Mak99]. Who [Ski99, Mat14b, Mat14a]. Wickham [Mil17b]. Wide [FFMS94, Ros95a]. Wigderson [Gre21e]. Wil [Mar09a]. wild [VM85].

Wilde [Mou16]. Wiley [Cen03, Cha05, Náv04, Pet04, Wil10c]. Wilf [Kre00]. Wilke [MT04]. Will [Har95]. Willi [Pap05a]. Willi-Hans [Pap05a]. William [Azi12a, Bón12, For99, Gre22h, Gas20c]. Williams [Khu12]. Williamson [Cha12, Nag21a]. Willinsky [Aar07]. Wilson [Bón14]. Wim [Pan93]. Win [All23, Ski01]. Winfree [Ogi01, Ste12]. Winner [Sch18, Rou05]. Winthrop [Che80]. wireless [GK09, HN14, Kei10b, LP10]. Wiretapping [DL98, DL07, Jan08b, Lan10]. Without [Gas07c, AF14, FFB89, Lew06]. Witt [Her13]. WOBCATS [Lie90]. Wojciech [Mar15b]. Wolfgang [Sga13]. Wolsey [Cha05]. Women [Bay95, Fri10, Fri11, GL09, Ken05, Ang91, Ber92, Toy93, Ros97d]. Wood [Amo99]. WOPA [CS91, Thu90]. word [Ben73b]. Words [BS08, Lot05, Bón09a, Dav08a, Rau18]. work [MM99]. Working [Dwy03, Wan05, BB89]. Works [Gol06]. Workshop [AP98, CH14, CS91, KR09, KR12, LB99, Mor97, PD98, Pro90, Rad14, Rod96, RW99, Thu90, Wel17c, WG00, BS09, Gol00a, GM12, HN14, Paz13]. Workshops [Cun95]. World [Gre07, Gre16g, Gre22h, VGS05, Wil10d, Ros95a, Law06, Gas15d, She14, dV18]. Worlds [For21, HRZ95]. wormholes [Chr06b, Ver06]. worst [BHHR99, KW81]. worst-case [BHHR99]. Wright [Gre18e]. Write [Che20, Mar20, Bra95]. Writing [HS98, Kah67a, Kah67b, Kah96, She15, Gas01b]. Wrong [Gas22f, Dun00]. WSS’01 [Her02b]. WTTM [CH14, GM12, KR12, Wel14c]. Wu [Bor09a]. Wu-Sheng [Bor09a]. www.scottaaronson.com [Gas07h]. www.scottaaronson.com/blog/ [Gas07h]. Wyoming [Hil07].

X [Wil04, Pur93, Wan05, dW07]. Xi [Gre22f, JK92]. Xiaoling [Mel11]. xii [Gas94, Pan93]. xiv [Sis93, Vla94, Wil10d]. XTANGO [Sta92]. xvii [O’R92a]. xviii [Sta93]. xx [Rie93]. xxvi [Har10].

Yampolsky [Cal10]. Yan [Mou10, Cou12, Ira12]. Yandell [Gas13c]. Yaniv [You99]. Yanofsky [Cou09a]. yardstick [Hem93]. Yaschenko [Kat05]. Yde [Hes03]. Year [FC^{+}99, Gas15h, Gas22f, Gas22e, Dif82, Kei08c]. Years [Ali20b, Gas15g, Hem23b, RR20, BC96, She15]. Yehuda [Ful10, Kat12]. Yehudayoff [Cad21]. Yellen [Gas04a]. Yi [Gre22f]. Ying [Lib17]. Yoav [Kin14]. Yona [Pap13b]. Yoram [Dek97]. Yorick [Pap05a]. York [Aro98, Zha05]. Yourself [Ali21c, Har21]. Yttri [Pru00]. Yu [dW07]. Yuri [Sim04]. Yves [Azi14, Ful12, Kam13, Pop12, Ter11].

Zampolli [Sub78]. Zax [Gas20d]. Zbigniew [Mas01b]. Zealand [All00]. Zehavi [JH20]. Zeilberger [Kre00]. Zelle [Jan10b]. Zero [Ban14a, Bel11, Kap99, Rei06, RG06, Bla06, Upa10]. Zero-Knowledge [RG06, Upa10]. Zhang [Fab11, Gre16g]. Zhu [Cen03]. Ziegler [Bón11c]. Zimand [Oli06]. Zimmermann [Tuc66, Yan12]. Zohar [Fur11, Sis93, Vla94]. Zomaya [Cha09, Ste11b]. Zumel [Mil17a]. zur [Hom17].

References

Aggarwal:2008:TRG

[AAC^{+}08] Gagan Aggarwal, Nir Ailon, Florin Constantin, Eyal Even-Dar, Jon Feldman, Gereon Frahling, Monika Rauch Henzinger, S. Muthukrishnan, Noam Nisan, Martin Pál, Mark Sandler, and Anastasios Sidiropoulos. Theory research at Google. ACM SIGACT News, 39(2):10–28, June 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Abraham:2011:DCM

[AAH11] Ittai Abraham, Lorenzo Alvisi, and Joseph Y. Halpern. Distributed computing meets game theory: combining insights from two fields. ACM SIGACT News, 42(2):69–76, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Aaronson:2005:GCN

[Aar05] Scott Aaronson. Guest column: NP-complete problems and physical reality. ACM SIGACT News, 36(1):30–52, March 2005. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Aaronson:2007:BRB

[Aar07] Scott Aaronson. Book review: The Access Principle, by John Willinsky, MIT Press, 2005. ACM SIGACT News, 38(4):19–23, December 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Wil06].

Aaronson:2020:BBF

[Aar20] Scott Aaronson. The busy beaver frontier. ACM SIGACT News, 51(3):32–54, September 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3427361.3427369.

Aharonov:2013:GCQ

[AAV13] Dorit Aharonov, Itai Arad, and Thomas Vidick. Guest column: the quantum PCP conjecture. ACM SIGACT News, 44(2):47–79, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Aazami:2018:RGT

[Aaz18] Amir Babak Aazami. Review of Game Theory, Alive. ACM SIGACT News, 49(4):11–12, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Arora:2009:CCM

[AB09a] Sanjeev Arora and Boaz Barak. Computational Complexity: a Modern Approach. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-42426-7. xxiv + 579 pp. LCCN QA267.7 .A76 2009.

Atallah:2009:ATC

[AB09b] Mikhail J. Atallah and Marina Blanton, editors. Algorithms and Theory of Computation Handbook: General Concepts and Techniques. Chapman and Hall/CRC applied algorithms and data structures series. Taylor and Francis, Boca Raton, FL, USA, second edition, 2009. ISBN 1-58488-822-9 (vol. 1). ???? pp. LCCN QA76.9.A43 A432 2009.

Atallah:2010:ATC

[AB10] Mikhail J. Atallah and Marina Blanton, editors. Algorithms and Theory of Computation Handbook: Special Topics and Techniques. Chapman and Hall/CRC applied algorithms and data structures series. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2010. ISBN 1-58488-820-2 (v. 2). ???? pp. LCCN QA76.9.A43 A433 2010.

Applegate:2006:TSP

[ABCC06] David L. Applegate, Robert E. Bixby, Vasek Chvátal, and William J. Cook. The Traveling Salesman Problem: a Computational Study. Princeton series in applied mathematics. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 1-4008-4110-0 (e-book), 0-691-12993-2 (hardcover). ix + 593 pp. LCCN QA164 .A67 2006.

Abiteboul:1995:RP

[Abi95] Serge Abiteboul. Report on PODS’95. ACM SIGACT News, 26 (4):16–18, December 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Amir:1995:CTC

[ABL^{+}95] Amihood Amir, Manuel Blum, Michael Loui, John Savage, and Carl Smith. Contributions of theoretical computer science. ACM SIGACT News, 26(4):2–4, December 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Alistarh:2023:BSP

[ABLO23] Dan Alistarh, Alkida Balliu, Dimitrios Los, and Sean Ovens. A brief summary of PODC 2022. ACM SIGACT News, 54(1):106– 112, March 2023. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3586165.3586179.

Adjeroh:2008:BWT

[ABM08] Donald Adjeroh, Tim Bell, and Amar Mukherjee. The Burrows– Wheeler Transform: Data Compression, Suffix Arrays, and Pattern Matching. Springer Science+Business Media, LLC, Boston, MA, 2008. ISBN 0-387-78908-1, 0-387-78909-X. xxii + 351 pp. LCCN QA76.9.D33 A35 2008.

Abramson:2011:BRB

[Abr11] Myriam Abramson. Book review: Combinatorics: A Problem Oriented Approach, by Daniel A. Marcus. ACM SIGACT News, 42 (3):40–41, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Mar98].

Abramson:2013:RRT

[Abr13] Myriam Abramson. Review of Resources for teaching discrete mathematics: classroom projects, history modules, and articles edited by Brian Hopkins. ACM SIGACT News, 44(4):31–34, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Abadi:1996:TO

[AC96] Martín Abadi and Luca Cardelli. A Theory of Objects. Monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94775-2 (hardcover). xiii + 396 pp. LCCN QA76.64

.A22 1996. URL http://www.loc.gov/catdir/enhancements/
fy0812/96017038-d.html.

Aceto:1992:ARP

[Ace92] Luca Aceto. Action Refinement in Process Algebras. Distinguished dissertations in computer science. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-43111-5. viii + 273 pp. LCCN QA76.58 .A23 1992. URL http://www.loc.gov/

catdir/description/cam025/92249890.html;   http://www.
loc.gov/catdir/toc/cam029/92249890.html. Thesis (Ph.D.),

University of Sussex, 1990.

ACM:2012:PAS

[ACM12] ACM, editor. Proceedings of the 2012 ACM Symposium on Principles of Distributed Computing. ACM Press, New York, NY 10036, USA, 2012. ISBN 1-4503-1450-3. LCCN ????

Adams:1974:LTC

[Ada74] William J. Adams. The Life and Times of the Central Limit Theorem. Kaedmon Pub. Co., New York, NY, USA, 1974. ISBN 0-913002-00-3. 119 + 9 pp. LCCN QA273.67 .A3.

Adams:2009:RCE

[Ada09a] Colin Conrad Adams. Riot at the Calc Exam and Other Mathematically Bent Stories. American Mathematical Society, Providence, RI, USA, 2009. ISBN 0-8218-4817-8. xvi + 271 pp.

LCCN QA99 .A235 2009. URL http://www.loc.gov/catdir/
toc/fy1001/2009009835.html.

Adams:2009:LTC

[Ada09b] William J. Adams. The Life and Times of the Central Limit Theorem, volume 35 of History of mathematics. American Mathematical Society, Providence, RI, USA, second edition, 2009. ISBN 0-8218-4899-2. xii + 195 pp. LCCN QA273.67 .A3 2009.

Adamopoulos:2011:BRB

[Ada11] Yiorgos Adamopoulos. Book review: Algorithms on Strings, by Crochemore, Hanchart and Lecroq. Published by Cambridge 2007, 392 pages, Hardcover, $60.00. ACM SIGACT News, 42(1):36–37, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [CHL07].

Apt:2009:VSC

[AdBO09] Krzysztof R. Apt, Frank S. de Boer, and E.-R. Olderog. Verification of Sequential and Concurrent Programs. Texts in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2009. ISBN 1-84882-744-X (hardcover). xxiii + 502 pp. LCCN QA76.76.V47 A67 2009.

Abrego:2015:CGC

[ÁDFT15] Bernardo Ábrego, Adrian Dumitrescu, Silvia Fernández, and Csaba D. Tóth. Computational geometry column 61. ACM SIGACT News, 46(2):65–77, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Arunachalam:2017:GCS

[AdW17] Srinivasan Arunachalam and Ronald de Wolf. Guest column: a survey of quantum learning theory. ACM SIGACT News, 48(2): 41–67, June 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Arfaoui:2014:WCC

[AF14] Heger Arfaoui and Pierre Fraigniaud. What can be computed without communications? ACM SIGACT News, 45(3):82–104, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Aksenov:2020:RHC

[AF20] Vitaly Aksenov and Alexey Fyodorov. Review of the Hydra Conference on Concurrent and Distributed Computing Hydra 2020. ACM SIGACT News, 51(4):85–88, December 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3444815.3444829.

Allender:1996:FCC

[AFG^{+}96] E. Allender, J. Feigenbaum, J. Goldsmith, T. Pitassi, and S. Rudich. The future of computational complexity theory: part II. ACM SIGACT News, 27(4):3–7, December 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Afrati:2005:RP

[Afr05] Foto N. Afrati. Report on PODS 2005. ACM SIGACT News, 36 (3):39–40, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Afshar:1982:CPN

[Afs82] Siroos K. Afshar. On the complexity of permutation network design. ACM SIGACT News, 14(3):31–35, Summer 1982. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Avni:2015:DR

[AG15] Hillel Avni and Rati Gelashvili. DISC 2015 review. ACM SIGACT News, 46(4):103–106, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Alistarh:2018:RAA

[AG18] Dan Alistarh and Rati Gelashvili. Recent algorithmic advances in population protocols. ACM SIGACT News, 49(3):63–73, September 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Agarwal:1991:IDA

[Aga91] Pankaj K. Agarwal. Intersection and Decomposition Algorithms for Planar Arrangements. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-40446-0. xvii + 277 pp. LCCN QA167 .A33

1991. URL http://www.loc.gov/catdir/description/cam025/
91155565.html;   http://www.loc.gov/catdir/toc/cam028/
91155565.html.

Aguilera:2004:PST

[Agu04] Marcos Kawazoe Aguilera. A pleasant stroll through the land of infinitely many creatures. ACM SIGACT News, 35(2):36–59, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Aguilera:2012:DCH

[Agu12] Marcos K. Aguilera, editor. Distributed Computing: 26th International Symposium, DISC 2012, Salvador, Brazil, October 16–18, 2012. Proceedings, volume 7611 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-33651-5. LCCN TK5105.5-5105.9.

Amirazizi:1983:TMP

[AH83] Hamid R. Amirazizi and Martin E. Hellman. Time-memory-processor tradeoffs. ACM SIGACT News, 15(1):18–19, Winter– Spring 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Alur:1991:SNL

[AH91] Rajeev Alur and Thomas A. Henzinger. Sigact news logic column 1: Time for logic. ACM SIGACT News, ??(??):??, ???? 1991. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Anderson:2006:ATM

[AH06] James A. (James Andrew) Anderson and Thomas J. Head. Automata Theory with Modern Applications. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-84887-3, 0-521-61324-8 (paperback). viii + 255 pp. LCCN QA267 .A53 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0618/2006281694-d.
html;   http://www.loc.gov/catdir/enhancements/fy0618/
2006281694-t.html;   http://www.loc.gov/catdir/enhancements/
fy0733/2006281694-b.html.

Aigner:2007:CE

[Aig07] Martin Aigner. A Course in Enumeration, volume 238 of Graduate texts in mathematics, 0072-5285. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-39032-4 (hardcover). x + 561 pp. LCCN QA164.8

.A36 2007. URL http://www.loc.gov/catdir/enhancements/
fy0825/2007928344-d.html.

Aho:1997:EOT

[AJK^{+}97] Alfred V. Aho, David S. Johnson, Richard M. Karp, S. Rao Kosaraju, Catherine C. McGeoch, Christos H. Papadimitriou, and Pavel Pevzner. Emerging opportunities for theoretical computer science. ACM SIGACT News, 28(3):65–74, September 1997. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Abiteboul:1990:DTC

[AK90] Serge Abiteboul and Paris C. Kanellakis. Database theory column: Query languages for complex object databases. ACM SIGACT News, 21(3):9–18, Summer 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Aksenov:2020:RTS

[AK20] Vitaly Aksenov and Petr Kuznetsov. Review of the Third Summer School on the Practice and Theory of Distributed Computing SPTDC 2020. ACM SIGACT News, 51(4):82–84, December 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3444815.
3444828.

Anta:2018:FID

[AKGN18] Antonio Fernández Anta, Kishori Konwar, Chryssis Georgiou, and Nicolas Nicolaou. Formalizing and implementing distributed ledger objects. ACM SIGACT News, 49(2):58–76, June 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Akman:1992:USE

[Akm92] Varol Akman. Undaunted sets (extended abstract). ACM SIGACT News, 23(1):47–48, Winter 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Amirazizi:1983:CKP

[AKR83] Hamid R. Amirazizi, Ehud D. Karnin, and Justin M. Reyneri. Compact knapsacks are polynomially solvable. ACM SIGACT News, 15(1):20–22, Winter–Spring 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Antoniou:2007:POA

[AL07] Andreas Antoniou and Wusheng Lu. Practical Optimization: Algorithms and Engineering Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-71106-6, 1-4419-4383-8. xvii + 669 pp. LCCN

QA402.5 .A68 2007. URL http://site.ebrary.com/lib/
stanford/Doc?id=10230218;   http://www.myilibrary.com?
id=133757;   http://www.springerlink.com/openurl.asp?genre=
book\%26isbn=978-0-387-71106-5.

Alagarsamy:2003:SMA

[Ala03] K. Alagarsamy. Some myths about famous mutual exclusion algorithms. ACM SIGACT News, 34(3):94–103, September 2003. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Albert:2003:TSU

[Alb03] Jim Albert. Teaching Statistics Using Baseball. Classroom resource materials. Mathematical Association of America, Washington, DC, USA, 2003. ISBN 0-88385-727-8. xi + 288 pp. LCCN QA276.18 .A44 2003.

Alistarh:2019:DCC

[Ali19] Dan Alistarh. Distributed computing column 76: Annual review 2019. ACM SIGACT News, 50(4):31–32, December 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3374857.3374865.

Alistarh:2020:DCCa

[Ali20a] Dan Alistarh. Distributed computing column 77 consensus dynamics: an overview. ACM SIGACT News, 51(1):57, March 2020. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

URL https://dl.acm.org/doi/10.1145/3388392.3388402.

Alistarh:2020:DCCb

[Ali20b] Dan Alistarh. Distributed computing column 78: 60 years of mastering concurrent computing through sequential thinking. ACM SIGACT News, 51(2):58, June 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.

org/doi/10.1145/3406678.3406689.

Alistarh:2020:DCCc

[Ali20c] Dan Alistarh. Distributed computing column 79: Using round elimination to understand locality. ACM SIGACT News, 51(3): 62, September 2020. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3427361.3427373.

Alistarh:2020:DCCd

[Ali20d] Dan Alistarh. Distributed computing column 80: Annual review 2020. ACM SIGACT News, 51(4):73–74, December 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3444815.3444826.

Alistarh:2021:DCCa

[Ali21a] Dan Alistarh. Distributed computing column 81: Byzantine agreement with less communication: Recent advances. ACM SIGACT News, 52(1):70, March 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3457588.3457599.

Alistarh:2021:DCCb

[Ali21b] Dan Alistarh. Distributed computing column 82 distributed computability: a few results masters students should know. ACM SIGACT News, 52(2):91, June 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.

org/doi/10.1145/3471469.3471483.

Alistarh:2021:DCCc

[Ali21c] Dan Alistarh. Distributed computing column 83 five ways not to fool yourself: Designing experiments for understanding performance. ACM SIGACT News, 52(3):60, September 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3494656.3494667.

Alistarh:2021:DCC

[Ali21d] Dan Alistarh. Distributed computing column 84: Perspectives on the paper “CCS Expressions, Finite State Processes, and Three Problems of Equivalence”. ACM SIGACT News, 52(4):74–75, December 2021. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3510382.3510393. See [BB21].

Alistarh:2022:DCC

[Ali22] Dan Alistarh. Distributed computing column 85: Elastic consistency: a consistency criterion for distributed optimization. ACM SIGACT News, 53(2):63, June 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.

org/doi/10.1145/3544979.3544990.

Alistarh:2023:DCCa

[Ali23a] Dan Alistarh. Distributed computing column 86: a summary of PODC 2022. ACM SIGACT News, 54(1):105, March 2023. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

URL https://dl.acm.org/doi/10.1145/3586165.3586178.

Alistarh:2023:DCCc

[Ali23b] Dan Alistarh. Distributed computing column 86: The environmental cost of our conferences. ACM SIGACT News, 54(4):92– 93, December 2023. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3639528.3639536.

Alistarh:2023:DCCb

[Ali23c] Dan Alistarh. Distributed computing column 87: Recent advances in multi-pass graph streaming lower bounds. ACM SIGACT News, 54(3):46–47, September 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3623800.3623807.

Allen:1972:DAD

[All72] C. David Allen. Derivation of axiomatic definitions of programming languages from algorithmic definitions. ACM SIGACT News, 4(14):15–26, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Allender:1997:MCC

[All97] Eric Allender. Making computation count: arithmetic circuits in the nineties. ACM SIGACT News, 28(4):2–15, December 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Allender:1998:BRB

[All98] Eric Allender. Book review: Complexity Theory Retrospective II. Edited by: Lane A. Hemaspaandra and Alan L. Selman (Springer Verlag). ACM SIGACT News, 29(1):2–5, March 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [HS97].

Allender:2000:RAS

[All00] Eric Allender. Report on the annual summer meeting of the New Zealand mathematics research institute. ACM SIGACT News, 31(2):60–61, June 1, 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Abelson:2008:BBY

[ALL08] Harold Abelson, Ken Ledeen, and Harry R. Lewis. Blown to Bits: Your Life, Liberty, and Happiness After the Digital Explosion. Addison-Wesley, Reading, MA, USA, 2008. ISBN 0-13-713559-9 (hardcover). xiv + 366 pp. LCCN QA76.9.C66 A245 2008.

Allender:2023:GCP

[All23] Eric Allender. Guest column: Parting thoughts and parting shots (read on for details on how to win valuable prizes!). ACM SIGACT News, 54(1):63–81, March 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/

doi/10.1145/3586165.3586175.

Alpern:2018:EWD

[Alp18] Bowen Alpern. Edsger W. Dijkstra: The man behind the prize. ACM SIGACT News, 49(4):64–65, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Alt:1975:LOS

[AM75] Helmut Alt and Kurt Mehlhorn. A language over a one symbol alphabet requiring only O(log log n) space. ACM SIGACT News, 7(4):31–33, November 1975. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ambainis:2004:QSA

[Amb04] Andris Ambainis. Quantum search algorithms. ACM SIGACT News, 35(2):22–35, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ambuhl:2017:SNO

[Amb17] Christoph Ambühl. SIGACT News online algorithms column 31. ACM SIGACT News, 48(3):68–82, September 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Alistarh:2022:ECC

[AMN22] Dan Alistarh, Ilia Markov, and Giorgi Nadiradze. Elastic consistency: a consistency criterion for distributed optimization. ACM SIGACT News, 53(2):64–82, June 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.

org/doi/10.1145/3544979.3544991.

Amos:1999:BRB

[Amo99] Martyn Amos. Book review: DNA Based Computers III, edited by Harvey Rubin and David Harlan Wood. ACM SIGACT News, 30(4):10–12, December 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [RW99].

Amos:2005:TED

[Amo05] Martyn Amos. Theoretical and Experimental DNA Computation. Natural computing series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-65773-

8. xiii + 172 pp. LCCN QP625.N89 A56 2005. URL http:/

/site.ebrary.com/lib/stanford/Doc?id=10143901;   http:
//www.loc.gov/catdir/enhancements/fy0662/2004116588-d.
html;   http://www.loc.gov/catdir/enhancements/fy0812/
2004116588-t.html;   http://www.myilibrary.com?id=30647;
 http://www.springerlink.com/openurl.asp?genre=book\%26isbn=
978-3-540-65773-6.

Amoretti:2020:REP

[Amo20] Michele Amoretti. Review of Elements of Parallel Computing. ACM SIGACT News, 51(3):10–13, September 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3427361.3427365.

Albert:2009:GNC

[AN09] Michael H. Albert and Richard J. Nowakowski, editors. Games Of No Chance 3, volume 56 of Mathematical Sciences Research Institute publications. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-86134-9 (hardcover), 0-521-67854-4 (paperback). x + 575 pp. LCCN QA269 .G374 2009.

Angluin:1976:FRA

[Ang76] Dana Angluin. The Four Russians’ Algorithm for Boolean matrix multiplication is optimal in its class. ACM SIGACT News, 8(1): 29–33, January 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). In 8(2) 1976 of this journal, the author writes that the work in this paper is subsumed by a more general treatment in [Sav74].

Angier:1991:GCT

[Ang91] Natalie Angier. The glass ceiling: in theory women swell ranks of science, but remain invisible at the top. ACM SIGACT News, 22(3):38–40, Summer 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Anonymous:2013:DPD

[Ano13] Anonymous. 2013 Dijkstra Prize in Distributed Computing to Nati Linial: 2013 Principles of Distributed Computing Doctoral Dissertation Award to Shiri Chechik and Danupon Nanongkai. ACM SIGACT News, 44(4):81–82, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Apt:1991:VSC

[AO91] Krzysztof R. Apt and E.-R. Olderog. Verification of Sequential and Concurrent Programs. Texts and monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97532-2 (Springer-Verlag New York Berlin Heidelberg), 3-540-97532-2 (Springer-Verlag Berlin Heidelberg New York). xvi + 441 pp. LCCN QA76.76.V47 A67 1991.

Apt:1997:VSC

[AO97] Krzysztof R. Apt and E.-R. Olderog. Verification of Sequential and Concurrent Programs. Graduate texts in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1997.

ISBN 0-387-94896-1 (hardcover). xvii + 364 pp. LCCN

QA76.76.V47 A67 1997. URL http://www.loc.gov/catdir/
enhancements/fy0812/96029771-d.html;   http://www.loc.
gov/catdir/enhancements/fy0812/96029771-t.html.

Agarwal:1998:CG

[AO98] Pankaj K. Agarwal and Joseph O’Rourke. Computational geometry. ACM SIGACT News, 29(3):27–32, September 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Aldous:1998:MDP

[AP98] D. J. (David J.) Aldous and James Propp, editors. Microsurveys in Discrete Probability: DIMACS Workshop, June 2–6, 1997, volume 41 of DIMACS series in discrete mathematics and theoretical computer science, 1052-1798. American Mathematical Society, Providence, RI, USA, 1998. ISBN 0-8218-0827-3 (hardcover). LCCN QA273.A1 M53 1998.

Alevras:2001:LOE

[AP01] Dimitris Alevras and Manfred W. Padberg. Linear Optimization and Extensions: Problems and Solutions. Universitext. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-41744-3 (paperback). ix + 449 pp. LCCN T57.74 .A44 2001. URL http:

//www.loc.gov/catdir/enhancements/fy0816/2001020974-d.
html;   http://www.loc.gov/catdir/enhancements/fy0816/
2001020974-t.html.

Apon:2010:JRB

[Apo10] Daniel Apon. Joint review of Computational Complexity: a Conceptual Perspective, by Oded Goldreich, published by Cambridge University Press, 2008 and Computational Complexity: a Modern Approach, by Sanjeev Arora and Boaz Barak, published by Cambridge University Press, 2009. ACM SIGACT News, 41(4):15–21, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Gol08b, AB09a].

Apon:2011:BRB

[Apo11] Daniel Apon. Book review: Games, Puzzles, & Computation, by Robert A. Hearn and Erik D. Demaine. ACM SIGACT News, 42 (3):45–49, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [HD09].

Apon:2012:BRB

[Apo12] Daniel Apon. Book review: Theory of Computation, by Dexter C. Kozen. ACM SIGACT News, 43(1):23–27, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Koz06].

Apon:2014:RSPb

[Apo14a] Daniel Apon. Review of Selected papers on design of algorithms by Donald E. Knuth. ACM SIGACT News, 45(2):14–16, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Apon:2014:RSPa

[Apo14b] Daniel Apon. Review of Selected papers on discrete mathematics by Donald E. Knuth. ACM SIGACT News, 45(2):10–13, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Apon:2016:RAB

[Apo16] Daniel Apon. Review of: Analysis of Boolean Functions by Ryan O’Donnell. ACM SIGACT News, 47(4):9–12, December 2016. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Augustine:2016:DAF

[APR16] John Augustine, Gopal Pandurangan, and Peter Robinson. Distributed algorithmic foundations of dynamic networks. ACM SIGACT News, 47(1):69–98, March 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Arenas:2018:DTC

[Are18] Marcelo Arenas. Database theory column report on PODS 2018. ACM SIGACT News, 49(4):55–57, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Arfaoui:2013:RS

[Arf13] Heger Arfaoui. A review of SIROCCO 2012. ACM SIGACT News, 44(2):113–118, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Arora:1998:BRB

[Aro98] Anish Arora. Book review: Verification of Sequential and Concurrent Programs, by Krzysztof R. Apt and Ernst-Rüdiger Olderog (Springer-Verlag New York, 1997). ACM SIGACT News, 29(3):46– 48, September 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [AO91, AO97, AdBO09].

Aurenhammer:1991:PVD

[AS91] F. Aurenhammer and G. Stöckl. On the Peeper’s Voronoi diagram. ACM SIGACT News, 22(4):50–59, Fall 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Allouche:2003:AST

[AS03] Jean-Paul Allouche and Jeffrey Outlaw Shallit. Automatic sequences: theory, applications, generalizations. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-82332-3. xvi + 571 pp. LCCN QA267.5.S4 A55 2003. URL http://www.myilibrary.

com?id=43653.

Ashcroft:1972:PCM

[Ash72] E. A. Ashcroft. Program correctness methods and language definition. ACM SIGACT News, 4(14):51–57, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Assadi:2017:SR

[Ass17] Sepehr Assadi. SPAA 2017 review. ACM SIGACT News, 48(4):87– 90, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Assadi:2023:RAM

[Ass23] Sepehr Assadi. Recent advances in multi-pass graph streaming lower bounds. ACM SIGACT News, 54(3):48–75, September 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3623800.
3623808.

Atkinson:1986:ICB

[ASU86] D. M. Atkinson, N. Santoro, and J. Urrutia. On the integer complexity of Boolean matrix multiplication. ACM SIGACT News, 18 (1):53, Summer 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Amouzandeh:2023:SNOa

[ASvS23a] Aflatoun Amouzandeh, Soren Schmitt, and Rob van Stee. SIGACT news online algorithms column 40: 2022 in review. ACM SIGACT News, 54(1):90–104, March 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/

doi/10.1145/3586165.3586177.

Amouzandeh:2023:SNOb

[ASvS23b] Aflatoun Amouzandeh, Sören Schmitt, and Rob van Stee. SIGACT news online algorithms column 41: 2023 in review. ACM SIGACT News, 54(4):108–125, December 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.

org/doi/10.1145/3639528.3639538.

Alur:2017:ASD

[AT17] Rajeev Alur and Stavros Tripakis. Automatic synthesis of distributed protocols. ACM SIGACT News, 48(1):55–90, March 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Atallah:1999:ATC

[Ata99] Mikhail J. Atallah, editor. Algorithms and Theory of Computation Handbook. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1999. ISBN 0-8493-2649-4. ???? pp. LCCN QA76.9.A43 A43 1999. URL http://www.loc.gov/

catdir/enhancements/fy0742/98038016-d.html.

Attiya:2008:NFT

[Att08] Hagit Attiya. Needed: foundations for transactional memory. ACM SIGACT News, 39(1):59–61, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Attiya:2011:SMR

[Att11] Hagit Attiya. Sharing memories, robustly. ACM SIGACT News, 42(4):79–82, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Avigad:2001:BRB

[Avi01] Jeremy Avigad. Book review: Basic Proof Theory: second edition, by A. S. Troelstra and H. Schwichtenberg. Cambridge University Press. ACM SIGACT News, 32(2):15–19, June 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [TS96, TS00].

Attiya:1998:DCF

[AW98] Hagit Attiya and Jennifer Welch. Distributed Computing: Fundamentals, Simulations and Advanced Topics. McGraw-Hill, New York, NY, USA, 1998. ISBN 0-07-709352-6 (paperback). xii + 451 pp. LCCN QA76.9.D5 A77 1998.

Attiya:2004:DCF

[AW04] Hagit Attiya and Jennifer Welch. Distributed Computing: Fundamentals, Simulations, and Advanced Topics. Wiley series on parallel and distributed computing. John Wiley, New York, NY, USA, second edition, 2004. ISBN 0-471-45324-2 (hardcover). xv + 414 pp. LCCN QA76.9.D5 A75 2004. URL http://www.

loc.gov/catdir/bios/wiley046/2004040728.html;   http:/
/www.loc.gov/catdir/description/wiley041/2004040728.html;
 http://www.loc.gov/catdir/toc/wiley041/2004040728.html.

Anzala-Yamajako:2012:RBC

[AY12] Alexandre Anzala-Yamajako. Review of Algorithmic Cryptanalysis, by Antoine Joux. ACM SIGACT News, 43(4):13–16, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Aycock:2006:CVM

[Ayc06] John Daniel Aycock. Computer Viruses and Malware, volume 22 of Advances in information security. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-30236-0, 0-387-34188-9 (e-book). xvi + 227 pp. LCCN QA76.76.C68 A96 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0824/2006925091-d.
html;   http://www.loc.gov/catdir/enhancements/fy0824/
2006925091-t.html.

Aycock:2011:SA

[Ayc11] John Daniel Aycock. Spyware and Adware, volume 50 of Advances in information security, 1568-2633. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 0-387-77740-7 , 0-387-77741-5 (e-book). xiv + 145 pp. LCCN QA76.76.S69 A93 2011.

Ayub:2012:BRB

[Ayu12] Abu Mohammad Omar Shehab Uddin Ayub. Book review: The Cryptoclub: Using Mathematics to Make and Break Secret Codes, by Janet Beissinger and Vera Pless. ACM SIGACT News, 43(1): 9–14, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BP06].

Aigner:1998:PB

[AZ98] Martin Aigner and Günter M. Ziegler. Proofs from The Book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-63698-6 (hardcover). viii + 199 pp. LCCN QA36 .A36 1998.

Aigner:1999:PB

[AZ99] Martin Aigner and Günter M. Ziegler. Proofs from The Book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second corrected printing edition, 1999. ISBN 3-540-63698-6. viii + 199 pp. LCCN QA36 .A36 1999.

Aigner:2001:PB

[AZ01] Martin Aigner and Günter M. Ziegler. Proofs from The Book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2001. ISBN 3-540-67865-4. viii + 215 pp. LCCN QA36 .A36 2001.

Aigner:2004:PB

[AZ04] Martin Aigner and Günter M. Ziegler. Proofs from The Book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2004. ISBN 3-540-40460-0. viii + 239

pp. LCCN QA36 .A36 2004. URL http://www.loc.gov/catdir/
enhancements/fy0817/2003060832-d.html.

Aigner:2010:PB

[AZ10] Martin Aigner and Günter M. Ziegler. Proofs from The Book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2010. ISBN 3-642-00855-0. viii + 274 pp. LCCN QA36 .A36 2010.

Aziz:2010:BRBa

[Azi10a] Haris Aziz. Book review: Multiagent Systems: Algorithmic, Game-Theoretic, and Logical Foundations, by Y. Shoham and K. Leyton-Brown, Cambridge University Press, 2008. ACM SIGACT News, 41(1):34–37, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [SLB09].

Aziz:2010:BRBb

[Azi10b] Haris Aziz. Book review: The Princeton Companion to Mathematics, by Timothy Gowers, June Barrow-Green and Imre Leader, Princeton University Press, 2008. ACM SIGACT News, 41(1): 41–43, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [GBGL08].

Aziz:2011:BRB

[Azi11] Haris Aziz. Book review: Graph Theory: a Problem Oriented Approach, by Daniel Marcus. ACM SIGACT News, 42(4):31–32, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Mar08a].

Aziz:2012:RBP

[Azi12a] Haris Aziz. Review of In Pursuit of the Traveling Salesman, by William J. Cook. ACM SIGACT News, 43(3):51–53, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Aziz:2012:RBT

[Azi12b] Haris Aziz. Review of Matching Theory, by László Lovász and Michael D. Plummer. ACM SIGACT News, 43(3):38–40, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Aziz:2014:RBF

[Azi14] Haris Aziz. Review of boolean functions: theory, algorithms, and applications by Yves Crama and Peter L. Hammer. ACM SIGACT News, 45(1):20–23, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Aziz:2016:RNC

[Azi16] Haris Aziz. Review of: The Nature of Computation by Cristopher Moore and Stephan Mertens. ACM SIGACT News, 47(1):6–8, March 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Aznaveh:2013:RBB

[Azn13] Mohsen Mahmoudi Aznaveh. Review of Introduction to Bio-Ontologies by Peter N. Robinson and Sebastian Bauer. ACM SIGACT News, 44(1):40–42, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Aznaveh:2014:RCB

[Azn14] Mohsen Mahmoudi Aznaveh. Review of Clustering in bioinformatics and drug design by John D. MacCuish and Norah E. Mac-Cuish. ACM SIGACT News, 45(3):15–17, September 2014. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ben-Amram:1991:SNN

[BA91] Amir M. Ben-Amram. Some notions on notations. ACM SIGACT News, 22(4):60–62, Fall 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ben-Amram:1995:WPM

[BA95] Amir M. Ben-Amram. What is a “pointer machine”? ACM SIGACT News, 26(2):88–95, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ben-Amram:2004:CTP

[BA04] Amir M. Ben-Amram. A complexity-theoretic proof of a recursiontheoretic theorem. ACM SIGACT News, 35(2):111–112, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ben-Amram:2005:CTT

[BA05] Amir M. Ben-Amram. The Church–Turing thesis and its lookalikes. ACM SIGACT News, 36(3):113–114, September 2005. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Baase:1978:CAI

[Baa78] Sara Baase. Computer Algorithms: Introduction to Design and Analysis. Addison-Wesley series in computer science. Addison-Wesley, Reading, MA, USA, 1978. ISBN 0-201-00327-9. xvii + 286 pp. LCCN QA76.6 .B25.

Baase:1988:CAI

[Baa88] Sara Baase. Computer Algorithms: Introduction to Design and Analysis. Addison-Wesley series in computer science. Addison-Wesley, Reading, MA, USA, second edition, 1988. ISBN 0-201-06035-3. xv + 415 pp. LCCN QA76.6 .B25 1988. US$36.95.

Ballard:1997:INC

[Bal97] Dana H. Ballard. An Introduction to Natural Computation. Complex Adaptive Systems. MIT Press, Cambridge, MA, USA, 1997.

ISBN 0-262-02420-9 (hardcover), 0-262-52258-6 (paperback), 0-585-08766-0 (e-book). xxii + 307 pp. LCCN QP356 .B345 1997.

Bane:2014:RZI

[Ban14a] John Tucker Bane. Review of From Zero to Infinity by Constance Reid. ACM SIGACT News, 45(4):21–23, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bane:2014:RMT

[Ban14b] John Tucker Bane. Review of Mathematical Treks by Ivars Peterson. ACM SIGACT News, 45(4):36–38, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bane:2015:RMB

[Ban15] John Tucker Bane. Review of: Math Bytes 5 of Math Bytes by Tim Chartier. ACM SIGACT News, 46(4):12–13, December 2015. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bard:2009:AC

[Bar09] Gregory V. Bard. Algebraic Cryptanalysis. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-88756-3, 0-387-88757-1 (e-book). xxxiii + 356 pp.

LCCN Z104 .B314 2009. URL http://www.loc.gov/catdir/
enhancements/fy1202/2009929845-d.html.

Barenboim:2010:RP

[Bar10] Leonid Barenboim. A review of PODC 2010. ACM SIGACT News, 41(4):100–105, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Barbay:2012:BRB

[Bar12] Jérémy Barbay. Book review: Understanding and Applying Cryptography and Data Security, by Adam J. Elbirt. ACM SIGACT News, 43(1):18–21, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Elb09].

Barenboim:2015:DSB

[Bar15] Leonid Barenboim. Distributed symmetry: Breaking from a local point of view. ACM SIGACT News, 46(4):84–87, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Batini:1993:AGD

[Bat93] Carlo Batini. Applications of graph drawing to software engineering (abstract). ACM SIGACT News, 24(1):57, Winter 1993. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bauer:2010:OFC

[Bau10] Friedrich Ludwig Bauer, editor. Origins and foundations of computing: in cooperation with Heinz Nixdorf MuseumsForum. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-02991-4 (hardcover), 3-642-02992-2 (e-book). viii + 142 pp. LCCN QA76.17 .B38 2010.

Baumeister:2017:GCC

[Bau17] D. Baumeister. Guest column: Complexity of verification in abstract argumentation frameworks. ACM SIGACT News, 48(4):43– 56, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Baylor:1995:GIW

[Bay95] Sandra Johnson Baylor. Graduate information for women in CS&E. ACM SIGACT News, 26(1):27–30, March 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bennett:1987:QPK

[BB87] Charles H. Bennett and Gilles Brassard. Quantum public key distribution reinvented. ACM SIGACT News, 18(4):51–53, Summer 1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bennett:1989:EQC

[BB89] C. H. Bennett and G. Brassard. Experimental quantum cryptography: the dawn of a new era for quantum cryptography: the experimental prototype is working. ACM SIGACT News, 20(4):78–80, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bartocci:2021:PCE

[BB21] Ezio Bartocci and Michael A. Bender. A perspective on “CCS Expressions, Finite State Processes, and Three Problems of Equivalence”. ACM SIGACT News, 52(4):76–77, December 2021. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

URL https://dl.acm.org/doi/10.1145/3510382.3510394. See

[Ali21d].

Barga:2009:CCA

[BBAGP09] Roger Barga, Jose Bernabeu-Auban, Dennis Gannon, and Christophe Poulain. Cloud computing architecture and application programming: DISC’09 tutorial, half day, Sept. 22nd 2009. ACM SIGACT News, 40(2):94–95, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Baig:2020:PR

[BBDD20] Ahad Mirza Baig, Alkida Balliu, Peter Davies, and Michal Dory. PODC 2020 review. ACM SIGACT News, 51(4):75–81, December 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3444815.
3444827.

Bauer:1973:NDF

[BBF^{+}73] M. Bauer, D. Brand, M. Fischer, A. Meyer, and M. Paterson. A note on disjunctive form tautologies. ACM SIGACT News, 5(2): 17–20, April 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bloch:1994:HHH

[BBG94] Stephen A. Bloch, Jonathan F. Buss, and Judy Goldsmith. How hard are n^{2}-hard problems? ACM SIGACT News, 25(2):83–85, June 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bollobas:1991:PCA

[BC^{+}91] Béla Bollobás, Fan R. K. Chung, et al., editors. Probabilistic Combinatorics and its Applications, volume 44 of Proceedings of symposia in applied mathematics; AMS short course lecture notes. American Mathematical Society, Providence, RI, USA, 1991. ISBN 0-8218-5500-X. xv + 196 pp. LCCN QA273.45 .P76 1992.

Brassard:1996:YQC

[BC96] Gilles Brassard and Claude Crépeau. 25 years of quantum cryptography. ACM SIGACT News, 27(3):13–24, September 1996. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Brualdi:2009:CAM

[BC09] Richard A. Brualdi and Dragoš M. Cvetković. A Combinatorial Approach to Matrix Theory and Its Applications. Discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, 2009. ISBN 1-4200-8223-X (hardcover). xiii + 267 pp.

LCCN QA188 .B778 2009. URL http://www.loc.gov/catdir/
toc/ecip0815/2008014627.html.

Buttcher:2010:IRI

[BCC10] Stefan Büttcher, Charles L. A. Clarke, and Gordon V. Cormack. Information Retrieval: Implementing and Evaluating Search Engines. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-02651-1 (hardcover). xxiv + 606 pp. LCCN TK5105.884 .B98 2010.

Bakali:2022:GCP

[BCG^{+}22] Eleni Bakali, Aggeliki Chalki, Andreas Göbel, Aris Pagourtzis, and Stathis Zachos. Guest column: a panorama of counting problems the decision version of which is in P^{3}. ACM SIGACT News, 53 (3):46–68, September 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3561064.3561072.

Becchetti:2020:CDO

[BCN20] Luca Becchetti, Andrea Clementi, and Emanuele Natale. Consensus dynamics: an overview. ACM SIGACT News, 51(1):58– 104, March 2020. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3388392.3388403.

Bohler:2003:PBB

[BCRV03] E. Böhler, N. Creignou, S. Reith, and H. Vollmer. Playing with Boolean blocks, part I: Post’s lattice with applications to complexity theory. ACM SIGACT News, 34(4):38–52, December 2003. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bohler:2004:PBB

[BCRV04] E. Böhler, N. Creignou, S. Reith, and H. Vollmer. Playing with Boolean blocks, part II: constraint satisfaction problems. ACM SIGACT News, 35(1):22–35, March 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Burgisser:1997:ACT

[BCS97] Peter Bürgisser, Michael Clausen, and Mohammad Amin Shokrollahi. Algebraic Complexity Theory, volume 315 of Grundlehren der mathematischen Wissenschaften, 0072-7830. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-60582-7 (Berlin: hardcover). xxiii + 618 pp.

LCCN QA267.7 .B87 1997. URL http://www.loc.gov/catdir/
enhancements/fy0812/96029754-d.html;   http://www.loc.
gov/catdir/enhancements/fy0812/96029754-t.html.

Blum:1998:CRC

[BCSS98] Lenore Blum, Felipe Cucker, Michael Shub, and Steve Smale. Complexity and Real Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98281-7 (hardcover). xvi + 453 pp. LCCN QA76

.C5474 1998. URL http://www.loc.gov/catdir/enhancements/
fy0815/97022859-d.html;   http://www.loc.gov/catdir/enhancements/

fy0815/97022859-t.html. Foreword by Richard M. Karp.

Birman:2009:TCC

[BCvR09] Ken Birman, Gregory Chockler, and Robbert van Renesse. Toward a cloud computing research agenda. ACM SIGACT News, 40(2): 68–80, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ben-David:2018:PR

[BD18] Naama Ben-David. PODC 2018 review. ACM SIGACT News, 49 (4):83–88, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ben-David:2019:PR

[BDCDL19] Naama Ben-David, Yi-Jun Chang, Michal Dory, and Dean Leitersdorf. PODC 2019 review. ACM SIGACT News, 50(4):33–45, December 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-

5827 (electronic). URL https://dl.acm.org/doi/abs/10.1145/
3374857.3374866.

Best:2001:PNA

[BDK01] Eike Best, Raymond Devillers, and Maciej Koutny. Petri Net Algebra. Monographs in theoretical computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-67398-9. xi + 378 pp. LCCN QA76.58

.B48 2001. URL http://www.loc.gov/catdir/enhancements/
fy0816/2001017045-d.html.

Blackburn:2001:ML

[BdRV01] Patrick Blackburn, Maarten de Rijke, and Yde Venema. Modal Logic, volume 53 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-80200-8 (hardcover). xxii + 554 pp. LCCN QA9.46

.B58 2001. URL http://www.loc.gov/catdir/description/
cam021/00054667.html;   http://www.loc.gov/catdir/toc/
fy02/00054667.html.

Beaumont:2007:BRB

[Bea07] James C. Beaumont. Book review: Symbolic Asymptotics, by John R. Shackell, Springer Verlag, 243 pages, $79.95. ACM SIGACT News, 38(2):11–16, June 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sha04].

Beck:2008:CGT

[Bec08] József Beck. Combinatorial games: tic-tac-toe theory, volume 114 of Encyclopedia of mathematics and its applications. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-46100-6 (hardcover). xiv + 732 pp. LCCN QA269 .B335 2008. URL

http://www.loc.gov/catdir/toc/fy0805/2008275067.html.

Behrends:2008:FMM

[Beh08] Ehrhard Behrends. Five-Minute Mathematics. American Mathematical Society, Providence, RI, USA, 2008. ISBN 0-8218-4348-6. xxvii + 380 pp. LCCN QA93 .B4413 2008.

Bella:2007:FCS

[Bel07] Giampaolo Bella. Formal Correctness of Security Protocols, volume 4752 of Information security and cryptography. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2007. ISBN 3-540-68136-1, 3-540-68134-5. xvi + 367 pp. LCCN TK5105.59 .B45 2007eb.

Belle:2011:BRB

[Bel11] Vaishak Belle. Book review: From Zero to Infinity: What Makes Numbers Interesting, by Constance Reid. ACM SIGACT News, 42 (2):10–11, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Rei06].

Belle:2014:RPH

[Bel14] Vaishak Belle. Review of Programming with higher-order logic by Dale Miller and Gopalan Nadathur. ACM SIGACT News, 45(2): 32–35, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Benson:1973:APA

[Ben73a] David B. Benson. An alternate proof of Aho & Ullman’s LR(k) viable prefix theorem. ACM SIGACT News, 5(4):11–14, October 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Benson:1973:ACN

[Ben73b] David B. Benson. Another comment on the null word definition problem. ACM SIGACT News, 5(1):14–17, January 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bennison:1974:STS

[Ben74] Victor L. Bennison. Saving tapes in the simulation of multihead Turing machines. ACM SIGACT News, 6(2):23–26, April 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Benson:1998:BRB

[Ben98] Gary Benson. Book review: Algorithms on Strings, Trees, and Sequences: Computer Science and Computational Biology, by Dan Gusfield (Cambridge University Press, Cambridge, England, 1997). ACM SIGACT News, 29(3):43–46, September 1998. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Gus97a].

Bender:2006:BRB

[Ben06] Adam Bender. Book review: Codes: The Guide to Secrecy from Ancient to Modern Times, by Richard A. Mollin, Chapman & Hall/CRC, 2005. ACM SIGACT News, 37(1):18–21, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Mol05].

Benedikt:2012:RP

[Ben12] Michael Benedikt. Report on PODS 2012. ACM SIGACT News, 43(3):85–86, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bennett:2023:CSV

[Ben23] Huck Bennett. The complexity of the shortest vector problem. ACM SIGACT News, 54(1):37–61, March 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3586165.3586172.

Berry:1972:EMT

[Ber72] Daniel M. Berry. The equivalence of models of tasking. ACM SIGACT News, 4(14):170–190, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bernstein:1986:TS

[Ber86] R. Bernstein. Testing for semilattices. ACM SIGACT News, 18(1): 49–50, Summer 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Berman:1992:CSW

[Ber92] Francine Berman. CRA status of women column. ACM SIGACT News, 23(1):14–17, Winter 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Berlekamp:2000:DBG

[Ber00] Elwyn R. Berlekamp. The Dots-and-Boxes Game: Sophisticated Child’s Play. A. K. Peters, Ltd., Wellesley, MA, USA, 2000. ISBN 1-56881-129-2 (paperback). xii + 131 pp. LCCN QA269 .B39 2000.

Berstel:2004:BRB

[Ber04] Jean Berstel. Book review: Automatic Sequences: Theory, Applications, Generalizations, by Jean-Paul Allouche and Jeffrey Shallit. Cambridge University Press. ACM SIGACT News, 35(1):12– 16, March 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [AS03].

Borodin:2005:OCC

[BEY05] Allan Borodin and Ran El-Yaniv. Online Computation and Competitive Analysis. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-61946-7 (paperback). xviii + 414 pp. LCCN QA76.9.A43 B67 2005. URL http:

//www.loc.gov/catdir/enhancements/fy0632/2005280521-d.
html;   http://www.loc.gov/catdir/enhancements/fy0632/
2005280521-t.html.

Belton:2017:RSS

[BF17] Robin Belton and Brittany Terese Fasy. Review of The Structure and Stability of Persistence Modules by Frédéric Chazal, Vin de Silva, Marc Glisse, and Steve Oudot. ACM SIGACT News, 48 (2):7–11, June 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Borodin:1979:TST

[BFK^{+}79] Allan Borodin, Michael J. Fischer, David G. Kirkpatrick, Nancy A. Lynch, and Martin Tompa. A time-space tradeoff for sorting and related non-oblivious computations. ACM SIGACT News, 11(2): 24, Fall 1979. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Boyar:2016:OAA

[BFK^{+}16] Joan Boyar, Lene M. Favrholdt, Christian Kudahl, Kim S. Larsen, and Jesper W. Mikkelsen. Online algorithms with advice: a survey. ACM SIGACT News, 47(3):93–129, September 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bender:2015:RCA

[BFM^{+}15] Michael A. Bender, Jeremy T. Fineman, Mahnush Movahedi, Jared Saia, Varsha Dani, Seth Gilbert, Seth Pettie, and Maxwell Young. Resource-competitive algorithms. ACM SIGACT News, 46(3):57– 71, September 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Brandt:2019:SR

[BFU19] Sebastian Brandt, Manuela Fischer, and Jara Uitto. SIROCCO 2019 review. ACM SIGACT News, 50(4):46–47, December 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3374857.
3374867.

Borger:1997:CDP

[BGG97] Egon Börger, Erich Grädel, and Yuri Gurevich. The Classical Decision Problem. Perspectives in mathematical logic. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-57073-X (hardcover). xii + 492 pp. LCCN QA9.65 .B64 1997.

Bera:2007:SDQ

[BGH07] Debajyoti Bera, Frederic Green, and Steven Homer. Small depth quantum circuits. ACM SIGACT News, 38(2):35–50, June 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bhattacharyya:2013:GCT

[Bha13] Arnab Bhattacharyya. Guest column: On testing affine-invariant properties over finite fields. ACM SIGACT News, 44(4):53–72, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Beygelzimer:1999:OWF

[BHHR99] Alina Beygelzimer, Lane A. Hemaspaandra, Christopher M. Homan, and Jörg Rothe. One-way functions in worst-case cryptography: algebraic and security properties are on the house. ACM SIGACT News, 30(4):25–40, December 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bundy:2005:RML

[BHI05] Alan Bundy, David Basin Dieter Hutter, and Andrew Ireland. Rippling: meta-level guidance for mathematical reasoning, volume 56 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-83449-X (hardcover). xiv + 202 pp. LCCN QA76.9.A96 B84 2005.

Bentley:1980:GMS

[BHS80] Jon Louis Bentley, Dorothea Haken, and James B. Saxe. A general method for solving divide-and-conquer recurrences. ACM SIGACT News, 12(3):36–44, Fall 1980. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Brassard:1997:QCH

[BHT97] Gilles Brassard, Peter Høyer, and Alain Tapp. Quantum cryptanalysis of hash and claw-free functions. ACM SIGACT News, 28 (2):14–19, June 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Buss:2008:CFP

[BI08] Jonathan F. Buss and Tarique Mesbaul Islam. The complexity of fixed-parameter problems: guest column. ACM SIGACT News, 39(1):33–46, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bierbrauer:2005:ICT

[Bie05] Juergen Bierbrauer. Introduction to Coding Theory. Discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, 2005. ISBN 1-58488-421-5. xxiii + 390 pp.

LCCN QA268 .B48 2004. URL http://www.loc.gov/catdir/
enhancements/fy0646/2004049447-d.html.

Biggs:2008:CII

[Big08] Norman Biggs. Codes: An introduction to Information Communication and Cryptography. Springer undergraduate mathematics series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-84800-273-4 (e-book), 1-84800-272-6 (paperback). x + 273 pp. LCCN QA268 .B496 2008eb.

Birget:1991:IRL

[Bir91] Jean-Camille Birget. Intersection of Regular Languages and State Complexity. ACM SIGACT News, 22(2):49, Spring 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bloem:2016:DPV

[BJK^{+}16] Roderick Bloem, Swen Jacobs, Ayrat Khalimov, Igor Konnov, Sasha Rubin, Helmut Veith, and Josef Widder. Decidability in parameterized verification. ACM SIGACT News, 47(2):53–64, June 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bonakdarpour:2012:AMR

[BK12] Borzoo Bonakdarpour and Sandeep S. Kulkarni. Automated model repair for distributed programs. ACM SIGACT News, 43(2):85– 107, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Biradavolu:2018:DR

[BK18] Aditya Biradavolu and Saptaparni Kumar. DISC 2018 review. ACM SIGACT News, 49(4):89–95, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bezem:2003:TRS

[BKdV03] M. (Marc) Bezem, J. W. Klop, and Roel de Vrijer, editors. Term rewriting systems, volume 55 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-39115-6. xxii + 884 pp. LCCN QA267 .T43 2003. URL http://www.loc.

gov/catdir/description/cam022/2002067658.html;   http:
//www.loc.gov/catdir/samples/cam033/2002067658.html;
http://www.loc.gov/catdir/toc/cam031/2002067658.html.

Banyassady:2018:CGC

[BKM18] Bahareh Banyassady, Matias Korman, and Wolfgang Mulzer. Computational geometry column 67. ACM SIGACT News, 49 (2):77–94, June 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Blum:2006:MPE

[BL06] E. K. (Edward K.) Blum and Sergey V. Lototsky. Mathematics of Physics and Engineering. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2006. ISBN 981-256-621-X. xiv + 482 pp. LCCN QA401 .B58 2006. URL

http://www.loc.gov/catdir/toc/fy0707/2006284392.html.

Blatner:1997:JP

[Bla97] David Blatner. The Joy of π. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 1997. ISBN 0-8027-1332-7 (hardcover), 0-8027-7562-4 (paperback). xiii + 129 pp. LCCN

QA484 .B55 1997. URL http://www.walkerbooks.com/books/
catalog.php?key=4.

Blank:2006:JRB

[Bla06] Brian Blank. A joint review of A History of Pi, by Petr Beckmann, St. Martins’s Press, 1976, Barnes and Noble Books, 1989; The Joy of Pi, by David Blatner, Walker & Co., 1997; The Nothing That Is, by Robert Kaplan, Oxford University Press, 1999; e: The Story of a Number, by Eli Maor, Princeton University Press, 1998; An Imaginary Tale, by Paul Nahin, Princeton University Press, 1998; Zero: The Biography of a Dangerous Idea, by Charles Seife, Viking Press, 2000. ACM SIGACT News, 37(3):19–26, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Bla97, Kap99, Mao94, Nah98].

Blanchette:2012:BPC

[Bla12] Jean-François Blanchette. Burdens of Proof: Cryptographic Culture and Evidence Law in the Age of Electronic Documents. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01751-2 (hardcover). 276 pp. LCCN K2269.5 .B58 2012.

Blum:1983:CFT

[Blu83] Manuel Blum. Coin flipping by telephone — a protocol for solving impossible problems. ACM SIGACT News, 15(1):23–27, Winter– Spring 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Barwise:1996:VCM

[BM96] Jon Barwise and Lawrence Stuart Moss. Vicious Circles: On the Mathematics of Non-Wellfounded Phenomena, volume 60 of CSLI lecture notes. CSLI Publications, Stanford, CA, USA, 1996. ISBN 1-57586-009-0 , 1-57586-008-2 (paperback). x + 390 pp. LCCN QA248 .B376 1996. URL http://www.loc.gov/

catdir/description/cam027/96000809.html;   http://www.
loc.gov/catdir/toc/cam023/96000809.html.

Bradley:2007:CCD

[BM07] Aaron R. Bradley and Zohar Manna. The Calculus of Computation: Decision Procedures with Applications to Verification. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-74113-5. xv + 366 pp. LCCN QA76.9.L63 B73 2007.

Birnbaum:2008:LBM

[BM08] Benjamin E. Birnbaum and Claire Mathieu. On-line bipartite matching made simple. ACM SIGACT News, 39(1):80–87, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Broda:2020:GCA

[BMMR20] Sabine Broda, Antonio Machiavelo, Nelma Moreira, and Rogerio Reis. Guest column: Analytic combinatorics and descriptional complexity of regular languages on average. ACM SIGACT News, 51(1):38–56, March 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3388392.3388401.

Britton:1977:NCI

[BMO77] Dianne E. Britton, Ralph B. McLaughlin, and Richard J. Orgass. A note concerning intermittent assertions. ACM SIGACT News, 9(2):30–35, Summer 1977. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Brass:2005:RPD

[BMP05] Peter Brass, W. O. J. Moser, and János Pach. Research Problems in Discrete Geometry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-29929-7. xii + 499 pp. LCCN QA640.7 .B73 2005. URL http://www.loc.

gov/catdir/toc/fy0608/2005924022.html.

Baader:1998:TRA

[BN98] Franz Baader and Tobias Nipkow. Term Rewriting and All That. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-45520-0 (hardcover). xii + 301 pp. LCCN QA267 .B314

1998. URL http://www.loc.gov/catdir/description/cam028/
97028286.html;   http://www.loc.gov/catdir/toc/cam027/
97028286.html.

Brabazon:2006:BIA

[BO06] Anthony Brabazon and Michael O’Neill. Biologically Inspired Algorithms for Financial Modelling. Natural computing series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-26252-0 (hardcover). xv + 275 pp. LCCN HG106 .B73 2006. URL http://www.loc.gov/

catdir/enhancements/fy0663/2005936099-d.html;   http:/
/www.loc.gov/catdir/toc/fy0608/2005936099.html.

Bollobas:1985:RG

[Bol85] Béla Bollobás. Random Graphs. Academic Press, New York, USA, 1985. ISBN 0-12-111755-3 , 0-12-111756-1 (paperback). xvi + 447 pp. LCCN QA166.17 .B66 1985.

Bollobas:1998:MGT

[Bol98] Béla Bollobás. Modern Graph Theory, volume 184 of Graduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98491-7 , 0-387-98488-7 (paperback). xiii + 394 pp.

LCCN QA166 .B663 1998. URL http://www.loc.gov/catdir/
enhancements/fy0816/98011960-d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/98011960-t.html.

Bollobas:2001:RG

[Bol01] Béla Bollobás. Random Graphs, volume 73 of Cambridge studies in advanced mathematics. Cambridge University Press, Cambridge, UK, second edition, 2001. ISBN 0-521-80920-7, 0-521-79722-5 (paperback). xviii + 498 pp. LCCN QA166.17 .B66 2001. URL

http://www.loc.gov/catdir/description/cam021/00068952.
html;   http://www.loc.gov/catdir/enhancements/fy0731/
00068952-b.html;   http://www.loc.gov/catdir/toc/cam027/
00068952.html.

Bona:2008:CP

[Bón08a] Miklós Bóna. Combinatorics of permutations. ACM SIGACT News, 39(4):21–25, December 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bonato:2008:CWG

[Bon08b] Anthony Bonato. A Course on the Web Graph, volume 89 of Graduate studies in mathematics. American Mathematical Society, Providence, RI, USA, 2008. ISBN 0-8218-4467-9. xi + 184 pp. LCCN TK5105.888 .B667 2008.

Bona:2009:BRBb

[Bón09a] Miklós Bóna. Book review: Algorithmic Combinatorics on Partial Words, by Francine Blanchet-Sadri. ACM SIGACT News, 40 (3):39–41, September 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BS08].

Bona:2009:BRBa

[Bón09b] Miklós Bóna. Book review: Combinatorics and Graph Theory: (second edition) by John M. Harris, Jeffry L. Hirst and Michael J. Mossinghoff. ACM SIGACT News, 40(3):37–39, September 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [HHM00, HHM08].

Bona:2009:BRBc

[Bón09c] Miklós Bóna. Book review: Random Graphs: (second edition) by Béla Bollobás. ACM SIGACT News, 40(3):46–48, September 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Bol85].

Bona:2010:BRBb

[Bón10a] Miklós Bóna. Book review: A Combinatorial Approach to Matrix Theory and Its Applications, by Richard Brualdi and Dragoš Cvetković, Published by Cambridge Press, 2009, 824 pages, hardcover. ACM SIGACT News, 41(2):19–22, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BC09].

Bona:2010:BRBa

[Bón10b] Miklós Bóna. Book review: Analytic Combinatorics, by Philippe Flajolet and Robert Sedgewick, published by Cambridge Press, 2009, 824 pages, hardcover. ACM SIGACT News, 41(2):11–14, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [FS09].

Bonnet:2010:RD

[Bon10c] François Bonnet. Review of DISC 2010. ACM SIGACT News, 41 (4):106–108, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bona:2011:BRBa

[Bón11a] Miklós Bóna. Book review: Applied Combinatorics, by Fred S. Roberts and Barry Tesman. ACM SIGACT News, 42(3):32–34, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Rob84, RT05].

Bona:2011:BRBb

[Bón11b] Miklós Bóna. Book review: Probability: Theory and Examples, 4th edition, by Rick Durrett. ACM SIGACT News, 42(3):42–45, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Dur10].

Bona:2011:BRBc

[Bón11c] Miklós Bóna. Book review: Proofs from THE BOOK (4th edition) by Martin Aigner and Günter M. Ziegler. ACM SIGACT News, 42(4):32–37, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [AZ98, AZ99, AZ01, AZ04, AZ10].

Bona:2012:BRB

[Bón12] Miklós Bóna. Book review: The Life and Times of the Central Limit Theorem, by William J. Adams. ACM SIGACT News, 43(1): 32–33, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Ada74, Ada09b].

Bona:2013:RBCa

[Bón13a] Miklós Bóna. Review of Bijective Combinatorics by Nicholas Loehr. ACM SIGACT News, 44(1):12–14, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bona:2013:RBCb

[Bón13b] Miklós Bóna. Review of Extremal Combinatorics with Applications to Computer Science (2nd edition) by Stasys Jukna. ACM SIGACT News, 44(3):24–27, September 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bona:2014:RAC

[Bón14] Miklós Bóna. Review of Analytic combinatorics in several variables by Robin Pemantle and Mark Wilson. ACM SIGACT News, 45 (3):32–33, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Booth:1977:BMM

[Boo77] Kellogg S. Booth. Boolean matrix multiplication using only O(n^{log_{2} 7} log n) bit operations. ACM SIGACT News, 9(3):23, Fall 1977. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Book:1984:RCCa

[Boo84a] Ronald V. Book. Relativizations of complexity classes. ACM SIGACT News, 15(4):61, Winter–Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Book:1984:RCCb

[Boo84b] Ronald V. Book. Relativizations of complexity classes. ACM SIGACT News, 16(1):61, Winter–Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Boothe:2010:BRB

[Boo10] Peter Boothe. Book review: A Course in Enumeration, by Martin Aigner, Springer, 2007 555 pages, hardcover. ACM SIGACT News, 41(2):17–19, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Aig07].

Borchers:2007:BRB

[Bor07a] Brian Borchers. Book review: Computational Techniques of the Simplex Method, by István Maros, Kluwer Academic Publishers, 2003. ACM SIGACT News, 38(3):27–30, September 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Mar03b].

Bortnikov:2007:RD

[Bor07b] Edward Bortnikov. Review of DISC ’07. ACM SIGACT News, 38 (4):49–53, December 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Borcher:2008:BRB

[Bor08] Brian Borcher. Book review: Combinatorial Optimization: Packing and Covering, by Gérard Cornuéjols. ACM SIGACT News, 39 (3):16–18, September 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Cor01].

Borchers:2009:BRB

[Bor09a] Brian Borchers. Book review: Practical Optimization: Algorithms and Engineering Applications, by Andreas Antoniou and Wu-Sheng Lu (Springer Verlag, 2007). ACM SIGACT News, 40(1): 20–22, March 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [AL07].

Bortnikov:2009:OSG

[Bor09b] Edward Bortnikov. Open-source grid technologies for web-scale computing. ACM SIGACT News, 40(2):87–93, June 2009. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Berman:2005:ASP

[BP05] Kenneth A. Berman and Jerome L. Paul. Algorithms: sequential, parallel, and distributed. Thomson, Boston, MA, USA, 2005. ISBN 0-534-42057-5. xxix + 962 pp. LCCN QA76.9.A43 B47 2005.

Beissinger:2006:CUM

[BP06] Janet Beissinger and Vera Pless. The Cryptoclub: Using Mathematics to Make and Break Secret Codes. A. K. Peters, Ltd., Wellesley, MA, USA, 2006. ISBN 1-56881-223-X. xvi + 199 pp.

LCCN QA40.5 .B45 2006. URL http://www.loc.gov/catdir/
toc/ecip067/2006002743.html.

Benjamin:2003:PRC

[BQ03] Arthur T. Benjamin and Jennifer J. Quinn. Proofs That Really Count: the Art of Combinatorial Proof, volume 27 of Dolciani mathematical expositions. Mathematical Association of America, Washington, DC, USA, 2003. ISBN 0-88385-333-7. xiv + 194 pp. LCCN QA164.8 B45 2003.

Blum:1998:TCS

[BR98] Avrim Blum and Prabhakar Raghavan. On a theory of computing symposia. ACM SIGACT News, 29(3):104–111, September 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Berlekamp:1999:MPP

[BR99] Elwyn R. Berlekamp and Tom Rodgers, editors. The Mathemagician and Pied Puzzler: a Collection in Tribute to Martin Gardner. A. K. Peters, Ltd., Wellesley, MA, USA, 1999. ISBN 1-56881-075-X. x + 266 pp. LCCN QA95 .M366 1999.

Brady:1977:TCS

[Bra77] J. M. Brady. The Theory of Computer Science: a Programming Approach. Chapman and Hall, Ltd., London, UK, 1977. ISBN 0-470-99103-8. xiii + 287 pp. LCCN QA76 .B697.

Brassard:1983:OSR

[Bra83] Giles Brassard. An optimally secure relativized cryptosystem. ACM SIGACT News, 15(1):28–33, Winter–Spring 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Brassard:1985:CBN

[Bra85] Gillea Brassard. Crusade for a better notation. ACM SIGACT News, 17(1):60–64, June 1985. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Knu76a, VM85].

Brassard:1989:CCb

[Bra89a] G. Brassard. Cryptology — column 2. ACM SIGACT News, 20(4): 13, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Brassard:1989:CCa

[Bra89b] Gilles Brassard. Cryptology column. ACM SIGACT News, 20(3): 15–19, July 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Brasard:1990:CCH

[Bra90a] Gilles Brasard. Cryptology — column 3 hot news on interactive protocols. ACM SIGACT News, 21(1):7, January 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Brassard:1990:CCH

[Bra90b] Gilles Brassard. Cryptology — column 4: hiding information from oracles. ACM SIGACT News, 21(2):5, Spring 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Brassard:1991:HCY

[Bra91] Gilles Brassard. How convincing is your protocol? ACM SIGACT News, 22(1):5–12, Winter 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bradfield:1992:VTP

[Bra92] Julian Charles Bradfield. Verifying Temporal Properties of Systems. Progress in theoretical computer science. Birkhäuser Verlag, Basel, Switzerland, 1992. ISBN 0-8176-3625-0 (hard), 3-7643-3625-0. vi + 113 pp. LCCN QA76.9.A96 B73 1992.

Brassard:1994:CCQ

[Bra94a] Gilles Brassard. Cryptology column – quantum computing: The end of classical cryptography? ACM SIGACT News, 25(4):15– 21, December 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Brassard:1994:QCE

[Bra94b] Gilles Brassard. Quantum computing: the end of classical cryptography? ACM SIGACT News, 25(4):15–21, December 1994. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Brassard:1995:CCB

[Bra95] Gilles Brassard. Cryptology Column: The Book I’ve Always Wanted To Write (almost). ACM SIGACT News, 26(2):18–20, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Brass:2008:ADS

[Bra08] Peter Brass. Advanced Data Structures. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-88037-8 (hardcover). xvi + 456 pp. LCCN QA76.9.A43 B73 2008. URL http://www.

loc.gov/catdir/toc/ecip0818/2008021408.html.

Brule:2013:RBA

[Bru13] Joshua Brulé. Review of Probably Approximately Correct by Leslie Valiant. ACM SIGACT News, 44(3):38–39, September 2013. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bentley:1979:AVS

[BS79] Jon Louis Bentley and James B. Saxe. Algorithms on vector sets. ACM SIGACT News, 11(2):36–39, Fall 1979. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Broder:1984:PAS

[BS84] Andrei Broder and Jorge Stolfi. Pessimal algorithms and simplexity analysis. ACM SIGACT News, 16(3):49–53, Fall 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bhatia:1998:BRB

[BS98] Randeep Bhatia and Yoram J. Sussmann. Book review: Approximation Algorithms for NP-hard Problems. Edited by Dorit S. Hochbaum (PWS, 1997). ACM SIGACT News, 29(2):17–20, June 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Hoc97b].

Blanchet-Sadri:2008:ACP

[BS08] Francine Blanchet-Sadri. Algorithmic Combinatorics on Partial Words. Discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, 2008. ISBN 1-4200-6092-9 (hardcover). 385 pp. LCCN QA76.9.A43

B53 2008. URL http://www.loc.gov/catdir/enhancements/
fy0808/2007031486-d.html;   http://www.loc.gov/catdir/
toc/ecip0724/2007031486.html.

Baldoni:2009:TAD

[BS09] Roberto Baldoni and Alexander A. Shvartsman. Theoretical aspects of dynamic distributed systems: report on the workshop, Elche, Spain, September 26, 2009. ACM SIGACT News, 40(4):87– 89, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Braud-Santoni:2013:PR

[BS13] Nicolas Braud-Santoni. PODC 2013 review. ACM SIGACT News, 44(4):83–86, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Basin:2011:AIS

[BSS11] David Basin, Patrick Schaller, and Michael Schläpfer. Applied information security: a hands-on approach. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-24473-4 (hardcover). xiv + 202 pp. LCCN QA76.9.A25 B37 2011.

Broadbent:2008:CQM

[BT08] Anne Broadbent and Alain Tapp. Can quantum mechanics help distributed computing? ACM SIGACT News, 39(3):67–76, September 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bun:2020:GCA

[BT20] Mark Bun and Justin Thaler. Guest column: Approximate degree in classical and quantum computing. ACM SIGACT News, 51 (4):48–72, December 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3444815.3444825.

Buchmann:2001:IC

[Buc01] Johannes Buchmann. Introduction to Cryptography. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-95034-6 (hardcover). xi + 281 pp. LCCN QA268 .B83 2001.

Buchmann:2004:IC

[Buc04] Johannes Buchmann. Introduction to Cryptography. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2004. ISBN 0-387-20756-2. xvi + 335 pp. LCCN QA268 .B83 2004. URL http:

//www.loc.gov/catdir/enhancements/fy0818/2004041657-b.
html;   http://www.loc.gov/catdir/enhancements/fy0818/
2004041657-d.html;   http://www.loc.gov/catdir/enhancements/
fy0818/2004041657-t.html.

Buhrman:1996:SNS

[Buh96] Harry Buhrman. A short note on Shor’s factoring algorithm. ACM SIGACT News, 27(1):89–90, March 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Burstall:1972:ADP

[Bur72] Rod M. Burstall. An algebraic description of programs with assertions, verification and simulation. ACM SIGACT News, 4(14): 7–14, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Burns:1978:MEL

[Bur78] James E. Burns. Mutual exclusion with linear waiting using binary shared variables. ACM SIGACT News, 10(2):42–47, Summer 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Burkholder:1987:HP

[Bur87] L. Burkholder. The halting problem. ACM SIGACT News, 18(3): 48–60, April 1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Bury:2015:SR

[Bur15] Marc Bury. SIROCCO 2015 review. ACM SIGACT News, 46 (4):88–93, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Buttelmann:1971:SSS

[But71] H. W. Buttelmann. Syntax-semantics systems as structure manipulation systems: phrase structure grammars and generalized finite automata. ACM SIGACT News, 3(9):10, April 1971. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Baase:2000:CAI

[BV00] Sara Baase and Allen Van Gelder. Computer Algorithms: Introduction to Design and Analysis. Addison-Wesley, Reading, MA, USA, third edition, 2000. ISBN 0-201-61244-5. xix + 688 pp. LCCN QA76.9.A43 B33 2000.

Buchmann:2007:BQF

[BV07] Johannes Buchmann and Ulrich Vollmer. Binary Quadratic Forms: an Algorithmic Approach, volume 20 of Algorithms and computation in mathematics, 1431-1550. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-46367-4 (hardcover). xiv + 318 pp. LCCN

QA243 .B9 2007. URL http://site.ebrary.com/lib/stanford/
Doc?id=10187632;   http://www.myilibrary.com?id=94925;
 http://www.springerlink.com/openurl.asp?genre=book\%26isbn=
978-3-540-46367-2.

Blackburn:2007:HML

[BvBW07] Patrick Blackburn, J. F. A. K. van Benthem, and Frank Wolter, editors. Handbook of Modal Logic, volume 3 of Studies in logic and practical reasoning. Elsevier, Amsterdam, The Netherlands, 2007. ISBN 0-444-51690-5. ???? pp. LCCN BC199.M6

H36 2007. URL http://www.loc.gov/catdir/enhancements/
fy0665/2006050759-d.html.

Baeza-Yates:1995:TA

[BY95] Ricardo A. Baeza-Yates. Teaching algorithms. ACM SIGACT News, 26(4):51–59, December 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Braverman:2009:CJS

[BY09] Mark Braverman and Michael Yampolsky. Computability of Julia Sets, volume 23 of Algorithms and computation in mathematics, 1431-1550. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-68546-4 (hardcover), 3-540-68547-2 (e-book). xiii + 151 pp. LCCN QA614.86 .B73 2009.

Brent:2011:MCA

[BZ11] Richard P. Brent and Paul Zimmermann. Modern Computer Arithmetic, volume 18 of Cambridge monographs on applied and computational mathematics. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-19469-5 (hardcover). xvi + 221 pp. LCCN

QA76.9.C62 BRE 2011. URL http://www.loria.fr/~zimmerma/
mca/pub226.html.

Cadilhac:2011:BRB

[Cad11] Michaël Cadilhac. Book review: Combinatorics: a Guided Tour, by David R. Mazur. ACM SIGACT News, 42(3):34–36, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Maz10].

Cadilhac:2012:RBW

[Cad12] Michaël Cadilhac. Review of Handbook of Weighted Automata, edited by Manfred Droste, Werner Kuich and Heiko Vogler. ACM SIGACT News, 43(3):32–37, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cadilhac:2013:RPA

[Cad13] Michaël Cadilhac. Review of Proofs and algorithms by Gilles Dowek (translation by Maribel Fernandez). ACM SIGACT News, 44(4):35–37, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cadilhac:2014:RGS

[Cad14] Michaël Cadilhac. Review of Graph structure and monadic secondorder logic: a language-theoretic approach by Bruno Courcelle and Joost Engelfriet. ACM SIGACT News, 45(3):24–25, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cadilhac:2016:RGR

[Cad16] Michaël Cadilhac. Review of: The Golden Ratio and Fibonacci Numbers by Richard A. Dunlap. ACM SIGACT News, 47(4):15– 17, December 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cadilhac:2021:RCC

[Cad21] Michael Cadilhac. Review of Communication Complexity and Applications by Anup Rao and Amir Yehudayoff. ACM SIGACT News, 52(3):11–13, September 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.

org/doi/10.1145/3494656.3494660.

Calude:1983:CIP

[Cal83] Cristian Calude. On a class of independent problems related to Rice theorem. ACM SIGACT News, 15(2):53–57, Spring–Summer 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Calvert:2010:BRB

[Cal10] Wesley Calvert. Book review: Computability of Julia Sets, by Mark Braverman and Michael Yampolsky, Publisher: Springer, 2009. ACM SIGACT News, 41(1):51–53, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BY09].

Calvert:2012:RBA

[Cal12] Wesley Calvert. Review of Process Algebra: Equational Theories of Communicating Processes by J. C. M. Baeten, T. Basten, and M. A. Reniers. ACM SIGACT News, 43(2):36–38, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Calvert:2013:RBS

[Cal13] Wesley Calvert. Review of Universal Semantic Communication by B. Juba. ACM SIGACT News, 44(3):12–15, September 2013. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Calvanese:2015:RP

[Cal15] Diego Calvanese. Report on PODS 2015. ACM SIGACT News, 46(4):79–81, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Canetti:2006:SCC

[Can06] Ran Canetti. Security and composition of cryptographic protocols: a tutorial (part I). ACM SIGACT News, 37(3):67–92, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cao:2013:RBN

[Cao13] Yixin Cao. Review of Flows in Networks by L. R. Ford, Jr. and D. R. Fulkerson. ACM SIGACT News, 44(2):28–30, June 2013. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Carter:1977:FG

[Car77] Larry Carter. A four-gadget. ACM SIGACT News, 9(2):36, Summer 1977. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Carpenter:1998:TLS

[Car98] Bob Carpenter. Type-logical semantics. Language, speech, and communication. MIT Press, Cambridge, MA, USA, 1998. ISBN 0-262-03248-1 (hardcover), 0-262-53149-6 (paperback). xxi + 575 pp. LCCN P325 .C324 1997. URL http://cognet.mit.edu/

library/books/view?isbn=0262531496.

Cardinal:2015:CGC

[Car15] Jean Cardinal. Computational geometry column 62. ACM SIGACT News, 46(4):69–78, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Casanova:2002:DCR

[Cas02] Henri Casanova. Distributed computing research issues in grid computing. ACM SIGACT News, 33(3):50–70, September 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Castaneda:2008:RP

[Cas08] Armando Castañeda. A review of PODC 2008. ACM SIGACT News, 39(4):55–59, December 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cassuto:2013:WCC

[Cas13] Yuval Cassuto. What can coding theory do for storage systems? ACM SIGACT News, 44(1):80–88, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Charron-Bost:2007:HDF

[CBS07] Bernadette Charron-Bost and André Schiper. Harmful dogmas in fault tolerant distributed computing. ACM SIGACT News, 38(1): 53–61, March 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Clark:1976:PMC

[CC76] K. L. (Keith L.) Clark and Don Cowell. Programs, Machines, and Computation: an Introduction to the Theory of Computing. Mc-Graw-Hill, New York, NY, USA, 1976. ISBN 0-07-084067-9. xi + 176 pp. LCCN QA267 .C54.

Colbourn:1978:GIS

[CC78] Marlene Jones Colbourn and Charles J. Colbourn. Graph isomorphism and self-complementary graphs. ACM SIGACT News, 10(1):25–29, Spring 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Calude:2015:GCA

[CCD15] Cristian S. Calude, Elena Calude, and Michael J. Dinneen. Guest column: Adiabatic quantum computing challenges. ACM SIGACT News, 46(1):40–61, March 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chung:1998:EGH

[CEG98] Fan R. K. Chung, Paul Erdős, and Ronald L. Graham. Erdős on Graphs: His Legacy of Unsolved Problems. A. K. Peters, Ltd., Wellesley, MA, USA, 1998. ISBN 1-56881-079-2. xiii + 142 pp. LCCN QA166 .C485 1998.

Clarke:2010:QAT

[CEG^{+}10] Dave Clarke, David Eppstein, Kaveh Ghasemloo, Lev Reyzin, András Salamon, Peter Shor, Aaron Sterling, and Suresh Venkatasubramanian. Questions answered. in theory: http://cstheory.stackexchange.com. ACM SIGACT News, 41(4):58–60, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cenek:2000:BRB

[Cen00] E. W. Cenek. Book review: Modern Graph Theory, by Béla Bollobás. ACM SIGACT News, 31(4):15–18, December 2000. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Bol98].

Cenek:2002:CCT

[Cen02] E. W. Cenek. Computability and complexity theory and the complexity theory companion. ACM SIGACT News, 33(3):17– 19, September 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cenek:2003:BRB

[Cen03] E. W. Cenek. Book review: Theory of Computational Complexity, by Ding-Zhu Du and Ker-I Ko. John Wiley & Sons. ACM SIGACT News, 34(3):11–12, September 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [DK00].

Cenek:2009:CEP

[Cen09] Eowyn Cenek. Chases and escapes by Paul J. Nahin. ACM SIGACT News, 40(3):48–50, September 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cenek:2011:BRB

[Cen11] Eowyn Cenek. Book review: Roots to Research, by Judith D. Sally and Paul J. Sally, Jr. ACM SIGACT News, 42(3):20–22, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [SS07].

Cenek:2014:RSS

[Cen14] Eowyn Cenek. Review of Six Sources of Collapse by Charles R. Hadlock. ACM SIGACT News, 45(4):38–40, December 2014. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cenek:2015:RKI

[Cen15] Eowyn Cenek. Review of: The King of Infinite Space: Euclid and his Elements by David Berlinski. ACM SIGACT News, 46(1):27, March 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cohen:2006:HEH

[CFA^{+}06] Henri Cohen, Gerhard Frey, Roberto Avanzi, et al., editors. Handbook of Elliptic and Hyperelliptic Curve Cryptography. Discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, 2006. ISBN 1-58488-518-1. xxxiv + 808 pp. LCCN QA567.2.E44 H36 2006. URL http://www.loc.gov/

catdir/enhancements/fy0647/2005041841-d.html.

Condon:1997:SDR

[CFF^{+}97] Anne Condon, Faith Fich, Greg N. Frederickson, Andrew V. Goldberg, David S. Johnson, Michael C. Loui, Steven Mahaney, Prabhakar Raghavan, John E. Savage, Alan L. Selman, and David B. Shmoys. Strategic directions in research in theory of computing. ACM SIGACT News, 28(3):75–93, September 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chaudhuri:1993:SST

[CH93] R. Chaudhuri and H. Höft. Splaying a search tree in preorder takes linear time. ACM SIGACT News, 24(2):88–93, Spring 1993. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Crary:2006:HOA

[CH06] Karl Crary and Robert Harper. Higher-order abstract syntax: setting the record straight. ACM SIGACT News, 37(3):93–96, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Crama:2010:BMM

[CH10] Yves Crama and P. L. Hammer, editors. Boolean Models and Methods in Mathematics, Computer Science, and Engineering, volume 134 of Encyclopedia of mathematics and its applications. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-84752-4. xviii + 759 pp. LCCN QA10.3 .B658 2010.

URL http://assets.cambridge.org/97805218/47520/cover/
9780521847520.jpg;   http://www.loc.gov/catdir/enhancements/
fy1009/2010017816-b.html;   http://www.loc.gov/catdir/
enhancements/fy1009/2010017816-d.html;   http://www.loc.
gov/catdir/enhancements/fy1009/2010017816-t.html.

Capdevielle:2014:WFW

[CH14] Claire Capdevielle and Sandeep Hans. WTTM 2013, the Fifth Workshop on the Theory of Transactional Memory. ACM SIGACT News, 45(3):75–81, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Censor-Hillel:2015:DAC

[CH15] Keren Censor-Hillel. Distributed algorithms as combinatorial structures. ACM SIGACT News, 46(1):63–76, March 2015. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chaitin:1970:MDL

[Cha70] G. J. Chaitin. To a mathematical definition of “life”. ACM SIGACT News, 2(4):12–18, January 1970. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chaitin:1971:CCG

[Cha71] G. J. Chaitin. Computational complexity and Gödel’s incompleteness theorem. ACM SIGACT News, 3(9):11–12, April 1971. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chaitin:1973:SAC

[Cha73a] G. J. Chaitin. Some abstracts from The Computer Science Conference. ACM SIGACT News, 5(2):24–25, April 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chaitin:1973:SPI

[Cha73b] G. J. Chaitin. Some philosophical implications of informationtheoretic computational complexity. ACM SIGACT News, 5(2):

21–23, April 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chaitin:1998:LMC

[Cha98] Gregory J. Chaitin. The Limits of Mathematics: a Course on Information Theory and Limits of Formal Reasoning. Springer series in discrete mathematics and theoretical computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 981-3083-59-X. ix + 148 pp. LCCN

QA76.9.M35 C48 1998. URL http://www.loc.gov/catdir/
enhancements/fy0902/97036103-d.html;   http://www.loc.
gov/catdir/enhancements/fy0902/97036103-t.html.

Chapin:2005:BRB

[Cha05] E. Jonathan Chapin. Book review: Integer Programming, by Laurence A. Wolsey, Wiley and Sons 1998. ACM SIGACT News, 36 (3):10–12, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Wol98].

Charalambides:2008:EC

[Cha08] Charalambos A. Charalambides. Enumerative combinatorics. ACM SIGACT News, 39(4):25–27, December 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chakrabarti:2009:RBH

[Cha09] Kushal Chakrabarti. Book review: The Handbook of Bioinspired Algorithms and Applications, edited by Stephan Olariu and Albert Y. Zomaya, Chapman & Hall / CRC, 679 pages, $139.95. ACM SIGACT News, 40(4):23–35, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [OZ06].

Chakrabarty:2012:RBA

[Cha12] Deeparnab Chakrabarty. Review of Design of Approximation Algorithms, by David P. Williamson and David B. Shmoys. ACM SIGACT News, 43(4):52–54, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chattopadhyay:2020:GCR

[Cha20] Eshan Chattopadhyay. Guest column: a recipe for constructing two-source extractors. ACM SIGACT News, 51(2):38–57, June 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3406678.
3406688.

Cheriton:1974:ELM

[Che74] David R. Cheriton. An extension to on-line multiplication lower bound results. ACM SIGACT News, 6(4):24–31, October 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cherniavsky:1977:BRD

[Che77] John C. Cherniavsky. Book review department. ACM SIGACT News, 9(1):15, January 1977. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cherniavsky:1979:BRBb

[Che79a] John C. Cherniavsky. Book review: Automated Theorem Proving: a Logical Basis, by D. W. Loveland. North-Holland Publishing Co. 1977. ACM SIGACT News, 11(1):18, Summer 1979. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Lov78].

Cherniavsky:1979:BRBa

[Che79b] John C. Cherniavsky. Book review: The Theory of Computer Science: a Programming Approach, by J. M. Brady. Chapman and Hall. ACM SIGACT News, 11(1):17, Summer 1979. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Bra77].

Cherniavsky:1979:BRBc

[Che79c] John C. Cherniavsky. Book review: TORIX: a Programming System for Operations on Vectors and Matrices over Arbitrary Fields and of Variable Size, vol 1, by S. G. Van Der Meulen and M. Veldhorst. Mathematisch Centrum 1978. ACM SIGACT News, 11(2):6, Fall 1979. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [VV78].

Cherniavsky:1980:BRB

[Che80] John Cherniavsky. Book reviews: Checking Landau’s “Grundlagen” in the Automath system, by L. S. Van Benthem Jutting. Mathematical Centre 1979: First Order Dynamic Logic, by David Harel. Springer-Verlag 1979. And A Programming Logic, by Robert L. Constable and Michael J. O’Donnell. Winthrop Publishers 1978. ACM SIGACT News, 12(3):14–16, Fall 1980. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [COJ78, van79, Har79].

Cherniavsky:1982:BRB

[Che82] John C. Cherniavsky. Book review: Unsolvable Classes of Quantificational Formulas, by Harry R. Lewis. Addison-Wesley 1979. and The Decision Problem: Solvable Classes of Quantificational Formulas, by Burton Dreben and Warren D. Goldfarb. Addison-Wesley 1979. ACM SIGACT News, 14(4):70–71, Fall–Winter 1982. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Lew79].

Cherniavsky:1985:NN

[Che85] John C. Cherniavsky. NSF news. ACM SIGACT News, 16(4): 46–47, April 1985. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chen:2006:RLC

[Che06] Hubie Chen. A rendezvous of logic, complexity, and algebra. ACM SIGACT News, 37(4):85–114, December 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chekuri:2007:RND

[Che07] Chandra Chekuri. Routing and network design with robustness to changing or uncertain traffic demands. ACM SIGACT News, 38 (3):106–129, September 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chen:2010:BRB

[Che10] David Chen. Book review: Elliptic Curves, by Lawrence C. Washington 2008, Chapman and Hall. ACM SIGACT News, 41(4):36– 38, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Was08b].

Chen:2011:EF

[Che11a] Li Chen. Education forum. ACM SIGACT News, 42(2):77–81, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chen:2011:GCC

[Che11b] X. Chen. Guest column: Complexity dichotomies of counting problems. ACM SIGACT News, 42(4):54–76, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chen:2013:EFD

[Che13] Li Chen. Education forum: digital geometry and its algorithms: an introduction. ACM SIGACT News, 44(2):119–124, June 2013. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chen:2020:HWG

[Che20] Li Chen. How to write a good project report. ACM SIGACT News, 51(1):118–119, March 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3388392.3388405.

Chen:2021:IVR

[Che21] Li Chen. Iteration vs. recursion: Two basic algorithm design methodologies. ACM SIGACT News, 52(1):81–86, March 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3457588.
3457601.

Chen:2024:BIQ

[Che24] Li Chen. Brief introduction to quantum computing for undergraduate students: Lecture one. ACM SIGACT News, 55(1): 52–65, March 2024. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3654780.3654786.

Chibnik:1986:AES

[Chi86] Mara Chibnik. Algorithmic elimination of spurious nondeterminism from Mealy machines. ACM SIGACT News, 17(3):30–34, Winter 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chitnis:2015:RFP

[Chi15] Rajesh Chitnis. Review of: Fundamentals of Parameterized Complexity by Rodney G. Downey and Michael R. Fellows. ACM SIGACT News, 46(1):23–26, March 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Condon:2009:AB

[CHK^{+}09] Anne Condon, David Harel, Joost N. Kok, Arto Salomaa, and Erik Winfree, editors. Algorithmic Bioprocesses. Natural Computing Series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-88868-3, 3-540-88869-1 (ebook). xx + 742 pp. LCCN QA76.9. URL http://public.eblib.

com/EBLPublic/PublicView.do?ptiID=478040.

Crochemore:2007:AS

[CHL07] Maxime Crochemore, Christophe Hancart, and Thierry Lecroq. Algorithms on Strings. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-84899-7 (hardcover). viii + 383 pp. LCCN QA76.9.A43 C757 2007. URL http:

//www.loc.gov/catdir/enhancements/fy0703/2006039263-d.
html;   http://www.loc.gov/catdir/enhancements/fy0703/
2006039263-t.html;   http://www.loc.gov/catdir/enhancements/
fy0803/2006039263-b.html.

Choo:2012:RBT

[Cho12] Kim-Kwang Raymond Choo. Review of Insider Threats in Cyber Security by Probst, Hunker, Gollman, Bishop. ACM SIGACT News, 43(2):38–40, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2004:ASN

[Chr04a] Marek Chrobak. ACM SIGACT news distributed computing column 15. ACM SIGACT News, 35(3):46–57, September 2004. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2004:CGC

[Chr04b] Marek Chrobak. Computational geometry column 45. ACM SIGACT News, 35(2):71–73, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2004:DAS

[Chr04c] Marek Chrobak. Distributed approximation: a survey. ACM SIGACT News, 35(4):40–57, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2004:PSF

[Chr04d] Marek Chrobak. A princess swimming in the fog looking for a monster cow. ACM SIGACT News, 35(2):74–78, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2004:SNOa

[Chr04e] Marek Chrobak. SIGACT news online algorithms column 2. ACM SIGACT News, 35(1):38–48, March 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2004:SNOb

[Chr04f] Marek Chrobak. SIGACT news online algorithms column 4. ACM SIGACT News, 35(3):58–66, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2005:AFI

[Chr05a] Marek Chrobak. Algorithmic foundations of the Internet. ACM SIGACT News, 36(2):45–62, June 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2005:SIF

[Chr05b] Marek Chrobak. A short introduction to failure detectors for asynchronous distributed systems. ACM SIGACT News, 36(1):53–70, March 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2005:SNL

[Chr05c] Marek Chrobak. SIGACT news logic column 13. ACM SIGACT News, 36(3):51–66, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2005:SNO

[Chr05d] Marek Chrobak. SIGACT news online algorithms column 8. ACM SIGACT News, 36(3):67, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2006:OP

[Chr06a] Marek Chrobak. 2005: an offline persepctive. ACM SIGACT News, 37(1):82–98, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2006:TTW

[Chr06b] Marek Chrobak. Travelling through wormholes: a new look at distributed systems models. ACM SIGACT News, 37(1):66–81, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2007:CPD

[Chr07] Marek Chrobak. Competitiveness via primal-dual. ACM SIGACT News, 38(3):100–105, September 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2008:DTC

[Chr08a] Marek Chrobak. Database theory column. ACM SIGACT News, 39(1):77–79, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2008:SNO

[Chr08b] Marek Chrobak. SIGACT news online algorithms column 13: 2007 — an offine perspective. ACM SIGACT News, 39(3):96– 121, September 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2009:ISN

[Chr09a] Marek Chrobak. Introduction to the SIGACT news online algorithms column. ACM SIGACT News, 40(4):98, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2009:SNO

[Chr09b] Marek Chrobak. SIGACT news online algorithms column 14. ACM SIGACT News, 40(3):86–98, September 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2010:SNOa

[Chr10a] Marek Chrobak. SIGACT news online algorithms column 16. ACM SIGACT News, 41(1):99, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2010:SNOb

[Chr10b] Marek Chrobak. SIGACT news online algorithms column 17. ACM SIGACT News, 41(4):114–121, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2011:SNO

[Chr11] Marek Chrobak. SIGACT news online algorithms column 19. ACM SIGACT News, 42(2):82, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2014:OAP

[Chr14] Marek Chrobak. Online aggregation problems. ACM SIGACT News, 45(1):91–102, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chung:1997:SGT

[Chu97] Fan R. K. Chung, editor. Spectral Graph Theory, volume 92 of Regional conference series in mathematics, 0160-7642. Published for the Conference Board of the Mathematical Sciences by the American Mathematical Society, Providence, RI, USA, 1997. ISBN 0-8218-0315-8 (softcover). LCCN QA1 .R33 no. 92.

Ciesinger:1988:NN

[Cie88] Joachim Ciesinger. Neural nets. ACM SIGACT News, 19(3–4): 46–47, Fall 1988. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:1989:LBS

[CK89] M. Chrobak and H. Karloff. A lower bound on the size of universal sets for planar graphs. ACM SIGACT News, 20(4):83–86, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Crescenzi:1998:HFB

[CK98] Pierluigi Crescenzi and Viggo Kann. How to find the best approximation results. ACM SIGACT News, 29(4):90–97, December 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Clote:2002:BFC

[CK02] Peter Clote and Evangelos Kranakis. Boolean Functions and Computation Models. Texts in theoretical computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-59436-1. xiv + 601 pp. LCCN QA267.7

.C58 2002. URL http://www.loc.gov/catdir/enhancements/
fy0817/2001031128-d.html;   http://www.loc.gov/catdir/
enhancements/fy0817/2001031128-t.html.

Chrobak:2004:CMC

[CK04] Marek Chrobak and Elias Koutsoupias. Coordination mechanisms for congestion games. ACM SIGACT News, 35(4):58–71, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Czumaj:2018:HAH

[CK18] Artur Czumaj and Robert Krauthgamer. 3rd highlights of algorithms (HALG 2018). ACM SIGACT News, 49(3):74–77, September 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chrobak:2006:SNO

[CKM06] Marek Chrobak and Claire Kenyon-Mathieu. SIGACT news online algorithms column 10: competitiveness via doubling. ACM SIGACT News, 37(4):115–126, December 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cohen:2021:BAL

[CKN21] Shir Cohen, Idit Keidar, and Oded Naor. Byzantine agreement with less communication: Recent advances. ACM SIGACT News, 52(1):71–80, March 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3457588.3457600.

Cachin:2009:TC

[CKS09] Christian Cachin, Idit Keidar, and Alexander Shraer. Trusting the cloud. ACM SIGACT News, 40(2):81–86, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Censor:2009:RP

[CL09] Keren Censor and Christoph Lenzen. A review of PODC 2009. ACM SIGACT News, 40(4):71–74, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Czumaj:2017:HHA

[CL17] Artur Czumaj and Stefano Leonardi. HALG: Highlights of algorithms. ACM SIGACT News, 48(3):83–86, September 2017. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Clement:2008:DCS

[Cle08] Allen Clement. Distributed computing in SOSP and OSDI. ACM SIGACT News, 39(2):84–91, June 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cadiou:1972:RDP

[CM72] J. M. Cadiou and Zohar Manna. Recursive definitions of partial functions and their computations. ACM SIGACT News, 4(14): 58–65, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cohen:1993:CCA

[Coh93a] Henri Cohen. A Course in Computational Algebraic Number Theory, volume 138 of Graduate texts in mathematics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-55640-0 (Springer-Verlag Berlin Heidelberg New York), 0-387-55640-0 (Springer-Verlag New York Berlin Heidelberg). xxi + 534 pp. LCCN QA247 .C55 1993.

Cohen:1993:DGD

[Coh93b] Robert F. Cohen. Dynamic graph drawing (abstract). ACM SIGACT News, 24(1):60, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cohen:1995:CCA

[Coh95] Henri Cohen. A Course in Computational Algebraic Number Theory, volume 138 of Graduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second corrected printing edition, 1995. ISBN 3-540-55640-0 (Berlin), 0-387-55640-0 (New York). xxii + 534 pp. LCCN QA247 .C55 1995.

Cohen:1996:CCA

[Coh96] Henri Cohen. A Course in Computational Algebraic Number Theory, volume 138 of Graduate texts in mathematics, 0072-5285. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1996. ISBN 3-540-55640-0 (Berlin), 0-387-55640-0 (New York). xx + 545 pp. LCCN QA247

.C55 1996. URL http://www.loc.gov/catdir/enhancements/
fy0812/96044745-d.html.

Cohen:2007:BRB

[Coh07] Jonathan Cohen. Book review: Introduction to Lattices and Order, by B. A. Davey and H. A. Priestley, Cambridge University Press. ACM SIGACT News, 38(1):17–23, March 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [DP90, DP02].

Cohen:2008:JRB

[Coh08] Jonathan A. Cohen. Joint review of General Lattice Theory (second edition) and The Congruences of a Finite Lattice: a Proof-by-Picture Approach, by George Grätzer. ACM SIGACT News, 39 (3):22–26, September 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Grä78a, Grä78b, Grä98].

Constable:1978:PLI

[COJ78] R. L. (Robert L.) Constable, Michael J. O’Donnell, and Scott D. Johnson. A Programming Logic: with an Introduction to the PL/CV Verifier. Winthrop Publishers, Cambridge, MA, USA, 1978. ISBN 0-87626-726-6. x + 389 pp. LCCN QA76.73.P252 C66.

Colton:1993:RPL

[Col93] Don Colton. A restated pumping lemma for context-free languages. ACM SIGACT News, 24(2):87, Spring 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Coleman:2010:BRB

[Col10] Graham Coleman. Book review: Binary Quadratic Forms: an Algorithmic Approach, by Johannes Buchmann and Ulrich Vollmer Springer 2007. ACM SIGACT News, 41(4):34–35, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BV07].

Condon:1995:ASP

[Con95] Anne Condon. Approximate solutions to problems in PSPACE. ACM SIGACT News, 26(2):4–13, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Constantinides:2003:BRB

[Con03] George A. Constantinides. Book review: Computer Arithmetic Algorithms, by Israel Koren. A.K. Peters. ACM SIGACT News, 34(3):13–15, September 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Kor93, Kor02].

Constanda:2009:DCY

[Con09] C. Constanda. Dude, Can You Count?: Stories, Challenges, and Adventures in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-84882-538-2. xviii + 294 pp. LCCN QA109 2009.

Cook:1976:SPP

[Coo76] Stephen A. Cook. A short proof of the pigeon hole principle using extended resolution. ACM SIGACT News, 8(4):28–32, October 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cooke:1983:NHS

[Coo83] D. J. Cooke. On non-hierarchical systems of binary operators. ACM SIGACT News, 15(3):36–44, Summer–Fall 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Coonce:2004:CSM

[Coo04] Harry B. Coonce. Computer science and the mathematics genealogy project. ACM SIGACT News, 35(4):117, December 2004. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cook:2012:PTS

[Coo12] William Cook. In pursuit of the traveling salesman: mathematics at the limits of computation. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-15270-5. xiii + 228 pp. LCCN QA164 .C69 2012.

Cantone:2001:STC

[COP01] Domenico Cantone, Eugenio Omodeo, and Alberto Policriti. Set Theory for Computing: From Decision Procedures to Declarative Programming with Sets. Monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-95197-0. xviii + 409 pp. LCCN QA248 .C1945 2001. Foreword by Jacob T. Schwartz.

Coren:1983:GA

[Cor83] Stanley Coren. The game of academe. ACM SIGACT News, 15(2): 58–62, Spring–Summer 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cornuejols:2001:COP

[Cor01] Gérard Cornuéjols. Combinatorial Optimization: Packing and Covering, volume 74 of CBMS-NSF regional conference series in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2001. ISBN 0-89871-481-8 (paperback). xi + 132 pp. LCCN QA166.7

.C67 2001. URL http://www.loc.gov/catdir/enhancements/
fy0708/00066073-d.html;   http://www.loc.gov/catdir/enhancements/
fy0708/00066073-t.html.

Coutinho:2009:BRBb

[Cou09a] S. C. Coutinho. Book review: Quantum Computing for Computer Scientists, by Noson S. Yanofsky and Mirco A. Mannucci, Cambridge University Press, 2008. ACM SIGACT News, 40(4):14–17, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [YM08].

Coutinho:2009:BRBa

[Cou09b] S. C. Coutinho. Book review: Solving Polynomial Equation Systems II: Macaulay’s Paradigm and Gröbner Technology, by Teo Mora (Cambridge University Press 2005). ACM SIGACT News, 40(1):14–17, March 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Mor03, Mor05].

Coutinho:2010:BRB

[Cou10] S. C. Coutinho. Book review: Essays in Constructive Mathematics, by Harold M. Edwards, published by Springer, 2005 211 pages, hardcover. ACM SIGACT News, 41(2):33–36, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Edw05].

Coutinho:2011:BRB

[Cou11] S. C. Coutinho. Book review: Not Always Buried Deep: A Second Course in Elementary Number Theory, by Paul Pollack. ACM SIGACT News, 42(2):34–37, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Pol09].

Coutinho:2012:RBT

[Cou12] S. C. Coutinho. Review of Primality Testing and Integer Factorization in Public Key Cryptography by Song Y. Yan. ACM SIGACT News, 43(2):33–35, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Coutinho:2014:RPP

[Cou14] S. C. Coutinho. Review of Perspectives on projective geometry by Jürgen Richter-Gebert. ACM SIGACT News, 45(1):34–37, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Coutinho:2015:RGM

[Cou15] S. C. Coutinho. Review of: Games and Mathematics: Subtle Connections by David Wells. ACM SIGACT News, 46(3):7–10, September 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Constable:1974:SIS

[CP74] Robert Constable and David Park. Special issue on semantics and program schemas: SIAM Journal on Computing. ACM SIGACT News, 6(4):32, October 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chattopadhyay:2010:SSD

[CP10] Arkadev Chattopadhyay and Toniann Pitassi. The story of set disjointness. ACM SIGACT News, 41(3):59–85, September 2010. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chung:2013:GCP

[CP13] Kai-Min Chung and Rafael Pass. Guest column: parallel repetition theorems for interactive arguments. ACM SIGACT News, 44(1): 50–69, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Codenotti:2004:CME

[CPV04] Bruno Codenotti, Sriram V. Pemmaraju, and Kasturi R. Varadarajan. The computation of market equilibria. ACM SIGACT News, 35(4):23–37, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cook:1974:CLP

[CR74] Stephen Cook and Robert Reckhow. Corrections for “On the lengths of proofs in the propositional calculus preliminary version”. ACM SIGACT News, 6(3):15–22, July 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). Original paper published in Proceedings of the Sixth Annual Symposium on the Theory of Computing, Seattle 1974, 135–148 (1974).

Cogliati:2003:MWP

[CR03] Josh Cogliati and Rocky Ross. Mathematics on the Web: part II. ACM SIGACT News, 34(2):61–68, June 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Crampton:2003:PMCa

[Cra03a] Jeremy W. Crampton. The Political Mapping of Cyberspace. Edinburgh University Press, Edinburgh, Scotland, 2003. ISBN 0-7486-1412-5, 0-7486-1413-3 (paperback). vii + 214 pp. LCCN HM851 .C73 2003b.

Crampton:2003:PMCb

[Cra03b] Jeremy W. Crampton. The Political Mapping of Cyberspace. University of Chicago Press, Chicago, IL, USA, 2003. ISBN 0-226-11746-4 (hardcover.), 0-226-11746-4. vii + 214 pp. LCCN HM851

.C73 2003. URL http://www.loc.gov/catdir/bios/uchi051/
2003066280.html;   http://www.loc.gov/catdir/description/
uchi051/2003066280.html;   http://www.loc.gov/catdir/toc/
uchi051/2003066280.html.

Cole:1991:RSW

[CS91] Richard Cole and Alan Siegel. Report on the Second Workshop on Parallel Algorithms (WOPA). ACM SIGACT News, 22(3):20– 23, Summer 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chen:2023:GCN

[CT23] Lijie Chen and Roei Tell. Guest column: New ways of studying the BPP = P conjecture. ACM SIGACT News, 54(2):44–69, June 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3604943.
3604950.

Cudia:1969:DDU

[Cud69] Dennis F. Cudia. The degree diagram of undecidable problems of formal grammars. ACM SIGACT News, 1(3):7–8, 1969. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Culik:1982:TCA

[Cul82] Karel Culik. Theory of computation on abstract/concrete computer automata. ACM SIGACT News, 14(1):21–35, Winter 1982. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cull:1999:TAF

[Cul99] Paul Cull. Table-automata/ finite co-finite languages. ACM SIGACT News, 30(1):41, March 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cuny:1995:WOM

[Cun95] Janice E. Cuny. Workshops offer mentoring opportunities. ACM SIGACT News, 26(3):42–44, September 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cheatham:1972:LSA

[CW72] T. E. Cheatham, Jr. and Ben Wegbreit. On a laboratory for the study of automating programming. ACM SIGACT News, 4 (14):208–211, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Chartrand:2009:CGT

[CZ09] Gary Chartrand and Ping Zhang. Chromatic Graph Theory. Discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, 2009. ISBN 1-58488-800-8 (hardcover). xiii + 483 pp. LCCN QA166.247 .C43 2009. URL http://www.loc.

gov/catdir/toc/fy0903/2008044200.html.

Dani:2006:BRB

[Dan06] Varsha Dani. Book review: Fair Division and Collective Welfare, by Hervé Moulin, MIT Press, 2003. ACM SIGACT News, 37(2): 12–17, June 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Mou03].

Das:1992:BRB

[Das92] Sajal K. Das. Book review: Introduction to Parallel Algorithms and Architectures: Arrays, Trees, Hypercubes, by F. T. Leighton (Morgan Kauffman Pub, 1992). ACM SIGACT News, 23(3):31– 32, Summer 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Lei92].

Das:2009:RS

[Das09] Shantanu Das. Review of SIROCCO 2009. ACM SIGACT News, 40(4):93–97, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Davison:1993:VTC

[Dav93] Andrew Davison. Vague text compression. ACM SIGACT News, 24(1):68–74, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dave:2002:BRB

[Dav02] Maulik Dave. Book review: The Clausal Theory of Types. ACM SIGACT News, 33(4):8–9, December 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Wol93].

Dave:2005:BRB

[Dav05] Maulik A. Dave. Book review: Data Structures, Near Neighbor Searches, and Methodology: Fifth and Sixth DIMACS Implementation Challenge, by Michael H. Goldwasser, David S. Johnson, Catherine C. McGeoch, American Mathematical Society 2002. ACM SIGACT News, 36(4):5–8, December 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [GJM02].

Dave:2006:BRBa

[Dav06a] Maulik A. Dave. Book review: Secure Communicating Systems: Design, analysis, and implementation, by Michael R. A. Huth, Cambridge University Press, 2001. ACM SIGACT News, 37(1): 26–27, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Hut01].

Dave:2006:BRBb

[Dav06b] Maulik A. Dave. Book review: Verification of Reactive Systems: Formal Methods and Algorithms, by Klaus Schneider, Springer-Verlag Berlin Heidelberg, 2004. ACM SIGACT News, 37(4):36– 37, December 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sch04].

Dave:2008:BRBc

[Dav08a] Maulik A. Dave. Book review: Applied Combinatorics on Words. ACM SIGACT News, 39(3):28–30, September 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Lot05].

Dave:2008:BRBa

[Dav08b] Maulik A. Dave. Book review: Semantic Integration of Heterogeneous Software Specifications, by Martin Große-Rhode, Springer-Verlag, 2004. ACM SIGACT News, 39(1):28–29, March 2008. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [GR04].

Dave:2008:BRBb

[Dav08c] Maulik A. Dave. Book review: Theoretical and Experimental DNA Computation, by M. Amos, Springer-Verlag Berlin Heidelberg, 2005. ACM SIGACT News, 39(2):45–46, June 2008. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Amo05].

Dave:2009:RTE

[Dav09] Maulik A. Dave. Review 17 of Theoretical and Experimental DNA Computation, by M. Amos. Published in 2005 by Springer-Verlag, Berlin, Heidelberg. ACM SIGACT News, 40(4):35–36, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Amo05].

Dave:2011:BRB

[Dav11] Maulik A. Dave. Book review: Rippling: Meta-Level Guidance for Mathematical Reasoning Cambridge Tracts in Theoretical Computer Science 56 by Alan Bundy, David Basin, Dieter Hutter, and Andrew Ireland. ACM SIGACT News, 42(4):21–23, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BHI05].

Davies:2017:PR

[Dav17] Peter Davies. PODC 2017 review. ACM SIGACT News, 48(4):91– 93, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

deBerg:1997:CGA

[dB^{+}97] Mark de Berg et al. Computational Geometry: Algorithms and Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-61270-X. xii + 365 pp. LCCN QA448.D38 C65 1997.

deBerg:2000:CGA

[dB^{+}00] Mark de Berg et al. Computational Geometry: Algorithms and Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 3-540-65620-0. xii + 367 pp. LCCN QA448.D38 C65 2000.

denBussche:2006:DTC

[dB06] Jan Van den Bussche. Database theory column: report on PODS 2006. ACM SIGACT News, 37(4):56–57, December 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

deBerg:2008:CGA

[dB^{+}08] Mark de Berg et al. Computational Geometry: Algorithms and Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2008. ISBN 3-540-77973-6 (hardcover), 3-540-77974-4 (ebook). xii + 386 pp. LCCN QA448.D38 C65 2008. URL http://site.ebrary.com/

lib/stanford/Doc?id=10231750.

Denny-Brown:1994:SMA

[DBHHT94] Derek Denny-Brown, Yenjo Han, Lane A. Hemaspaandra, and Leen Torenvliet. Semi-membership algorithms: some recent advances. ACM SIGACT News, 25(3):12–23, September 1994. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Davis:1982:RMG

[DCV82] Natalie Zemon Davis, Jean-Claude Carrière, and Daniel Vigne. Le retour de Martin Guerre. R. Laffont, Paris, France, 1982. ISBN 2-221-00744-1. 269 pp. LCCN KJV130.D8 D38 1982.

Dean:2009:BRB

[Dea09] Alice M. Dean. Book review: Visibility Algorithms in the Plane, by Subir Kumar Ghosh (Cambridge University Press, 2007). ACM SIGACT News, 40(2):33–35, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Gho07].

Dekhtyar:1997:BRB

[Dek97] Alexander Dekhtyar. Book review: Reasoning About Knowledge, by Ronald Fagin, Joseph Halpern, Yoram Moses and Moshe Verdi. ACM SIGACT News, 28(4):20–23, December 1997. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [FHMV95, FHMV03].

Dekhtyar:1998:BRB

[Dek98] Alexander Dekhtyar. Book review: Vicious Circles, by Jon Barwise and Lawrence Moss (Cambridge University Press 1996). ACM SIGACT News, 29(4):10–14, December 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BM96].

Dekhtyar:1999:BRBb

[Dek99a] Alexander Dekhtyar. Book review: Logic for Applications, Second Edition, by Anil Nerode and Richard A. Shore (Springer Verlag 1997). ACM SIGACT News, 30(3):15–18, September 1999. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [NS93, NS97].

Dekhtyar:1999:BRBa

[Dek99b] Michael Dekhtyar. Book review: Computational Geometry in C, Second Edition, by Joseph O’Rourke (Cambridge University Press 1988). ACM SIGACT News, 30(3):8–13, September 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [O’R94d].

Demontis:2009:SNH

[Dem09] R. Demontis. A simple NP-hard problem. ACM SIGACT News, 40 (2):45–48, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Denysyuk:2014:RP

[Den14] Oksana Denysyuk. Review of PODC 2014. ACM SIGACT News, 45(4):89–93, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dewdney:1979:LCP

[Dew79] A. K. Dewdney. Logic circuits in the plane: minimal crossovers. ACM SIGACT News, 10(3):38–48, Winter 1979. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dewdney:1986:EGR

[Dew86] A. K. Dewdney. Experiments with a generic reduction computer. ACM SIGACT News, 18(1):75–79, Summer 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dey:2007:CSR

[Dey07] Tamal K. (Tamal Krishna) Dey. Curve and Surface Reconstruction: Algorithms with Mathematical Analysis, volume 23 of The Cambridge monographs on applied and computational mathematics. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-86370-8 (hardcover). xiii + 214 pp. LCCN QA565

.D49 2007. URL http://www.loc.gov/catdir/enhancements/
fy0659/2006017359-d.html;   http://www.loc.gov/catdir/
enhancements/fy0659/2006017359-t.html;   http://www.loc.
gov/catdir/enhancements/fy0733/2006017359-b.html.

Downey:1999:PC

[DF99] R. G. (Rod G.) Downey and M. R. (Michael Ralph) Fellows. Parameterized Complexity. Monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-94883-X (hardcover). xv + 533 pp.

LCCN QA267.7 .D68 1999. URL http://www.loc.gov/catdir/
enhancements/fy0815/97022882-d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/97022882-t.html.

Dorst:2007:GAC

[DFM07] Leo Dorst, Daniel Fontijne, and Stephen Mann. Geometric Algebra for Computer Science: an Object-Oriented Approach to Geometry. Morgan Kaufmann series in computer graphics. Elsevier, Amsterdam, The Netherlands, 2007. ISBN 0-12-369465-5, 0-12-374942-5 (hardcover). xxxv + 626 pp. LCCN QA564

.D67 2007. URL http://www.loc.gov/catdir/enhancements/
fy1011/2010481601-d.html.

deFraysseix:1993:POP

[dFR93] Hubert de Fraysseix and Pierre Rosenstiehl. Partial orders for planarity and drawings (abstract). ACM SIGACT News, 24(1):58, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Danek:2008:RD

[DGW08] Robert Danek, Wojciech Golab, and Wojciech Wawrzyniak. Review of DISC 2008. ACM SIGACT News, 39(4):60–65, December 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Derus:1979:LTD

[DH79a] Kenneth H. Derus and John C. Hansen. Logics of truth and disposition. ACM SIGACT News, 11(1):36–43, Summer 1979. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Derus:1979:PMD

[DH79b] Kenneth H. Derus and John C. Hansen. Propositions with multiple dispositions and multiple truth values. ACM SIGACT News, 11 (2):30–35, Fall 1979. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Downey:2010:ARC

[DH10] R. G. (Rod G.) Downey and Denis Roman Hirschfeldt. Algorithmic Randomness and Complexity. Theory and applications of computability, 2190-619X. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 0-387-95567-4, 0-387-68441-7 (e-book). xxviii + 855 pp. LCCN QA267.7

.D67 2010. URL http://www.loc.gov/catdir/enhancements/
fy1202/2011377427-d.html;   http://www.loc.gov/catdir/
enhancements/fy1202/2011377427-t.html.

Dhulipa:2019:SR

[Dhu19] Laxman Dhulipa. SPAA 2018 review. ACM SIGACT News, 50(1): 61–64, March 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

DiBattista:1993:ARA

[Di 93] Giuseppe Di Battista. Area requirements (abstract). ACM SIGACT News, 24(1):58–59, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

DiCosmo:1995:ITL

[Di 95] Roberto Di Cosmo. Isomorphisms of Types: From λ-Calculus to Information Retrieval and Language Design. Progress in theoretical computer science. Birkhäuser Verlag, Basel, Switzerland, 1995. ISBN 0-8176-3763-X , 3-7643-3763-X. viii + 235 pp. LCCN QA76.7 .D5 1995.

Diderich:1993:BMT

[Did93] Claude G. Diderich. A bibliography on minimax trees. ACM SIGACT News, 24(4):82–89, December 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL http:/

/diwww.epfl.ch/~diderich/research/minimax/bib/minimax.
html.

Dietzfelbinger:2004:PTP

[Die04] Martin Dietzfelbinger. Primality Testing in Polynomial Time: From Randomized Algorithms to “Primes” Is in P, volume 3000 of LNCS, 0302-9743; Tutorial. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-40344-2. x + 147 pp. LCCN QA246 .D54 2004. URL http:

//www.loc.gov/catdir/enhancements/fy0817/2004107785-b.
html;   http://www.loc.gov/catdir/enhancements/fy0817/
2004107785-d.html.

Diffie:1982:CTF

[Dif82] Whitfield Diffie. Cryptographic technology: fifteen year forecast. ACM SIGACT News, 14(4):38–57, Fall–Winter 1982. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dinur:2008:PSS

[Din08] Irit Dinur. PCPs with small soundness error. ACM SIGACT News, 39(3):41–57, September 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dinitz:2013:RAM

[Din13] Michael Dinitz. Recent advances on the matroid secretary problem. ACM SIGACT News, 44(2):126–142, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dumitrescu:2013:CGCa

[DJ13a] Adrian Dumitrescu and Minghui Jiang. Computational geometry column 56. ACM SIGACT News, 44(2):80–87, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dumitrescu:2013:CGCb

[DJ13b] Adrian Dumitrescu and Minghui Jiang. Computational geometry column 58. ACM SIGACT News, 44(4):73–78, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dumitrescu:2014:CGCb

[DJ14] Adrian Dumitrescu and Minghui Jiang. Computational geometry column 60. ACM SIGACT News, 45(4):76–82, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dumitrescu:2019:CGC

[DJ19] Adrian Dumitrescu and Minghui Jiang. Computational geometry column 69. ACM SIGACT News, 50(3):75–90, September 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

deJong:1975:IBN

[dJvL75] Lieuwe de Jong and Jan van Leeuwen. An improved bound on the number of multiplications and divisions necessary to evaluate a polynomial and all its derivatives. ACM SIGACT News, 7(3): 32–34, Summer 1975. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Du:2000:TCC

[DK00] Dingzhu Du and Ker-I. Ko. Theory of Computational Complexity. John Wiley, New York, NY, USA, 2000. ISBN 0-471-34506-7. xiii + 491 pp. LCCN QA267.7 .D8 2000. URL

http://www.loc.gov/catdir/bios/wiley042/99053576.html;
 http://www.loc.gov/catdir/description/wiley034/99053576.
html;   http://www.loc.gov/catdir/toc/onix05/99053576.
html.

Delfs:2002:ICP

[DK02] Hans Delfs and Helmut Knebl. Introduction to Cryptography: Principles and Applications. Information security and cryptography. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-42278-1 (hardcover). xiv + 310 pp. LCCN QA76.9.A25 D447 2002.

Delfs:2007:ICP

[DK07] Hans Delfs and Helmut Knebl. Introduction to Cryptography: Principles and Applications. X-9218-2100-8. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2007. ISBN 3-540-49243-7 (hardcover). xvi + 367 pp. LCCN QA76.9.A25 D447 2007.

deKoninck:2009:FN

[dK09] J. M. de Koninck. Those Fascinating Numbers. American Mathematical Society, Providence, RI, USA, 2009. ISBN 0-8218-4807-0. xviii + 426 pp. LCCN QA241 .K685913 2009.

Droste:2009:HWA

[DKV09] Manfred Droste, Werner Kuich, and Heiko Vogler, editors. Handbook of Weighted Automata. Monographs in theoretical computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-642-01491-7. xvii + 608 pp. LCCN QA267 .H34 2009. URL http:

//www.loc.gov/catdir/enhancements/fy1109/2009926056-d.

html;   http://www.loc.gov/catdir/enhancements/fy1109/
2009926056-t.html.

Dantsin:2005:QVR

[DKW05] Evgeny Dantsin, Vladik Kreinovich, and Alexander Wolpert. On quantum versions of record-breaking algorithms for SAT. ACM SIGACT News, 36(4):103–108, December 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

DeMillo:1979:BRB

[DL79] Richard A. DeMillo and Richard J. Lipton. Book review: Proofs and Refutations: the Logic of Mathematical Discovery, by Imre Lakatos. Cambridge University Press 1976. ACM SIGACT News, 11(2):7–9, Fall 1979. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Lak76].

Diffie:1998:PLP

[DL98] Whitfield Diffie and Susan Eva Landau. Privacy on the Line: the Politics of Wiretapping and Encryption. MIT Press, Cambridge, MA, USA, 1998. ISBN 0-262-04167-7 (hardcover). ix + 342 pp. LCCN KF9670 .D54 1998.

Diffie:2007:PLP

[DL07] Whitfield Diffie and Susan Eva Landau. Privacy on the Line: the Politics of Wiretapping and Encryption. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-04240-1 (hardcover). xvii + 472 pp.

LCCN UB256.U6 D54 2007. URL http://www.loc.gov/catdir/
toc/ecip073/2006035514.html.

Durand:2007:SPP

[DLM07] Arnaud Durand, Clemens Lautemann, and Malika More. A simple proof of the polylog counting ability of first-order logic: guest column. ACM SIGACT News, 38(4):40–45, December 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dorrigiv:2005:SPM

[DLO05] Reza Dorrigiv and Alejandro López-Ortiz. A survey of performance measures for on-line algorithms. ACM SIGACT News, 36(3):67– 81, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dorrigiv:2009:DNM

[DLO09] Reza Dorrigiv and Alejandro López-Ortiz. On Developing New Models, with Paging as a Case Study. ACM SIGACT News, 40 (4):98–123, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dwibedy:2022:OLS

[DM22] Debasis Dwibedy and Rakesh Mohanty. Online list scheduling for makespan minimization: a review of the state-of-the-art results, research challenges and open problems. ACM SIGACT News, 53 (2):84–105, June 2022. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3544979.3544993.

Demaine:1999:CGC

[DO99] Erik D. Demaine and Joseph O’Rourke. Computational geometry column 37. ACM SIGACT News, 30(3):39–42, September 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Demaine:2007:GFA

[DO07] Erik D. Demaine and Joseph O’Rourke. Geometric Folding Algorithms: Linkages, Origami, Polyhedra. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-85757-

0. xiii + 472 pp. LCCN QA491 .D46 2007. URL http:

//www.loc.gov/catdir/enhancements/fy0703/2006038156-d.
html;   http://www.loc.gov/catdir/enhancements/fy0703/
2006038156-t.html;   http://www.loc.gov/catdir/enhancements/
fy0803/2006038156-b.html.

Datta:2013:OCT

[DO13] Anwitaman Datta and Frédérique Oggier. An overview of codes tailor-made for better repairability in networked distributed storage systems. ACM SIGACT News, 44(1):89–105, March 2013. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Guimaraes:2010:BRB

[dOG10] José de Oliveira Guimarães. Book review: Is Mathematics Inevitable? A Miscellany, by Underwood Dudley (editor), The Mathematical Association of America, 2008 324 pages, hardcover. ACM SIGACT News, 41(2):36–37, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Dud08].

Guimaraes:2012:RBC

[dOG12] José de Oliveira Guimarães. Review of Models of Computation: an Introduction to Computability Theory, by Maribel Fernández.

ACM SIGACT News, 43(3):28–31, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dolev:2007:RDC

[Dol07] Shlomi Dolev. A review of the DISC 2006 conference. ACM SIGACT News, 38(1):46–52, March 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

DosReis:1987:RLD

[Dos87a] A. J. Dos Reis. Regular languages do not form a lattice under GSM mappings. ACM SIGACT News, 18(3):46–47, April 1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

DosReis:1987:RLU

[Dos87b] A. J. Dos Reis. Regular languages under F-gsm mappings. ACM SIGACT News, 18(3):41–45, April 1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Davey:1990:ILO

[DP90] B. A. Davey and H. A. (Hilary A.) Priestley. Introduction to Lattices and Order. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-36584-8, 0-521-36766-2 (paperback). viii + 248 pp. LCCN QA171.5 .D38 1990. URL http://www.loc.gov/

catdir/description/cam023/89009753.html;   http://www.
loc.gov/catdir/toc/cam029/89009753.html.

Davey:2002:ILO

[DP02] B. A. Davey and H. A. (Hilary A.) Priestley. Introduction to Lattices and Order. Cambridge University Press, Cambridge, UK, second edition, 2002. ISBN 0-521-78451-4 (paperback). xii + 298 pp. LCCN QA171.5 .D38 2002. URL http://www.loc.

gov/catdir/description/cam022/2001043910.html;   http:
//www.loc.gov/catdir/toc/cam026/2001043910.html.

Denchev:2008:DQC

[DP08] Vasil S. Denchev and Gopal Pandurangan. Distributed quantum computing: a new frontier in distributed systems or science fiction? ACM SIGACT News, 39(3):77–95, September 2008. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dubhashi:2009:CMA

[DP09] Devdatt Dubhashi and Alessandro Panconesi. Concentration of Measure for the Analysis of Randomized Algorithms. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-88427-6 (hardcover). xiv + 196 pp. LCCN QA273 .D765 2009.

Doxiadis:2009:LES

[DPP09] Apostolos Doxiadis, Alecos Papadatos, and Christos H. Papadimitriou. Logicomix: an Epic Search for Truth. Bloomsbury, New York, NY, USA, 2009. ISBN 1-59691-452-1. 347 pp. LCCN PA5615.O87 L6413 2009.

Dasgupta:2008:A

[DPV08] Sanjoy Dasgupta, Christos H. Papadimitriou, and Umesh Virkumar Vazirani. Algorithms. McGraw-Hill Higher Education, Boston, MA, USA, 2008. ISBN 0-07-352340-2. x + 320 pp. LCCN QA9.58

.D37 2008. URL http://www.loc.gov/catdir/enhancements/
fy0665/2006049014-t.html;   http://www.loc.gov/catdir/
enhancements/fy0701/2006049014-d.html.

Dershowitz:2012:MIP

[DR12] Nachum Dershowitz and Edward M. Reingold. Modulo intervals: a proposed notation. ACM SIGACT News, 43(3):60–64, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Costa:1994:IC

[dRCR94] Antônio Carlos da Rocha Costa and Vanderlei Moraes Rodrigues. Inspecting continuations. ACM SIGACT News, 25(2):90–91, June 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

deRezende:2022:GCP

[dRGR22] S. F. de Rezende, M. Göös, and R. Robere. Guest column: Proofs, circuits, and communication. ACM SIGACT News, 53(1):59– 82, March 2022. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3532737.3532746.

Dumitrescu:2012:CGCb

[DT12] Adrian Dumitrescu and Csaba D. Tóth. Computational geometry column 54. ACM SIGACT News, 43(4):90–97, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dumitrescu:2014:CGCa

[DT14] Adrian Dumitrescu and Csaba D. Tóth. Computational geometry column 59. ACM SIGACT News, 45(2):68–72, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dudley:2008:MIM

[Dud08] Underwood Dudley, editor. Is Mathematics Inevitable?: a Miscellany. Spectrum series. Mathematical Association of America, Washington, DC, USA, 2008. ISBN 0-88385-566-6 (hardcover). x + 325 pp. LCCN QA99 .I7 2008.

Dudley:2009:GEN

[Dud09] Underwood Dudley. A Guide to Elementary Number Theory, volume 41; # 5 of The Dolciani mathematical expositions; MAA guides. Mathematical Association of America, Washington, DC, USA, 2009. ISBN 0-88385-347-7. x + 141 pp. LCCN QA241 .D793 2009.

Dumitrescu:2012:CGCa

[Dum12] Adrian Dumitrescu. Computational geometry column 53. ACM SIGACT News, 43(2):78–83, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dumitrescu:2016:CGC

[Dum16] Adrian Dumitrescu. Computational geometry column 64. ACM SIGACT News, 47(4):44–47, December 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dumitrescu:2018:CGC

[Dum18] Adrian Dumitrescu. Computational geometry column 68. ACM SIGACT News, 49(4):46–54, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dunham:1984:SUF

[Dun84] Charles B. Dunham. A sequence of uncomputable functions. ACM SIGACT News, 16(3):48, Fall 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dunham:1986:CBP

[Dun86a] C. B. Dunham. The cycle burning problem. ACM SIGACT News, 18(1):51, Summer 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dunham:1986:SAB

[Dun86b] Charles B. Dunham. A simpler approach to the Busy Beaver Problem. ACM SIGACT News, 17(3):29, Winter 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dunham:1987:PU

[Dun87] C. B. Dunham. Pessimization is unsolvable. ACM SIGACT News, 18(3):37, April 1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dunham:2000:PWC

[Dun00] Charles B. Dunham. Partially Wrong Completions. ACM SIGACT News, 31(2):69, June 1, 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Durrett:2010:PTE

[Dur10] Richard Durrett. Probability: Theory and Examples. Cambridge University Press, Cambridge, UK, fourth edition, 2010. ISBN 0-534-42441-4, 0-521-76539-0. x + 428 pp. LCCN QA273 .D865 2005.

Vera:2016:RDS

[dV16] Ramon de Vera, Jr. Review of: Distributed Systems: an Algorithmic Approach (2nd Edition) by Sukumar Ghosh. ACM SIGACT News, 47(4):13–14, December 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Vera:2018:RRW

[dV18] Ramon de Vera Jr. Review of Real-World Algorithms: a Beginner’s Guide by Panos Louridas. ACM SIGACT News, 49(1):16–19, March 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dvir:2009:GCR

[Dvi09] Z. Dvir. Guest column: from randomness extraction to rotating needles. ACM SIGACT News, 40(4):46–61, December 2009. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

deWolf:2007:JRB

[dW07] Ronald de Wolf. Joint review of An Introduction to Quantum Computing Algorithms, by Arthur O. Pittenger, Birkhäuser, ISBN 0-8176-4127-0; Quantum Computing, by Mika Hirvensalo, Springer, ISBN 3-540-66783-0; and Classical and Quantum Computation, by A. Yu. Kitaev, A. Shen, and M. N. Vyalyi, American Mathematical Society, ISBN 0-8218-2161-X. ACM SIGACT News, 38(1):14–17, March 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Hir01, KSV02, Pit00] and review [Vio04].

DiPaula:2013:RBG

[DW13] Cynthia DiPaula and Andrew Wonnacott. Review of The Golden Ticket: P, NP, and the Search for the Impossible by Lance Fortnow. ACM SIGACT News, 44(3):35–37, September 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Du:2019:MKK

[DW19] Ding-Zhu Du and Jie Wang. In memoriam: Ker-I Ko (1950–2018). ACM SIGACT News, 50(1):79, March 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dwork:1989:DCC

[Dwo89] C. Dwork. Distributed computing column. ACM SIGACT News, 20(4):14–16, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dwork:1990:DCC

[Dwo90] Cynthia Dwork. Distributed computing column. ACM SIGACT News, 21(2):12, Spring 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dwork:1995:DCC

[Dwo95] Cynthia Dwork. Distributed computing column: Lotus Notes security and authentication. ACM SIGACT News, 26(1):17–19, March 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dwork:1996:DCC

[Dwo96] Cynthia Dwork. Distributed computing column. ACM SIGACT News, 27(3):50–54, September 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dwyer:1990:KGA

[Dwy90] Rex A. Dwyer. Kinder, gentler average-case analysis for convex hulls and maximal vectors. ACM SIGACT News, 21(2):64, Spring 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Dwyer:2003:GPB

[Dwy03] Rex A. Dwyer. Genomic Perl: From Bioinformatics Basics to Working Code. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-80177-X. xvii + 334 pp. LCCN

QA76.73.P22 D88 2003. URL http://www.loc.gov/catdir/toc/
cam031/2003268666.html.

Dyer:2011:BRB

[Dye11] Jason Dyer. Book review: Creative Mathematics, by H. S. Wall. ACM SIGACT News, 42(1):24–25, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Wal63, Wal09].

El-Atawy:2009:BRB

[EA09] Adel El-Atawy. Book review: An Introduction to Difference Equations: third edition by Saber Elaydi. ACM SIGACT News, 40 (3):42–46, September 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Ela96, Ela99, Ela05].

El-Atawy:2010:BRB

[EA10] Adel El-Atawy. Book review: An Introduction to Data Structures and Algorithms, by James A. Storer, Birkhäuser. ACM SIGACT News, 41(1):15–19, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sto02].

Eades:1993:ADT

[Ead93] Peter Eades. Algorithms for drawing trees (abstract). ACM SIGACT News, 24(1):57, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Elbassioni:2017:CGC

[ED17] Khaled Elbassioni and Adrian Dumitrescu. Computational geometry column 66. ACM SIGACT News, 48(4):57–74, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Edmonds:2008:HTA

[Edm08] Jeff Edmonds. How to Think about Algorithms. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-84931-4 (hardcover), 0-521-61410-4 (paperback). xiii + 448 pp. LCCN QA9.58

.E36 2008. URL http://www.loc.gov/catdir/enhancements/
fy0808/2008001238-b.html;   http://www.loc.gov/catdir/
enhancements/fy0808/2008001238-d.html;   http://www.loc.
gov/catdir/enhancements/fy0808/2008001238-t.html.

Edwards:2005:ECM

[Edw05] Harold M. Edwards. Essays in Constructive Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-21978-1. xix + 211 pp. LCCN QA9.56

.E39 2005. URL http://www.loc.gov/catdir/enhancements/
fy0662/2004049156-d.html;   http://www.loc.gov/catdir/
enhancements/fy0818/2004049156-b.html;   http://www.loc.
gov/catdir/enhancements/fy0818/2004049156-t.html.

Edwards:2008:HAA

[Edw08] Harold M. Edwards. Higher Arithmetic: an Algorithmic Introduction to Number Theory, volume 45 of Student mathematical library. American Mathematical Society, Providence, RI, USA, 2008. ISBN 0-8218-4439-3. ISSN 1520-9121. xii + 210 pp. LCCN QA241 .E39 2008.

Eilenberg:1976:BRB

[Eil76] Samuel Eilenberg. Book review: Algebraic and Automata-Theoretic Properties of Formal Languages, by Seymour Ginsburg. North Holland, 1975. ACM SIGACT News, 8(4):11–12, October 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Gin75].

Ehrig:1979:AIA

[EKP79] H. Ehrig, H. J. Kreowski, and P. Padawitz. Algebraic implementation of abstract data types: an announcement. ACM SIGACT News, 11(2):25–29, Fall 1979. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Elaydi:1996:IDE

[Ela96] Saber Elaydi. An Introduction to Difference Equations. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94582-2 (hardcover). xiii + 389 pp. LCCN QA431 .E43 1996.

Elaydi:1999:IDE

[Ela99] Saber Elaydi. An Introduction to Difference Equations. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1999. ISBN 0-387-98830-0 (hardcover). xviii + 427 pp. LCCN QA431 .E43 1999.

Elaydi:2005:IDE

[Ela05] Saber Elaydi. An Introduction to Difference Equations. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2005. ISBN 0-387-23059-9. xxii + 539 pp. LCCN QA431

.E43 2005. URL http://www.loc.gov/catdir/enhancements/
fy0662/2004058916-d.html;   http://www.loc.gov/catdir/
enhancements/fy0813/2004058916-t.html.

Elbirt:2009:UAC

[Elb09] Adam J. Elbirt. Understanding and Applying Cryptography and Data Security. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-6160-7. xxvii + 637 pp. LCCN QA76.9.A25 E43 2009. URL http://www.loc.gov/

catdir/toc/ecip0821/2008028154.html.

Elkin:2004:DAS

[Elk04] Michael Elkin. Distributed approximation: a survey. ACM SIGACT News, 35(4):40–57, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ellis:1970:NCF

[Ell70] Clarence A. Ellis. Non-context-free probabilistic languages. ACM SIGACT News, 2(6):9–14, 1970. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Elrod:2011:BRB

[Elr11] Hal C. Elrod. Book review: Algorithms and Data Structures: the Basic Toolbox, by Kurt Mehlhorn and Peter Sanders. ACM SIGACT News, 42(4):26–29, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [MS08].

Engelfriet:1989:CCP

[Eng89] Joost Engelfriet. The complexity of the circularity problem for attribute grammars: a note on a counterexample for a simpler construction. ACM SIGACT News, 20(3):57–59, July 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Englert:2018:RBP

[Eng18] Matthias Englert. The reordering buffer problem on the line revisited. ACM SIGACT News, 49(1):67–72, March 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Epstein:1987:NCC

[Eps87] D. Epstein. On the NP-completeness of cryptarithms. ACM SIGACT News, 18(3):38–40, April 1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Eqvhazy:1985:GSF

[ER85] C. J. Eqvhazy and T. B. Ramsay. Generating a set of fundamental cycles in a graph. ACM SIGACT News, 17(1):78–82, June 1985. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Erickson:2010:PDM

[Eri10] Martin J. Erickson. Pearls of Discrete Mathematics. Discrete mathematics and its applications. Taylor and Francis, Boca Raton, FL, USA, 2010. ISBN 1-4398-1616-6 (paperback). x + 270 pp. LCCN QA164 .E745 2010.

Escardo:1993:LNN

[Esc93] Martín Hötzel Escardó. On lazy natural numbers with applications to computability theory and functional programming. ACM SIGACT News, 24(1):61–67, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Essl:2005:BRB

[Ess05] Georg Essl. Book review: Computational Line Geometry, by H. Pottmann and J. Wallner, Springer Verlag, 2001. ACM SIGACT News, 36(3):13–17, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [PW01, PW09].

Even:1983:PSC

[Eve83] Shimon Even. A protocol for signing contracts. ACM SIGACT News, 15(1):34–39, Winter–Spring 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Eyal:2013:FTT

[Eya13] Ittay Eyal. Fault tolerant transaction architectures. ACM SIGACT News, 44(3):78–84, September 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fruhwirth:2003:ECP

[FA03] Thom Frühwirth and Slim Abdennadher. Essentials of Constraint Programming. Cognitive technologies. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-67623-6. ix + 144 pp. LCCN QA76.612

.F78 2003. URL http://www.loc.gov/catdir/enhancements/
fy0812/2002044502-d.html;   http://www.loc.gov/catdir/
enhancements/fy0812/2002044502-t.html.

Faber:2011:BRB

[Fab11] Vance Faber. Book review: Chromatic Graph Theory, by Gary Chartrand and Ping Zhang. ACM SIGACT News, 42(3):23–28, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [CZ09].

Fahle:2004:BRB

[Fah04] William Fahle. Book review: Introduction to the Design and Analysis of Algorithms, by Ananay Levitin, Addison-Wesley. ACM SIGACT News, 35(4):18–19, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Lev07, Lev12].

Fahle:2005:BRB

[Fah05] William Fahle. Book review: Graphs, Networks, and Algorithms (second edition) by Dieter Jungnickel, Springer-Verlag 2005. ACM SIGACT News, 36(4):12–14, December 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Jun99, Jun05, Jun08].

Faliszewski:2007:BRB

[Fal07] Piotr Faliszewski. Book review: Complexity Theory and Cryptology: An Introduction to Cryptocomplexity, by Jörg Rothe, Springer, 2005, 484 pages. ACM SIGACT News, 38(2):20–22, June 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Rot05].

Falke:2012:RBM

[Fal12a] Stephan Falke. Review of Introduction to Mathematics of Satisfiability, by Victor W. Marek. ACM SIGACT News, 43(3):41– 44, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Falke:2012:RBT

[Fal12b] Stephan Falke. Review of Transitions and Trees: an Introduction to Structural Operational Semantics, by Hans Hüttel. ACM SIGACT News, 43(4):34–37, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fan:2013:RP

[Fan13] Wenfei Fan. Report on PODS 2013. ACM SIGACT News, 44 (3):69–71, September 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Farr:2019:UGT

[Far19] Graham Farr. Using Go in teaching the theory of computation. ACM SIGACT News, 50(1):65–78, March 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fass:1983:LCF

[Fas83] Leona F. Fass. Learning context-free languages from their structured sentences. ACM SIGACT News, 15(3):24–35, Summer–Fall 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fass:1986:ICC

[Fas86] L. F. Fass. On the inference of canonical context-free grammars. ACM SIGACT News, 17(4):55–60, Spring 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fass:2004:LMO

[Fas04] Leona F. Fass. Language, mathematics and other dangerous things. ACM SIGACT News, 35(1):74–79, March 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Faulon:2010:HCA

[FB10] Jean-Loup Faulon and Andreas Bender, editors. Handbook of Chemoinformatics Algorithms. Chapman and Hall/CRC mathematical and computational biology series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2010. ISBN 1-4200-8292-2 (hardcover). xii + 440 pp. LCCN QD39.3.E46 H357 2010.

Fletez-Brant:2014:RBP

[FB14] Kipper Fletez-Brant. Review of Basic phylogenetic combinatorics by Andreas Dress, Katharina T. Huber, Jacobus Koolen, Vincent Moulton and Andreas Spillner. ACM SIGACT News, 45(3):26– 28, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Findler:1983:AIT

[FC83] Nicholas V. Findler and Robert F. Cromp. An artificial intelligence technique to generate self-optimizing experimental designs. ACM SIGACT News, 15(2):77, Spring–Summer 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Farach-Colton:1999:MSM

[FC^{+}99] Martin Farach-Colton et al., editors. Mathematical Support for Molecular Biology: Papers Related to the Special Year in Mathematical Support for Molecular Biology, 1994–1998, volume 47 of DIMACS series in discrete mathematics and theoretical computer science, 1052-1798. American Mathematical Society, Providence, RI, USA, 1999. ISBN 0-8218-0826-5. xv + 288 pp. LCCN QH506 .M385 1999.

Feuilloley:2023:ECO

[FdV23] Laurent Feuilloley and Tijn de Vos. The environmental cost of our conferences: The CO_{2} emissions due to travel at PODC and DISC. ACM SIGACT News, 54(4):94–107, December 2023. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

URL https://dl.acm.org/doi/10.1145/3639528.3639537.

Feigenbaum:1989:RDS

[Fei89] J. Feigenbaum. Report on DIMACS seminar series. ACM SIGACT News, 20(4):48–49, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Feigenbaum:1990:RDS

[Fei90] Joan Feigenbaum. Report on DIMACS seminar series. ACM SIGACT News, 21(2):25, Spring 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fekete:2009:TAT

[Fek09] Alan D. Fekete. Teaching about threading: where and what? ACM SIGACT News, 40(1):51–57, March 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fenner:2021:RA

[Fen21a] Stephen Fenner. Remembrances of Alan. ACM SIGACT News, 52(1):87–93, March 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3457588.3457603.

Fenner:2021:RTQ

[Fen21b] Stephen A. Fenner. Review of The Theory of Quantum Information, John Watrous. ACM SIGACT News, 52(1):16–24, March 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3457588.
3457593.

Ferreira:2005:SRC

[Fer05] Ulisses Ferreira. The sets of real and complex numbers are denumerable. ACM SIGACT News, 36(2):126–130, June 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fernandez:2009:MCI

[Fer09] Maribel Fernández. Models of computation: an introduction to computability theory. Undergraduate topics in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-84882-433-5 (paperback). xi + 184 pp. LCCN QA267 .F47 2009.

Feferman:2004:ATL

[FF04] Anita Burdman Feferman and Solomon Feferman. Alfred Tarski: Life and Logic. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-80240-7. vi + 425 pp. LCCN QA29.T32 F44

2004. URL http://www.loc.gov/catdir/description/cam041/
2004045748.html;   http://www.loc.gov/catdir/toc/cam041/
2004045748.html.

Fischer:1989:PRI

[FFB89] M. J. Fischer, R. P. Fischer, and R. Beigel. Primitive recursion without implicit predecessor. ACM SIGACT News, 20(4):87–91, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fenner:1996:CTN

[FFG96] Stephen A. Fenner, Lance J. Fortnow, and William J. Gasarch. Complexity theory newsflash. ACM SIGACT News, 27(3):126, September 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Felber:2008:TBT

[FFGH08] Pascal Felber, Christof Fetzer, Rachid Guerraoui, and Tim Harris. Transactions are back—but are they the same?: Le Retour de Martin Guerre (Sommersby). ACM SIGACT News, 39(1):48–58, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [DCV82].

French:1994:WAT

[FFMS94] Jim French, Edward Fox, Kurt Maly, and Alan L. Selman. Wide area technical report service-technical reports online. ACM SIGACT News, 25(3):124–127, September 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fenner:2018:FPE

[FGH18] Stephen Fenner, Frederic Green, and Steven Homer. Fixedparameter extrapolation and aperiodic order: Open problems. ACM SIGACT News, 49(3):35–47, September 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ferson:2002:CVI

[FGK^{+}02] Scott Ferson, Lev Ginzburg, Vladik Kreinovich, Luc Longpré, and Monica Aviles. Computing variance for interval data is NPhard. ACM SIGACT News, 33(2):108–118, June 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

http://www.cs.utep.edu/vladik/2002/tr02-13b.pdf.

Fenner:2017:GCP

[FGT17] Stephen Fenner, Rohit Gurjar, and Thomas Thierauf. Guest column: Parallel algorithms for perfect matching. ACM SIGACT News, 48(1):102–109, March 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Faliszewski:2006:OQT

[FH06] Piotr Faliszewski and Lane A. Hemaspaandra. Open questions in the theory of semifeasible computation. ACM SIGACT News, 37(1):47–65, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fagin:1995:RAK

[FHMV95] Ronald Fagin, Joseph Y. Halpern, Yoram Moses, and Moshe Y. Vardi. Reasoning about Knowledge. MIT Press, Cambridge, MA, USA, 1995. ISBN 0-262-06162-7 (hardcover). xiii + 477 pp. LCCN BD181 .R38 1995.

Fagin:2003:RAK

[FHMV03] Ronald Fagin, Joseph Y. Halpern, Yoram Moses, and Moshe Y. Vardi. Reasoning about Knowledge. MIT Press, Cambridge, MA, USA, 2003. ISBN 0-262-56200-6. xiii + 517 pp. LCCN BD181 .R38 2003.

Fich:1997:IIR

[Fic97] Faith Fich. Infrastructure issues related to theory of computing research. ACM SIGACT News, 28(3):94–99, September 1997. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fischer:1972:LCS

[Fis72] Michael J. Fischer. Lambda calculus schemata. ACM SIGACT News, 4(14):104–109, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fisher:2008:RPS

[Fis08] Len Fisher. Rock, Paper, Scissors: Game Theory in Everyday Life. Basic Books, New York, NY, USA, 2008. ISBN 0-465-00938-7. x + 265 pp. LCCN QA269 .F49 2008. URL http:

//www.loc.gov/catdir/enhancements/fy1003/2008025592-b.
html;   http://www.loc.gov/catdir/enhancements/fy1003/
2008025592-d.html;   http://www.loc.gov/catdir/toc/ecip0820/
2008025592.html.

Fitzsimons:2011:BRB

[Fit11] Joseph Fitzsimons. Book review: Algebraic Cryptanalysis, by Gregory V. Bard. ACM SIGACT News, 42(2):14–18, June 2011. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Bar09].

Farmer:1989:MAL

[FK89] W. M. Farmer and A. J. Kfoury. Minutes of the 4th annual LICS business meeting. ACM SIGACT News, 20(4):43–47, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Friedman:2008:MFC

[FKR08] Roy Friedman, Anne-Marie Kermarrec, and Michel Raynal. Modularity: a first class concept to address distributed systems. ACM SIGACT News, 39(2):91–110, June 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fortnow:1994:IVO

[FKW94] Lance Fortnow, Stuart Kurtz, and Duke Whang. The infinite version of an open communication complexity problem is independent of the axioms of set theory. ACM SIGACT News, 25(1): 87–89, March 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Findler:1983:NFE

[FL83] Nicholas V. Findler and Ron Lo. A note on the functional estimation of values of hidden variables: an extended module for expert systems. ACM SIGACT News, 15(2):78, Spring–Summer 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Flannery:2005:SRD

[Fla05] David Flannery. The square root of 2: a dialogue concerning a number and a sequence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-20220-X. x + 256 pp. LCCN QA247.5 .F536 2006.

Feitelson:2005:DLJ

[FM05] Dror G. Feitelson and Ahuva W. Mu’alem. On the definition of “on-line” in job scheduling problems. ACM SIGACT News, 36(1): 122–131, March 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Feigenbaum:2006:TTN

[FM06] Joan Feigenbaum and Michael Mitzenmacher. Towards a theory of networked computation. ACM SIGACT News, 37(4):22–26, December 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fasy:2008:BRB

[FM08] Brittany Terese Fasy and David L. Millman. Book review: Geometric Algebra for Computer Science, by Leo Dorst, Daniel Fontijne, and Stephen Mann (Morgan Kaufmann Publishers, 2007). ACM SIGACT News, 39(4):27–30, December 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [DFM07].

Fasy:2009:BRB

[FM09] Brittany Terese Fasy and David L. Millman. Book review: Higher Arithmetic: an Algorithmic Introduction to Number Theory, by H. M. Edwards (American Mathematical Society Student Mathematical Library Vol. 45, 2008). ACM SIGACT News, 40(2):38–41, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Edw08].

Fasy:2011:BRB

[FM11] Brittany Terese Fasy and David L. Millman. Book review: Geometric Folding Algorithms, by E. D. Demaine and J. O’Rourke. ACM SIGACT News, 42(1):43–46, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [DO07].

Fasy:2013:RBF

[FM13] Brittany Terese Fasy and David L. Millman. Review of How to Fold It by J. O’Rourke. ACM SIGACT News, 44(3):17–19, September 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fasy:2014:RDC

[FM14] Brittany Terese Fasy and David L. Millman. Review of Discrete and computational geometry by Satyan L. Devadoss and Joseph O’Rourke. ACM SIGACT News, 45(1):27–30, March 2014. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fasy:2015:RPA

[FM15] Brittany Terese Fasy and David L. Millman. Review of: Polyhedral and Algebraic Methods in Computational Geometry by Michael Joswig and Thorsten Theobald. ACM SIGACT News, 46(3):17– 20, September 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fischer:2017:DR

[FM17] Manuela Fischer and Yannic Maus. DISC 2017 review. ACM SIGACT News, 48(4):94–99, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fischer:1970:NIR

[FMOP70] M. Fischer, A. Meyer, P. O’Neil, and M. Paterson. A note on independence of a regularity-preserving operator. ACM SIGACT News, 2(4):3–7, 1970. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fortnow:1999:BRB

[For99] Lance Fortnow. Book review: Bounded Queries in Recursion Theory, by William A. Gasarch and Georgia A. Martin (Birkhäuser. Boston, Basel, Berlin, 1999). ACM SIGACT News, 30(3):13–15, September 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [GM99].

Fortnow:2004:BRB

[For04] Lance Fortnow. Book review: Theory of Semi-Feasible Algorithms, by Lane Hemaspaandra and Leen Torenvliet. Springer. ACM SIGACT News, 35(2):16–18, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [HT03].

Fortnow:2021:WHO

[For21] Lance Fortnow. Worlds to die harder for open oracle questions for the 21st century. ACM SIGACT News, 52(3):26–36, September 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3494656.
3494663.

Fotakis:2011:OIA

[Fot11] Dimitris Fotakis. Online and incremental algorithms for facility location. ACM SIGACT News, 42(1):97–131, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Frost:2004:PPP

[FPE04] Christopher Frost, Michael Peck, and David Evans. Pancakes, puzzles, and polynomials: cracking the cracker barrel. ACM SIGACT News, 35(1):80–84, March 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Franco:1984:DFIa

[Fra84a] John Franco. Duality, finite improvement and efficiently solved optimization problems. ACM SIGACT News, 15(4):62, Winter– Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Franco:1984:DFIb

[Fra84b] John Franco. Duality, finite improvement and efficiently solved optimization problems. ACM SIGACT News, 16(1):62, Winter– Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Franco:1985:SPR

[Fra85] J. Franco. Sensitivity of probabilistic results on algorithms for NPcomplete problems to input distributions. ACM SIGACT News, 17 (1):40–59, June 1985. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fredriksson:2001:CSP

[Fre01] Einar H. Fredriksson, editor. A Century of Science Publishing: a Collection of Essays. IOS Press, Amsterdam, The Netherlands, 2001. ISBN 1-58603-148-1. viii + 312 pp. LCCN Z286.S4 C46 2001.

Friedler:2010:BRB

[Fri10] Sorelle A. Friedler. Book review: Change is Possible: Stories of Women and Minorities in Mathematics, by Patricia Clark Kenschaft, published by AMS, 2005 212 pages, softcover. ACM SIGACT News, 41(2):47–50, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Ken05].

Friedler:2011:BRB

[Fri11] Sorelle A. Friedler. Book review: Pioneering Women in American Mathematics: the pre-1940 PhD’s, by Judy Green and Jeanne LaDuke. ACM SIGACT News, 42(2):37–41, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [GL09].

Fekete:2017:AFP

[FRRS17] Sándor P. Fekete, Andréa Richa, Kay Römer, and Christian Scheideler. Algorithmic foundations of programmable matter Dagstuhl Seminar 16271. ACM SIGACT News, 48(2):87–94, June 2017. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fakcharoenphol:2004:AMT

[FRT04] Jittat Fakcharoenphol, Satish Rao, and Kunal Talwar. Approximating metrics by tree metrics. ACM SIGACT News, 35(2):60–70, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Flajolet:2009:AC

[FS09] Philippe Flajolet and Robert Sedgewick. Analytic Combinatorics. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-89806-4. xiii + 810 pp. LCCN QA164 .F57 2009. URL http:

//www.loc.gov/catdir/enhancements/fy0916/2009280635-b.
html;   http://www.loc.gov/catdir/enhancements/fy0916/
2009280635-d.html;   http://www.loc.gov/catdir/enhancements/
fy0916/2009280635-t.html.

Forbes:2015:CTC

[FS15] Michael A. Forbes and Amir Shpilka. Complexity theory column 88: Challenges in polynomial factorization. ACM SIGACT News, 46(4):32–49, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Foerster:2019:SRD

[FS19] Klaus-Tycho Foerster and Stefan Schmid. Survey of reconfigurable data center networks: Enablers, algorithms, complexity. ACM SIGACT News, 50(2):62–79, June 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Fulton:2010:BRB

[Ful10] Ben Fulton. Book review: Introduction to Modern Cryptography, by Jonathan Katz and Yehuda Lindell, Publisher: Chapman & Hall-CRC 2008 1-58488-551-3. ACM SIGACT News, 41(4):44–47, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [KL08].

Fulton:2012:RBS

[Ful12] Ben Fulton. Review of Introduction to Scheduling, by Yves Robert and Frédéric Vivien. ACM SIGACT News, 43(4):41–43, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Furia:2011:BRB

[Fur11] Carlo A. Furia. Book review: The Calculus of Computation: Decision Procedures with Applications to Verification, by Aaron R. Bradley and Zohar Manna. ACM SIGACT News, 42(1):32–35, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BM07].

Gabow:1975:HGN

[Gab75] Harold N. Gabow. How to gracefully number certain symmetric trees. ACM SIGACT News, 7(4):33–36, November 1975. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gabizon:2010:DEW

[Gab10] Ariel Gabizon. Deterministic Extraction from Weak Random Sources. Monographs in theoretical computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-14902-2, 3-642-14903-0 (e-book). 148 pp. LCCN QA564 .G33 2011.

Galil:1974:SDE

[Gal74] Zvi Galil. On some direct encodings of nondeterministic Turing machines operating in polynomial time into p-complete problems. ACM SIGACT News, 6(1):19–24, January 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Galil:1975:CLA

[Gal75] Zvi Galil. On converting on-line algorithms into real-time and on real-time algorithms for string-matching and palindrome recognition. ACM SIGACT News, 7(4):26–30, November 1975. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Galil:1978:KTB

[Gal78] Zvi Galil. Killing two birds with one stone. ACM SIGACT News, 10(1):22–24, Spring 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Galil:1992:RC

[Gal92] Zvi Galil. Renato Capocelli. ACM SIGACT News, 23(2):104, Spring 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Galle:2013:RBI

[Gal13] Matthias Gallé. Review of Grammatical Inference: Learning Automata and Grammars by Colin de la Higuera. ACM SIGACT News, 44(2):12–14, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Galle:2014:RBR

[Gal14] Matthias Gallé. Review of Bayesian reasoning and machine learning by David Barber. ACM SIGACT News, 45(2):27–29, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Galle:2015:RAS

[Gal15] Matthias Gallé. Review of: Algorithms on Strings by Maxime Crochemore, Christophe Hancart and Thierry Lecroq. ACM SIGACT News, 46(3):15–16, September 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Galinkin:2021:RFC

[Gal21] Erick Galinkin. Review of The Foundations of Computability Theory (Second Edition) by Borut Robic. ACM SIGACT News, 52 (2):7–9, June 2021. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3471469.3471472.

Galinkin:2022:RGU

[Gal22] Erick Galinkin. Review of The Geometry of Uncertainty by Fabio Cuzzolin. ACM SIGACT News, 53(2):11–14, June 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3544979.3544983.

Gardner:2006:ATV

[Gar06] Martin Gardner. Aha! A Two Volume Collection: Aha! Gotcha, Aha! Insight. Spectrum series. Mathematical Association of America, Washington, DC, USA, 2006. ISBN 0-88385-551-8. xvii + 164 (vol. 1), ix + 181 (vol. 2) pp. LCCN QA95 .G37 2006.

Gardner:2008:OES

[Gar08] Martin Gardner. Origami, Eleusis, and the Soma cube: Martin Gardner’s Mathematical Diversions, volume 2 of The new Martin Gardner mathematical library. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-75610-3 (hardcover), 0-521-73524-6 (paperback). xi + 234 pp. LCCN QA95 .G2975 2008. URL http:

//www.loc.gov/catdir/enhancements/fy0834/2008012534-b.
html;   http://www.loc.gov/catdir/enhancements/fy0834/
2008012534-d.html;   http://www.loc.gov/catdir/enhancements/
fy0834/2008012534-t.html.

Gasarch:1994:BRB

[Gas94] William Gasarch. Book review: Finite Automata, Formal Logic, and Circuit Complexity. By Howard Straubing. (Birkhäuser. 1994. xii + 226pp. ISBN 0-8176-3719-2. $39.50). ACM SIGACT News, 25(3):28–32, September 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Str94].

Gasarch:1997:BRB

[Gas97a] William Gasarch. Book review: An Introduction to Kolmogorov Complexity and its Applications Second Edition, 1997 by Ming Li and Paul Vitanyi (Springer (Graduate Text Series)). ACM SIGACT News, 28(3):37–40, September 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [LV93, LV97, LV08].

Gasarch:1997:BRC

[Gas97b] William Gasarch. The book review column. ACM SIGACT News, 28(4):16, December 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:1998:BRCa

[Gas98a] William Gasarch. The book review column. ACM SIGACT News, 29(1):2, March 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:1998:BRCb

[Gas98b] William Gasarch. The book review column. ACM SIGACT News, 29(2):17, June 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:1998:BRCc

[Gas98c] William Gasarch. The book review column. ACM SIGACT News, 29(3):43, September 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:1998:BRCd

[Gas98d] William Gasarch. The book review column. ACM SIGACT News, 29(4):2–4, December 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:1999:BRBa

[Gas99a] William Gasarch. Book review: Algorithms and Theory of Computation Handbook, edited by Mikhail Atallah. ACM SIGACT News, 30(2):3, June 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Ata99, AB09b, AB10].

Gasarch:1999:BRBb

[Gas99b] William Gasarch. Book review: Handbook of Combinatorics (in two volumes), edited by R. L. Graham, M. Grötschel, L. Lovász. ACM SIGACT News, 30(2):7, June 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [GGL95].

Gasarch:1999:BRCa

[Gas99c] William Gasarch. The book review column. ACM SIGACT News, 30(1):2, March 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:1999:BRCb

[Gas99d] William Gasarch. The book review column. ACM SIGACT News, 30(2):2, June 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:1999:BRCd

[Gas99e] William Gasarch. The book review column. ACM SIGACT News, 30(4):3, December 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2000:BRB

[Gas00a] William Gasarch. Book review: Indiscrete Thoughts, by Gina-Carlo Rota (Birkhäuser, 1996). ACM SIGACT News, 31(2):9–11, June 1, 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [RP97].

Gasarch:2000:BRCa

[Gas00b] William Gasarch. The book review column. ACM SIGACT News, 31(1):2–3, March 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2000:BRCb

[Gas00c] William Gasarch. The book review column. ACM SIGACT News, 31(2):2, June 1, 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2000:BRCc

[Gas00d] William Gasarch. The book review column. ACM SIGACT News, 31(3):10, September 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2000:BRCd

[Gas00e] William Gasarch. The book review column. ACM SIGACT News, 31(4):3–4, December 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2000:RTB

[Gas00f] William Gasarch. Reviews of THREE books on fair division of resources. ACM SIGACT News, 31(4):4–10, December 2000. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2001:BRBb

[Gas01a] William Gasarch. Book review: Proofs and Refutations, by Imre Lakatos. ACM SIGACT News, 32(4):6–8, December 2001. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Lak76].

Gasarch:2001:BRBa

[Gas01b] William Gasarch. Book review: The Codebreakers: the Story of Secret Writing, by David Kahn. Scribner. ACM SIGACT News, 32(2):5–6, June 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Kah67a, Kah67b, Kah74, Kah96].

Gasarch:2001:BRCa

[Gas01c] William Gasarch. The book review column. ACM SIGACT News, 32(1):2–3, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2001:BRCb

[Gas01d] William Gasarch. The book review column. ACM SIGACT News, 32(2):3–4, June 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2001:BRCc

[Gas01e] William Gasarch. The book review column. ACM SIGACT News, 32(4):2–5, December 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2002:BML

[Gas02a] William Gasarch. Bioinformatics: the machine learning approach. ACM SIGACT News, 33(4):5–8, December 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2002:BRB

[Gas02b] William Gasarch. Book review: Calculated Bets. ACM SIGACT News, 33(2):19–20, June 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Ski01].

Gasarch:2002:BRCa

[Gas02c] William Gasarch. The book review column. ACM SIGACT News, 33(1):3–20, March 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2002:BRCb

[Gas02d] William Gasarch. The book review column. ACM SIGACT News, 33(2):2–9, June 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2002:BRCc

[Gas02e] William Gasarch. The book review column. ACM SIGACT News, 33(3):6, September 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2002:BRCd

[Gas02f] William Gasarch. The book review column. ACM SIGACT News, 33(4):4–5, December 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2003:BRCa

[Gas03a] William Gasarch. The book review column. ACM SIGACT News, 34(1):3, March 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2003:BRCb

[Gas03b] William Gasarch. The book review column. ACM SIGACT News, 34(2):2, June 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2003:BRCc

[Gas03c] William Gasarch. The book review column. ACM SIGACT News, 34(3):3–4, September 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2003:BRCd

[Gas03d] William Gasarch. The book review column. ACM SIGACT News, 34(4):14–15, December 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2004:BRB

[Gas04a] William I. Gasarch. Book review: Handbook of Graph Theory, edited by Gross and Yellen. CRC, 2004. ACM SIGACT News, 35(3):5–8, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [GY04].

Gasarch:2004:BRCa

[Gas04b] William I. Gasarch. The book review column. ACM SIGACT News, 35(1):3–4, March 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2004:BRCb

[Gas04c] William I. Gasarch. The book review column. ACM SIGACT News, 35(2):3–4, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2004:BRCc

[Gas04d] William I. Gasarch. The book review column. ACM SIGACT News, 35(4):4–5, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2005:BRBb

[Gas05a] William I. Gasarch. Book review: Cryptological Mathematics, by Robert Lewand; MAA, 2000, $33.95, Softcover. ACM SIGACT News, 36(2):4–7, June 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Lew00].

Gasarch:2005:BRBa

[Gas05b] William I. Gasarch. Book review: Proofs that Really Count: The Art of Combinatorial Proof, by Arthur T. Benjamin and Jennifer J. Quinn; MAA, 2003. ACM SIGACT News, 36(1):12–14, March 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BQ03].

Gasarch:2005:BRCa

[Gas05c] William I. Gasarch. The book review column. ACM SIGACT News, 36(2):3–4, June 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2005:BRCb

[Gas05d] William I. Gasarch. The book review column. ACM SIGACT News, 36(3):4–5, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2005:BRCc

[Gas05e] William I. Gasarch. The book review column. ACM SIGACT News, 36(4):4–5, December 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2006:BRCa

[Gas06a] William I. Gasarch. The book review column. ACM SIGACT News, 37(1):9–11, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2006:BRCb

[Gas06b] William I. Gasarch. The book review column. ACM SIGACT News, 37(2):10–12, June 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2006:BRCc

[Gas06c] William I. Gasarch. The book review column. ACM SIGACT News, 37(3):14–17, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2006:BRCd

[Gas06d] William I. Gasarch. The book review column. ACM SIGACT News, 37(4):27–29, December 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2006:JRB

[Gas06e] William I. Gasarch. A joint review of Reality Conditions: Short Mathematical Fiction, by Alex Kasman, MAA 2005; Numb3rs, TV show. CBS, Free. Currently running Fridays at 10:00PM; Mathematical Apocryphia: Stories and Anecdotes of Mathematicians and the Mathematical, by Steven Kranz, MAA, 2002; Mathematical Apocryphia Redux: More Stories and Anecdotes of Mathematicians and the Mathematical, by Steven Kranz, MAA, 1999. ACM SIGACT News, 37(3):17–19, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Kas05].

Gasarch:2007:JBN

[Gas07a] William Gasarch. The joys of being an NSF program director. ACM SIGACT News, 38(1):7–8, March 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2007:BRBb

[Gas07b] William I. Gasarch. Book review: A Century of Scientific Publishing: A collection of essays, edited by Fredriksson, IOS Press. ACM SIGACT News, 38(4):23–24, December 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Fre01].

Gasarch:2007:BRBa

[Gas07c] William I. Gasarch. Book review: Excellence Without a Soul: How a Great University Forgot Education, by Harry Lewis, Public Affairs, 290 pages. ACM SIGACT News, 38(1):9–13, March 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Lew06].

Gasarch:2007:BRBc

[Gas07d] William I. Gasarch. Book review: Research Problems in Discrete Geometry, by Brass, Moser, Pach, Springer-Verlag. ACM SIGACT News, 38(4):31–34, December 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BMP05].

Gasarch:2007:BRCa

[Gas07e] William I. Gasarch. The book review column. ACM SIGACT News, 38(2):8–10, June 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2007:BRCb

[Gas07f] William I. Gasarch. The book review column. ACM SIGACT News, 38(3):14–16, September 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [MU05].

Gasarch:2007:BRCc

[Gas07g] William I. Gasarch. The book review column. ACM SIGACT News, 38(4):16–18, December 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Wil06, Fre01].

Gasarch:2007:JRB

[Gas07h] William I. Gasarch. Joint review of Three Blogs by Theorists: Computational Complexity (weblog.fortnow.com), by Lance Fortnow, Shtetl-Optimized (www.scottaaronson.com/blog/), by Scott Aaronson, In theory (in-theory.blogspot.com) by Luca Trevisan. ACM SIGACT News, 38(2):23–25, June 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2008:BRCc

[Gas08a] William Gasarch. The book review column. ACM SIGACT News, 39(4):15–17, December 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2008:TRG

[Gas08b] William Gasarch. Theory research at Google. ACM SIGACT News, 39(2):10–28, June 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2008:BRCa

[Gas08c] William I. Gasarch. The book review column. ACM SIGACT News, 39(1):9–12, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2008:BRCb

[Gas08d] William I. Gasarch. The book review column. ACM SIGACT News, 39(2):29–31, June 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2009:BRBa

[Gas09a] William Gasarch. Book review: Blown to Bits: Your Life, Liberty, and Happiness After the Digital Explosion, by Hal Abelson, Ken Ledeen, and Harry Lewis (Addison Wesley, 2008). ACM SIGACT News, 40(1):10–13, March 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [ALL08].

Gasarch:2009:BRBb

[Gas09b] William Gasarch. Book review: Rock, Paper, Scissors: Game Theory for Everyday Life, by Len Fisher (Basic Books, 2008). ACM SIGACT News, 40(1):22–23, March 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Fis08].

Gasarch:2009:BRBc

[Gas09c] William Gasarch. Book review: The Mathematical Coloring Book: Mathematics of Coloring and the Colorful Life of Its Creators, by Alexander Soifer. ACM SIGACT News, 40(3):24–31, September 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Soi09].

Gasarch:2009:BRCa

[Gas09d] William Gasarch. The book review column. ACM SIGACT News, 40(1):8–10, March 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2009:BRCb

[Gas09e] William Gasarch. The book review column. ACM SIGACT News, 40(2):10–13, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2009:BRCc

[Gas09f] William Gasarch. The book review column. ACM SIGACT News, 40(3):21–24, September 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2009:BRCd

[Gas09g] William Gasarch. The book review column. ACM SIGACT News, 40(4):8–10, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2009:JRP

[Gas09h] William Gasarch. Joint review of Professor Stewart’s Cabinet of Mathematical Curiosities, by Ian Stewart and Five Minute Mathematics, by Ehrhard Behrends and Aha Gotcha!-Aha Insight! by Martin Gardner and Origami, Eleusis, and the Soma Cube, by Martin Gardner and Hexaexagons, Probability Paradoxes, and the Tower of Hanoi, by Martin Gardner and Group Theory in the Bedroom and Other Mathematical Diversions, by Brian Hayes. ACM SIGACT News, 40(3):32–37, September 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Gar06, Beh08, Gar08, Hay08, Ste09].

Gasarch:2010:BRBd

[Gas10a] William Gasarch. Book review: Games of No Chance (1998, edited by Richard Nowakowski) and More Games of No Chance (2002, edited by Richard Nowakowski) and Games of No Chance III (2009, edited by Michael Albert and Richard Nowakowski published by Cambridge Press). ACM SIGACT News, 41(3):26–28, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [AN09, Now02].

Gasarch:2010:BRBa

[Gas10b] William Gasarch. Book review: Logicomix Text, by Apostolos Doxiadis and Christos Papadimitriou Art by Alecos Papadatos and Annie di Donna, published by Bloomsbury, 2009 314 pages, softcover. comic book! ACM SIGACT News, 41(2):26–28, June 2010. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [DPP09].

Gasarch:2010:BRBe

[Gas10c] William Gasarch. Book review: Mathematical Treks: from Surreal Numbers to Magic Circles, by Ivars Peterson published by the MAA, 2002 170 pages. ACM SIGACT News, 41(3):29–30, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Pet02].

Gasarch:2010:BRBc

[Gas10d] William Gasarch. Book review: Random Curves: Journeys of a Mathematician, by Neal Koblitz, published by Springer, 2008, 390 pages. ACM SIGACT News, 41(3):18–25, September 2010. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Kob08].

Gasarch:2010:BRBb

[Gas10e] William Gasarch. Book review: Riot at the Calc Exam and Other Mathematically Bent Stories, by Colin Adams, published by the AMS, 2009 271 pages, softcover and The Great Debate: Which Is the Best Number?, by Colin Adams VS Thomas Garrity, moderated by Edward Burger, published by the MAA, 2006 and The United States of Mathematics Presidential Debate, by Colin Adams VS Thomas Garrity, moderated by Edward Burger, published by the MAA, 2009. ACM SIGACT News, 41(2):50–51, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Ada09a].

Gasarch:2010:BRBf

[Gas10f] William Gasarch. Book review: The P = NP question and Gödel’s lost letter, by Richard J. Lipton, Springer, 2010. ACM SIGACT News, 41(4):25–29, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Lip10].

Gasarch:2010:BRBg

[Gas10g] William Gasarch. Book review: The Pea and the Sun: a Mathematical Paradox, by Leonard Wapner Published by A. K. Peters, 2005. ACM SIGACT News, 41(4):30–32, December 2010. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Wap05].

Gasarch:2010:BRCa

[Gas10h] William Gasarch. The book review column. ACM SIGACT News, 41(1):10, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2010:BRCb

[Gas10i] William Gasarch. The book review column. ACM SIGACT News, 41(2):7–10, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2010:BRCc

[Gas10j] William Gasarch. The book review column. ACM SIGACT News, 41(3):15–17, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2010:BRCd

[Gas10k] William Gasarch. The book review column. ACM SIGACT News, 41(4):12–15, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2011:BRBa

[Gas11a] William Gasarch. Book review: Dude, Can You Count?: Stories, Challenges, and Adventures in Mathematics, by Christian Constanda. ACM SIGACT News, 42(1):49–54, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Con09].

Gasarch:2011:BRBb

[Gas11b] William Gasarch. Book review: Those Fascinating Numbers, by Jean-Marie De Koninck. ACM SIGACT News, 42(2):24–27, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [dK09].

Gasarch:2011:BRCa

[Gas11c] William Gasarch. The book review column. ACM SIGACT News, 42(1):12–15, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2011:BRCb

[Gas11d] William Gasarch. The book review column. ACM SIGACT News, 42(2):7–10, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2011:BRCc

[Gas11e] William Gasarch. The book review column. ACM SIGACT News, 42(3):14–15, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2011:BRCd

[Gas11f] William Gasarch. The book review column. ACM SIGACT News, 42(4):15–16, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2011:JRB

[Gas11g] William Gasarch. Joint review of The Mathemagician and Pied Puzzler: a Collection in Tribute to Martin Gardner. ACM SIGACT News, 42(1):16–22, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BR99].

Gasarch:2012:BRB

[Gas12a] William Gasarch. Book review: An Introduction to the History of Algebra Solving Equations from Mesopotamian Times to the Renaissance, by Jacques Sesiano. ACM SIGACT News, 43(1):39, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Ses09].

Gasarch:2012:BRCa

[Gas12b] William Gasarch. The book review column. ACM SIGACT News, 43(1):7–9, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2012:BRCb

[Gas12c] William Gasarch. The book review column. ACM SIGACT News, 43(2):6–8, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2012:BRCc

[Gas12d] William Gasarch. The book review column. ACM SIGACT News, 43(3):15–18, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2012:RBG

[Gas12e] William Gasarch. Review of Combinatorial Games: Tic-Tac-Toe Theory, by Jozsef Beck. ACM SIGACT News, 43(3):19–21, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2012:GCS

[Gas12f] William I. Gasarch. Guest Column: the second P =? NP poll. ACM SIGACT News, 43(2):53–77, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2013:BRCa

[Gas13a] William Gasarch. The book review column. ACM SIGACT News, 44(1):7–9, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2013:BRC

[Gas13b] William Gasarch. The book review column. ACM SIGACT News, 44(2):7–9, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2013:JRH

[Gas13c] William Gasarch. Joint review of the honor class: Hilbert’s problems and their solver by Ben Yandell and: Mathematical developments arising from Hilbert’s problems edited by Felix Browder. ACM SIGACT News, 44(4):18–24, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2013:RAP

[Gas13d] William Gasarch. Review of Algorithmic puzzles by Anany Levitin and Maria Levitin. ACM SIGACT News, 44(4):47–48, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2013:RBF

[Gas13e] William Gasarch. Review of Boolean Function Complexity: Advances and Frontiers by Stasys Jukna. ACM SIGACT News, 44 (2):39–41, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2013:RBC

[Gas13f] William Gasarch. Review of Theoretical Computer Science: Introduction to Automata, Computability, Complexity, Algorithmics, Randomization, Communication, and Cryptography by Juraj Hromkovic. ACM SIGACT News, 44(3):7–8, September 2013. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2014:BRCa

[Gas14a] William Gasarch. The book review column. ACM SIGACT News, 45(2):7–9, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2014:BRCb

[Gas14b] William Gasarch. The book review column. ACM SIGACT News, 45(3):7–9, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2014:BRCc

[Gas14c] William Gasarch. The book review column. ACM SIGACT News, 45(4):14–16, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2014:JRF

[Gas14d] William Gasarch. Joint reviews of four articles. ACM SIGACT News, 45(2):22–27, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2014:RCP

[Gas14e] William Gasarch. Review of Companion to the papers of Donald Knuth by Donald E. Knuth. ACM SIGACT News, 45(2):19–21, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2014:RPP

[Gas14f] William Gasarch. Review of People, problems, and proofs by Richard Lipton and Ken Regan. ACM SIGACT News, 45(2):36–39, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2014:RSPa

[Gas14g] William Gasarch. Review of Selected papers on fun & games by Donald E. Knuth. ACM SIGACT News, 45(2):17–19, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2014:RED

[Gas14h] William Gasarch. Review of The Erdős distance problem by Julia Garibaldi, Alex Iosevich and Steven Senger. ACM SIGACT News, 45(3):13–14, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2014:RSPb

[Gas14i] William Gasarch. Review of The Satisfiability Problem: Algorithms and Analyses by Uwe Schöning and Jacobo Torán. ACM SIGACT News, 45(4):45–47, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2015:BRC

[Gas15a] William Gasarch. The book review column. ACM SIGACT News, 46(1):8–9, March 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2015:RWT

[Gas15b] William Gasarch. Review of: A Walk Through Combinatorics by Miklós Bóna. ACM SIGACT News, 46(1):13–14, March 2015. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2015:RAB

[Gas15c] William Gasarch. Review of: Algorithmic Barriers Falling: P=NP? by Donald E. Knuth and Edgar G. Daylight and The Essential Knuth by Donald E. Knuth and Edgar G. Daylight. ACM SIGACT News, 46(2):21–22, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2015:RIH

[Gas15d] William Gasarch. Review of: Infinitesimal: How a dangerous mathematical theory shaped the modern world by Amir Alexander. ACM SIGACT News, 46(2):16–18, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2015:RLM

[Gas15e] William Gasarch. Review of: Love and Math: The Heart of Hidden Reality by Edward Frenkel. ACM SIGACT News, 46(2):23–24, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2015:RMGa

[Gas15f] William Gasarch. Review of: Martin Gardner in the Twenty-First Century: Edited by Michael Henle and Brian Hopkins. ACM SIGACT News, 46(2):19–20, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2015:RMGb

[Gas15g] William Gasarch. Review of Mathematics Galore: The First Five Years of the St. Marks’ Institute of Mathematics by James Tanton. ACM SIGACT News, 46(4):9–11, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2015:RSR

[Gas15h] William Gasarch. Review of: Structure and Randomness: Pages from Year One of a Mathematical Blog by Terence Tao. ACM SIGACT News, 46(2):25–27, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2015:RCP

[Gas15i] William Gasarch. Review of: The Cult of Pythagoras: Math and Myths by Alberto A. Martinez. ACM SIGACT News, 46(2):13–15, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2016:OPA

[Gas16a] William Gasarch. Open problems about grid coloring and the complexity of grid colorings. ACM SIGACT News, 47(3):39–43, September 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2016:RAJ

[Gas16b] William Gasarch. Review of: Asymptopia by Joel Spencer and Laura Florescu. ACM SIGACT News, 47(2):12–13, June 2016. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2016:RRT

[Gas16c] William Gasarch. Review of: Ramsey Theory over the Integers (Second Edition) by Bruce M. Landman and Aaron Robertson. ACM SIGACT News, 47(2):14–17, June 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2016:RST

[Gas16d] William Gasarch. Review of Slicing the Truth: On the Computability Theoretic and Reverse Mathematical Analysis of Combinatorial Principles by Denis Hirschfeldt. ACM SIGACT News, 47(1): 21–24, March 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2016:RJF

[Gas16e] William Gasarch. Review of: The Joy of Factoring by Samuel Wagstaff. ACM SIGACT News, 47(2):10–11, June 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2016:RSS

[Gas16f] William Gasarch. Review of The Scholar and the State: In Search of Van der Waerden by Alexander Soifer. ACM SIGACT News, 47(1):25–28, March 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2016:RTC

[Gas16g] William Gasarch. Review of: Turing Computability: Theory and Applications by Robert Soare. ACM SIGACT News, 47(4):6–8, December 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2016:RWC

[Gas16h] William Gasarch. Review of What is College For?: The Public Purpose of Higher Education by Ellen Condliffe Lagemann and Harry Lewis. ACM SIGACT News, 47(1):16–20, March 2016. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2017:OPCa

[Gas17a] William Gasarch. Open problems column. ACM SIGACT News, 48(1):38, March 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2017:OPCb

[Gas17b] William Gasarch. Open problems column. ACM SIGACT News, 48(2):34–39, June 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2017:OPCd

[Gas17c] William Gasarch. Open problems column. ACM SIGACT News, 48(4):35–41, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2017:RRT

[Gas17d] William Gasarch. Review of Ramsey Theory for Discrete Structures by Hans Jürgen Prömel. ACM SIGACT News, 48(4):18–21, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2018:OPCa

[Gas18a] William Gasarch. Open problems column. ACM SIGACT News, 49(1):40–54, March 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2018:OPCb

[Gas18b] William Gasarch. Open problems column. ACM SIGACT News, 49(2):29–31, June 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2018:OPCc

[Gas18c] William Gasarch. Open problems column. ACM SIGACT News, 49(3):34, September 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2018:OPCd

[Gas18d] William Gasarch. Open problems column. ACM SIGACT News, 49(4):28, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2019:OPCa

[Gas19a] William Gasarch. Open problems column. ACM SIGACT News, 50(1):28–34, March 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2019:OPCb

[Gas19b] William Gasarch. Open problems column. ACM SIGACT News, 50(2):28, June 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2019:OPCc

[Gas19c] William Gasarch. Open problems column. ACM SIGACT News, 50(3):24, September 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2019:OPC

[Gas19d] William Gasarch. Open problems column. ACM SIGACT News, 50(4):26–30, December 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3374857.3374863.

Gasarch:2019:RIR

[Gas19e] William Gasarch. Review of An Introduction to Ramsey Theory: Fast Functions, Infinity, and Metamathematics by Matthew Katz and Jan Reimann. ACM SIGACT News, 50(2):9–11, June 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2019:RFM

[Gas19f] William Gasarch. Review of Factor Man by Matt Ginsberg. ACM SIGACT News, 50(4):7–8, December 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3374857.3374860.

Gasarch:2019:RQQ

[Gas19g] William Gasarch. Review of Q is for Quantum by Terry Rudolph. ACM SIGACT News, 50(2):6–8, June 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2019:GCT

[Gas19h] William I. Gasarch. Guest column: The third P =? NP poll. ACM SIGACT News, 50(1):38–59, March 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2020:OPCb

[Gas20a] William Gasarch. Open problems column. ACM SIGACT News, 51(2):27–35, June 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3406678.3406685.

Gasarch:2020:OPCc

[Gas20b] William Gasarch. Open problems column. ACM SIGACT News, 51(3):31, September 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3427361.3427368.

Gasarch:2020:OPCd

[Gas20c] William Gasarch. Open problems column edited by William Gasarch. This issue’s column! ACM SIGACT News, 51(4):30– 46, December 2020. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3444815.3444822.

Gasarch:2020:RED

[Gas20d] William Gasarch. Review of Essential Discrete Mathematics for Computer Science by Harry Lewis and Rachel Zax. ACM SIGACT News, 51(2):6–8, June 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3406678.3406681.

Gasarch:2020:RTC

[Gas20e] William Gasarch. Review of Theorems of the 21st Century: Volume I. ACM SIGACT News, 51(3):14–19, September 2020. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

URL https://dl.acm.org/doi/10.1145/3427361.3427366.

Gasarch:2020:RWC

[Gas20f] William Gasarch. Review of What Can Be Computed: a Practical Guide to the Theory of Computation by John MacCormick. ACM SIGACT News, 51(1):9–11, March 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.

org/doi/10.1145/3388392.3388395.

Gasarch:2021:HTP

[Gas21a] William Gasarch. Hilbert’s Tenth Problem: Refinements and variants. ACM SIGACT News, 52(2):36–44, June 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3471469.3471476.

Gasarch:2021:OPCb

[Gas21b] William Gasarch. Open problems column. ACM SIGACT News, 52(3):25, September 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3494656.3494662.

Gasarch:2021:RBB

[Gas21c] William Gasarch. Review of Blown to Bits: Your Life, Liberty, and Happiness after the Digital Explosion by Hal Abelson, Ken Ledeen, Harry Lewis, and Wendy Seltzer. ACM SIGACT News, 52 (2):18–23, June 2021. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3471469.3471474.

Gasarch:2021:RIC

[Gas21d] William Gasarch. Review of Ideas that Created the Future: Classic Papers of Computer Science, Edited by Harry Lewis. ACM SIGACT News, 52(2):10–17, June 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.

org/doi/10.1145/3471469.3471473.

Gasarch:2022:OPCa

[Gas22a] William Gasarch. Open problems column. ACM SIGACT News, 53(2):26, June 2022. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3544979.3544985.

Gasarch:2022:OPCb

[Gas22b] William Gasarch. Open problems column. ACM SIGACT News, 53(3):36–40, September 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3561064.3561070.

Gasarch:2022:OPCc

[Gas22c] William Gasarch. Open problems column. ACM SIGACT News, 53(4):11–31, December 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3577971.3577975.

Gasarch:2022:RMR

[Gas22d] William Gasarch. Review of A Map that Reflects the Territory: Essays by the LessWrong Community Author: LessWrong. ACM SIGACT News, 53(1):13–24, March 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://

dl.acm.org/doi/10.1145/3532737.3532741.

Gasarch:2022:RTI

[Gas22e] William Gasarch. Review of Tales of Impossibility: The 2000-Year Quest to Solve the Mathematical Problems of Antiquity Author: David Richeson. ACM SIGACT News, 53(1):9–12, March 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3532737.
3532740.

Gasarch:2022:REC

[Gas22f] William Gasarch. Review of “The Engines of Cognition: Essays by the Less Wrong Community by Less Wrong Less Wrong Press, 720 pages, Year: 2019 $30.00”. ACM SIGACT News, 53(3):6– 16, September 2022. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3561064.3561066.

Gasarch:2023:JRM

[Gas23a] William Gasarch. Joint review of The Mathematics of Various Entertaining Subjects: Research in Recreational Math and The Mathematics of Various Entertaining Subjects: Vol. 2. Research in Games, Graphs, Counting, and Complexity and The Mathematics of Various Entertaining Subjects Vol. 3. The Magic of Mathematics. ACM SIGACT News, 54(1):9–14, March 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3586165.3586168.

Gasarch:2023:OPCa

[Gas23b] William Gasarch. Open problems column. ACM SIGACT News, 54(2):26–42, June 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3604943.3604947.

Gasarch:2023:OPCb

[Gas23c] William Gasarch. Open problems column. ACM SIGACT News, 54(3):33–41, September 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3623800.3623804.

Gasarch:2023:OPCc

[Gas23d] William Gasarch. Open problems column. ACM SIGACT News, 54(4):41–52, December 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3639528.3639532.

Gasarch:2024:OPC

[Gas24] William Gasarch. Open problems column. ACM SIGACT News, 55(1):32–51, March 2024. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3654780.3654784.

Gavoille:2001:RDN

[Gav01] Cyril Gavoille. Routing in distributed networks: Overview and open problems. ACM SIGACT News, 32(1):36–52, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2020:OPCa

[GB20] William Gasarch and Aarav Bajaj. Open problems column. ACM SIGACT News, 51(1):27–36, March 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://

dl.acm.org/doi/10.1145/3388392.3388398.

Gowers:2008:PCM

[GBGL08] Timothy Gowers, June Barrow-Green, and Imre Leader, editors. The Princeton Companion to Mathematics. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-11880-9 (hardcover). xx + 1034 pp. LCCN QA11.2 .P745 2008. URL http://www.loc.

gov/catdir/toc/ecip0818/2008020450.html.

Gerety:1986:TCT

[GC86] C. Gerety and P. Cull. Time complexity of the Towers of Hanoi problem. ACM SIGACT News, 18(1):80–87, Summer 1986. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Geerts:2017:DTC

[Gee17] Floris Geerts. Database theory column report on PODS 2017. ACM SIGACT News, 48(4):75–77, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Georgiou:2009:GTA

[Geo09] Chryssis Georgiou. Game-theoretic aspects of distributed computing. ACM SIGACT News, 40(4):89–92, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2017:OPCc

[GFW17] William Gasarch, Brittany Terese Fasy, and Bei Wang. Open problems in computational topology. ACM SIGACT News, 48(3):32– 36, September 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Galil:1986:ISM

[GG86] Z. Galil and R. Giancarlo. Improved string matching with k mismatches. ACM SIGACT News, 17(4):52–54, Spring 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gassama:2021:RSC

[GG21] Abdulai Gassama and Frederic Green. Review of A Short Course in Computational Geometry and Topology by Herbert Edelsbrunner. ACM SIGACT News, 52(4):11–14, December 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3510382.3510386.

Graham:1995:HC

[GGL95] Ronald L. Graham, Martin Grötschel, and László Lovász, editors. Handbook of Combinatorics. Elsevier, Amsterdam, The Netherlands, 1995. ISBN 0-444-88002-X (set: Elsevier), 0-262-07169-X (set: MIT Press), 0-444-82346-8 (v. 1.: Elsevier), 0-262-07170-3 (v. l.: MIT), 0-444-82351-4 (v. 2.: Elsevier), 0-262-07171-1 (v. 2.: MIT). cii + 2198 (2 volumes) pp. LCCN QA164 .H33 1995.

Glasser:2000:MPC

[GH00] Christian Glaßer and Lane A. Hemaspaandra. A moment of perfect clarity II: consequences of sparse sets hard for NP with respect to weak reductions. ACM SIGACT News, 31(4):39–51, December 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gillman:2009:MCC

[GH09] Rick Gillman and David Housman. Models of Conflict and Cooperation. American Mathematical Society, Providence, RI, USA, 2009. ISBN 0-8218-4872-0. xi + 417 pp. LCCN QA269 .G487 2009.

Gharibian:2023:GCF

[Gha23] Sevag Gharibian. Guest column: The 7 faces of quantum NP. ACM SIGACT News, 54(4):54–91, December 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://

dl.acm.org/doi/10.1145/3639528.3639535.

Gasarch:2019:GCM

[GHMP19] William Gasarch, Scott Huddleston, Erik Metz, and Jacob Prinz. Guest column: The muffin problem. ACM SIGACT News, 50(2): 31–60, June 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ghosh:2007:VAP

[Gho07] Subir Kumar Ghosh. Visibility Algorithms in the Plane. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-87574-9. xiii + 318 pp. LCCN TA1634 .G485 2007. URL http:

//www.loc.gov/catdir/enhancements/fy0803/2007297923-b.
html;   http://www.loc.gov/catdir/enhancements/fy0803/
2007297923-d.html;   http://www.loc.gov/catdir/toc/fy0713/
2007297923.html.

Gasarch:2022:CCN

[GHOP22] William Gasarch, Nathan Hayes, Anthony Ostuni, and Davin Park. The complexity of chromatic number when restricted to graphs with bounded genus or bounded crossing number. ACM SIGACT News, 53(2):27–38, June 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.

org/doi/10.1145/3544979.3544987.

Glasser:2014:DNP

[GHSW14] C. Glaßer, A. Hughes, A. L. Selman, and N. Wisiol. Disjoint NPpairs and propositional proof systems. ACM SIGACT News, 45 (4):59–75, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Galil:1991:REC

[GI91] Zvi Galil and Giuseppe F. Italiano. Reducing edge connectivity to vertex connectivity. ACM SIGACT News, 22(1):57–61, Winter 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ginsburg:1975:AAT

[Gin75] Seymour Ginsburg. Algebraic and Automata-Theoretic Properties of Formal Languages, volume 2 of Fundamental studies in computer science. North-Holland Publishing Co., Amsterdam, The Netherlands, 1975. ISBN 0-444-10586-7 (American Elsevier). xii + 313 pp. LCCN QA267.3 .G54.

Garey:1978:NNS

[GJ78] M. R. Garey and D. S. Johnson. A note on “A note on ’Some simplified NP-complete graph problems”’. ACM SIGACT News, 9 (4):17, Winter 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Kri77].

Goldwasser:2002:DSN

[GJM02] Michael H. Goldwasser, David S. Johnson, and Catherine C. Mc-Geoch, editors. Data structures, near neighbor searches, and methodology: fifth and sixth DIMACS implementation challenges: papers related to the DIMACS challenge on dictionaries and priority queues (1995–1996) and the DIMACS challenge on near neighbor searches (1998–1999), volume 59 of DIMACS series in discrete mathematics and theoretical computer science. American Mathematical Society, Providence, RI, USA, 2002. ISBN 0-8218-2892-4 (hardcover). LCCN QA76.9.D35 D385 2002.

Gottlieb:1980:NSI

[GK80] Allan Gottlieb and Clyde P. Kruskal. A note on sorting integers from a bounded range. ACM SIGACT News, 12(3):66–67, Fall 1980. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Greene:1981:MAA

[GK81] Daniel H. Greene and Donald Ervin Knuth. Mathematics for the Analysis of Algorithms, volume 1 of Progress in computer science. Birkhäuser Verlag, Basel, Switzerland, 1981. ISBN 3-7643-3046-5. 107 pp. LCCN QA76.6 .G7423.

Greene:1982:MAA

[GK82] Daniel H. Greene and Donald E. Knuth. Mathematics for the Analysis of Algorithms. Birkhäuser Verlag, Basel, Switzerland, second edition, 1982. ISBN 3-7643-3102-X. 123 pp. LCCN QA76.6 .G7423 1982.

Greene:1990:MAA

[GK90] Daniel H. Greene and Donald Ervin Knuth. Mathematics for the Analysis of Algorithms, volume 1 of Progress in computer science and applied logic. Birkhäuser Verlag, Basel, Switzerland, third edition, 1990. ISBN 0-8176-3515-7 (U.S.), 3-7643-3515-7 (Germany). viii + 132 pp. LCCN QA76.6 .G7423 1990.

Gasarch:1999:BRCc

[GK99] William Gasarch and Samir Khuller. The book review column. ACM SIGACT News, 30(3):8, September 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Greene:2007:MAA

[GK07] Daniel H. Greene and Donald Ervin Knuth. Mathematics for the Analysis of Algorithms. Modern Birkhäuser Classics. Birkhäuser Verlag, Basel, Switzerland, third edition, 2007. ISBN 0-8176-4728-7, 0-8176-4729-5. viii + 132 pp. LCCN QA76.6 .G7423 2008.

Gilbert:2009:RSW

[GK09] Seth Gilbert and Dariusz R. Kowalski. Reliability and security in wireless networks. ACM SIGACT News, 40(4):86–87, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2006:BRB

[GKKK06] William I. Gasarch, Alexander Kruskal, Justin Kruskal, and Rebecca Kruskal. Book review: The Square Root of 2: A Dialogue Concerning a Number and a Sequence, by David Flannery, Copernicus Books, 2006. ACM SIGACT News, 37(3):27–32, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Fla05].

Graham:1988:TAB

[GKP88] Ronald L. Graham, Donald E. Knuth, and Oren Patashnik. Textbook announcement: Concrete Mathematics: a Foundation for Computer Science. ACM SIGACT News, 20(??):??, ???? 1988. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [GKP89].

Graham:1989:CM

[GKP89] Ronald L. Graham, Donald E. Knuth, and Oren Patashnik. Concrete Mathematics. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-14236-8. xiii + 625 pp. LCCN QA39.2 .G7331 1989.

Gasarch:1996:FLS

[GKP96] William Gasarch, Wayne Kelly, and William Pugh. Finding the i th largest of n for small i, n. ACM SIGACT News, 27(2):88–96, June 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Garland:1972:TRS

[GL72] Stephen J. Garland and David C. Luckham. Translating recursion schemes into program schemes. ACM SIGACT News, 4(14):83– 96, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gens:1980:CAA

[GL80] Georgii Gens and Evgenii Levner. Complexity of approximation algorithms for combinatorial problems: a survey. ACM SIGACT News, 12(3):52–65, Fall 1980. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gurari:1984:DDCa

[GL84a] Eitan M. Gurari and Ten Hwang Lai. Deadlock detection in communicating finite state machines. ACM SIGACT News, 15(4): 63–64, Winter–Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gurari:1984:DDCb

[GL84b] Eitan M. Gurari and Ten Hwang Lai. Deadlock detection in communicating finite state machines. ACM SIGACT News, 16(1): 63–64, Winter–Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gilbert:2002:BCF

[GL02] Seth Gilbert and Nancy Lynch. Brewer’s conjecture and the feasibility of consistent, available, partition-tolerant Web services. ACM SIGACT News, 33(2):51–59, June 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2009:PWA

[GL09] Judy Green and Jeanne LaDuke. Pioneering Women in American Mathematics: the pre-1940 PhD’s, volume 34 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2009. ISBN 0-8218-4376-1 (hardcover). xviii + 349 pp. LCCN QA28 .G74 2009.

Glenn:2005:BRB

[Gle05] James Glenn. Book review: Teaching Statistics Using Baseball, by Jim Albert, MAA, 2003. ACM SIGACT News, 36(3):19–21, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Alb03].

Goldsmith:1996:LN

[GLM96] Judy Goldsmith, Matthew A. Levy, and Martin Mundhenk. Limited nondeterminism. ACM SIGACT News, 27(2):20–29, June 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Grumbach:1996:QLB

[GLMW96] Stéphane Grumbach, Leonid Libkin, Tova Milo, and Limsoon Wong. Query languages for bags: expressive power and complexity. ACM SIGACT News, 27(2):30–37, June 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:1999:BQR

[GM99] William I. Gasarch and Georgia A. Martin. Bounded Queries in Recursion Theory, volume 16 of Progress in computer science and applied logic. Birkhäuser Verlag, Basel, Switzerland, 1999. ISBN 0-8176-3966-7 (hardcover). xiii + 353 pp. LCCN QA267.7 .G38 1999.

Gay:2010:STQ

[GM10] Simon Gay and I. (Ian) Mackie, editors. Semantic Techniques in Quantum Computation. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-51374-X (hardcover). xiv + 478 pp.

LCCN QA76.889 .S36 2010. URL http://assets.cambridge.
org/97805215/13746/cover/9780521513746.jpg.

Gramoli:2012:WFW

[GM12] Vincent Gramoli and Alessia Milani. WTTM 2012, the fourth workshop on the theory of transactional memory. ACM SIGACT News, 43(4):116–122, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gasarch:2021:OPCa

[GM21] William Gasarch and Erik Metz. Open problems column. ACM SIGACT News, 52(1):36–40, March 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://

dl.acm.org/doi/10.1145/3457588.3457595.

Gallaire:1986:LDR

[GMN87] H. Gallaire, J. Minker, and J. M. Nicolas. Logic and databases: a response. ACM SIGACT News, 18(2):52–56, September/March 1986–1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Germano:1974:LAC

[GMS74] G. Germano and A. Maggiolo-Schettini. Loops in Algol 60 and in category theory. ACM SIGACT News, 6(4):21–23, October 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Guo:2007:IDR

[GN07] Jiong Guo and Rolf Niedermeier. Invitation to data reduction and problem kernelization. ACM SIGACT News, 38(1):31–45, March 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Goldschlager:1977:MPC

[Gol77] Leslie M. Goldschlager. The monotone and planar circuit value problems are log space complete for P. ACM SIGACT News, 9(2): 25–29, Summer 1977. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Goldreich:1992:CST

[Gol92] Oded Goldreich. Critique of some trends in the TCS community in light of two controversies. ACM SIGACT News, 23(1):44–46, Winter 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Goldreich:1993:TPS

[Gol93] Oded Goldreich. A taxonomy of proof systems (part 1). ACM SIGACT News, 24(4):2–13, December 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Goldreich:1994:TPS

[Gol94] Oded Goldreich. A taxonomy of proof systems (part 2). ACM SIGACT News, 25(1):22–30, March 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Goldwasser:1995:SLP

[Gol95] Michael Goldwasser. A survey of linear programming in randomized subexponential time. ACM SIGACT News, 26(2):96–104, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Goldreich:1999:MCP

[Gol99] Oded Goldreich. Modern cryptography, probabilistic proofs, and pseudorandomness, volume 17 of Algorithms and combinatorics, 0937-5511. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-64766-X. xv + 182 pp. LCCN QA76.9.A25 G64 1999. URL http://www.loc.gov/

catdir/enhancements/fy0815/98050548-d.html;   http://www.
loc.gov/catdir/enhancements/fy0815/98050548-t.html.

Goldengorin:2000:BRB

[Gol00a] Boris Goldengorin. Book review: Network Design: Connectivity and Facilities Location. Proceedings from DIMACS workshop in April 1997 (AMS 1998). ACM SIGACT News, 31(2):5–9, June 1, 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [PD98].

Goldsmith:2000:BRB

[Gol00b] Judy Goldsmith. Book review: Theory of Computing: A Gentle Introduction, by Kinber and Smith (Prentice-Hall, 2001). ACM SIGACT News, 31(3):19–22, September 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [KS01].

Goldreich:2001:FCB

[Gol01] Oded Goldreich. Foundations of Cryptography: Basic Tools. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-79172-3, 0-521-83084-2 (vol. 2). xix + 372 pp. LCCN QA268 .G5745 2001.

Goldreich:2006:GCB

[Gol06] Oded Goldreich. Guest column: Bravely, moderately — a common theme in four recent works. ACM SIGACT News, 37(2):31–46, June 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Goldreich:2008:CCCa

[Gol08a] Oded Goldreich. Computational complexity: a conceptual perspective. ACM SIGACT News, 39(3):35–39, September 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Goldreich:2008:CCCb

[Gol08b] Oded Goldreich. Computational Complexity: a Conceptual Perspective. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-88473-X (hardcover), 0-511-64992-4 (e-book), 0-511-40077-2 (ebook). xxiv + 606 pp. LCCN QA267.7 .G65 2008.

Goldreich:2009:ODS

[Gol09] Oded Goldreich. On our duties as scientists. ACM SIGACT News, 40(3):53–59, September 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Goldreich:2010:PNN

[Gol10a] Oded Goldreich. P, NP, and NP-completeness: the basics of computational complexity. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19248-X (hardcover), 0-521-12254-6 (paperback). xxix + 184 pp. LCCN QA267.7 .G652

2010. URL http://assets.cambridge.org/97805211/92484/
cover/9780521192484.jpg.

Goldwasser:2010:SBM

[Gol10b] Michael H. Goldwasser. A survey of buffer management policies for packet switches. ACM SIGACT News, 41(1):100–128, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Goldreich:2012:IIV

[Gol12a] Oded Goldreich. On intellectual and instrumental values in science. ACM SIGACT News, 43(2):45–50, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Goldreich:2012:SCS

[Gol12b] Oded Goldreich. On struggle and competition in scientic fields. ACM SIGACT News, 43(1):43–60, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Golab:2018:PP

[Gol18] Wojciech Golab. Proving PACELC. ACM SIGACT News, 49(1): 73–81, March 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Goodrich:1993:PAC

[Goo93] Michael T. Goodrich. Parallel algorithms column 1: models of computation. ACM SIGACT News, 24(4):16–21, December 1993. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Goos:2012:RB

[Göö12] Mika Göös. Review of DISC 2012. ACM SIGACT News, 43(4): 112–115, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gopalakrishnan:2008:RCE

[Gop08] Ganesh Lalitha Gopalakrishnan. Review computation engineering: applied automata theory and logic. ACM SIGACT News, 39(3):30– 32, September 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gottlob:2000:RP

[Got00] Georg Gottlob. Report on PODS 2000. ACM SIGACT News, 31 (3):43–46, September 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Guessarian:1989:LLI

[GP89] I. Guessarian and D. Perrin. LITP: Laboratoire D’Informatique théorique et programmation Paris: Presentation of scientific activity. ACM SIGACT News, 20(4):50–53, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Guessarian:1983:IVR

[GPP83] Irène Guessarian and Francesco Parisi-Presicce. Iterative vs. regular factor algebras. ACM SIGACT News, 15(2):32–44, Spring– Summer 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Grosse-Rhode:2004:SIH

[GR04] Martin Große-Rhode. Semantic Integration of Heterogeneous Software Specifications. Monographs in theoretical computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-40257-8. ix + 327 pp. LCCN QA76.76.D47 G79 2004. URL http:

//www.loc.gov/catdir/enhancements/fy0817/2003069097-b.
html;   http://www.loc.gov/catdir/enhancements/fy0817/
2003069097-d.html;   http://www.loc.gov/catdir/enhancements/
fy0817/2003069097-t.html.

Gratzer:1978:GLTa

[Grä78a] George A. Grätzer. General Lattice Theory. Lehrbücher und Monographien aus dem Gebiete der exakten Wissenschaften: Mathematische Reihe; Bd. 52. Birkhäuser Verlag, Basel, Switzerland, 1978. ISBN 3-7643-0813-3. xiii + 381 pp. LCCN QA171.5 .G7. 78.00F.

Gratzer:1978:GLTb

[Grä78b] George A. Grätzer. General Lattice Theory, volume 75 of Pure and applied mathematics: a series of monographs and textbooks. Academic Press, New York, USA, 1978. ISBN 0-12-295750-4. xiii + 381 pp. LCCN QA3 .P8 vol. 75.

Grandjean:1982:NUQ

[Gra82] Etienne Grandjean. Number of universal quantifiers and computational complexity: preliminary report. ACM SIGACT News, 14 (3):30, Summer 1982. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gratzer:1998:GLT

[Grä98] George A. Grätzer. General Lattice Theory. Birkhäuser Verlag, Basel, Switzerland, second edition, 1998. ISBN 3-7643-5239-6 (Basel: hardcover), 0-8176-5239-6 (Boston: hardcover). xix + 663 pp. LCCN QA171.5 .G7 1998. With new appendices with B. A. Davey and others.

Gratzer:2008:BRJ

[Grä08] George Grätzer. A brief response to J. A. Cohen’s joint review. ACM SIGACT News, 39(3):26–28, September 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Coh08].

Gramoli:2009:WTT

[Gra09] Vincent Gramoli. What theory for transactional memory? ACM SIGACT News, 40(4):79–81, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Greibach:1981:CRT

[Gre81] S. A. Greibach. Comments on the roots of theorems and languages both easy and hard. ACM SIGACT News, 13(1):26–29, Winter 1981. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2007:BRB

[Gre07] Frederic Green. Book review: Mathematics of Physics and Engineering, by Edward K. Blum and Sergey V. Lototsky, World Scientific. ACM SIGACT News, 38(4):25–30, December 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BL06].

Green:2013:RQC

[Gre13] Frederic Green. Review: Quantum computing since Democritus by Scott Aaronson. ACM SIGACT News, 44(4):42–47, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2015:BRCa

[Gre15a] Frederic Green. The book review column. ACM SIGACT News, 46 (2):10–12, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2015:BRCb

[Gre15b] Frederic Green. The book review column. ACM SIGACT News, 46(3):5–6, September 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2015:BRCc

[Gre15c] Frederic Green. The book review column. ACM SIGACT News, 46(4):4–5, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2015:RIN

[Gre15d] Frederic Green. Review of: Incredible Numbers by Ian Stewart. ACM SIGACT News, 46(4):6–8, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2016:BRCa

[Gre16a] Frederic Green. The book review column. ACM SIGACT News, 47 (1):4–5, March 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2016:BRCb

[Gre16b] Frederic Green. The book review column. ACM SIGACT News, 47(2):4–5, June 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2016:BRCc

[Gre16c] Frederic Green. The book review column. ACM SIGACT News, 47(3):5–6, September 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2016:BRCd

[Gre16d] Frederic Green. The book review column. ACM SIGACT News, 47(4):4–5, December 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2016:RPT

[Gre16e] Frederic Green. Review of: Primality Testing for Beginners by Lasse Rempe-Gillen and Rebecca Waldecker. ACM SIGACT News, 47(2):6–9, June 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2016:RQA

[Gre16f] Frederic Green. Review of: Quantum Algorithms via Linear Algebra by Richard J. Lipton and Kenneth W. Regan. ACM SIGACT News, 47(3):7–11, September 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2016:RFW

[Gre16g] Frederic Green. Review of: The Fascinating World of Graph Theory by Arthur Benjamin, Gary Chartrand and Ping Zhang. ACM SIGACT News, 47(4):18–21, December 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2016:RNC

[Gre16h] Frederic Green. Review of: The Nature of Computation by Cristopher Moore and Stephan Mertens. ACM SIGACT News, 47(1):

9–11, March 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2017:BRCa

[Gre17a] Frederic Green. The book review column. ACM SIGACT News, 48 (1):8–9, March 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2017:BRCb

[Gre17b] Frederic Green. The book review column. ACM SIGACT News, 48(2):4–6, June 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2017:BRCc

[Gre17c] Frederic Green. The book review column. ACM SIGACT News, 48(3):4–6, September 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2017:BRCd

[Gre17d] Frederic Green. The book review column. ACM SIGACT News, 48(4):4–6, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2017:JRM

[Gre17e] Frederic Green. Joint review of The Magic of Math by Arthur Benjamin and How to Bake by Eugenia Cheng. ACM SIGACT News, 48(2):16–20, June 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2017:RAL

[Gre17f] Frederic Green. Review of Art in the Life of Mathematicians by Anna Kepes Szemerédi. ACM SIGACT News, 48(1):10–17, March 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2017:RQW

[Gre17g] Frederic Green. Review of Quantum Walks and Search Algorithms by Renato Portugal. ACM SIGACT News, 48(2):12–15, June 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2017:RST

[Gre17h] Frederic Green. Review of Set Theory: a First Course by Daniel W. Cunningham. ACM SIGACT News, 48(3):7–9, September 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2018:BRCa

[Gre18a] Frederic Green. The book review column. ACM SIGACT News, 49 (1):9–11, March 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2018:BRCb

[Gre18b] Frederic Green. The book review column. ACM SIGACT News, 49(2):4–6, June 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2018:BRCc

[Gre18c] Frederic Green. The book review column. ACM SIGACT News, 49(3):6–8, September 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2018:BRCd

[Gre18d] Frederic Green. The book review column. ACM SIGACT News, 49(4):4–6, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2018:JRQ

[Gre18e] Frederic Green. Joint review of Quadratic Residues and Non-Residues by Steve Wright and The Quadratic Reciprocity Law by Oswald Baumgart. Edited and translated by Franz Lemmermeyer. ACM SIGACT News, 49(1):20–28, March 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2019:BRCa

[Gre19a] Frederic Green. The book review column. ACM SIGACT News, 50 (1):7–8, March 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2019:BRCb

[Gre19b] Frederic Green. The book review column. ACM SIGACT News, 50(2):4–5, June 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2019:BRCc

[Gre19c] Frederic Green. The book review column. ACM SIGACT News, 50(3):4–5, September 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2019:BRC

[Gre19d] Frederic Green. The book review column. ACM SIGACT News, 50(4):4–6, December 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3374857.3374859.

Green:2019:RMC

[Gre19e] Frederic Green. Review of Modern Cryptography and Elliptic Curves, A Beginner’s Guide by Thomas R. Shemanske. ACM SIGACT News, 50(2):12–14, June 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2019:RNT

[Gre19f] Frederic Green. Review of Number Theory: an Introduction via the Density of Primes, second edition. ACM SIGACT News, 50 (1):9–13, March 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2019:RHG

[Gre19g] Frederic Green. Review of handbook of graph theory, combinatorial optimization, and algorithms. ACM SIGACT News, 50(3): 6–11, September 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Green:2020:BRCa

[Gre20a] Frederic Green. The book review column. ACM SIGACT News, 51 (1):7–8, March 2020. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3388392.3388394.

Green:2020:BRCb

[Gre20b] Frederic Green. The book review column. ACM SIGACT News, 51(2):4–5, June 2020. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3406678.3406680.

Green:2020:BRCc

[Gre20c] Frederic Green. The book review column. ACM SIGACT News, 51(3):5–6, September 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3427361.3427363.

Green:2020:BRCd

[Gre20d] Frederic Green. The book review column1. ACM SIGACT News, 51(4):4–5, December 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3444815.3444817.

Green:2020:RFC

[Gre20e] Frederic Green. Review of Forbidden Configurations in Discrete Geometry by David Eppstein. ACM SIGACT News, 51(4):15– 17, December 2020. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3444815.3444820.

Green:2021:BRCa

[Gre21a] Frederic Green. The book review column. ACM SIGACT News, 52 (1):7–8, March 2021. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3457588.3457590.

Green:2021:BRCb

[Gre21b] Frederic Green. The book review column. ACM SIGACT News, 52(2):4–6, June 2021. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3471469.3471471.

Green:2021:BRCc

[Gre21c] Frederic Green. The book review column. ACM SIGACT News, 52(3):3–5, September 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3494656.3494658.

Green:2021:BRC

[Gre21d] Frederic Green. The book review column. ACM SIGACT News, 52(4):3–5, December 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3510382.3510384.

Green:2021:RMC

[Gre21e] Frederic Green. Review of Mathematics and Computation by Avi Wigderson. ACM SIGACT News, 52(3):6–10, September 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3494656.
3494659.

Green:2021:RTL

[Gre21f] Frederic Green. Review of Three Lectures on Complexity and Black Holes by Leonard Susskind. ACM SIGACT News, 52(4): 6–10, December 2021. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3510382.3510385.

Grechuk:2022:SOD

[Gre22a] Bogdsan Grechuk. On the smallest open Diophantine equations. ACM SIGACT News, 53(1):36–57, March 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3532737.3532743.

Green:2022:BRCa

[Gre22b] Frederic Green. The book review column. ACM SIGACT News, 53 (1):6–8, March 2022. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3532737.3532739.

Green:2022:BRCb

[Gre22c] Frederic Green. The book review column. ACM SIGACT News, 53(2):3–5, June 2022. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3544979.3544981.

Green:2022:BRCc

[Gre22d] Frederic Green. The book review column. ACM SIGACT News, 53(3):3–5, September 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3561064.3561065.

Green:2022:BRCd

[Gre22e] Frederic Green. The book review column. ACM SIGACT News, 53 (4):6, December 2022. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3577971.3577973.

Green:2022:RCD

[Gre22f] Frederic Green. Review of Complexity Dichotomies for Counting Problems Volume 1: Boolean Domain by Jin-Yi Cai and Xi Chen. ACM SIGACT News, 53(2):6–10, June 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://

dl.acm.org/doi/10.1145/3544979.3544982.

Green:2022:RFC

[Gre22g] Frederic Green. Review of Feasible Computations and Provable Complexity Properties by Juris Hartmanis. ACM SIGACT News, 53(4):7–10, December 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3577971.3577974.

Green:2022:RMM

[Gre22h] Frederic Green. Review of “Mathematical Muffin Morsels — Nobody wants a small piece by William Gasarch, Erik Metz, Jacob Prinz, and Daniel Smolyak World Scientific, 2021, 210 pages, Softcover, $59.99”. ACM SIGACT News, 53(3):17–20, September 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3561064.
3561067.

Griffin:2010:BRB

[Gri10] John S. Griffin. Book review: The modern algebra of information retrieval by Sandor Dominich, Springer-Verlag, Berlin / Heidelberg. ACM SIGACT News, 41(1):30–34, March 2010. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Grohe:2014:DTC

[Gro14] Martin Grohe. Database theory column report on PODS 2014. ACM SIGACT News, 45(4):83–85, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gaifman:1980:RHC

[GS80] H. Gaifman and E. Shamir. Roots of the hardest context free language and other constructs. ACM SIGACT News, 12(3):45–51, Fall 1980. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gold:2008:POD

[GS08] Bonnie Gold and Roger A. Simons, editors. Proof and other dilemmas: mathematics and philosophy. Spectrum series. Mathematical Association of America, Washington, DC, USA, 2008. ISBN 0-88385-567-4 (hardcover). xxxii + 346 pp. LCCN QA8.4 .P755 2008.

Goodrich:1997:DSA

[GT97] Michael T. Goodrich and Roberto Tamassia. Data Structures and Algorithms in Java. John Wiley, New York, NY, USA, 1997. ISBN 0-471-19308-9 (paperback). xviii + 738 pp. LCCN

QA76.73.J38 G66 1997. URL http://www.loc.gov/catdir/
enhancements/fy0706/97039794-d.html;   http://www.loc.
gov/catdir/toc/onix01/97039794.html.

Goodrich:2001:DSA

[GT01] Michael T. Goodrich and Roberto Tamassia. Data Structures and Algorithms in Java. John Wiley, New York, NY, USA, second edition, 2001. ISBN 0-471-38367-8 (hardcover). xiii + 641 pp. LCCN QA76.73.J38 G66 2001. URL

http://www.loc.gov/catdir/bios/wiley043/00043294.html;
 http://www.loc.gov/catdir/description/wiley035/00043294.
html;   http://www.loc.gov/catdir/toc/onix07/00043294.
html.

Goodrich:2002:ADF

[GT02] Michael T. Goodrich and Roberto Tamassia. Algorithm Design: Foundations, Analysis, and Internet Examples. John Wiley, New York, NY, USA, 2002. ISBN 0-471-38365-1. xii + 708 pp. LCCN QA76.9.A43 G67 2002. URL http://www.loc.gov/

catdir/description/wiley034/2002265982.html;   http://
www.loc.gov/catdir/toc/wiley022/2002265982.html.

Goodrich:2004:DSA

[GT04] Michael T. Goodrich and Roberto Tamassia. Data Structures and Algorithms in Java. John Wiley, New York, NY, USA, third edition, 2004. ISBN 0-471-46983-1 (hardcover). xvii + 681 pp. LCCN QA76.73.J38 G66 2004. URL http://www.

loc.gov/catdir/bios/wiley046/2003071070.html;   http:/
/www.loc.gov/catdir/description/wiley041/2003071070.html;
 http://www.loc.gov/catdir/toc/wiley041/2003071070.html.

Goodrich:2006:DSA

[GT06] Michael T. Goodrich and Roberto Tamassia. Data Structures and Algorithms in Java. John Wiley, New York, NY, USA, fourth edition, 2006. ISBN 0-471-73884-0. xxii + 696 pp. LCCN QA76.73.J38 G66 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0627/2005282681-b.
html;   http://www.loc.gov/catdir/enhancements/fy0627/
2005282681-d.html;   http://www.loc.gov/catdir/enhancements/
fy0627/2005282681-t.html.

Goodrich:2010:DSA

[GT10] Michael T. Goodrich and Roberto Tamassia. Data Structures and Algorithms in Java. John Wiley, New York, NY, USA, fifth edition, 2010. ISBN 0-470-38326-7 (hardcover), 0-470-39880-9 (paperback). xxii + 714 pp. LCCN QA76.73.J38 G66 2010.

Gradel:2002:ALI

[GTW02] Erich Grädel, Wolfgang Thomas, and Thomas Wilke. Automata, Logics, and Infinite Games, volume 2500 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-00388-6. viii + 385 pp. LCCN QA76.9.M35 A85 2002. URL http://www.loc.gov/

catdir/enhancements/fy0817/2002044587-d.html.

Guessarian:1983:SCI

[Gue83] Irène Guessarian. Survey on classes of interpretations and some of their applications. ACM SIGACT News, 15(3):45–71, Summer– Fall 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gunter:1992:SPL

[Gun92] Carl A. Gunter. Semantics of Programming Languages: Structures and Techniques. Foundations of computing. MIT Press, Cambridge, MA, USA, 1992. ISBN 0-262-07143-6. xviii + 419 pp. LCCN QA76.7 .G86 1992.

Gupta:1998:BRB

[Gup98] Sanjay Gupta. Book review: Theories of Computability, by Nicholas Pippenger (Cambridge University Press 1997). ACM SIGACT News, 29(4):5–10, December 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Pip97].

Gupta:2002:CCF

[Gup02] Vijay Shivshanker Gupta. Communication complexity for file synchronization is undecidable. ACM SIGACT News, 33(3):110–112, September 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gurevich:1986:WDM

[Gur86] Y. Gurevich. What does O(n) mean. ACM SIGACT News, 17(4): 61–63, Spring 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gurevich:1994:LAE

[Gur94] Yuri Gurevich. Logic activities in Europe. ACM SIGACT News, 25 (2):11–24, June 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Guruswami:2004:GCE

[Gur04] Venkatesan Guruswami. Guest column: error-correcting codes and expander graphs. ACM SIGACT News, 35(3):25–41, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gusfield:1997:AST

[Gus97a] Dan Gusfield. Algorithms on Strings, Trees, and Sequences: Computer Science and Computational Biology. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-58519-8 (hardcover). xviii + 534 pp. LCCN QA76.9.A43 G87 1997. URL http://

www.loc.gov/catdir/description/cam028/96046612.html;
http://www.loc.gov/catdir/toc/cam023/96046612.html.

Gusfield:1997:BST

[Gus97b] Dan Gusfield. Algorithms on Strings, Trees, and Sequences: Computer Science and Computational Biology. ACM SIGACT News, 28(4):41–60, December 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Gus97a].

Guz:2008:RS

[Guz08] Zvika Guz. Review of SPAA’08. ACM SIGACT News, 39(4):66– 68, December 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Goldreich:1997:TCS

[GW97] Oded Goldreich and Avi Wigderson. Theory of computing: a scientific perspective (extended abstract). ACM SIGACT News, 28 (3):100–102, September 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Grochow:2018:BNB

[GW18] Joshua A. Grochow and David H. Wolpert. Beyond number of bit erasures: New complexity questions raised by recently discovered thermodynamic costs of computation. ACM SIGACT News, 49 (2):33–56, June 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Gross:2004:HGT

[GY04] Jonathan L. Gross and Jay Yellen, editors. Handbook of Graph Theory. Discrete mathematics and its applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2004. ISBN 1-58488-090-2. 1167 pp. LCCN QA166

.H36 2004. URL http://www.loc.gov/catdir/enhancements/
fy0646/2003065270-d.html.

Hacken:2009:QCQ

[Hac09] George Hacken. Quantum computation and quantum communication: theory and experiments Author of book: Mladen Pavicic. ACM SIGACT News, 40(4):10–14, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hadzilacos:2015:QCW

[Had15] Vassos Hadzilacos. A quarter-century of wait-free synchronization. ACM SIGACT News, 46(2):79–88, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Haglin:1999:TRC

[Hag99] David J. Haglin. Technical report column. ACM SIGACT News, 30(4):22–24, December 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Haglin:2000:TRC

[Hag00] David J. Haglin. Technical Report Column. ACM SIGACT News, 31(1):14–15, March 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Haglin:2002:TRC

[Hag02] David J. Haglin. Technical report column. ACM SIGACT News, 33 (2):29–30, June 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Halpern:2003:RAU

[Hal03] Joseph Y. Halpern. Reasoning About Uncertainty. MIT Press, Cambridge, MA, USA, 2003. ISBN 0-262-08320-5 (hardcover). xiv + 483 pp. LCCN Q375 .H35 2003.

Hamlet:1976:ADP

[Ham76] Richard Hamlet. Application of “DOVETAILING’’ to program testing. ACM SIGACT News, 8(2):25–26, April/June 1976. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hambrusch:1990:RMT

[Ham90] Susanne E. Hambrusch. Report of the 21st Midwest Theory Consortium. ACM SIGACT News, 21(2):30, Spring 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hame:2013:RBC

[Ham13] Angèle M. Hame. Review of Enumerative Combinatorics, volume 1, second edition by Richard P. Stanley. ACM SIGACT News, 44 (3):28–31, September 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hankin:1994:LCG

[Han94] Chris Hankin. Lambda calculi: a guide for computer scientists, volume 3 of Graduate texts in computer science. Clarendon Press, New York, NY, USA, 1994. ISBN 0-19-853841-3, 0-19-853840-5. xiii + 162 pp. LCCN QA9.5 .H36 1994.

Harel:1979:FOD

[Har79] David Harel. First-Order Dynamic Logic, volume 68 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. ISBN 0-387-09237-4. 133 pp. LCCN QA76.6 .H34 1979.

Harel:1980:FT

[Har80] David Harel. On folk theorems. ACM SIGACT News, 12(3):68–80, Fall 1980. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Harel:1986:LDC

[Har86] D. Harel. Logic and databases: a critique. ACM SIGACT News, 18(1):68–74, Summer 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Harel:1995:WPW

[Har95] David Harel. Will I be pretty, will I be rich?: some thoughts on theory vs. practice in systems engineering. ACM SIGACT News, 26 (2):21–25, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Haralambous:2010:BRB

[Har10] Yannis Haralambous. Book review: History of Mathematics, Princeton University Press, 2004, xxvi + 372 pages. ACM SIGACT News, 41(3):46–50, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Nah04].

Haralambous:2011:BRB

[Har11] Yannis Haralambous. Book review: A View from the Top: Analysis, Combinatorics and Number Theory, by Alex Iosevich. ACM SIGACT News, 42(1):38–42, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Ios07].

Harris:2015:HTC

[Har15] Tim Harris. Hardware trends: Challenges and opportunities in distributed computing. ACM SIGACT News, 46(2):89–95, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Harris:2021:FWF

[Har21] Tim Harris. Five ways not to fool yourself: Designing experiments for understanding performance. ACM SIGACT News, 52(3):61– 68, September 2021. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3494656.3494668.

Hastings:1985:CSA

[Has85a] Harold M. Hastings. Convergence of simulated annealing. ACM SIGACT News, 17(2):52–63, September 1985. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hastings:1985:MIP

[Has85b] Harold M. Hastings. Maps of the interval, polynomial time, and polynomial space. ACM SIGACT News, 17(2):44–51, September 1985. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hautamaki:2012:RBC

[Hau12] Ville Hautamäki. Review of A Concise Introduction to Data Compression by David Salomon. ACM SIGACT News, 43(2):9–10, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Havel:1971:WCM

[Hav71] Ivan M. Havel. Weak complexity measures. ACM SIGACT News, 3(8):21–30, January 1971. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Havel:1974:ATM

[Hav74] Ivan M. Havel. Automata theory motivated by problem solving. ACM SIGACT News, 6(2):18–23, April 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hayes:2008:GTB

[Hay08] Brian Hayes. Group Theory in the Bedroom and Other Mathematical Diversions. Hill and Wang, New York, NY, USA, 2008. ISBN 0-8090-5219-9 (hardcover). xi + 269 pp. LCCN T185

.H39 2008. URL http://www.loc.gov/catdir/enhancements/
fy0740/2007023337-b.html;   http://www.loc.gov/catdir/
enhancements/fy0740/2007023337-d.html.

Hearn:2009:GPC

[HD09] Robert A. Hearn and Erik D. Demaine. Games, Puzzles, and Computation. A. K. Peters, Ltd., Wellesley, MA, USA, 2009. ISBN 1-56881-322-8 (hardcover). ix + 237 pp. LCCN QA63 .H35 2009.

Helmi:2011:RP

[Hel11] Maryam Helmi. A review of PODC 2011. ACM SIGACT News, 42(4):83–86, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:1993:LYN

[Hem93] Lane A. Hemaspaandra. Lowness: a yardstick for NP-P. ACM SIGACT News, 24(2):10–14, Spring 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:1994:CTC

[Hem94a] Lane A. Hemaspaandra. Complexity theory column 5: the notready-for-prime-time conjectures. ACM SIGACT News, 25(2):5– 10, June 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:1994:TCC

[Hem94b] Lane A. Hemaspaandra. Teaching Computational Complexity: Resources to Treasure. ACM SIGACT News, 25(4):2–11, December 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:1995:SNCb

[Hem95] Lane A. Hemaspaandra. SIGACT News Complexity Theory Column 10. ACM SIGACT News, 26(3):2–12, September 1995. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:1996:SNC

[Hem96] Lane A. Hemaspaandra. SIGACT News Complexity Theory Column 12. ACM SIGACT News, 27(1):2–13, March 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:1997:J

[Hem97a] Lane A. Hemaspaandra. Journals to Die For. ACM SIGACT News, 28(1):2, March 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:1997:SNC

[Hem97b] Lane A. Hemaspaandra. SIGACT News complexity theory column 18. ACM SIGACT News, 28(3):2–11, September 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:1998:THC

[Hem98] Lane A. Hemaspaandra. Take-home complexity. ACM SIGACT News, 29(2):9–13, June 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:1999:BCR

[Hem99] Lane A. Hemaspaandra. Biomolecular computing: recent theoretical and experimental advances. ACM SIGACT News, 30(2):22–30, June 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2001:CT

[Hem01a] Lane A. Hemaspaandra. Complexity theory. ACM SIGACT News, 32(3):40–52, September 1, 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2001:SNCa

[Hem01b] Lane A. Hemaspaandra. SIGACT news complexity theory column 31. ACM SIGACT News, 32(1):21–31, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2001:SNCb

[Hem01c] Lane A. Hemaspaandra. SIGACT News complexity theory column 32. ACM SIGACT News, 32(2):32–43, June 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2001:SNCc

[Hem01d] Lane A. Hemaspaandra. SIGACT news complexity theory column 34. ACM SIGACT News, 32(4):24–33, December 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2002:SNCa

[Hem02a] Lane A. Hemaspaandra. SIGACT news complexity theory column 35. ACM SIGACT News, 33(1):32–45, March 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2002:SNCb

[Hem02b] Lane A. Hemaspaandra. SIGACT news complexity theory column 36. ACM SIGACT News, 33(2):34–47, June 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2002:SNCd

[Hem02c] Lane A. Hemaspaandra. SIGACT news complexity theory column 38. ACM SIGACT News, 33(4):22–36, December 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2002:SNCc

[Hem02d] Lane A. Hemaspaandra. SIGACT news complexity theory comun 37. ACM SIGACT News, 33(3):32–49, September 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2003:BRB

[Hem03a] Lane A. Hemaspaandra. Book review: The π-Calculus: a Theory of Mobile Processes. ACM SIGACT News, 34(1):17–24, March 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [SW01].

Hemaspaandra:2003:SNCa

[Hem03b] Lane A. Hemaspaandra. SIGACT news complexity theory column 40. ACM SIGACT News, 34(2):27–41, June 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2003:SNCb

[Hem03c] Lane A. Hemaspaandra. SIGACT News complexity theory column 41. ACM SIGACT News, 34(3):26–39, September 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2003:SNCc

[Hem03d] Lane A. Hemaspaandra. SIGACT news complexity theory column 42. ACM SIGACT News, 34(4):38–52, December 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2004:SNC

[Hem04] Lane A. Hemaspaandra. SIGACT news complexity theory column 43. ACM SIGACT News, 35(1):22–35, March 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2005:TRC

[Hem05a] Lane Hemaspaandra. Technical report column. ACM SIGACT News, 36(3):22–23, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2005:SNCa

[Hem05b] Lane A. Hemaspaandra. SIGACT news complexity theory column 48. ACM SIGACT News, 36(3):24–38, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2005:SNCb

[Hem05c] Lane A. Hemaspaandra. SIGACT news complexity theory column 49. ACM SIGACT News, 36(4):24–35, December 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2006:AAS

[Hem06a] Lane Hemaspaandra. Approximation algorithms for 2-stage stochastic optimization problems. ACM SIGACT News, 37(1): 33–46, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2006:TRC

[Hem06b] Lane Hemaspaandra. Technical report column. ACM SIGACT News, 37(3):33–35, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2006:SNCa

[Hem06c] Lane A. Hemaspaandra. SIGACT news complexity theory column 51. ACM SIGACT News, 37(2):31, June 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2006:SNCb

[Hem06d] Lane A. Hemaspaandra. SIGACT news complexity theory column 52. ACM SIGACT News, 37(3):36–54, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2006:SNCc

[Hem06e] Lane A. Hemaspaandra. SIGACT news complexity theory column 53. ACM SIGACT News, 37(4):47–55, December 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2007:Ia

[Hem07a] Lane A. Hemaspaandra. Introduction. ACM SIGACT News, 38 (3):34–38, September 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2007:Ib

[Hem07b] Lane A. Hemaspaandra. Introduction. ACM SIGACT News, 38 (4):39–40, December 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2007:TRCa

[Hem07c] Lane A. Hemaspaandra. Technical report column. ACM SIGACT News, 38(1):28–30, March 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2007:TRCb

[Hem07d] Lane A. Hemaspaandra. Technical report column. ACM SIGACT News, 38(2):26–28, June 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2008:RNI

[Hem08a] Lane A. Hemaspaandra. Research Notices: The injectivity of the global function of a cellular automaton in the hyperbolic plane is undecidable. ACM SIGACT News, 39(3):40, September 2008. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2008:SNCa

[Hem08b] Lane A. Hemaspaandra. SIGACT news complexity theory column 59: introduction. ACM SIGACT News, 39(2):50, June 2008. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2008:SNCb

[Hem08c] Lane A. Hemaspaandra. SIGACT news complexity theory column 61. ACM SIGACT News, 39(4):35–36, December 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2008:TRC

[Hem08d] Lane A. Hemaspaandra. Technical report column. ACM SIGACT News, 39(1):30–32, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2009:SNCa

[Hem09a] Lane A. Hemaspaandra. SIGACT news complexity theory column 62. ACM SIGACT News, 40(1):26, March 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2009:SNCb

[Hem09b] Lane A. Hemaspaandra. SIGACT news complexity theory column 63. ACM SIGACT News, 40(2):49, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2009:SNCc

[Hem09c] Lane A. Hemaspaandra. SIGACT news complexity theory column 64. ACM SIGACT News, 40(3):60–76, September 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2009:SNCd

[Hem09d] Lane A. Hemaspaandra. SIGACT news complexity theory column 65. ACM SIGACT News, 40(4):45, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2010:SNCa

[Hem10a] Lane A. Hemaspaandra. SIGACT News Complexity Theory Column 67. ACM SIGACT News, 41(3):58, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2010:SNCb

[Hem10b] Lane A. Hemaspaandra. SIGACT news complexity theory column 68. ACM SIGACT News, 41(4):73–94, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2010:TRCa

[Hem10c] Lane A. Hemaspaandra. Technical report column. ACM SIGACT News, 41(2):53–56, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2010:TRCb

[Hem10d] Lane A. Hemaspaandra. Technical report column. ACM SIGACT News, 41(3):56–58, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2011:SNCa

[Hem11a] Lane A. Hemaspaandra. SIGACT news complexity theory column 69. ACM SIGACT News, 42(1):58, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2011:SNCb

[Hem11b] Lane A. Hemaspaandra. SIGACT news complexity theory column 70. ACM SIGACT News, 42(2):51, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2011:SNCc

[Hem11c] Lane A. Hemaspaandra. SIGACT news complexity theory column 71. ACM SIGACT News, 42(3):53–54, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2011:SNCd

[Hem11d] Lane A. Hemaspaandra. SIGACT news complexity theory column 72. ACM SIGACT News, 42(4):53, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2012:SNCa

[Hem12a] Lane A. Hemaspaandra. SIGACT news complexity theory column 73. ACM SIGACT News, 43(1):61, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2012:SNCb

[Hem12b] Lane A. Hemaspaandra. SIGACT news complexity theory column 74. ACM SIGACT News, 43(2):51–52, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2012:SNCc

[Hem12c] Lane A. Hemaspaandra. SIGACT news complexity theory column 75. ACM SIGACT News, 43(3):65–66, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2013:SNCa

[Hem13a] Lane A. Hemaspaandra. SIGACT news complexity theory column 77. ACM SIGACT News, 44(1):49, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2013:SNCb

[Hem13b] Lane A. Hemaspaandra. SIGACT News complexity theory column 78. ACM SIGACT News, 44(2):45–46, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2013:SNC

[Hem13c] Lane A. Hemaspaandra. SIGACT news complexity theory column 80. ACM SIGACT News, 44(4):52, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2014:BSA

[Hem14a] Lane A. Hemaspaandra. Beautiful structures: an appreciation of the contributions of Alan Selman. ACM SIGACT News, 45(3):54– 70, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2014:SNCa

[Hem14b] Lane A. Hemaspaandra. SIGACT news complexity theory column 81. ACM SIGACT News, 45(1):47, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2014:SNCb

[Hem14c] Lane A. Hemaspaandra. SIGACT news complexity theory column 82. ACM SIGACT News, 45(2):46, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2014:SNCc

[Hem14d] Lane A. Hemaspaandra. SIGACT news complexity theory column 83. ACM SIGACT News, 45(3):53, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2014:SNCd

[Hem14e] Lane A. Hemaspaandra. SIGACT News complexity theory column 84. ACM SIGACT News, 45(4):58, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2015:SNCa

[Hem15a] Lane A. Hemaspaandra. SIGACT News complexity theory column 85. ACM SIGACT News, 46(1):39, March 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2015:SNCb

[Hem15b] Lane A. Hemaspaandra. SIGACT News complexity theory column 86: Introduction to complexity theory column 86. ACM SIGACT News, 46(2):40, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2015:SNCc

[Hem15c] Lane A. Hemaspaandra. SIGACT News complexity theory column 87: Introduction to complexity theory column 87. ACM SIGACT News, 46(3):36, September 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2015:SNCd

[Hem15d] Lane A. Hemaspaandra. SIGACT News complexity theory column 88/89: Introduction to complexity theory column 88/89. ACM SIGACT News, 46(4):31, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2016:SNC

[Hem16a] Lane A. Hemaspaandra. SIGACT News complexity theory column 93. ACM SIGACT News, 47(3):44–45, September 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2016:SNCa

[Hem16b] Lane A. Hemaspaandra. SIGACT news complexity theory column 90: Introduction to complexity theory column 90. ACM SIGACT News, 47(1):41, March 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2016:SNCb

[Hem16c] Lane A. Hemaspaandra. SIGACT news complexity theory column 91. ACM SIGACT News, 47(2):65, June 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2017:SNC

[Hem17a] Lane Hemaspaandra. SIGACT News complexity theory column 96. ACM SIGACT News, 48(4):42, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2017:SNCa

[Hem17b] Lane A. Hemaspaandra. SIGACT News complexity theory column 93. ACM SIGACT News, 48(1):101, March 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2017:SNCb

[Hem17c] Lane A. Hemaspaandra. SIGACT News complexity theory column 94. ACM SIGACT News, 48(2):40, June 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2018:SNCa

[Hem18a] Lane A. Hemaspaandra. SIGACT news complexity theory column 97. ACM SIGACT News, 49(1):54, March 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2018:SNCb

[Hem18b] Lane A. Hemaspaandra. SIGACT news complexity theory column 98. ACM SIGACT News, 49(2):32, June 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2018:SNC

[Hem18c] Lane A. Hemaspaandra. SIGACT news complexity theory column 99. ACM SIGACT News, 49(3):48–50, September 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2019:SNCa

[Hem19a] Lane A. Hemaspaandra. SIGACT news complexity theory column 100. ACM SIGACT News, 50(1):35–37, March 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2019:SNCb

[Hem19b] Lane A. Hemaspaandra. SIGACT news complexity theory column 101. ACM SIGACT News, 50(2):29–30, June 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2019:SNCc

[Hem19c] Lane A. Hemaspaandra. SIGACT news complexity theory column 102. ACM SIGACT News, 50(3):51, September 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2019:SNC

[Hem19d] Lane A. Hemaspaandra. SIGACT news complexity theory column 103. ACM SIGACT News, 50(4):56, December 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3374857.3374869.

Hemaspaandra:2020:SNCa

[Hem20a] Lane A. Hemaspaandra. SIGACT news complexity theory column 104. ACM SIGACT News, 51(1):37, March 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3388392.3388400.

Hemaspaandra:2020:SNCb

[Hem20b] Lane A. Hemaspaandra. SIGACT news complexity theory column 105. ACM SIGACT News, 51(2):36–37, June 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3406678.3406687.

Hemaspaandra:2020:SNCc

[Hem20c] Lane A. Hemaspaandra. SIGACT news complexity theory column 106: Teaching models, computability, and complexity in time of COVID-19. ACM SIGACT News, 51(3):55–58, September 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3427361.
3427371.

Hemaspaandra:2020:SNCd

[Hem20d] Lane A. Hemaspaandra. SIGACT news complexity theory column 107. ACM SIGACT News, 51(4):47, December 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3444815.3444824.

Hemaspaandra:2021:SNCa

[Hem21a] Lane A. Hemaspaandra. SIGACT News complexity theory column 108. ACM SIGACT News, 52(1):41–46, March 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3457588.3457597.

Hemaspaandra:2021:SNCb

[Hem21b] Lane A. Hemaspaandra. SIGACT News complexity theory column 109. ACM SIGACT News, 52(2):45, June 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3471469.3471478.

Hemaspaandra:2021:SNCc

[Hem21c] Lane A. Hemaspaandra. SIGACT News complexity theory column 110. ACM SIGACT News, 52(3):37, September 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3494656.3494665.

Hemaspaandra:2021:SNC

[Hem21d] Lane A. Hemaspaandra. SIGACT News complexity theory column 111. ACM SIGACT News, 52(4):55, December 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3510382.3510391.

Hemaspaandra:2022:SNCa

[Hem22a] Lane A. Hemaspaandra. SIGACT News complexity theory column 112. ACM SIGACT News, 53(1):58, March 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3532737.3532745.

Hemaspaandra:2022:SNCb

[Hem22b] Lane A. Hemaspaandra. SIGACT news complexity theory column 113. ACM SIGACT News, 53(2):39, June 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3544979.3544988.

Hemaspaandra:2022:SNCc

[Hem22c] Lane A. Hemaspaandra. SIGACT news complexity theory column 114. ACM SIGACT News, 53(3):41–45, September 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3561064.3561071.

Hemaspaandra:2022:SNCd

[Hem22d] Lane A. Hemaspaandra. SIGACT news complexity theory column 115: Juris Hartmanis and two Golden Rules. ACM SIGACT News, 53(4):35–40, December 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3577971.3577977.

Hemaspaandra:2023:SNCa

[Hem23a] Lane A. Hemaspaandra. SIGACT news complexity theory column 116. ACM SIGACT News, 54(1):62, March 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3586165.3586174.

Hemaspaandra:2023:SNCb

[Hem23b] Lane A. Hemaspaandra. SIGACT news complexity theory column 117: Thirty years of complexity theory (columns). ACM SIGACT News, 54(1):82–89, March 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/

doi/10.1145/3586165.3586176.

Henze:1983:SGE

[Hen83] Ernst Henze. The solution of a general equation for the public key system. ACM SIGACT News, 15(1):45–49, Winter–Spring 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Henzinger:1993:BRB

[Hen93] Thomas A. Henzinger. Book review: Verifying Temporal Properties of Systems, by Julian Charles Bradfield. (Birkhäuser, 1992 viii + 113pp. ISBN 0-8176-3625-0. $49.50). ACM SIGACT News, 24(1):27–28, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Bra92].

Hendler:2009:BRB

[Hen09] Danny Hendler. Book review: Synchronization Algorithms and Concurrent Programming, by Gadi Taubenfeld (Pearson/Prentice Hall, 2006). ACM SIGACT News, 40(1):47–50, March 2009. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Tau06].

Herman:1977:BRB

[Her77] Gabor Herman. Book review: Automata, Languages, Development, by A. Lindenmayer and G. Rozenberg. North Holland Publishing Co. 1976. ACM SIGACT News, 9(1):16–18, January 1977. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [LR76].

Herlestam:1983:FCD

[Her83] Tore Herlestam. On the feasibility of computing discrete logarithms using Adleman’s subexponential algorithm. ACM SIGACT News, 15(1):50–55, Winter–Spring 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Herlihy:2000:BRB

[Her00a] Maurice Herlihy. Book review: Distributed Computing, by Attiya and Welch. ACM SIGACT News, 31(1):3, March 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [AW98, AW04].

Hertrampf:2000:AAM

[Her00b] Ulrich Hertrampf. Algebraic acceptance mechanisms for polynomial time machines. ACM SIGACT News, 31(2):22–33, June 1, 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Herbrich:2002:LKC

[Her02a] Ralf Herbrich. Learning Kernel Classifiers: Theory and Algorithms. Adaptive computation and machine learning. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-08306-X (hardcover). xx + 364 pp. LCCN Q325.5 .H48 2002.

Herman:2002:SSW

[Her02b] Ted Herman. Self-stabilization at WSS’01 and DISC’01. ACM SIGACT News, 33(1):54–57, March 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Herrmann:2013:RBC

[Her13] Sven Herrmann. Review of Bioinspired Computation in Combinatorial Optimization by Frank Neumann and Carsten Witt. ACM SIGACT News, 44(2):22–26, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hull:1972:CNA

[HES72] T. E. Hull, W. H. Enright, and A. E. Sedgwick. The correctness of numerical algorithms. ACM SIGACT News, 4(14):66–73, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hesselink:1992:PRU

[Hes92] Wim H. Hesselink. Programs, Recursion, and Unbounded Choice: Predicate-Transformation Semantics and Transformation Rules, volume 27 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-40436-3. xii + 223 pp. LCCN QA76.6 .H483 1992. URL http://

www.loc.gov/catdir/description/cam025/92232354.html;
http://www.loc.gov/catdir/toc/cam029/92232354.html.

Hestand:2002:MTD

[Hes02] P. Daniel Hestand. Mathematical theory of domains. ACM SIGACT News, 33(3):19–22, September 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hestand:2003:BRB

[Hes03] P. Daniel Hestand. Book review: Modal Logic: Cambridge tracts in theoretical computer science #53 by Patrick Blackburn, Maarten de Rijke, and Yde Venema. Cambridge University Press 2001. ACM SIGACT News, 34(2):15–19, June 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BdRV01, BvBW07].

Hemaspaandra:2000:MPC

[HG00] Lane A. Hemaspaandra and Christian Glaßer. A moment of perfect clarity I: the parallel census technique. ACM SIGACT News, 31 (3):37–42, September 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hartmanis:1976:IRC

[HH76] J. Hartmanis and J. E. Hopcroft. Independence results in computer science. ACM SIGACT News, 8(4):13–24, October 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hatami:2024:GCS

[HH24] Hamed Hatami and Pooya Hatami. Guest column: Structure in communication complexity and constant-cost complexity classes. ACM SIGACT News, 55(1):67–93, March 2024. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3654780.3654788.

Hemaspaandra:1998:WDC

[HHH98] Edith Hemaspaandra, Lane A. Hemaspaandra, and Harald Hempel. What’s up with downward collapse: using the easyhard technique to link Boolean and polynomial hierarchy collapses. ACM SIGACT News, 29(3):10–22, September 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hankerson:2000:CTC

[HHL00] Darrel R. Hankerson, D. G. Hoffman, and D. A. Leonard. Coding Theory and Cryptography: the Essentials, volume 234 of Monographs and textbooks in pure and applied mathematics; Pure and applied mathematics. Marcel Dekker, Inc., New York, NY, USA, second edition, 2000. ISBN 0-8247-0465-7. x + 350 pp. LCCN QA268 .C675 2000.

Harris:2000:CGT

[HHM00] John M. (John Michael) Harris, Jeffry L. Hirst, and Michael J. Mossinghoff. Combinatorics and Graph Theory. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-98736-3. xiii + 225 pp. LCCN QA165 .H37 2000. URL http://www.loc.

gov/catdir/enhancements/fy0816/99049806-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0816/99049806-t.html.

Harris:2008:CGT

[HHM08] John M. (John Michael) Harris, Jeffry L. Hirst, and Michael J. Mossinghoff. Combinatorics and Graph Theory. Undergraduate texts in mathematics, 0172-6056. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2008. ISBN 0-387-79710-6. xv + 381 pp. LCCN QA165

.H37 2008. URL http://www.loc.gov/catdir/enhancements/
fy0907/2008934034-d.html;   http://www.loc.gov/catdir/
enhancements/fy0907/2008934034-t.html.

Hemaspaandra:1997:RNL

[HHR97] Edith Hemaspaandra, Lane A. Hemaspaandra, and Jörg Rothe. Raising NP lower bounds to parallel NP lower bounds. ACM SIGACT News, 28(2):2–13, June 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Higham:2001:RD

[Hig01] Lisa Higham. Report on DISC’00. ACM SIGACT News, 32(1): 32–36, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hill:2007:HCD

[Hil07] Robin K. Hill. How close did kurt Gödel get to the university of wyoming? ACM SIGACT News, 38(2):87–90, June 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hindley:1997:BST

[Hin97] J. Roger Hindley. Basic Simple Type Theory, volume 42 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-46518-4. xi + 186 pp. LCCN QA76.7 .H55 1997. URL http://www.loc.gov/

catdir/description/cam027/95009058.html;   http://www.
loc.gov/catdir/samples/cam034/95009058.html;   http://
www.loc.gov/catdir/toc/cam027/95009058.html.

Hinek:2010:CRV

[Hin10] M. Jason Hinek. Cryptanalysis of RSA and Its Variants. Chapman and Hall/CRC cryptography and network security. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4200-7518-7 (hardcover). xviii + 268 pp. LCCN TK5102.94 .H56 2010.

Hirvensalo:2001:QC

[Hir01] Mika Hirvensalo. Quantum Computing. Natural computing series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-66783-0. xi + 190 pp. LCCN QA76.889 .H57 2001.

Hakimi:1984:MTCa

[HK84a] S. Louis Hakimi and Oded Kariv. Midwest theory of computation symposium: on a generalization of edge-coloring in graphs. ACM SIGACT News, 15(4):60, Winter–Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hakimi:1984:MTCb

[HK84b] S. Louis Hakimi and Oded Kariv. Midwest theory of computation symposium: on a generalization of edge-coloring in graphs. ACM SIGACT News, 16(1):60, Winter–Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hodge:2005:MVE

[HK05] Jonathan K. Hodge and Richard E. Klima. The Mathematics of Voting and Elections: a Hands-On Approach, volume 22 of Mathematical world, 1055-9426. American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3798-2. xiv + 226 pp. LCCN JF1001 .H63 2005.

Hellman:1983:NCE

[HKR83] Martin E. Hellman, Ehud D. Karnin, and Justin Reyneri. On the necessity of cryptanalytic exhaustive search. ACM SIGACT News, 15(1):40–44, Winter–Spring 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Harel:2000:DL

[HKT00] David Harel, Dexter Kozen, and Jerzy Tiuryn. Dynamic Logic. Foundations of computing. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-08289-6 (hardcover). xv + 459 pp. LCCN QA76.9.L63 H37 2000.

Harel:2001:DL

[HKT01] David Harel, Dexter Kozen, and Jerzy Tiuryn. Dynamic logic. ACM SIGACT News, 32(1):66–69, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Herlihy:2008:DCM

[HL08] Maurice Herlihy and Victor Luchangco. Distributed computing and the multicore revolution. ACM SIGACT News, 39(1):62–72, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hsu:2009:GTI

[HL09] Lih-Hsing Hsu and Cheng-Kuan Lin. Graph theory and interconnection networks. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-4481-8 (hardcover). xiii + 706 pp. LCCN QA166 .H78 2009. URL http://

www.loc.gov/catdir/toc/ecip0817/2008019917.html.

Hazay:2010:EST

[HL10] Carmit Hazay and Yehuda Lindell. Efficient Secure Two-Party Protocols: Techniques and Constructions. Information security and cryptography: texts and monographs. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-14302-4 (hardcover), 3-642-14303-2 (e-book). ISSN 1619-7100. xiii + 263 pp. LCCN Z103 .H39 2010eb.

Hakimi:2011:RD

[HM11] Michael Hakimi and Adam Morrison. A review of DISC 2011. ACM SIGACT News, 42(4):87–91, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hopcroft:2001:IATa

[HMU01a] John E. Hopcroft, Rajeev Motwani, and Jeffrey D. Ullman. Introduction to Automata Theory, Languages, and Computation, 2nd edition. ACM SIGACT News, 32(1):60–65, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [HMU01b, HMU07].

Hopcroft:2001:IATb

[HMU01b] John E. Hopcroft, Rajeev Motwani, and Jeffrey D. Ullman. Introduction to Automata Theory, Languages, and Computation. Addison-Wesley, Reading, MA, USA, second edition, 2001. ISBN 0-201-44124-1. xiv + 521 pp. LCCN QA267 .H56 2001.

Hopcroft:2007:IAT

[HMU07] John E. Hopcroft, Rajeev Motwani, and Jeffrey D. Ullman. Introduction to Automata Theory, Languages, and Computation. Pearson/Addison Wesley, Boston, MA, USA, third edition, 2007. ISBN 0-321-45536-3, 0-321-46225-4, 0-321-45537-1 (student access kit). xvii + 535 pp. LCCN QA267 .H56 2007. URL http://www.loc.

gov/catdir/toc/ecip0613/2006014263.html.

Halldorsson:2014:MWA

[HN14] Magnús Halldórsson and Calvin Newport. Making wireless algorithm theory more useful: five ideas from the 2013 workshop on realistic models for algorithms in wireless networks. ACM SIGACT News, 45(3):72–74, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2001:CTC

[HO01] Lane A. Hemaspaandra and Mitsunori Ogihara. The complexity theory companion. ACM SIGACT News, 32(4):66–68, December 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hochba:1997:AAN

[Hoc97a] Dorit S. Hochba. Approximation Algorithms for NP-hard Problems. ACM SIGACT News, 28(2):40–52, June 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hochbaum:1997:AAN

[Hoc97b] Dorit S. Hochbaum, editor. Approximation Algorithms for NP-Hard Problems. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1997. ISBN 0-534-94968-1 (hardcover). xxii + 596 pp. LCCN T57.7 .A68 1997.

Hofmann:2000:PLC

[Hof00] Martin Hofmann. Programming languages capturing complexity classes. ACM SIGACT News, 31(1):31–42, March 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Homer:2017:RCS

[Hom17] Steve Homer. Review of Crypto School by Joachim von zur Gathen. ACM SIGACT News, 48(3):10–13, September 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hoffstein:2008:IMC

[HPS08] Jeffrey Hoffstein, Jill Catherine Pipher, and Joseph H. Silverman. An Introduction to Mathematical Cryptography, volume 666 of Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-77993-0 (hardcover). xv + 523 pp. LCCN QA268 .H64 2008.

Harfst:2000:PBA

[HR00] Gregory C. Harfst and Edward M. Reingold. A potential-based amortized analysis of the union-find data structure. ACM SIGACT News, 31(3):86–95, September 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hamburger:2002:LLM

[HR02] Henry Hamburger and Dana Richards. Logic and language models for computer science. ACM SIGACT News, 33(1):67–70, March 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hromkovic:2003:AHP

[Hro03] Juraj Hromkovič. Algorithmics for Hard Problems: Introduction to Combinatorial Optimization, Randomization, Approximation, and Heuristics. Texts in theoretical computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2003. ISBN 3-540-44134-4. xiii + 544 pp. LCCN QA76.58 .H76 2003. URL http:

//www.loc.gov/catdir/enhancements/fy0817/2002032405-d.
html;   http://www.loc.gov/catdir/enhancements/fy0817/
2002032405-t.html.

Hromkovic:2009:AAH

[Hro09] Juraj Hromkovic. Algorithmic adventures: from knowledge to magic. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-85986-1. xiii + 363 pp. LCCN QA76.9.A43 H76 2009. URL http://site.ebrary.com/

lib/upcatalunya/docDetail.action?docID=10313472.

Herlihy:2012:CDC

[HRR12] Maurice Herlihy, Sergio Rajsbaum, and Michel Raynal. Computability in distributed computing: a Tutorial. ACM SIGACT News, 43(3):88–110, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:1995:W

[HRZ95] Lane A. Hemaspaandra, Ajit Ramachandran, and Marius Zimand. Worlds to die for. ACM SIGACT News, 26(4):5–15, December 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:1997:CTR

[HS97] Lane A. Hemaspaandra and Alan L. Selman, editors. Complexity Theory Retrospective II. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94973-9 (hardcover). xi + 339 pp. LCCN QA267.7

.C674 1997. URL http://www.loc.gov/catdir/enhancements/
fy0815/96029772-d.html.

Hemaspaandra:1998:WEC

[HS98] Lane A. Hemaspaandra and Alan L. Selman. Writing and editing complexity theory: tales and tools. ACM SIGACT News, 29(4):20– 27, December 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hardy:2001:CQC

[HS01] Yorick Hardy and W.-H. Steeb. Classical and Quantum Computing: with C++ and Java Simulations. Birkhäuser Verlag, Basel, Switzerland, 2001. ISBN 3-7643-6610-9. xxiii + 589 pp. LCCN

QA76.73.C153 H367 2001. URL http://www.loc.gov/catdir/
enhancements/fy0812/2001052712-d.html.

Herlihy:2012:TMB

[HS12] Maurice Herlihy and Nir Shavit. Transactional memory: beyond the first two decades. ACM SIGACT News, 43(4):101–103, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2018:CTD

[HS18] Lane A. Hemaspaandra and Holger Spakowski. Column: Team diagonalization. ACM SIGACT News, 49(3):51–61, September 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hohne:2019:SNO

[HSvS19] Felix Höhne, Sören Schmitt, and Rob van Stee. SIGACT news online algorithms column 35: 2019 in review. ACM SIGACT News, 50(4):77–92, December 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3374857.3374871.

Hohne:2020:SNO

[HSvS20] Felix Höhne, Sören Schmitt, and Rob van Stee. SIGACT news online algorithms column 36: 2020 in review. ACM SIGACT News, 51(4):89–107, December 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3444815.3444830.

Hohne:2021:SNO

[HSvS21] Felix Hohne, Soren Schmitt, and Rob van Stee. SIGACT News online algorithms column 38: 2021 in review. ACM SIGACT News, 52(4):80–96, December 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3510382.3510396.

Hemaspaandra:2003:TSF

[HT03] Lane A. Hemaspaandra and Leen Torenvliet. Theory of Semi-Feasible Algorithms. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-42200-5. x + 148 pp. LCCN QA76.9.A43 H46 2003. URL http://www.loc.

gov/catdir/enhancements/fy0816/2002029158-d.html.

Hopcroft:1979:IAT

[HU79] John E. Hopcroft and Jeffrey D. Ullman. Introduction to Automata Theory, Languages, and Computation. Addison-Wesley series in computer science. Addison-Wesley, Reading, MA, USA, 1979. ISBN 0-201-02988-X. x + 418 pp. LCCN QA267 .H56. US$20.00.

Huang:1998:CKM

[Hua98] Andy Huang. Computation of the Knuth–Morris–Pratt skip tables. ACM SIGACT News, 29(3):59–61, September 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Huang:2003:PTA

[Hua03] Xiaofei Huang. A polynomial-time algorithm for solving NP-hard problems in practice. ACM SIGACT News, 34(1):101–108, March 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Huang:2014:SNO

[Hua14] Zhiyi Huang. SIGACT News online algorithms column 25: Online primal dual: Beyond linear programs. ACM SIGACT News, 45 (4):105–119, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hull:1996:RP

[Hul96] Richard Hull. Report on PODS ’96. ACM SIGACT News, 27(4): 8–10, December 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hutchins:1970:SSC

[Hut70] Sandra E. Hutchins. Stochastic sources for context-free languages. ACM SIGACT News, 2(4):19, 1970. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Huth:2001:SCS

[Hut01] Michael Huth. Secure communicating systems: design, analysis, and implementation. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-80731-X. x + 283 pp. LCCN TK5102.85

.H88 2001. URL http://site.ebrary.com/lib/stanford/Doc?
id=10006816;   http://www.myilibrary.com?id=41927.

Huttel:2010:TTI

[Hüt10] Hans Hüttel. Transitions and Trees: an Introduction to Structural Operational Semantics. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19746-5 (hardcover), 0-521-14709-3 (paperback). xvii + 272 pp. LCCN QA76.7 .H88 2010. URL http:

//www.loc.gov/catdir/enhancements/fy1116/2010282941-b.
html;   http://www.loc.gov/catdir/enhancements/fy1116/
2010282941-d.html;   http://www.loc.gov/catdir/enhancements/
fy1116/2010282941-t.html.

Hemaspaandra:1995:SNCa

[HV95] Lane A. Hemaspaandra and Heribert Vollmer. The satanic notations: counting classes beyond #P and other definitional adventures. ACM SIGACT News, 26(1):2–13, March 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hallett:1994:CPC

[HW94] Michael T. Hallett and H. Todd Wareham. A compendium of parameterized complexity results. ACM SIGACT News, 25(3): 122–123, September 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hemaspaandra:2012:SNCd

[HW12] Lane A. Hemaspaandra and Ryan Williams. SIGACT News Complexity Theory Column 76: an atypical survey of typical-case heuristic algorithms. ACM SIGACT News, 43(4):70–89, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hardy:2009:AAC

[HWR09] Darel W. Hardy, Carol L. Walker, and Fred Richman. Applied algebra: codes, ciphers, and discrete algorithms. Discrete mathematics and its applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2009. ISBN 1-4200-7142-4 (hardcover). 410 pp. LCCN QA268 .H365 2009.

Ibarra:1991:RD

[Iba91] Oscar H. Ibarra. On resetting DLBA’s. ACM SIGACT News, 22(1):62–63, Winter 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Im:2011:TAL

[IMP11] Sungjin Im, Benjamin Moseley, and Kirk Pruhs. A tutorial on amortized local competitiveness in online scheduling. ACM SIGACT News, 42(2):83–97, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ingemarsson:1983:AIK

[Ing83] Ingemar Ingemarsson. Are all injective knapsacks partly solvable after multiplication modulo q? ACM SIGACT News, 15(1):56–60, Winter–Spring 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Iosevich:2007:VTA

[Ios07] Alex Iosevich. A View From the Top: Analysis, Combinatorics and Number Theory, volume 39 of Student mathematical library. American Mathematical Society, Providence, RI, USA, 2007. ISBN 0-8218-4397-4 (paperback). xiii + 136 pp. LCCN QA295 .I57 2007.

Irani:2005:APP

[IP05] Sandy Irani and Kirk Pruhs. Algorithmic problems in power management. ACM SIGACT News, 36(2):63–76, June 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Irakliotis:2012:BRB

[Ira12] Leo Irakliotis. Book review: Cryptanalytic Attacks on RSA, by Song Y. Yan. ACM SIGACT News, 43(1):14–16, March 2012. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Yan07].

Irwin:2003:BRBb

[Irw03a] Robert J. Irwin. Book review: Coding Theory and Cryptography: the Essentials, second edition, revised and expanded by D.R. Hankerson, et al. Marcel Dekker, 2000. ACM SIGACT News, 34 (4):17–21, December 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [HHL00].

Irwin:2003:BRBa

[Irw03b] Robert J. Irwin. Book review: Set Theory for Computing: From Decision Procedures to Declarative Programming with Sets, by Domenico Cantone, Eugenio Omodeo and Alberto Policriti. Springer-Verlag 2001. ACM SIGACT News, 34(3):4–11, September 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [COP01].

Irwin:2008:BRB

[Irw08] Robert J. Irwin. Book review: Derivation and Computation: Taking the Curry–Howard Correspondence Seriously, by Harold Simmons, Cambridge University Press, 2000. ACM SIGACT News, 39 (2):42–44, June 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sim00].

Istrate:2000:BRB

[Ist00] Gabriel Istrate. Book review: Parameterized Complexity, by R. Downey and M. Fellows. ACM SIGACT News, 31(4):13–15, December 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [DF99].

Istrate:2012:RBL

[Ist12] Gabriel Istrate. Review of Handbook of Large-Scale Random Networks by Bela Bollobás, Robert Kozma and Deszö Miklós. ACM SIGACT News, 43(2):25–28, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ivanov:2002:BRB

[Iva02a] Ivelin Ivanov. Book review: Combinatorial Optimization — Theory and Algorithms. ACM SIGACT News, 33(2):14–16, June 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [KV00, KV02, Kor06, KV08, KV12].

Ivanov:2002:PPP

[Iva02b] Ivelin Ivanov. Parallel processing and parallel algorithms: theory and computation. ACM SIGACT News, 33(4):12–14, December 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Isard:1970:TOQ

[IZ70] Stephen D. Isard and Arnold M. Zwicky. Three open questions in the theory of one-symbol Smullyan systems. ACM SIGACT News, 2(7):11–19, 1970. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Jackson:2006:SAL

[Jac06] Daniel Jackson. Software Abstractions: Logic, Language, and Analysis. MIT Press, Cambridge, MA, USA, 2006. ISBN 0-262-10114-9. xvi + 350 pp. LCCN QA76.76.D47 J29 2006. URL

http://library.stanford.edu/bks24\_id=12936.

Jaffe:1978:NSP

[Jaf78] Jeffrey Jaffe. A necessary and sufficient pumping lemma for regular languages. ACM SIGACT News, 10(2):48–49, Summer 1978. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

James:1998:BRB

[Jam98] Jerry James. Book review: Algorithms and Programming: Problems and Solutions, by Alexander Shen (Birkhäuser Boston, 1997). ACM SIGACT News, 29(3):48–52, September 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [She08, She10].

Jankowski:2008:BRBa

[Jan08a] Richard Jankowski. Book review: Foundations of Computer Security, by David Salomon, Springer-Verlag, 2006. ACM SIGACT News, 39(2):40–41, June 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sal05b].

Jankowski:2008:BRBb

[Jan08b] Richard Jankowski. Book review: Privacy on the Line: The Politics of Wiretapping and Encryption, by Whitfield Diffie and Susan Landau. ACM SIGACT News, 39(4):30–32, December 2008. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [DL98, DL07].

Jankowski:2010:BRBb

[Jan10a] Richard Jankowski. Book review: Advanced Data Structures, by Peter Brass, Cambridge University Press 2008. ACM SIGACT News, 41(1):19–20, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Bra08].

Jankowski:2010:BRBa

[Jan10b] Richard Jankowski. Book review: Data Structures and Algorithms Using Python and C++, by David M. Reed and John Zelle Franklin, Beedle and Associates 2009. ACM SIGACT News, 41(1): 13–15, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [RZ09].

Jawor:2005:TDP

[Jaw05] Wojciech Jawor. Three dozen papers on online algorithms. ACM SIGACT News, 36(1):71–85, March 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Jensen:1992:CPN

[Jen97] K. (Kurt) Jensen. Coloured Petri Nets: Basic Concepts, Analysis Methods, and Practical Use, volume 26 of v. 1: EATCS monographs on theoretical computer science; v. 2–3: Monographs in theoretical computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992– 1997. ISBN 3-540-55597-8 (v. 1), 0-387-55597-8 (New York : v. 1), 3-540-58276-2 (v. 2 : acid-free), 0-387-58276-2 (New York : v. 2 : acid-free), 3-540-62867-3 (v. 3). ???? pp.

LCCN QA267 .J46 1992. URL http://www.loc.gov/catdir/
enhancements/fy0817/92227749-d.html;   http://www.loc.
gov/catdir/enhancements/fy0817/92227749-t.html.

Jennings:2001:BRB

[Jen01] Christopher G. Jennings. Book review: Analysis of Algorithms: An Active Learning Approach, by Jeffrey J. McConnell. ACM SIGACT News, 32(4):17–18, December 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [McC01].

Jackman:2020:RKT

[JH20] Tim Jackman and Steve Homer. Review of Kernelization: Theory of Parameterized Preprocessing by Fedor V. Fomin, Daniel Lokshtanov, Saket Saurabh, and Meirav Zehavi. ACM SIGACT News, 51(4):11–14, December 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3444815.3444819.

Jacob:1992:BRB

[JK92] Tom Jacob and Bill Kaizer. Book review: A Unifying Framework for Structured Analysis and Design Models. By T. H. Tse. (Cambridge University Press, 1991. xi + 179pp. ISBN 0-521-39196-2). ACM SIGACT News, 23(3):30, Summer 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Tse91].

Jiang:1999:SOP

[JKL99] Tao Jiang, Paul Kearney, and Ming Li. Some open problems in computational molecular biology. ACM SIGACT News, 30(3):43– 49, September 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Jakobsson:2007:DPD

[JM07] Markus Jakobsson and Steven Myers. Delayed password disclosure. ACM SIGACT News, 38(3):56–75, September 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Johnson:1984:GTC

[Joh84] David S. Johnson. The genealogy of theoretical computer science: a preliminary report. ACM SIGACT News, 16(2):36–49, Summer 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Johnson:1990:SFB

[Joh90] David S. Johnson. A STOC/FOCS bibliography: the last progress report. ACM SIGACT News, 21(2):4, Spring 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Johnson:1996:BHS

[Joh96a] David S. Johnson. A Brief History of SIGACT News and its Editors. ACM SIGACT News, 27(3):125, September 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Johnson:1996:HDE

[Joh96b] David S. Johnson. How to do experiments (extended advertisement). ACM SIGACT News, 27(2):87, June 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Johnson:2010:BRB

[Joh10] Samuel Johnson. Book review: Branching Programs and Binary Decision Diagrams: Theory and Applications, by Ingo Wegener, Society for Industrial and Applied Mathematics, 2000, 408 pages. ACM SIGACT News, 41(3):36–38, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Weg00].

Jones:1972:FDC

[Jon72] C. B. Jones. Formal development of correct algorithms: an example based on Earley’s recogniser. ACM SIGACT News, 4(14):150– 169, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Jones:1973:MOS

[Jon73] Louise H. Jones. Microprogramming: an opportunity for SIGACT. ACM SIGACT News, 5(4):9–11, October 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Jordan:1983:VPK

[Jor83] John P. Jordan. A variant of a public key cryptosystem based on Goppa Codes. ACM SIGACT News, 15(1):61–66, Winter–Spring 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Jain:1999:SLI

[JORS99] Sanjay Jain, Daniel Osherson, James S. Royer, and Arun Sharma. Systems That Learn: an Introduction to Learning Theory. Learning, development, and conceptual change. MIT Press, Cambridge, MA, USA, second edition, 1999. ISBN 0-262-10077-0 (hardcover). xii + 317 pp. LCCN BF318 .O83 1999.

Jouannaud:1994:BRB

[Jou94] Jean-Pierre Jouannaud. Book review: A Proof Theory for General Unification. By Wayne Snyder. (Birkhäuser, 1991. vi + 175 pages. ISBN 0-8176-3593-9. $28.00). ACM SIGACT News, 25(2):25, June 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sny91].

Joux:2009:AC

[Jou09] Antoine Joux. Algorithmic Cryptanalysis. Chapman and Hall/CRC cryptography and network security. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-7002-9 (hardcover). 501 pp. LCCN QA76.9.A43; QA76.9.A43 J693 2009.

Jarvinen:2000:TE

[JP00] J. Järvinen and T. Pasanen. Two exercises. ACM SIGACT News, 31(1):75–76, March 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Johnsonbaugh:2003:A

[JS03] Richard Johnsonbaugh and Marcus Schaefer. Algorithms. Pearson Education, Upper Saddle River, NJ, USA, 2003. ISBN 0-02-360692-4, 0-13-122853-6, 81-317-0868-3. xiii + 752 pp. LCCN Q A76.9.A43 J67 2003.

Julstrom:1992:BSF

[Jul92] Bryant A. Julstrom. A bound on the shift function in terms of the Busy Beaver function. ACM SIGACT News, 23(3):100–106, Summer 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Jungnickel:1999:GNA

[Jun99] D. (Dieter) Jungnickel. Graphs, Networks, and Algorithms, volume 5 of Algorithms and computation in mathematics, 1431-1550. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-63760-5. xii + 589 pp. LCCN QA402.5 .J8913 1999.

Jungnickel:2005:GNA

[Jun05] D. (Dieter) Jungnickel. Graphs, Networks, and Algorithms, volume 5 of Algorithms and computation in mathematics, 1431-1550. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2005. ISBN 3-540-21905-6. xvi + 611 pp. LCCN QA402.5 .J8913 2005. URL http:

//www.loc.gov/catdir/enhancements/fy0826/2004112299-d.
html;   http://www.loc.gov/catdir/enhancements/fy0826/
2004112299-t.html.

Jungnickel:2008:GNA

[Jun08] D. (Dieter) Jungnickel. Graphs, Networks, and Algorithms, volume 5 of Algorithms and computation in mathematics, 1431-1550. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2008. ISBN 3-540-72779-5. xix + 650 pp. LCCN QA402.5 .J8913 2008. URL http://www.loc.gov/

catdir/toc/fy0804/2007929271.html.

Kanellakis:1989:DBR

[KA89a] P. Kanellakis and S. Abiteboul. Deciding bounded recursion in database logic programs. ACM SIGACT News, 20:4(??), Fall 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kanellakis:1989:DTC

[KA89b] Paris C. Kanellakis and Serge Abiteboul. Database theory column: Deciding bounded recursion in database logic programs. ACM SIGACT News, 20(4):17–23, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kahn:1967:CSSa

[Kah67a] David Kahn. The Codebreakers: the Story of Secret Writing. Macmillan Publishing Company, New York, NY, USA, 1967. xvi

+ 1164 pp. LCCN Z103 .K28. URL http://www.cl.cam.ac.uk/
~fapp2/steganography/bibliography/1000.html.

Kahn:1967:CSSb

[Kah67b] David Kahn. The Codebreakers: the Story of Secret Writing. Weidenfeld and Nicolson, London, UK, 1967. xvi + 1164 pp. LCCN Z103 .K28 1967.

Kahn:1974:C

[Kah74] David Kahn. The Codebreakers. Weidenfeld and Nicolson, London, UK, abridged edition, 1974. ISBN 0-297-76785-2. xvi + 576 pp. LCCN Z103 .K28 1974.

Kahn:1996:CSS

[Kah96] David Kahn. The Codebreakers: the Story of Secret Writing. Scribner, New York, NY, USA, revised edition, 1996. ISBN 0-684-83130-9. xviii + 1181 pp. LCCN Z103 .K28 1996. See [Tuc66].

Kaminski:2013:RBM

[Kam13] Marcin Kamiński. Review of Boolean Models and Methods in Mathematics, Computer Science, and Engineering by Yves Crama and Peter L. Hammer. ACM SIGACT News, 44(1):21–24, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kantabutra:1988:LBP

[Kan88] V. Kantabutra. A lower bound on the path length of binary trees. ACM SIGACT News, 19(2):48–50, Summer 1988. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kant:1993:NMP

[Kan93] Goos Kant. A new method for planar graph drawings on a grid (abstract). ACM SIGACT News, 24(1):60, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kanellakis:1994:DQC

[Kan94] Paris Kanellakis. Database querying and constraint programming. ACM SIGACT News, 25(4):22–87, December 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kantorovitz:2005:NTM

[Kan05] Isaiah Pinchas Kantorovitz. A note on Turing machine computability of rule driven systems. ACM SIGACT News, 36(4):109–110, December 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kannan:2014:KPC

[Kan14] Ravindran Kannan. 14th Knuth prize: call for nominations. ACM SIGACT News, 45(1):7–8, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kao:1997:MSD

[Kao97] Ming-Yang Kao. Multiple-size divide-and-conquer recurrences. ACM SIGACT News, 28(2):67–69, June 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kaplan:1999:NNH

[Kap99] Robert Kaplan. The Nothing That Is: a Natural History of Zero. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-512842-7, 0-19-514237-3. xii + 225 pp. LCCN QA141 .K36 1999. US$22.00.

Kari:2000:HCA

[Kar00] Lila Kari. Half century of automata theory. ACM SIGACT News, 31(4):121–124, December 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kartha:2011:BRB

[Kar11] Neelakantan Kartha. Book review: The Algorithm Design Manual, second edition by Steven S. Skiena. ACM SIGACT News, 42 (4):29–31, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Ski98, Ski08].

Kasman:2005:RCS

[Kas05] Alex Kasman. Reality conditions: short mathematical fiction. Spectrum series. Mathematical Association of America, Washington, DC, USA, 2005. ISBN 0-88385-552-6. ix + 247 pp. LCCN

PS3611.A7854 R43 2005. URL http://www.loc.gov/catdir/
toc/fy0614/2005924781.html.

Katseff:1978:RFT

[Kat78] Howard P. Katseff. From the recursive function theory newsletter. ACM SIGACT News, 9(4):22–23, Winter 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Katz:2005:CBR

[Kat05] Jonathan Katz. Comparative book review: Cryptography: An Introduction, by V. V. Yaschenko (American Mathematical Society, 2002); Cryptanalysis of Number Theoretic Ciphers, by S. S. Wagstaff, Jr. (Chapman & Hall/CRC Press, 2003); RSA and Public-Key Cryptography, by R. A. Mollin (Chapman & Hall/CRC Press, 2003); Foundations of Cryptography, vol. 1: Basic Tools, by O. Goldreich, (Cambridge University Press, 2001). ACM SIGACT News, 36(2):14–19, June 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Gol01, Wag03, Mol03].

Katz:2006:BRBa

[Kat06a] Jonathan Katz. Book review: A Computational Introduction to Number Theory and Algebra, by Victor Shoup, Cambridge University Press, 2005. ACM SIGACT News, 37(1):12–13, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sho05, Sho09].

Katz:2006:BRBb

[Kat06b] Jonathan Katz. Book review: Primality Testing in Polynomial Time, by Martin Dietzfelbinger, Springer-Verlag, 2004. ACM SIGACT News, 37(1):14–15, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Die04].

Katz:2007:BRB

[Kat07] Jonathan Katz. Book review: Probability and Computing: Randomized Algorithms and Probabilistic Analysis, by Michael Mitzenmacher and Eli Upfal, Cambridge University Press, 2005. ACM SIGACT News, 38(3):20–22, September 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [MU05].

Katz:2012:BRB

[Kat12] Jonathan Katz. Book review: Efficient Secure Two-Party Protocols: Techniques and Constructions, by Carmit Hazay and Yehuda Lindell. ACM SIGACT News, 43(1):21–23, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [HL10].

Katz:2013:RIB

[Kat13a] Jon Katz. Review of Identity-based encryption by Sanjit Chattarjee and Palash Sarkar. ACM SIGACT News, 44(4):29–31, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Katz:2013:RBI

[Kat13b] Jonathan Katz. Review of Applied Information Security by David Basin, Patrick Schaller, and Michael Schläpfer. ACM SIGACT News, 44(1):38–40, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Katz:2014:RNL

[Kat14a] Jonathan Katz. Review of Networked life: 20 questions and answers by Mung Chiang. ACM SIGACT News, 45(3):21–23, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Katz:2014:RBC

[Kat14b] Jonathan Katz. Review of The block cipher companion by Lars R. Knudsen and Matthew J. B. Robshaw. ACM SIGACT News, 45 (3):18–20, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kaufmann:1993:ARS

[Kau93] Michael Kaufmann. Angular resolution of straight-line drawings (abstract). ACM SIGACT News, 24(1):59, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kayal:2018:GCP

[Kay18] Neeraj Kayal. Guest column: a paradigm for arithmetic circuit lower bounds. ACM SIGACT News, 49(1):55–65, March 2018. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Krishnamoorthy:1979:GTH

[KB79a] M. S. Krishnamoorthy and Somenath Biswas. The Generalized Towers of Hanoi. ACM SIGACT News, 10(3):49, Winter 1979. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Krishnamoorthy:1979:RTR

[KB79b] M. S. Krishnamoorthy and Somenath Biswas. Recent technical reports. ACM SIGACT News, 10(3):50–94, Winter 1979. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2007:I

[Kei07] Idit Keidar. Introduction. ACM SIGACT News, 38(4):46–49, December 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2008:ASNa

[Kei08a] Idit Keidar. ACM SIGACT News Distributed Computing Column 29. ACM SIGACT News, 39(1):47–48, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2008:ASNb

[Kei08b] Idit Keidar. ACM SIGACT news distributed computing column 31: quantum computers meet distributed computing. ACM SIGACT News, 39(3):66, September 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2008:ASNc

[Kei08c] Idit Keidar. ACM SIGACT news distributed computing column 32: the year in review. ACM SIGACT News, 39(4):53–54, December 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2008:CFP

[Kei08d] Idit Keidar. The complexity of fixed-parameter problems: guest column. ACM SIGACT News, 39(1):33–46, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2008:DCP

[Kei08e] Idit Keidar. On distributed computing principles in systems research: introduction. ACM SIGACT News, 39(2):82–83, June 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2009:ASNa

[Kei09a] Idit Keidar. ACM SIGACT news distributed computing column 33 teaching concurrency. ACM SIGACT News, 40(1):45–46, March 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2009:ASNb

[Kei09b] Idit Keidar. ACM SIGACT news distributed computing column 34: distributed computing in the clouds. ACM SIGACT News, 40 (2):67, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2009:ASNc

[Kei09c] Idit Keidar. ACM SIGACT news distributed computing column 35: theory and practice in large distributed systems. ACM SIGACT News, 40(3):77, September 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2009:DCC

[Kei09d] Idit Keidar. Distributed computing column 36: Distributed computing: 2009 edition. ACM SIGACT News, 40(4):64–67, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2010:DCCa

[Kei10a] Idit Keidar. Distributed computing column 37: Reconfiguring state machines … and the history of common knowledge. ACM SIGACT News, 41(1):57, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2010:DCCb

[Kei10b] Idit Keidar. Distributed Computing Column 38: Models for algorithm design in wireless networks. ACM SIGACT News, 41(2):73, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2010:DCCc

[Kei10c] Idit Keidar. Distributed Computing Column 39: Byzantine Generals: The Next Generation. ACM SIGACT News, 41(3):88, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2010:DCCd

[Kei10d] Idit Keidar. Distributed computing column 40: Annual review 2010. ACM SIGACT News, 41(4):95–99, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2011:DCCa

[Kei11a] Idit Keidar. Distributed computing column 41: Computing over dynamic networks. ACM SIGACT News, 42(1):81, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2011:DCCb

[Kei11b] Idit Keidar. Distributed computing column 42: Game theory and fault tolerance in distributed computing. ACM SIGACT News, 42 (2):68, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2011:DCCc

[Kei11c] Idit Keidar. Distributed computing column 43: Using social networks to overcome Sybil attacks. ACM SIGACT News, 42(3): 79, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2011:DCCd

[Kei11d] Idit Keidar. Distributed computing column 44: 2011 in review. ACM SIGACT News, 42(4):76–78, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2012:DCCa

[Kei12a] Idit Keidar. Distributed computing column 45: What theory for transactional memory? ACM SIGACT News, 43(1):86, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2012:DCCb

[Kei12b] Idit Keidar. Distributed computing column 46: synthesizing distributed and concurrent programs. ACM SIGACT News, 43(2):84, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2012:DCCc

[Kei12c] Idit Keidar. Distributed computing column 47: distributed computability. ACM SIGACT News, 43(3):87, September 2012. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2012:DCCd

[Kei12d] Idit Keidar. Distributed computing column 48: annual review 2012. ACM SIGACT News, 43(4):98–100, December 2012. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2013:DCCa

[Kei13a] Idit Keidar. Distributed computing column 49: coding for distributed storage. ACM SIGACT News, 44(1):79, March 2013. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2013:DCCb

[Kei13b] Idit Keidar. Distributed computing column 50: distributing trusted third parties, innovation prize, and SIROCCO review. ACM SIGACT News, 44(2):88, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2013:DCCc

[Kei13c] Idit Keidar. Distributed computing column 51: large-scale transaction replication. ACM SIGACT News, 44(3):72, September 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2005:TRC

[Kel05] Dean Kelley. Technical report column. ACM SIGACT News, 36 (3):22–23, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2006:TRCa

[Kel06a] Dean Kelley. Technical report column. ACM SIGACT News, 37(1): 29–32, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2006:TRCb

[Kel06b] Dean Kelley. Technical report column. ACM SIGACT News, 37 (2):28–30, June 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2006:TRCc

[Kel06c] Dean Kelley. Technical report column. ACM SIGACT News, 37 (3):33–35, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2006:TRCd

[Kel06d] Dean Kelley. Technical report column. ACM SIGACT News, 37 (4):44–46, December 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2007:TRCa

[Kel07a] Dean Kelley. Technical report column. ACM SIGACT News, 38(1): 28–30, March 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2007:TRCb

[Kel07b] Dean Kelley. Technical report column. ACM SIGACT News, 38 (2):26–28, June 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2007:TRCc

[Kel07c] Dean Kelley. Technical report column. ACM SIGACT News, 38 (3):31–33, September 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2007:TRCd

[Kel07d] Dean Kelley. Technical report column. ACM SIGACT News, 38 (4):35–38, December 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2008:TRCa

[Kel08a] Dean Kelley. Technical report column. ACM SIGACT News, 39(1): 30–32, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2008:TRCb

[Kel08b] Dean Kelley. Technical report column. ACM SIGACT News, 39 (2):47–49, June 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2008:TRCc

[Kel08c] Dean Kelley. Technical report column. ACM SIGACT News, 39 (3):33–34, September 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2008:TRCd

[Kel08d] Dean Kelley. Technical report column. ACM SIGACT News, 39 (4):33–34, December 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2008:BRB

[Kel08e] Timothy Kelley. Book review: A Course in Computational Algebraic Number Theory, by Henri Cohen, Springer, 2000. ACM SIGACT News, 39(2):36–39, June 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Coh93a, Coh95, Coh96].

Kelley:2009:BRB

[Kel09a] Dean Kelley. Book review: A Programmer’s Companion to Algorithm Analysis, by Ernst Leiss (Chapman & Hall/CRC, 2007). ACM SIGACT News, 40(2):18–22, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Lei07].

Kelley:2009:JRB

[Kel09b] Dean Kelley. Joint review of Algorithms, by Richard Johnsonbaugh and Marcus Schaefer (Pearson/Prentice-Hall, 2004) and Algorithms by Sanjoy Dasgupta, Christos Papadimitriou and Umesh Vazirani (McGraw-Hill, 2008). ACM SIGACT News, 40(2):23–25, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [JS03, DPV08].

Kelley:2009:TRCa

[Kel09c] Dean Kelley. Technical report column. ACM SIGACT News, 40(1): 24–25, March 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2009:TRCb

[Kel09d] Dean Kelley. Technical report column. ACM SIGACT News, 40 (2):42–44, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2009:TRCc

[Kel09e] Dean Kelley. Technical report column. ACM SIGACT News, 40 (3):51–52, September 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2009:TRCd

[Kel09f] Dean Kelley. Technical report column. ACM SIGACT News, 40 (4):42–44, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2010:TRCa

[Kel10a] Dean Kelley. Technical report column. ACM SIGACT News, 41(1): 54–56, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2010:TRCb

[Kel10b] Dean Kelley. Technical report column. ACM SIGACT News, 41 (4):55–57, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2011:TRCa

[Kel11a] Dean Kelley. Technical report column. ACM SIGACT News, 42(1): 55–57, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2011:TRCb

[Kel11b] Dean Kelley. Technical report column. ACM SIGACT News, 42 (2):47–50, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2011:TRCc

[Kel11c] Dean Kelley. Technical report column. ACM SIGACT News, 42 (3):50–52, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2011:TRCd

[Kel11d] Dean Kelley. Technical report column. ACM SIGACT News, 42 (4):50–52, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2012:TRCa

[Kel12a] Dean Kelley. Technical report column. ACM SIGACT News, 43(1): 40–43, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2012:TRCb

[Kel12b] Dean Kelley. Technical report column. ACM SIGACT News, 43 (2):41–44, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2012:TRCc

[Kel12c] Dean Kelley. Technical report column. ACM SIGACT News, 43 (3):56–59, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2012:TRCd

[Kel12d] Dean Kelley. Technical report column. ACM SIGACT News, 43 (4):55–56, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2013:TRCa

[Kel13a] Dean Kelley. Technical report column. ACM SIGACT News, 44(1): 46–48, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2013:TRCb

[Kel13b] Dean Kelley. Technical report column. ACM SIGACT News, 44 (2):42–45, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2013:TRCc

[Kel13c] Dean Kelley. Technical report column. ACM SIGACT News, 44 (3):40–41, September 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2013:TRC

[Kel13d] Dean Kelley. Technical report column. ACM SIGACT News, 44 (4):49–51, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2014:TRCa

[Kel14a] Dean Kelley. Technical report column. ACM SIGACT News, 45(1): 44–46, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2014:TRCb

[Kel14b] Dean Kelley. Technical report column. ACM SIGACT News, 45 (2):43–45, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2014:TRCc

[Kel14c] Dean Kelley. Technical report column. ACM SIGACT News, 45 (3):37–46, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2014:TRCd

[Kel14d] Dean Kelley. Technical report column. ACM SIGACT News, 45 (4):48–57, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2015:TRCa

[Kel15a] Dean Kelley. Technical report column. ACM SIGACT News, 46(1): 28–38, March 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2015:TRCb

[Kel15b] Dean Kelley. Technical report column. ACM SIGACT News, 46 (2):28–39, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2015:TRCc

[Kel15c] Dean Kelley. Technical report column. ACM SIGACT News, 46 (3):23–35, September 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2015:TRCd

[Kel15d] Dean Kelley. Technical report column. ACM SIGACT News, 46 (4):19–30, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelk:2016:RGS

[Kel16a] Steven Kelk. Review of: Genome-Scale Algorithm Design (Biological sequence analysis in the era of high-throughput sequencing) by Veli Mäkinen, Djamal Belazzougui, Fabio Cunial and Alexandru I. Tomescu. ACM SIGACT News, 47(3):15–18, September 2016. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelk:2016:RRA

[Kel16b] Steven Kelk. Review of ReCombinatorics: The Algorithmics of Ancestral Recombination Graphs and Explicit Phylogenetic Networks by Dan Gusfield. ACM SIGACT News, 47(1):12–15, March 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2016:TRCa

[Kel16c] Dean Kelley. Technical report column. ACM SIGACT News, 47(1): 29–40, March 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2016:TRCb

[Kel16d] Dean Kelley. Technical report column. ACM SIGACT News, 47 (2):21–33, June 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2016:TRCc

[Kel16e] Dean Kelley. Technical report column. ACM SIGACT News, 47 (3):25–38, September 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2016:TRCd

[Kel16f] Dean Kelley. Technical report column. ACM SIGACT News, 47 (4):22–32, December 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2017:TRCa

[Kel17a] Dean Kelley. Technical report column. ACM SIGACT News, 48(1): 27–37, March 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2017:TRCb

[Kel17b] Dean Kelley. Technical report column. ACM SIGACT News, 48 (2):21–33, June 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2017:TRCc

[Kel17c] Dean Kelley. Technical report column. ACM SIGACT News, 48 (3):20–31, September 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2017:TRCd

[Kel17d] Dean Kelley. Technical report column. ACM SIGACT News, 48 (4):22–34, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2018:TRCa

[Kel18a] Dean Kelley. Technical report column. ACM SIGACT News, 49(1): 29–39, March 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2018:TRCb

[Kel18b] Dean Kelley. Technical report column. ACM SIGACT News, 49 (2):18–28, June 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2018:TRCc

[Kel18c] Dean Kelley. Technical report column. ACM SIGACT News, 49 (3):23–33, September 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2018:TRCd

[Kel18d] Dean Kelley. Technical report column. ACM SIGACT News, 49 (4):17–27, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2019:TRCa

[Kel19a] Dean Kelley. Technical report column. ACM SIGACT News, 50(1): 17–27, March 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2019:TRCb

[Kel19b] Dean Kelley. Technical report column. ACM SIGACT News, 50 (2):15–27, June 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2019:TRCc

[Kel19c] Dean Kelley. Technical report column. ACM SIGACT News, 50 (3):12–23, September 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kelley:2019:TRC

[Kel19d] Dean Kelley. Technical report column. ACM SIGACT News, 50 (4):14–25, December 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3374857.3374862.

Kelley:2020:TRCa

[Kel20a] Dean Kelley. Technical report column. ACM SIGACT News, 51(1): 15–26, March 2020. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3388392.3388397.

Kelley:2020:TRCb

[Kel20b] Dean Kelley. Technical report column. ACM SIGACT News, 51 (2):15–26, June 2020. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3406678.3406684.

Kelley:2020:TRCc

[Kel20c] Dean Kelley. Technical report column. ACM SIGACT News, 51 (3):20–30, September 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3427361.3427367.

Kelley:2020:TRCd

[Kel20d] Dean Kelley. Technical report column. ACM SIGACT News, 51 (4):18–29, December 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3444815.3444821.

Kelley:2021:TRCa

[Kel21a] Dean Kelley. Technical report column. ACM SIGACT News, 52(1): 25–35, March 2021. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3457588.3457594.

Kelley:2021:TRCb

[Kel21b] Dean Kelley. Technical report column. ACM SIGACT News, 52 (2):24–35, June 2021. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3471469.3471475.

Kelley:2021:TRCc

[Kel21c] Dean Kelley. Technical report column. ACM SIGACT News, 52 (3):14–24, September 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3494656.3494661.

Kelley:2021:TRC

[Kel21d] Dean Kelley. Technical report column. ACM SIGACT News, 52 (4):18–30, December 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3510382.3510388.

Kelley:2022:TRCa

[Kel22a] Dean Kelley. Technical report column. ACM SIGACT News, 53(1): 25–35, March 2022. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3532737.3532742.

Kelley:2022:TRCb

[Kel22b] Dean Kelley. Technical report column. ACM SIGACT News, 53 (2):15–25, June 2022. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3544979.3544984.

Kelley:2022:TRCc

[Kel22c] Dean Kelley. Technical report column. ACM SIGACT News, 53 (3):24–35, September 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3561064.3561069.

Kelley:2023:TRCa

[Kel23a] Dean Kelley. Technical report column. ACM SIGACT News, 54(1): 22–36, March 2023. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3586165.3586171.

Kelley:2023:TRCb

[Kel23b] Dean Kelley. Technical report column. ACM SIGACT News, 54 (2):15–25, June 2023. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3604943.3604946.

Kelley:2023:TRCc

[Kel23c] Dean Kelley. Technical report column. ACM SIGACT News, 54 (3):22–32, September 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3623800.3623803.

Kelley:2023:TRCd

[Kel23d] Dean Kelley. Technical report column. ACM SIGACT News, 54 (4):27–40, December 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3639528.3639531.

Kelley:2024:TRC

[Kel24] Dean Kelley. Technical report column. ACM SIGACT News, 55(1): 20–31, March 2024. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3654780.3654783.

Kenschaft:2005:CPS

[Ken05] Patricia C. Kenschaft. Change is Possible: Stories of Women and Minorities in Mathematics. American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3748-6. ix + 212 pp. LCCN QA10.5 .K46 2005.

Kramer:2010:FDC

[KGO10] Simon Kramer, Rajeev Goré, and Eiji Okamoto. Formal definitions and complexity results for trust relations and trust domains fit for TTPs, the web of trust, PKIs, and ID-based cryptography. ACM SIGACT News, 41(1):75–98, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khot:2005:GCI

[Kho05] Subhash Khot. Guest column: inapproximability results via Long Code based PCPs. ACM SIGACT News, 36(2):25–42, June 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:1989:OP

[Khu89] S. Khuller. Open problems: 3. ACM SIGACT News, 20(4):24, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:1990:OP

[Khu90] Samir Khuller. Open problems. ACM SIGACT News, 21(1):12, January 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:1995:OP

[Khu95] Samir Khuller. Open Problems: 11. ACM SIGACT News, 26 (1):33, March 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:1996:OPa

[Khu96a] Samir Khuller. Open Problems: 13. ACM SIGACT News, 27(2): 52–54, June 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:1996:OPb

[Khu96b] Samir Khuller. Open Problems 14. ACM SIGACT News, 27(4): 11, December 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:1997:OP

[Khu97] Samir Khuller. Open problems: 15. ACM SIGACT News, 28 (3):33–36, September 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:1998:BRB

[Khu98a] Samir Khuller. Book review: Selected Papers on Computer Science, by Donald E. Knuth. ACM SIGACT News, 29(2):21–26, June 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Knu96].

Khuller:1998:OP

[Khu98b] Samir Khuller. Open problems: 16. ACM SIGACT News, 29(1): 15–17, March 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:2001:AC

[Khu01] Samir Khuller. Algorithms column. ACM SIGACT News, 32(2): 28–31, June 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:2002:ACV

[Khu02] Samir Khuller. Algorithms column: the vertex cover problem. ACM SIGACT News, 33(2):31–33, June 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:2004:PST

[Khu04] Samir Khuller. A pleasant stroll through the land of infinitely many creatures. ACM SIGACT News, 35(2):36–59, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:2005:FCS

[Khu05a] Samir Khuller. Four colors suffice! ACM SIGACT News, 36(2): 43–44, June 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:2005:GCI

[Khu05b] Samir Khuller. Guest column: inapproximability results via Long Code based PCPs. ACM SIGACT News, 36(2):25–42, June 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:2006:TRC

[Khu06] Samir Khuller. Technical report column. ACM SIGACT News, 37(1):29–32, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:2007:I

[Khu07] Samir Khuller. Introduction. ACM SIGACT News, 38(3):106, September 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khuller:2012:ACO

[Khu12] Samir Khuller. Algorithms column: an overview of the recent progress on matrix multiplication by Virginia Vassilevska Williams. ACM SIGACT News, 43(4):57–59, December 2012. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kilov:2012:RBF

[Kil12] Haim Kilov. Review of Origins and Foundations of Computing, by Friedrich L. Bauer. ACM SIGACT News, 43(4):38–40, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kilov:2014:RUCa

[Kil14a] Haim Kilov. Review of The universal computer: the road from Leibniz to Turing by Martin Davis. ACM SIGACT News, 45(3):29– 31, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kilov:2014:RUCb

[Kil14b] Haim Kilov. Review of The Universal Computer. The Road from Leibniz to Turing by Martin Davis. ACM SIGACT News, 45(4):17– 20, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kindahl:2005:BRB

[Kin05] Mats Kindahl. Book review: Types and Programming Languages, by Benjamin C. Pierce; The MIT Press, 2002. ACM SIGACT News, 36(1):15–20, March 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Pie02].

Kindahl:2006:BRB

[Kin06] Mats Kindahl. Book review: Types and Programming Languages, by Benjamin C. Pierce, MIT Press, 2002. ACM SIGACT News, 37(4):29–34, December 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Pie02].

Kingsford:2011:BRB

[Kin11] Carl Kingsford. Book review: Complex Social Networks, by Fernando Vega-Redondo. ACM SIGACT News, 42(1):29–31, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [VR07].

Kintali:2012:RBA

[Kin12] Shiva Kintali. Review of Elements of Automata Theory, by Jacques Sakarovitch, Translator (from French) Reuben Thomas. ACM SIGACT News, 43(3):45–47, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kintali:2014:RBF

[Kin14] Shiva Kintali. Review of Boosting: foundations and algorithms by Robert E. Schapire and Yoav Freund. ACM SIGACT News, 45(1): 41–43, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kintali:2015:RAU

[Kin15] Shiva Kintali. Review of Algorithms Unplugged by B. Vöcking, H. Alt, M. Dietzfelbinger, R. Reischuk, C. Scheideler, H. Vollmer, and D. Wagner: The Power of Algorithms by Giorgio Ausiello and Rossella Petreschi. ACM SIGACT News, 46(4):14–16, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kirchner:1991:RTA

[KL91] Hélène Kirchner and Pierre Lescanne. Rewriting techniques and applications, RTA’91. ACM SIGACT News, 22(3):24–30, Summer 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kreinovich:2005:KCL

[KL05] Vladik Kreinovich and Luc Longpré. Kolmogorov complexity leads to a representation theorem for idempotent probabilities (sigmamaxitive measures). ACM SIGACT News, 36(3):107–112, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Katz:2008:IMC

[KL08] Jonathan Katz and Yehuda Lindell. Introduction to Modern Cryptography. Chapman and Hall/CRC cryptography and network security. Chapman and Hall/CRC, Boca Raton, FL, USA, 2008. ISBN 1-58488-551-3. xviii + 534 pp. LCCN QA76.9.A25

K36 2008. URL http://www.loc.gov/catdir/enhancements/
fy0807/2007017861-d.html;   http://www.loc.gov/catdir/
toc/ecip0716/2007017861.html.

Koshelev:2010:AIT

[KL10] Misha Koshelev and Luc Longpré. Algorithmic information theory may explain the pathogenic number of DNA repeats in myotonic dystrophy type 1 (and in similar diseases). ACM SIGACT News, 41(4):61–64, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Klein:2016:CGC

[KL16] Rolf Klein and Elmar Langetepe. Computational geometry column 63. ACM SIGACT News, 47(2):34–39, June 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Klawe:1988:SSS

[Kla88] Maria Klawe. SIGACT salary survey. ACM SIGACT News, 19(3– 4):11–17, Fall 1988. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kliot:2008:RD

[Kli08] Gabriel Kliot. Review of DSN’08. ACM SIGACT News, 39(4):69– 73, December 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Knop:2021:GCM

[KLMY21] A. Knop, S. Lovett, S. McGuire, and W. Yuan. Guest column: Models of computation between decision trees and communication. ACM SIGACT News, 52(2):46–70, June 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://

dl.acm.org/doi/10.1145/3471469.3471479.

Kuller:2010:SNA

[KM10] Samir Kuller and Michael W. Mahoney. SIGACT news algorithms column: Computation in large-scale scientific and Internet data applications is a focus of MMDS 2010. ACM SIGACT News, 41 (4):65–72, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Khoussainov:2001:ATA

[KN01] Bakhadyr Khoussainov and Anil Nerode. Automata Theory and Its Applications, volume 21 of Progress in computer science and applied logic. Birkhäuser Verlag, Basel, Switzerland, 2001. ISBN 0-8176-4207-2 , 3-7643-4207-2. xiv + 430 pp. LCCN QA267

.K56 2001. URL http://www.loc.gov/catdir/enhancements/
fy0812/2001025819-d.html;   http://www.loc.gov/catdir/
enhancements/fy0812/2001025819-t.html.

Knuth:1974:PAN

[Knu74a] Donald E. Knuth. Postscript about NP-hard problems. ACM SIGACT News, 6(2):15–16, April 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Knuth:1974:TP

[Knu74b] Donald E. Knuth. A terminological proposal. ACM SIGACT News, 6(1):12–18, January 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Knuth:1976:BOB

[Knu76a] Donald E. Knuth. Big Omicron and big Omega and big Theta. ACM SIGACT News, 8(2):18–24, April/June 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See comments [Bra85, VM85].

Knuth:1976:MSR

[Knu76b] Donald E. Knuth. Mariages stables et leurs relations avec d’autres problèmes combinatoires. Collection de la Chaire Aisenstadt. Les Presses de l’Université de Montréal, Montréal, Québec, Canada, 1976. ISBN 0-8405-0342-3, 2-7606-0529-9. 106 pp. LCCN QA164 .K59 1976. Edition revue et corrigée, 1981. Currently available from Les Publications CRM / Centre de Recherches Mathématiques, Université de Montréal, Montréal, Québec.

Knuth:1977:CS

[Knu77a] Donald E. Knuth. The complexity of songs. ACM SIGACT News, 9(2):17–24, Summer 1977. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). Reprinted in [Knu84a]; errata (June 1984), 593. Reprinted in Metafolkloristica, edited by Franz Kinder and Boaz the Clown (Salt Lake City, Utah 84158-8183: Frank and Boaz, P.O. Box 58183), 63–65. Reprinted in Humour the Computer, edited by Andrew Davison (Cambridge, Massachusetts: MIT Press, 1994), to appear.

Knuth:1977:SMA

[Knu77b] Donald E. Knuth. The Samson–Mueller algorithm. ACM SIGACT News, 9(??):??, ???? 1977. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Knuth:1984:CS

[Knu84a] Donald E. Knuth. The complexity of songs. Communications of the ACM, 27(4):344–346, April 1984. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See erratum in [Knu84b].

Knuth:1984:CCS

[Knu84b] Donald E. Knuth. Corrigendum: “The complexity of songs”. Communications of the ACM, 27(6):593, June 1984. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Knu84a].

Knuth:1996:SPC

[Knu96] Donald Ervin Knuth. Selected Papers on Computer Science, volume 59 of CSLI lecture notes. Cambridge University Press, Cambridge, UK, 1996. ISBN 1-881526-91-7 (paperback), 1-881526-92-5 (hardcover). xii + 274 pp. LCCN QA76.K537 1996.

Knuth:1997:SMR

[Knu97] Donald E. Knuth. Stable Marriage and Its Relation to Other Combinatorial Problems. An Introduction to the Mathematical Analysis of Algorithms, volume 10 of CRM Proceedings & Lecture Notes. American Mathematical Society, Providence, RI, USA, 1997. ISBN 0-8218-0603-3. xiv + 74 pp. LCCN QA164 .K5913 1997. Translated from the French by Martin Goldstein and revised by the author.

Knuth:2000:SPA

[Knu00] Donald Ervin Knuth. Selected Papers on Analysis of Algorithms. CSLI Publications, Stanford, CA, USA, 2000. ISBN 1-57586-211-5, 1-57586-212-3 (paperback). xvi + 621 pp. LCCN QA9.58 .K65

2000. URL http://www.loc.gov/catdir/description/cam021/
00023847.html;   http://www.loc.gov/catdir/toc/cam023/
00023847.html.

Knuth:2003:RWF

[Knu03a] Donald E. Knuth. Robert W Floyd, In Memoriam. ACM SIGACT News, 34(4):3–13, December 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Knuth:2003:SPD

[Knu03b] Donald E. Knuth. Selected Papers on Discrete Mathematics, volume 106 of CSLI Lecture Notes. CSLI Publications, Stanford, CA, USA, 2003. ISBN 1-57586-248-4. xvi + 812 pp. LCCN QA39.3 .K59 2001. US$40.00.

Knuth:2011:ACP

[Knu11] Donald E. Knuth. The Art of Computer Programming: Volume 4, Combinatorial algorithms. Part 1, volume 4A of The art of computer programming. Addison-Wesley, Reading, MA, USA, 2011. ISBN 0-201-03804-8. xv + 883 pp. LCCN QA76.6 2005.

Kuhn:2011:DNM

[KO11] Fabian Kuhn and Rotem Oshman. Dynamic networks: models and algorithms. ACM SIGACT News, 42(1):82–96, March 2011. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Koblitz:1998:BRB

[Kob98] Neal Koblitz. Book review: Discrete Mathematics in the Schools. Edited by Joseph G. Rosenstein, Deborah S. Franzblau, and Fred S. Roberts (American Mathematical Society). ACM SIGACT News, 29(1):8–12, March 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [RFR97].

Koblitz:2008:RCJ

[Kob08] Neal Koblitz. Random Curves: Journeys of a Mathematician. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 3-540-74077-5 (hardcover), 3-540-74078-3 (e-book). 392 pp. LCCN QA29.K57 A3 2008.

Koch:2019:DTC

[Koc19] Christoph Koch. Database theory column report on PODS 2019. ACM SIGACT News, 50(4):93–94, December 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3374857.3374872.

Konrad:2017:SR

[Kon17] Christian Konrad. SIROCCO 2017 review. ACM SIGACT News, 48(4):80–86, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kopparty:2011:BRB

[Kop11] Swastik Kopparty. Book review: Algebraic Function Fields and Codes, by Henning Stichtenoth. ACM SIGACT News, 42(2):19–24, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sti93, Sti09].

Koren:1993:CAA

[Kor93] Israel Koren. Computer Arithmetic Algorithms. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-151952-2. xiii + 210 pp. LCCN QA76.9.C62 K67 1993.

Koren:2002:CAA

[Kor02] Israel Koren. Computer Arithmetic Algorithms. A. K. Peters, Ltd., Wellesley, MA, USA, second edition, 2002. ISBN 1-56881-160-8. xv + 281 pp. LCCN QA76.9.C62 K67 2002.

Korte:2006:COT

[Kor06] Bernhard Korte. Combinatorial optimization: theory and algorithms, volume 21 of Algorithms and combinatorics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2006. ISBN 3-540-25684-9. ISSN 0937-5511. xv + 597 pp. LCCN QA402.5 .K6665 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0662/2005931374-d.
html;   http://www.loc.gov/catdir/enhancements/fy0826/
2005931374-t.html.

Kosaraju:1974:CRP

[Kos74a] S. Rao Kosaraju. Correction to “Regularity preserving functions” (ACM SIGACT News 6 (1974), no. 2, 16–17). ACM SIGACT News, 6(3):22, July 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Kos74b].

Kosaraju:1974:RPF

[Kos74b] S. Rao Kosaraju. Regularity preserving functions. ACM SIGACT News, 6(2):16–17, April 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See correction [Kos74a].

Kostin:1997:NAD

[Kos97] A. E. Kostin. The novel algorithm for determining the reachability in acyclic Petri nets. ACM SIGACT News, 28(2):70–79, June 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kozen:1978:CPE

[Koz78] Dexter Kozen. A clique problem equivalent to graph isomorphism. ACM SIGACT News, 10(2):50–52, Summer 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kozen:1986:FPO

[Koz87] D. Kozen. Fast parallel orthogonalization. ACM SIGACT News, 18(2):47, September/March 1986–1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kozen:2006:TC

[Koz06] Dexter Kozen. Theory of Computation. Texts in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-84628-297-7. xiii + 418 pp. LCCN QA267.7 .K69 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0663/2005937504-d.
html;   http://www.loc.gov/catdir/enhancements/fy0818/
2005937504-t.html.

Kozma:2014:RTH

[Koz14] László Kozma. Review of The tower of Hanoi: myths and maths by Andreas M. Hinz, Sandi Klavzar, Uroz Milutinović and Ciril Petr. ACM SIGACT News, 45(3):34–36, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kozma:2018:RCD

[Koz18] László Kozma. Review of Compact Data Structures — a practical approach by Gonzalo Navarro. ACM SIGACT News, 49(3):9–13, September 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kozen:2022:RJ

[Koz22] Dexter Kozen. Reminiscences of Juris. ACM SIGACT News, 53 (4):32–34, December 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3577971.3577976.

Kosowski:2011:S

[KPS11] Adrian Kosowski, Dominik Pajak, and Zuzanna Stamirowska. SIROCCO 2011. ACM SIGACT News, 42(4):92–95, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Keidar:2001:CFT

[KR01] Idit Keidar and Sergio Rajsbaum. On the cost of fault-tolerant consensus when there are no faults: preliminary version. ACM SIGACT News, 32(2):45–63, June 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kumar:2003:ACS

[KR03] Ravi Kumar and Ronitt Rubinfeld. Algorithms column: sublinear time algorithms. ACM SIGACT News, 34(4):57–67, December 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kuznetsov:2009:BWW

[KR09] Petr Kuznetsov and Rodrigo Rodrigues. BFTW 3: Why? when? where? workshop on the theory and practice of Byzantine fault tolerance. ACM SIGACT News, 40(4):82–86, December 2009. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kuznetsov:2012:WTW

[KR12] Petr Kuznetsov and Srivatsan Ravi. WTTM 2011: The Third Workshop on the Theory of Transactional Memory. ACM SIGACT News, 43(1):87–92, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kramer:2007:LCC

[Kra07] Simon Kramer. Logical concepts in cryptography. ACM SIGACT News, 38(4):65–66, December 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kreinovich:2000:BRB

[Kre00] Vladik Kreinovich. Book review: A = B, by Marko Petkovšek, Herbert S. Wilf, and Doron Zeilberger. ACM SIGACT News, 31 (4):18–24, December 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [PWZ96].

Kreinovich:2001:INB

[Kre01] Vladik Kreinovich. Itanium’s new basic operation of fused multiply-add: Theoretical explanation and theoretical challenge. ACM SIGACT News, 32(1):115–117, March 2001. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

URL http://www.cs.utep.edu/vladik/2000/tr00-42.pdf;
http://www.cs.utep.edu/vladik/2000/tr00-42.ps.gz.

Kreinovich:2007:BRB

[Kre07] Vladik Kreinovich. Book review: Computational Techniques for the Summation of Series, by Anthony Sofo, Kluwer Academic Publishers, 2003. ACM SIGACT News, 38(1):24–27, March 2007. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sof03].

Krishnamoorthy:1975:NHP

[Kri75] M. S. Krishnamoorthy. An NP-hard problem in bipartite graphs. ACM SIGACT News, 7(1):26, January 1975. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Krishnamoorthy:1977:NSS

[Kri77] M. S. Krishnamoorthy. A note on “Some simplified NP-complete graph problems”. ACM SIGACT News, 9(3):24, Fall 1977. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [GJ78].

Krizanc:1999:BRB

[Kri99] Danny Krizanc. Book review: Probabilistic Combinatorics and Its Applications, edited by Béla Bollobás. ACM SIGACT News, 30 (2):12–14, June 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BC^{+}91, Bol01].

Krizanc:2000:BRB

[Kri00] Danny Krizanc. Book review: Gems of Theoretical Computer Science, by Uwe Schöning and Randall Pruim (Springer-Verlag, 1998). ACM SIGACT News, 31(2):2–5, June 1, 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [SP98].

Krizanc:2008:BRB

[Kri08] Danny Krizanc. Book review: The Game’s Afoot: Game Theory in Myth and Paradox, by Alexander Mehlmann, American Mathematical Society, 2000, 0-8218-2121-0. ACM SIGACT News, 39(1): 22–24, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Meh00].

Kung:2009:CRW

[KRY09] Joseph P. S. Kung, Gian-Carlo Rota, and Catherine H. Yan. Combinatorics: the Rota Way. Cambridge mathematical library. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-73794-X (paperback), 0-521-88389-X (hardcover). xii + 396 pp. LCCN QA164 .K86 2009.

Kreher:1999:CAG

[KS99] Donald L. Kreher and Douglas R. Stinson. Combinatorial algorithms: generation, enumeration, and search. ACM SIGACT News, 30(1):33–35, March 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kinber:2001:TCG

[KS01] Efim Kinber and Carl H. Smith. Theory of Computing: a Gentle Introduction. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-027961-7. xiv + 207 pp. LCCN QA76 .K475 2001.

Keidar:2006:WSC

[KS06a] Idit Keidar and Assaf Schuster. Want scalable computing?: speculate! ACM SIGACT News, 37(3):59–66, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kreinovich:2006:ANH

[KS06b] Vladik Kreinovich and Max Shpak. Aggregability is NP-hard. ACM SIGACT News, 37(3):97–104, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

King:2010:SBC

[KS10] Valerie King and Jared Saia. Scalable Byzantine computation. ACM SIGACT News, 41(3):89–104, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kopparty:2016:GCL

[KS16] Swastik Kopparty and Shubhangi Saraf. Guest column: Local testing and decoding of high-rate error-correcting codes. ACM SIGACT News, 47(3):46–66, September 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kemme:2014:DSR

[KSRS14] Bettina Kemme, André Schiper, G. Ramalingam, and Marc Shapiro. Dagstuhl seminar review: consistency in distributed systems. ACM SIGACT News, 45(1):67–89, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kitaev:2002:CQC

[KSV02] A. Yu. (Alexei Yu.) Kitaev, A. (Alexander) Shen, and M. N. (Mikhail N.) Vyalyi. Classical and Quantum Computation, volume 47 of Graduate studies in mathematics. American Mathematical Society, Providence, RI, USA, 2002. ISBN 0-8218-2161-X. ISSN 1065-7339. xiii + 257 pp. LCCN QA267 .K5713 2002.

Kostin:1998:YAR

[KT98] Alexander E. Kostin and Svetlana A. Tchoudaikina. Yet another reachability algorithm for Petri nets. ACM SIGACT News, 29 (4):98–110, December 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kudlek:1995:RI

[Kud95] Manfred Kudlek. Report on IFIP’94. ACM SIGACT News, 26(1): 92–98, March 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kugel:1975:T

[Kug75] Peter Kugel. Trees. ACM SIGACT News, 7(2):19, April 1975. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kugel:1976:DAC

[Kug76a] Peter Kugel. Digital to analog conversion: a speculation. ACM SIGACT News, 8(2):27–33, April/June 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kugel:1976:UT

[Kug76b] Peter Kugel. On uninteresting theorems. ACM SIGACT News, 8(1):27–29, January 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kulkarni:2013:GDT

[Kul13] Raghav Kulkarni. Gems in decision tree complexity revisited. ACM SIGACT News, 44(3):42–55, September 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kumar:2012:RBR

[Kum12] Akash Kumar. Review of Polynomia and Related Realms by Dan Kalman. ACM SIGACT News, 43(2):15–20, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kumar:2013:RBO

[Kum13] Akash Kumar. Review of Combinatorial Optimization by B. Korte and J. Vygen. ACM SIGACT News, 44(3):31–35, September 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kupcu:2013:DTT

[Küp13] Alptekin Küpçü. Distributing trusted third parties. ACM SIGACT News, 44(2):92–112, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kurz:2006:CTL

[Kur06] Alexander Kurz. Coalgebras and their logics. ACM SIGACT News, 37(2):57–77, June 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kuznetsov:2018:FSS

[Kuz18] Petr Kuznetsov. The First Summer School on Practice and Theory of Concurrent Computing SPTCC 2017. ACM SIGACT News, 49(1):81–90, March 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kuznetsov:2019:RSS

[Kuz19] Petr Kuznetsov. Review of the Second Sankt Petersburg Summer School on Practice and Theory of Distributed Computing (SPTDC 2019). ACM SIGACT News, 50(4):48–55, December 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3374857.
3374868.

Kearns:1995:CLT

[KV95] Michael J. Kearns and Umesh V. Vazirani. Computational Learning Theory. ACM SIGACT News, 26(1):43–45, March 1995. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Korte:2000:COT

[KV00] Bernhard H. (Bernhard H.) Korte and Jens Vygen. Combinatorial optimization: theory and algorithms, volume 21 of Algorithms and combinatorics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67226-5. ISSN 0937-5511. xi + 530 pp. LCCN QA402.5 .K6665 2000.

Korte:2002:COT

[KV02] Bernhard H. (Bernhard H.) Korte and Jens Vygen. Combinatorial optimization: theory and algorithms, volume 21 of Algorithms and combinatorics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2002. ISBN 3-540-43154-3. ISSN 0937-5511. xiii + 530 pp. LCCN QA402.5 .K6665

Korte:2008:COT

[KV08] Bernhard Korte and Jens Vygen. Combinatorial Optimization: Theory and Algorithms, volume 21 of Algorithms and combinatorics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2008. ISBN 3-540-71843-5 (print), 3-540-71844-3 (e-book). ISSN 0937-5511. xvii + 627 pp. LCCN QA402.5 .K6665 2008.

Korte:2012:COH

[KV12] Bernhard Korte and Jens Vygen. Combinatorial Optimization: Theory and Algorithms, volume 21 of Algorithms and combinatorics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fifth edition, 2012. ISBN 3-642-24487-4, 3-642-24488-2. ISSN 0937-5511. xix + 659 pp. LCCN QA402.5 .K6665 2012.

Kuperstein:2012:AIM

[KVY12] Michael Kuperstein, Martin Vechev, and Eran Yahav. Automatic inference of memory fences. ACM SIGACT News, 43(2):108–123, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kruskal:1981:NWC

[KW81] Clyde P. Kruskal and Elia Weixelbaum. A note on the worst case of heapsort. ACM SIGACT News, 13(1):35–38, Winter 1981. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kuhn:2007:TCC

[KW07] Michael Kuhn and Roger Wattenhofer. The theoretic center of computer science. ACM SIGACT News, 38(4):54–63, December 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Kyer:2009:BRB

[Kye09] Brad G. Kyer. Book review: Biologically Inspired Algorithms for Financial Modelling, by Anthony Brabazon, Michael O’Neill Springer-Verlag Berlin Heidelberg, 2006. ACM SIGACT News, 40 (4):17–23, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BO06].

Labarre:2012:RBP

[Lab12] Anthony Labarre. Review of Combinatorial Pattern Matching Algorithms in Computational Biology using Perl and R, by Gabriel Valiente. ACM SIGACT News, 43(3):48–50, September 2012. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Labarre:2013:RBA

[Lab13] Anthony Labarre. Review of Graph Algorithms (2nd edition) by Shimon Even, edited by Guy Even. ACM SIGACT News, 44(3):15– 16, September 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ladner:1975:CVP

[Lad75] Richard E. Ladner. The circuit value problem is log space complete for P. ACM SIGACT News, 7(1):18–20, January 1975. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lakatos:1976:PRL

[Lak76] Imre Lakatos, editor. Proofs and Refutations: the Logic of Mathematical Discovery. Cambridge University Press, Cambridge, UK, 1976. ISBN 0-521-21078-X, 0-521-29038-4. xii + 174 pp. LCCN

QA8.4 .L34. URL http://www.loc.gov/catdir/description/
cam032/75032478.html;   http://www.loc.gov/catdir/toc/

cam031/75032478.html. Edited by John Worrall and Elie Zahar after Lakatos’ death.

Lamport:2009:TC

[Lam09] Leslie Lamport. Teaching concurrency. ACM SIGACT News, 40(1):58–62, March 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lampis:2013:RBE

[Lam13] Michael Lampis. Review of Exact Exponential Algorithms by Fedor V. Fomin and Dieter Kratsch. ACM SIGACT News, 44(2):17–21, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lampis:2014:GCE

[Lam14] Michael Lampis. Guest column: the elusive inapproximability of the TSP. ACM SIGACT News, 45(1):48–65, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Landweber:1982:CTR

[Lan82] L. H. Landweber. CSNET: a tool for researchers in theoretical computer science. ACM SIGACT News, 14(3):21–22, Summer 1982. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Landau:2010:SSR

[Lan10] Susan Eva Landau. Surveillance or security?: The risks posed by new wiretapping technologies. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01530-7 (hardcover),0-262-29558-X (e-book). xvi + 383 pp. LCCN TK5102.85 .L36 2010.

Laplante:1991:HUP

[Lap91] P. A. Laplante. The Heisenberg uncertainty principle and the halting problem. ACM SIGACT News, 22(3):63–65, Summer 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Latch:1995:NATa

[Lat95a] Dana May Latch. NSF Announcements: Theory of Computing Program. ACM SIGACT News, 26(1):31–32, March 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Latch:1995:NATb

[Lat95b] Dana May Latch. NSF Announcements: Theory of Computing Program. ACM SIGACT News, 26(2):37–38, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Latch:1995:NATc

[Lat95c] Dana May Latch. NSF Announcements: Theory of Computing Program. ACM SIGACT News, 26(4):22–23, December 1995. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

LaTorra:2011:BRB

[LaT11] Sage LaTorra. Book review: Comprehensive Mathematics for Computer Scientists 1, by Guerino B. Mazzola, Gerard Milmeister, Jody Weissmann. ACM SIGACT News, 42(1):23–24, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [MMW06].

Law:2006:BRB

[Law06] James Law. Book review: Algorithmic Learning in a Random World, by Vovk, Gammerman and Shafer, Springer, 2005, ISBN 0-387-00152-2. ACM SIGACT News, 37(4):38–40, December 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [VGS05].

Landweber:1999:DBC

[LB99] Laura F. (Laura Faye) Landweber and Eric B. Baum, editors. DNA Based Computers II: DIMACS Workshop, June 10–12, 1996, volume 44 of DIMACS series in discrete mathematics and theoretical computer science. American Mathematical Society, Providence, RI, USA, 1999. ISBN 0-8218-0756-0. LCCN QA76.887 .D52 1999.

League:1997:BRB

[Lea97] Christopher League. Book review: Isomorphisms of Types: From λ-Calculus to Information Retrieval and Language Design, by Roberto Di Cosmo (Birkhäuser, 1995). ACM SIGACT News, 28 (4):24–27, December 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Di 95].

League:2000:BRB

[Lea00] Christopher League. Book review: Lambda Calculi: A Guide for Computer Scientists, by Chris Hankin. ACM SIGACT News, 31 (1):8–13, March 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Han94].

Ledin:2016:RME

[Led16] George Ledin, Jr. Review of: The Mathematics of Encryption: An Elementary Introduction by Margaret Cozzens and Steven J. Miller. ACM SIGACT News, 47(3):19–21, September 2016. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lee:1972:DVR

[Lee72] John A. N. Lee. The definition and validation of the radix sorting technique. ACM SIGACT News, 4(14):142–149, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lee:1989:NRC

[Lee89] D. T. Lee. NSF report — computer and computation research. ACM SIGACT News, 20(4):25, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lee:1990:NRCa

[Lee90a] D. T. Lee. NSF report — computer and computation research. ACM SIGACT News, 21(1):13, January 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lee:1990:NRCb

[Lee90b] D. T. Lee. NSF report — computer and computation research. ACM SIGACT News, 21(2):16, Spring 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lee:2003:BRBa

[Lee03a] Andrew C. Lee. Book review: Introduction to Cryptography, by Johannes A. Buchmann. Springer Verlag, 2001. ACM SIGACT News, 34(4):15–17, December 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Buc01, Buc04].

Lee:2003:BRBb

[Lee03b] Andrew C. Lee. Book review: Modern Cryptography, Probabilistic Proofs and Pseudorandomness Algorithms and Combinatorics, vol 17, by Oded Goldreich. Springer Verlag, 1999. ACM SIGACT News, 34(4):32–34, December 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Gol99].

Lee:2010:BRB

[Lee10] Andrew C. Lee. Book review: Software Abstractions: Logic, Language and Analysis, by Daniel Jackson, MIT Press, 2006. ACM SIGACT News, 41(4):50–54, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Jac06].

Leiss:1980:CFA

[Lei80] Ernst Leiss. Constructing a finite automaton for a given regular expression. ACM SIGACT News, 12(3):81–87, Fall 1980. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Leiss:1991:SCR

[Lei91] Ernst L. Leiss. Some comments on a recent note by Ravikumar. ACM SIGACT News, 22(1):64, Winter 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Leighton:1992:IPA

[Lei92] Frank Thomson Leighton. Introduction to Parallel Algorithms and Architectures: Arrays, Trees, Hypercubes. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1992. ISBN 1-55860-117-1. xvii + 831 pp. LCCN QA76.58 .L45 1992. US$54.95.

Leiss:2007:PCA

[Lei07] Ernst L. Leiss. A Programmer’s Companion to Algorithm Analysis. Chapman and Hall/CRC, Boca Raton, FL, USA, 2007. ISBN 1-58488-673-0. 255 pp. LCCN QA402.5

.L398 2007. URL http://www.loc.gov/catdir/enhancements/
fy0654/2006044552-d.html;   http://www.loc.gov/catdir/
toc/fy0705/2006044552.html.

Lenzerini:2008:DTC

[Len08] Maurizio Lenzerini. Database theory column: report on PODS 2008. ACM SIGACT News, 39(3):63–65, September 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lescanne:2006:BRB

[Les06] Pierre Lescanne. Book review: Alfred Tarski: Life and Logic, by Anita Burdman Feferman and Solomon Feferman, Cambridge University Press 2004. ACM SIGACT News, 37(1):27–28, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [FF04].

Levin:1991:TCH

[Lev91] Leonid Levin. Theory of computation: how to start. ACM SIGACT News, 22(1):47–56, Winter 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Levin:1995:SC

[Lev95] Leonid A. Levin. STOC Criteria. ACM SIGACT News, 26(2):77, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Levin:1996:FCC

[Lev96] Leonid A. Levin. Fundamentals of computing (a cheatlist). ACM SIGACT News, 27(3):89, September 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See errata [Lev97].

Levin:1997:EFC

[Lev97] Leonid Levin. Errata to “Fundamentals of Computing”. ACM SIGACT News, 28(2):80, June 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Lev96].

Levitin:2007:IDA

[Lev07] Anany Levitin. Introduction to the Design and Analysis of Algorithms. Pearson Addison-Wesley, Boston, MA, USA, second edition, 2007. ISBN 0-321-35828-7. xxiii + 562 pp. LCCN

QA76.9.A43 L48 2007. URL http://www.loc.gov/catdir/toc/
ecip063/2005033620.html.

Levin:2010:BRB

[Lev10] Dave Levin. Book review: Algorithmic Game Theory, edited by Noam Nisan, Tim Roughgarden, Eva Tardos, Vijay V. Vazirani. Cambridge University Press, ISBN 978-0-521-87282-9. ACM SIGACT News, 41(4):21–24, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [NRTV07].

Levitin:2012:IDA

[Lev12] Anany Levitin. Introduction to the Design and Analysis of Algorithms. Pearson, Boston, MA, USA, third edition, 2012. ISBN 0-13-231681-1. xxiv + 565 pp. LCCN QA76.9.A43 L48 2012.

Lewis:1978:BRB

[Lew78] Harry R. Lewis. Book review: Mariages stables et leurs relations avec d’autre problèmes combinatoires: introduction à l’analyze mathématique des algorithmes, by Donald E. Knuth. Les Presses de l’Université de Montréal. ACM SIGACT News, 9(4):13–14, Winter 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Knu76b].

Lewis:1979:UCQ

[Lew79] Harry R. Lewis. Unsolvable Classes of Quantificational Formulas. Addison-Wesley, Reading, MA, USA, 1979. ISBN 0-201-04069-7. xv + 198 pp. LCCN QA9.63 .L48.

Lewand:2000:CM

[Lew00] Robert Lewand. Cryptological Mathematics. Classroom resource materials. Mathematical Association of America, Washington, DC, USA, 2000. ISBN 0-88385-719-7. xiv + 199 pp. LCCN QA268 .L48

2000. URL http://www.loc.gov/catdir/description/cam041/
00105256.html;   http://www.loc.gov/catdir/toc/cam041/
00105256.html.

Lewis:2006:ESH

[Lew06] Harry R. Lewis. Excellence without a soul: how a great university forgot education. PublicAffairs, New York, NY, USA, 2006. ISBN 1-58648-393-5. xiv + 305 pp. LCCN LA227.4

.L49 2006. URL http://www.loc.gov/catdir/enhancements/
fy0828/2006002732-b.html;   http://www.loc.gov/catdir/
enhancements/fy0828/2006002732-d.html;   http://www.loc.
gov/catdir/toc/ecip068/2006002732.html.

Lewis:2013:RBP

[Lew13] Harry Lewis. Review of Burdens of Proof by Jean-François Blanchette. ACM SIGACT News, 44(1):19–21, March 2013. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lewis:2014:RUA

[Lew14] Harry Lewis. Review of Unauthorized access: the crisis in online privacy and security by Robert H. Sloan and Richard Warner. ACM SIGACT News, 45(1):16–19, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Li:2004:EAI

[Li04] Dafa Li. The equality axioms are not independent. ACM SIGACT News, 35(3):98–101, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Li:2008:DFL

[Li08] Fanzhang Li. Dynamic Fuzzy Logic and Its Applications. Nova Science Publishers, New York, NY, USA, 2008. ISBN 1-60021-428-2. x + 296 pp. LCCN QA9.64 .L5 2008. URL http://www.

loc.gov/catdir/toc/ecip0619/2006028308.html.

Li:2012:HFM

[Li12] Chong-Zhang Li. Hilbert’s formalistic method and its development in computer science. ACM SIGACT News, 43(2):124–126, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Li:2014:RIM

[Li14] Yang D. Li. Review of Iterative methods in combinatorial optimization by Lap Chi Lau, R. Ravi and Mohit Singh. ACM SIGACT News, 45(1):31–34, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Libkin:1999:QLA

[Lib99] Leonid Libkin. Query languages with arithmetic and constraint databases. ACM SIGACT News, 30(4):41–50, December 1999.

CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Liberatore:2017:RCN

[Lib17] Vincenzo Liberatore. Review of Communication Networks: an Optimization, Control, and Stochastic Networks Perspective, R. Srikant and L. Ying. ACM SIGACT News, 48(4):7–12, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lien:1975:PPS

[Lie75] Y. Edmund Lien. Periodic properties of strings. ACM SIGACT News, 7(1):21–25, January 1975. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Liestman:1990:RWM

[Lie90] Arthur L. Liestman. Report on WOBCATS meeting. ACM SIGACT News, 21(2):32, Spring 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lincoln:1992:LL

[Lin92a] Patrick Lincoln. Linear logic. ACM SIGACT News, 23(2):29–37, Spring 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lincoln:1992:LLG

[Lin92b] Patrick Lincoln. Linear logic: Guest logic column. ACM SIGACT News, 23(2):29–37, Spring 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL http://www.csl.sri.

com/lincoln/sigact92.dvi.Z.

Liotta:1993:ALF

[Lio93] Giuseppe Liotta. An automatic layout facility (abstract). ACM SIGACT News, 24(1):60, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lipton:2010:PQG

[Lip10] Richard J. Lipton. The P = NP question and Gödel’s lost letter. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-4419-7154-8, 1-4419-7155-6 (e-book). xiii + 239 pp. LCCN QA9.65. URL http://catdir.loc.gov/

catdir/enhancements/fy1208/2010932770-b.html;   http:/

/catdir.loc.gov/catdir/enhancements/fy1208/2010932770-
d.html;   http://catdir.loc.gov/catdir/enhancements/fy1208/
2010932770-t.html;   http://www.gbv.de/dms/bowker/toc/
9781441971548.pdf;   http://www.gbv.de/dms/ilmenau/toc/
629401365.PDF.

Litow:1992:R

[Lit92] B. Litow. On Rat_{∞}. ACM SIGACT News, 23(3):98–99, Summer 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Litow:2003:HCP

[Lit03] B. Litow. The Hamiltonian circuit problem and automaton theory. ACM SIGACT News, 34(2):91–94, June 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lipton:1975:QCN

[LL75] Richard J. Lipton and Nancy Lynch. A quantifier characterization for nondeterministic log space. ACM SIGACT News, 7(4):24–25, November 1975. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lind:1973:CLS

[LM73] John Lind and Albert R. Meyer. A characterization of log-space computable functions. ACM SIGACT News, 5(3):26–29, July 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lamport:2010:RSM

[LMZ10] Leslie Lamport, Dahlia Malkhi, and Lidong Zhou. Reconfiguring a state machine. ACM SIGACT News, 41(1):63–73, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lopez-Ortiz:1994:LPM

[LO94] Alejandro López-Ortiz. Linear pattern matching of repeated substrings. ACM SIGACT News, 25(3):114–121, September 1994. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lopez-Ortiz:2005:AFI

[LO05] Alejandro López-Ortiz. Algorithmic foundations of the internet. ACM SIGACT News, 36(2):45–62, June 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Loehr:2011:BC

[Loe11] Nicholas A. Loehr. Bijective combinatorics. Discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, 2011. ISBN 1-4398-4884-X (hardcover). xxii + 590 pp. LCCN QA164 .L64 2011.

Longo:1971:ATD

[Lon71] Giuseppe Longo. Axioms for time dependence of resource consumption in computing recursive functions. ACM SIGACT News, 3(12):14–24, 1971. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

London:1972:CCL

[Lon72] Ralph L. London. Correctness of a compiler for a Lisp subset. ACM SIGACT News, 4(14):121–127, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Longpre:1997:RC

[Lon97] Luc Longpré. Report on Complexity 1997. ACM SIGACT News, 28(4):80–83, December 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Longpre:1998:RC

[Lon98] Luc Longpré. Report on Complexity 1998. ACM SIGACT News, 29(3):92–93, September 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lothaire:2005:ACW

[Lot05] M. Lothaire. Applied Combinatorics on Words, volume 104 of Encyclopedia of mathematics and its applications. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-84802-4. xv + 610 pp. LCCN QA164 .L66 2005. URL http://www.loc.

gov/catdir/description/cam051/2004057012.html;   http:
//www.loc.gov/catdir/toc/cam051/2004057012.html.

Loui:1989:RCR

[Lou89] M. C. Loui. Reprint from Computing Reviews. ACM SIGACT News, 20(4):26–29, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Loui:1990:RCR

[Lou90] Michael C. Loui. Reprint from Computing Reviews. ACM SIGACT News, 21(1):14, January 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Loui:1991:TCA

[Lou91] Michael C. Loui. Theory of computing: achievements, challenges, and opportunities. ACM SIGACT News, 22(3):41–48, Summer 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Loui:1995:RCRa

[Lou95a] Michael C. Loui. Reprints from Computing Reviews. ACM SIGACT News, 26(1):24–26, March 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Loui:1995:RCRb

[Lou95b] Michael C. Loui. Reprints from Computing Reviews. ACM SIGACT News, 26(2):33–36, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Loui:1995:RCRc

[Lou95c] Michael C. Loui. Reprints from Computing Reviews. ACM SIGACT News, 26(3):32–41, September 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Loui:1996:RCRa

[Lou96a] Michael C. Loui. Reprint from Computing Reviews. ACM SIGACT News, 27(1):14–15, March 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Loui:1996:RCRb

[Lou96b] Michael C. Loui. Reprints from Computing Reviews. ACM SIGACT News, 27(2):45–46, June 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Loui:1997:RCR

[Lou97] Michael C. Loui. Reprints from Computing Reviews. ACM SIGACT News, 28(1):9, March 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Loulergue:2005:BRB

[Lou05] Frédéric Loulergue. Book review: Term Rewriting Systems, by Terese, Cambridge University Press 2003, 0-521-39115-6. ACM SIGACT News, 36(4):17–20, December 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BKdV03].

Louridas:2018:JRP

[Lou18a] Panos Louridas. Joint review of The Power of Networks: Six Principles that Connect our Lives. ACM SIGACT News, 49(4):7–10, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Louridas:2018:RNS

[Lou18b] Panos Louridas. Review of Network Science by Albert-Làaszlò Barabàasi. ACM SIGACT News, 49(2):10–13, June 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Loveland:1978:ATP

[Lov78] Donald W. Loveland. Automated Theorem Proving: a Logical Basis, volume 6 of Fundamental studies in computer science. North-Holland Publishing Co., Amsterdam, The Netherlands, 1978. ISBN 0-7204-0499-1. xii + 405 pp. LCCN QA76.9.A96 L68.

Low:2012:BRB

[Low12] Robert J. Low. Book review: Codes: an Introduction to Information Communication and Cryptography, by Norman L. Biggs. ACM SIGACT News, 43(1):27–29, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Big08].

Lovasz:1986:MT

[LP86] László Lovász and M. D. Plummer. Matching theory, volume 121 / 29 of North-Holland mathematics studies; Annals of discrete mathematics. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986. ISBN 0-444-87916-1 (paperback). xxxiii + 544 pp. LCCN QA164 .L7 1986.

Luccio:1992:SPC

[LP92] Fabrizio Luccio and Linda Pagli. The p-shovelers problem: (computing with time-varying data). ACM SIGACT News, 23(2):72–75, Spring 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lewis:1998:ETC

[LP98] Harry R. Lewis and Christos H. Papadimitriou. Elements of the Theory of Computation. ACM SIGACT News, 29(3):62–78, September 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lomuscio:2007:SMC

[LP07] Alessio Lomuscio and Wojciech Penczek. Symbolic model checking for temporal-epistemic logics. ACM SIGACT News, 38(3):77–99, September 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lovasz:2009:MT

[LP09] László Lovász and M. D. Plummer. Matching theory. AMS Chelsea Publishing, Providence, RI, USA, 2009. ISBN 0-8218-4759-7. xxxiii + 547 pp. LCCN QA164 .L7 2009.

Lotker:2010:SAS

[LP10] Zvi Lotker and David Peleg. Structure and algorithms in the SINR wireless model. ACM SIGACT News, 41(2):74–84, June 2010. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lindenmayer:1976:ALD

[LR76] Aristid Lindenmayer and Grzegorz Rozenberg, editors. Automata, Languages, Development. North-Holland Publishing Co., Amsterdam, The Netherlands, 1976. ISBN 0-7204-0474-6 (North Holland), 0-444-11118-2 (American Elsevier). LCCN QH491 .A95.

Lindner:1997:DT

[LR97] Charles C. Lindner and C. A. (Christopher Andrew) Rodger. Design Theory. The CRC Press series on discrete mathematics and its applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1997. ISBN 0-8493-3986-3. 198 pp.

LCCN QA166.25 .L56 1997. URL http://www.loc.gov/catdir/
enhancements/fy0729/97007965-d.html.

Linton:2010:PPP

[LRV10] Steve Linton, Nik Ruškuc, and Vincent Vatter, editors. Permutation Patterns: [Proceedings of the Fifth International Conference on Permutation Patterns, 11–15 June 2007, University of St. Andrews, Scotland], volume 376 of London Mathematical Society Lecture Note Series. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-72834-7 (paperback), 0-511-90249-2 (e-book). 352 pp. LCCN QA164 .P476 2007.

Lipton:1974:ARC

[LS74] Richard J. Lipton and Lawrence Snyder. On the Aanderaa– Rosenberg Conjecture. ACM SIGACT News, 6(1):30–31, January 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Litow:1978:NEO

[LS78] B. Litow and I. H. Sudborough. On non-erasing oracle tapes in space bounded reducibility. ACM SIGACT News, 10(2):53–57, Summer 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Li:2006:NIP

[LS06] Duan Li and Xiaoling Sun. Nonlinear Integer Programming, volume 84 of International series in operations research and management science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-29503-8, 0-387-32995-1 (e-book). xxi + 435 pp. LCCN T57.7 .L5 2006; MLCM 2006/41439 (T). URL http://www.loc.gov/

catdir/enhancements/fy0663/2006922525-d.html;   http:/
/www.loc.gov/catdir/toc/fy0712/2006922525.html.

Limaye:2022:GCL

[LST22] N. Limaye, S. Srinivasan, and S. Tavenas. Guest column: Lower bounds against constant-depth algebraic circuits. ACM SIGACT News, 53(2):40–62, June 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3544979.3544989.

Luby:1996:STR

[Lub96] Michael Luby. SICACT Treasurer’s Report. ACM SIGACT News, 27(2):5–6, June 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lurie:1999:BRB

[Lur99] Jacob Lurie. Book review: Spectral Graph Theory: by Fan R. K. Chung. ACM SIGACT News, 30(2):14–16, June 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Chu97].

Li:1993:IKC

[LV93] Ming Li and P. M. B. Vitányi. An Introduction to Kolmogorov complexity and Its Applications. Texts and monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-94053-7 (New York), 3-540-94053-7 (Berlin). xx + 546 pp. LCCN QA267.7 .L5 1993.

Li:1997:IKC

[LV97] Ming Li and P. M. B. Vitányi. An Introduction to Kolmogorov Complexity and Its Applications. Graduate texts in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1997. ISBN 0-387-94868-6 (hardcover). xx + 637 pp.

LCCN QA267.7 .L5 1997. URL http://www.loc.gov/catdir/
enhancements/fy0815/96042357-d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/96042357-t.html.

Li:2008:IKC

[LV08] Ming Li and P. M. B. Vitányi. An Introduction to Kolmogorov Complexity and Its Applications. Texts in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2008. ISBN 0-387-33998-1 (hardcover), 0-387-49820-6 (e-book). xxiii + 790 pp. LCCN QA267.7 .L5 2008.

Macarie:1995:SBP

[Mac95] Ioan I. Macarie. Space-bounded probabilistic computation: old and new stories. ACM SIGACT News, 26(3):1–12, September 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

MacCormick:2020:RPP

[Mac20] John MacCormick. Review of problems with a point: Exploring math and computer science. ACM SIGACT News, 51(1): 12–14, March 2020. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3388392.3388396.

Mahmoud:2008:PUM

[Mah08] Hosam M. (Hosam Mahmoud) Mahmoud. Pólya Urn Models. Chapman and Hall/CRC texts in statistical science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2008. ISBN 1-4200-5983-1. 290 pp. LCCN QA273

.M317 2008. URL http://www.loc.gov/catdir/toc/fy0805/
2008013347.html.

Mahoney:2019:DAI

[Mah19] Michael W. Mahoney. The difficulties of addressing interdisciplinary challenges at the foundations of data science. ACM SIGACT News, 50(3):91–95, September 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Maier:1980:BRB

[Mai80] David Maier. Book review: Introduction to Automata Theory, Languages and Computation, by John E. Hopcroft and Jeffrey D. Ullman. Addison-Wesley 1979. ACM SIGACT News, 12(3):13–14, Fall 1980. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [HU79, HMU01a, HMU07].

Makowsky:1993:BRB

[Mak93] J. A. Makowsky. Book review: Predicate Transformer Semantics. By Ernest G. Manes. (Cambridge University Press, 1992. 233pp. ISBN 0-521-42036-9. $39.95). ACM SIGACT News, 24(4):36–38, December 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Man92].

Maklevitch:1999:BRB

[Mak99] Joseph Maklevitch. Book review: Privacy on the Line, by Whitfield Diffie and Susan Landau (MIT Press 1998). ACM SIGACT News, 30(1):6–8, March 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [DL98, DL07].

Michiels:2007:TAL

[MAK07] Wil Michiels, E. H. L. (Emile H. L.) Aarts, and Jan Korst. Theoretical Aspects of Local Search. Monographs in theoretical computer science, 1431-2654. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-35853-6. viii + 235 pp. LCCN QA402.5 .M526 2007. URL http://www.loc.gov/

catdir/enhancements/fy0824/2006936842-d.html;   http:/
/www.loc.gov/catdir/toc/fy0707/2006936842.html.

Mandrioli:1982:TTFa

[Man82a] Dino Mandrioli. On teaching theoretical foundations of computer science. ACM SIGACT News, 14(3):36–53, Summer 1982. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mandrioli:1982:TTFb

[Man82b] Dino Mandrioli. On teaching theoretical foundations of Computer Science. ACM SIGACT News, 14(4):58–69, Fall–Winter 1982. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Manes:1992:PTS

[Man92] Ernest G. Manes. Predicate Transformer Semantics, volume 33 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-42036-9. 223 pp. LCCN QA76.7 .M342 1992. URL http://www.loc.gov/

catdir/description/cam025/92249094.html;   http://www.
loc.gov/catdir/toc/cam029/92249094.html.

Maor:1994:SN

[Mao94] Eli Maor. e: The Story of a Number. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03390-0. xiv + 223 pp.

LCCN QA247.5.M33 1994. URL http://www-gap.dcs.st-and.
ac.uk/~history/HistTopics/e.html. This book chronicles the

story of the ultimate version number of .

Marcus:1998:CPO

[Mar98] Daniel A. Marcus. Combinatorics: a Problem Oriented Approach. Classroom resource materials. Mathematical Association of America, Washington, DC, USA, 1998. ISBN 0-88385-710-3. x + 136 pp. LCCN QA164 .M345 1998. URL http://www.loc.gov/

catdir/description/cam029/98085594.html;   http://www.
loc.gov/catdir/toc/cam027/98085594.html.

Marchini:2003:BRB

[Mar03a] Andrea Marchini. Book review: An Introduction to Quantum Computing Algorithms, by Arthur O. Pittenger. Birkhäuser 2000. ACM SIGACT News, 34(3):16–18, September 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Pit00].

Maros:2003:CTS

[Mar03b] István Maros. Computational Techniques of the Simplex Method, volume 61 of International series in operations research and management science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-7332-1. xx + 325 pp. LCCN T57.76 .M37 2003. URL http:

//www.loc.gov/catdir/enhancements/fy0820/2002040603-d.

html;   http://www.loc.gov/catdir/enhancements/fy0820/
2002040603-t.html.

Marcus:2008:GTP

[Mar08a] Daniel A. Marcus. Graph Theory: a Problem Oriented Approach. MAA textbooks. Mathematical Association of America, Washington, DC, USA, 2008. ISBN 0-88385-753-

7. xvi + 205 pp. LCCN QA166 .M37 2008. URL http:

//www.loc.gov/catdir/enhancements/fy0902/2008922013-b.
html;   http://www.loc.gov/catdir/enhancements/fy0902/
2008922013-d.html;   http://www.loc.gov/catdir/enhancements/
fy0902/2008922013-t.html.

Margenstern:2008:RNI

[Mar08b] Maurice Margenstern. Research notices: The injectivity of the global function of a cellular automaton in the hyperbolic plane is undecidable. ACM SIGACT News, 39(3):40, September 2008. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Marecek:2009:BRB

[Mar09a] Jakub Marecek. Book review: Theoretical Aspects of Local Search, by Wil P. A. J. Michiels, Emile H. L. Aarts, and Jan H. M. Korst (Springer in the EATCS Series Monographs in Theoretical Computer Science, 007). ACM SIGACT News, 40(2):27–30, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [MAK07].

Marek:2009:IMS

[Mar09b] Wiktor Marek. Introduction to mathematics of satisfiability. Chapman and Hall/CRC studies in informatics series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4398-0167-3 (hardcover). xiii + 350 pp. LCCN QA9.3 .M37 2009.

Marcus:2015:RGE

[Mar15a] Shoshana Marcus. Review of: A Guide to Experimental Algorithmics by Catherine C. McGeoch. ACM SIGACT News, 46(1): 20–22, March 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Marcus:2015:RJS

[Mar15b] Shoshana Marcus. Review of: Jewels of Stringology by Maxime Crochemore and Wojciech Rytter. ACM SIGACT News, 46(3):11– 14, September 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Marcus:2020:RPS

[Mar20] Shoshana Marcus. Review of The Problem With Software: Why Smart Engineers Write Bad Code. ACM SIGACT News, 51(3): 7–9, September 2020. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3427361.3427364.

Masum:2000:BRBb

[Mas00a] Hassan Masum. Book review: Computational Geometry: Algorithms and Applications (2nd ed.) by Mark de Berg, Marc van Kreveld, Mark Overmars, and Otfried Schwarzkopf. ACM SIGACT News, 31(4):10–12, December 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [dB^{+}97, dB^{+}00, dB^{+}08].

Masum:2000:BRBa

[Mas00b] Hassan Masum. Book review: Microsurveys in Discrete Probability, edited by David Aldous and James Propp. (AMS 1998). ACM SIGACT News, 31(3):22–24, September 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [AP98].

Masum:2001:BRBa

[Mas01a] Hassan Masum. Book review: Data Structures and Algorithms in Java (2nd ed): Michael T Goodrich and Roberto Tamassia. ACM SIGACT News, 32(1):3–5, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [GT97, GT01, GT04, GT06, GT10].

Masum:2001:BRBb

[Mas01b] Hassan Masum. Book review: How to Solve It: Modern Heuristics, by Zbigniew Michalewicz and David B. Fogel. ACM SIGACT News, 32(1):8–12, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [MF02].

Masum:2002:BRB

[Mas02] Hassan Masum. Book review: A New Kind of Science. ACM SIGACT News, 33(4):15–21, December 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Wol02].

Masum:2003:BRB

[Mas03] Hassan Masum. Book review: Algorithmics for Hard Problems: Introduction to Combinatorial Optimization, Randomization, Approximation, and Heuristics, by Juraj Hromkovic. Springer 2001. ACM SIGACT News, 34(2):6–8, June 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Hro03].

Matiyasevich:1993:HTP

[Mat93] Yuri V. Matiyasevich. Hilbert’s Tenth Problem. Foundations of computing. MIT Press, Cambridge, MA, USA, 1993. ISBN 0-262-13295-8. xxii + 264 pp. LCCN QA242 .M399 1993. With a foreword by Martin Davis.

Matos:1994:IUP

[Mat94] Armando B. Matos. An introduction to ultimately periodic sets of integers. ACM SIGACT News, 25(1):90–96, March 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Matias:1997:PAC

[Mat97] Yossi Matias. Parallel algorithms column: on the search for suitable models. ACM SIGACT News, 28(3):21–29, September 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mattei:2014:RWS

[Mat14a] Nicholas Mattei. Review of Who’s #1?: the science of ranking and rating by Amy N. Langville and Carl D. Meyer. ACM SIGACT News, 45(1):38–40, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mattei:2014:RWB

[Mat14b] Nicholas Mattei. Review of Who’s bigger?: where historical figures really rank by Steven Skiena and Charles B. Ward. ACM SIGACT News, 45(2):40–42, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mavronicolas:2002:BRB

[Mav02] Marios Mavronicolas. Book review: Introduction to Distributed Algorithms. ACM SIGACT News, 33(2):16–18, June 2002. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Tel00].

Mavronicolas:2009:BRB

[Mav09] Marios Mavronicolas. Book review: Design and Analysis of Randomized Algorithms: Introduction to Design Paradigms, by Juraj Hromkovic (Published by Springer). ACM SIGACT News, 40(2): 25–27, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Hro03].

Mazur:2010:CGT

[Maz10] David R. Mazur. Combinatorics: a Guided Tour. MAA textbooks. Mathematical Association of America, Washington, DC, USA, 2010. ISBN 0-88385-762-6 (hardcover). xviii + 391 pp. LCCN QA164 .M398 2010.

Mazumdar:2014:RCP

[Maz14] Arya Mazumdar. Review of Classic Papers in Combinatorics Edited by Ira Gessel and Gian-Carlo Rota. ACM SIGACT News, 45(4):32–35, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Miller:2005:ASP

[MB05] Russ Miller and Laurence Boxer. Algorithms sequential and parallel: a unified approach. Charles River Media, Hingham, MA, USA, second edition, 2005. ISBN 1-58450-412-9, 1-58450-652-0. xix + 372 pp. LCCN QA76.9.A43 M55 2005.

McBeth:1985:TSE

[McB85] Rod McBeth. The tree structure of exponential calculations. ACM SIGACT News, 17(1):65–77, June 1985. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

McBeth:1986:TSE

[McB87a] R. McBeth. The tree structure of exponential calculations — addendum. ACM SIGACT News, 18(2):64, September/March 1986– 1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

McBeth:1987:PGI

[McB87b] R. McBeth. A proof of generalized induction. ACM SIGACT News, 18(3):31–36, April 1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

McBeth:1991:FTE

[McB91] M. A. McBeth. Finite types, exponential forms and complete trees. ACM SIGACT News, 22(3):55–62, Summer 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

McBeth:1992:TMA

[McB92] M. A. McBeth. The Turing machine of Ackermann’s function. ACM SIGACT News, 23(1):37–43, Winter 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

McConnell:2001:AAA

[McC01] Jeffrey J. McConnell. Analysis of Algorithms: an Active Learning Approach. Jones and Bartlett Publishers, Sudbury, MA, USA, 2001. ISBN 0-7637-1634-0. LCCN QA76.9.A43 M38 2001b.

McGowan:1972:CML

[McG72a] Clement McGowan. The Contour Model Lambda Calculus Machine. ACM SIGACT News, 4(14):110–115, January 1972. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

McGowan:1972:MRE

[McG72b] Clement L. McGowan. The “most recent” error: Its causes and correction. ACM SIGACT News, 4(14):191–202, January 1972. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

McNicholl:1999:BRB

[McN99] Timothy H. McNicholl. Book review: Stable Marriage and its Relation to Other Combinatorial Problems: An Introduction to Algorithm Analysis, by Donald E. Knuth (American Mathematical Society 1996). ACM SIGACT News, 30(1):2–4, March 1999. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Knu97].

McNicholl:2001:BRB

[McN01a] Timothy H. McNicholl. Book review: Complexity and Real Computation, by Blum, Cucker, Shub, and Smale. Springer-Verlag. ACM SIGACT News, 32(2):14–15, June 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BCSS98].

McNicholl:2001:RBP

[McN01b] Timothy H. McNicholl. A review of Selected Papers on Analysis of Algorithms: by Donald E. Knuth. ACM SIGACT News, 32 (1):5–8, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Knu00].

McNaughton:2004:BRB

[McN04] Robert McNaughton. Book review: Introduction to Languages, Machines and Logic: Computable Languages, Abstract Machines and Formal Logic, by Alan P. Parkes. Springer-Verlag 2002. ACM SIGACT News, 35(2):13–14, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Par02b].

Mehlmann:2000:GAG

[Meh00] Alexander Mehlmann. The Game’s Afoot!: Game Theory in Myth and Paradox, volume 5 of Student mathematical library. American Mathematical Society, Providence, RI, USA, 2000. ISBN 0-8218-2121-0. ISSN 1520-9121. xx + 159 pp. LCCN QA269 .M44513 2000. Translated by David Kramer.

Meiklejohn:2010:BRB

[Mei10] Sarah Meiklejohn. Book review: An Introduction to Mathematical Cryptography, by Jeffrey Hoffstein, Jill Pipher, and Joseph Silverman Springer-Verlag, 2008. ACM SIGACT News, 41(4):47– 50, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [HPS08].

Melvin:2011:BRB

[Mel11] Justin Melvin. Book review: Nonlinear Integer Programming, by Duan Li and Xiaoling Sun. ACM SIGACT News, 42(1):26–29, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [LS06].

Mendelson:2004:IGT

[Men04] Elliott Mendelson. Introducing Game Theory and Its Applications. Discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, 2004. ISBN 1-58488-300-6. 259 pp. LCCN QA269 .M455 2004. URL http://www.loc.gov/

catdir/enhancements/fy0646/2004042807-d.html.

Mermin:2007:QCS

[Mer07] N. David Mermin. Quantum Computer Science: an Introduction. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-87658-3 (hardcover). xiv + 220 pp. LCCN QA76.889

.M47 2007. URL http://site.ebrary.com/lib/stanford/Doc?
id=10202764.

Michalewicz:2002:HSI

[MF02] Zbigniew Michalewicz and David B. Fogel. How to Solve It: Modern Heuristics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-66061-5. xv + 467 pp. LCCN QA63 .M53 2000.

Miettinen:2011:BRB

[Mie11] Pauli Miettinen. Book review: Mathematical Tools for Data Mining: Set Theory, Partial Orders, Combinatorics, by Dan A. Simovici and Chabane Djeraba. ACM SIGACT News, 42(2):43–46, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [SD08].

Mikolajzak:1993:BRB

[Mik93] Boleslaw Mikolajzak. Book review: Coloured Petri Nets: Basic Concepts, Analysis Methods and Practical Use, Volume One. By Kurt Jensen. (Springer-Verlag, 1992. vii + 234 pages. ISBN 0-387-55597-8. $69.00). ACM SIGACT News, 24(3):31–33, October 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Jen97].

Mikolajczak:2002:BRB

[Mik02] Boleslaw Mikolajczak. Book review: Petri Net Algebra. ACM SIGACT News, 33(2):10–14, June 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BDK01].

Miksa:2011:BRB

[Mik11] Mladen Miksa. Book review: Mathematics for the Analysis of Algorithms, by Daniel H. Greene and Donald E. Knuth. ACM SIGACT News, 42(2):12–14, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [GK81, GK82, GK90, GK07].

Miller:1972:BBD

[Mil72a] Raymond E. Miller. A boundary between decidability and undecidability for parallel program schemata (extended abstract). ACM SIGACT News, 4(14):116–120, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Milner:1972:IAS

[Mil72b] Robin Milner. Implementation and applications of Scott’s logic for computable functions. ACM SIGACT News, 4(14):1–6, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Milner:1999:CMS

[Mil99] R. (Robin) Milner. Communicating and Mobile Systems: the π-Calculus. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-05216-4 (hardcover), 0-521-05216-5 (paperback). xii + 161 pp. LCCN QA76.59 .M55 1999.

Milner:2009:SMC

[Mil09] R. (Robin) Milner. The Space and Motion of Communicating Agents. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-49030-8 (hardcover), 0-521-73833-4 (paperback). xxi + 191 pp. LCCN TK5102.5 .M524 2009. URL http:

//www.loc.gov/catdir/enhancements/fy0729/2007014231-b.
html;   http://www.loc.gov/catdir/enhancements/fy0729/
2007014231-d.html;   http://www.loc.gov/catdir/toc/ecip0715/
2007014231.html.

Milchior:2013:RBF

[Mil13] Arthur Milchior. Review of Logical Foundation of Proof Complexity by Stephen Cook and Phuong Nguyen. ACM SIGACT News, 44 (2):14–17, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Miller:2017:RPD

[Mil17a] Allan M. Miller. Review of Practical Data Science with R by Nina Zumel and John Mount. ACM SIGACT News, 48(1):18–22, March 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Miller:2017:RRD

[Mil17b] Allan M. Miller. Review of R for Data Science: Import, Tidy, Transform, Visualize, and Model Data by Hadley Wickham and Garrett Grolemund. ACM SIGACT News, 48(3):14–19, September 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Miller:2018:RMD

[Mil18a] Allan M. Miller. Review of Modern Data Science. ACM SIGACT News, 49(4):13–16, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Miller:2018:SR

[Mil18b] Avery Miller. SIROCCO 2018 review. ACM SIGACT News, 49 (4):66–82, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mitzenmacher:2001:CSC

[Mit01a] Michael Mitzenmacher. Challenging students with creative assignments. ACM SIGACT News, 32(1):70–73, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mitzenmacher:2001:EAR

[Mit01b] Michael Mitzenmacher. An experimental assignment on random processes. ACM SIGACT News, 32(1):74–78, March 2001. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

McGeoch:1999:HPP

[MM99] Catherine C. McGeoch and Bernard M. E. Moret. How to present a paper on experimental work with algorithms. ACM SIGACT News, 30(4):85–90, December 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mullen:2007:FFA

[MM07] Gary L. Mullen and Carl Mummert. Finite Fields and Applications, volume 41 of Student mathematical library. American Mathematical Society, Providence, RI, USA, 2007. ISBN 0-8218-4418-0 (paperback). ix + 175 pp. LCCN QA247.3 .M85 2007.

Malkhi:2013:SCC

[MM13] Dahlia Malkhi and Jean-Philippe Martin. Spanner’s concurrency control. ACM SIGACT News, 44(3):73–77, September 2013. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mazzola:2006:CMC

[MMW06] G. (Guerino) Mazzola, Gérard Milmeister, and Jody Weissmann. Comprehensive Mathematics for Computer Scientists 1: Sets and Numbers, Graphs and Algebra, Logic and Machines, Linear Geometry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2006. ISBN 3-540-36873-6, 3-540-36874-4. xiv + 383 pp. LCCN QA267.7 .M39 2006.

Manna:1972:IMP

[MNV72] Zohar Manna, Stephen Ness, and Jean Vuillemin. Inductive methods for proving properties of programs. ACM SIGACT News, 4 (14):27–50, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mitchell:2001:CG

[MO01] Joseph S. B. Mitchell and Joseph O’Rourke. Computational geometry. ACM SIGACT News, 32(3):63–72, September 1, 2001. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Moffat:1983:EPM

[Mof83] Alistair Moffat. The effect of paged memory upon algorithm performance: short note. ACM SIGACT News, 15(2):45–52, Spring– Summer 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mohar:1993:CPR

[Moh93] Bojan Mohar. Circle packing representation in the plane and other surfaces. ACM SIGACT News, 24(1):59, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mollin:2003:RPK

[Mol03] Richard A. Mollin. RSA and Public-Key Cryptography. Discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, 2003. ISBN 1-58488-338-3. xii + 291 pp. LCCN QA268 .M655 2003. URL http://www.loc.gov/

catdir/enhancements/fy0646/2002031096-d.html;   http:/
/www.loc.gov/catdir/toc/fy036/2002031096.html.

Mollin:2005:CGS

[Mol05] Richard A. Mollin. Codes: The Guide to Secrecy from Ancient to Modern Times. Chapman and Hall/CRC, Boca Raton, FL, USA, 2005. ISBN 1-4200-3508-8 (e-book), 1-58488-470-3 (hardcover). xx + 678 pp. LCCN QA76.9.A25 M67 2005. URL http://marc.

crcnetbase.com/isbn/9781420035087.

Morgan:1976:PDR

[Mor76] Carroll Morgan. A prime decomposition result for parallel systems. ACM SIGACT News, 8(3):14–20, July 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Morrison:1983:SEA

[Mor83] D. R. Morrison. Subtractive encryptors: alternatives to the DES. ACM SIGACT News, 15(1):67–77, Winter–Spring 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Morgenstern:1985:HCF

[Mor85] Jacques Morgenstern. How to compute fast a function and all its derivatives, A variation on the theorem of Baur-Strassen. ACM SIGACT News, 16(4):60–62, April 1985. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Moret:1988:PNP

[Mor88] B. M. E. Moret. Planar NAE3SAT is in P. ACM SIGACT News, 19(2):51–54, Summer 1988. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Moret:1997:WAE

[Mor97] Bernard M. E. Moret. Workshop on algorithm engineering: a report. ACM SIGACT News, 28(4):77–79, December 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mora:2003:SPE

[Mor03] Teo Mora. Solving Polynomial Equation Systems. I. The Kronecker–Duval Philosophy, volume 88 of Encyclopedia of mathematics and its applications. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-81154-6. xiii + 423 pp. LCCN QA218 .M64 2003.

Mora:2005:SPE

[Mor05] Teo Mora. Solving Polynomial Equation Systems: II: Macaulay’s Paradigm and Gröbner Technology, volume 99 of Encyclopedia of mathematics and its applications. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-81156-2. xxii + 759 pp. LCCN QA218 .M64 2005.

Moran:2013:RBS

[Mor13] Tal Moran. Review of Surveillance or Security? by Susan Landau. ACM SIGACT News, 44(1):14–16, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mosses:1992:AS

[Mos92] Peter D. Mosses. Action Semantics, volume 26 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-40347-2. xx + 372 pp. LCCN QA76.7 .M68 1992. URL http://www.loc.gov/

catdir/description/cam025/92249086.html;   http://www.
loc.gov/catdir/toc/cam029/92249086.html.

Moss:2005:JRB

[Mos05] Lawrence S. Moss. Joint review of Introduction To Natural Computation, by Dana H. Ballard; MIT Press, 1997, ISBN 0-262-52258-6 and Mathematical Methods in Artificial Intelligence, by Edward A. Bender, IEEE Press, 1996, ISBN 0-8186-7200-5. ACM SIGACT News, 36(1):21–24, March 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Bal97].

Moses:2010:BSK

[Mos10] Yoram Moses. Behind the scenes of K&CK: the undelivered speech for the 2009 Dijkstra Prize. ACM SIGACT News, 41(1):58–62, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Moshkovitz:2012:GCA

[Mos12] Dana Moshkovitz. Guest column: Algebraic construction of projection PCPs. ACM SIGACT News, 43(1):62–81, March 2012. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Moshkovitz:2019:SSC

[Mos19] Dana Moshkovitz. Sliding scale conjectures in PCP. ACM SIGACT News, 50(3):25–33, September 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Moulin:2003:FDC

[Mou03] Hervé Moulin. Fair Division and Collective Welfare. MIT Press, Cambridge, MA, USA, 2003. ISBN 0-262-13423-3. vi + 289 pp. LCCN HB846 .M68 2003.

Mount:2010:BRB

[Mou10] John Mount. Book review: Combinatorics: the Rota Way, by Joseph P. S. Kung, Gian-Carlo Rota and Catherine H. Yan, published by Cambridge Press, 2009 396 pages, softcover. ACM SIGACT News, 41(2):14–17, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [KRY09].

Moulick:2014:RUC

[Mou14] Subhayan Roy Moulick. Review of Understanding cryptography: a textbook for students and practitioners by Christof Paar and Jan Pelzl. ACM SIGACT News, 45(3):10–12, September 2014. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Moulick:2015:RDS

[Mou15] Subhayan Roy Moulick. Review of: Digital Signatures by Jonathan Katz. ACM SIGACT News, 46(1):10–12, March 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Moulick:2016:RQI

[Mou16] Subhayan Roy Moulick. Review of: Quantum Information Theory by Mark M. Wilde. ACM SIGACT News, 47(3):12–14, September 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Manna:1992:TLR

[MP92] Zohar Manna and Amir Pnueli. The Temporal Logic of Reactive and Concurrent Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97664-7 (New York : v. 1), 3-540-97664-7 (Berlin : v. 1). xiv + 427 (vol.

1) pp. LCCN QA76.6 .M3564 1991. URL http://www.loc.gov/

catdir/enhancements/fy0815/91028181-d.html.

Mereghetti:2021:GCQ

[MP21] Carlo Mereghetti and Beatrice Palano. Guest column: Quantum finite automata: From theory to practice. ACM SIGACT News, 52(3):38–59, September 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3494656.3494666.

Motwani:1995:RA

[MR95] Rajeev Motwani and Prabhaka Raghavan. Randomized Algorithms. ACM SIGACT News, 26(3):48–50, September 1995. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mitchell:1997:APS

[MR97] John C. Mitchell and Jon G. Riecke. The analysis of programming structure. ACM SIGACT News, 28(2):24–31, June 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mehlhorn:2008:ADS

[MS08] Kurt Mehlhorn and Peter Sanders. Algorithms and Data Structures: the Basic Toolbox. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 3-540-77978-7, 3-540-77977-9. xii + 300 pp. LCCN QA76.9.D35 M47 2008eb.

Moss:2004:BRB

[MT04] Lawrence S. Moss and Hans-Jörg Tiede. Book review: Automata Theory and its Applications, by Bakhadyr Khoussainov and Anil Nerode. Birkhäuser Boston, Inc. 2001.: and Automata, Logics, and Infinite Games, by E. Grädel, W. Thomas, and T. Wilke. Springer-Verlag. ACM SIGACT News, 35(1):8–12, March 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [KN01, GTW02].

Mitzenmacher:2005:PCR

[MU05] Michael Mitzenmacher and Eli Upfal. Probability and Computing: Randomized Algorithms and Probabilistic Analysis. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-83540-2. LCCN QA274 .M574 2005.

Muller:1969:UBF

[Mul69] D. E. Muller. Universal Boolean functions. ACM SIGACT News, 1(1):8–11, 1969. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Murphy:2010:BRB

[Mur10] Cillian Murphy. Book review: Introduction to Cryptography, by Hans Delfs and Helmut Knebl, Publisher: Springer, 2007, ISBN 978-3-540-49243-6. ACM SIGACT News, 41(4):42–44, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [DK02, DK07].

Murphy:2013:RBS

[Mur13] Michael Murphy. Review of Triangulations: Structure for Algorithms and Applications by Jesús A. De Lorea, Jörg Rambau, and Francisco Santos. ACM SIGACT News, 44(2):26–28, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Manna:1993:DFC

[MW93] Zohar Manna and Richard Waldinger. The Deductive Foundations of Computer Programming: a One-Volume Version of the Logical Basis for Computer Programming. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-54886-0. xiv + 717 pp. LCCN QA76.6 .M35595 1993.

Nagaraj:2013:RHM

[Nag13a] S. V. Nagaraj. Review of History of mathematics: highways and byways by Amy Dahan-Dalmedico and Jeanne Peiffer. ACM SIGACT News, 44(4):27–28, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nagaraj:2013:RBN

[Nag13b] S. V. Nagaraj. Review of P, NP, and NP-Completeness by Oded Goldreich. ACM SIGACT News, 44(1):10–11, March 2013. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nagaraj:2014:RIM

[Nag14] S. V. Nagaraj. Review of Integrated methods for optimization, second edition, 2012 by John Hooker. ACM SIGACT News, 45 (2):30–32, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nagaraj:2015:RMO

[Nag15a] S. V. Nagaraj. Review of: A Mathematical Orchard — Problems and Solutions by Mark Krusemeyer, George Gilbert, and Loren Larson. ACM SIGACT News, 46(3):21–22, September 2015. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nagaraj:2015:RHF

[Nag15b] S. V. Nagaraj. Review of: Handbook of Finite Fields by Gary L. Mullen and Daniel Panario. ACM SIGACT News, 46(4):17– 18, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nagaraj:2016:RME

[Nag16] S. V. Nagaraj. Review of: Mathematics Everywhere by Martin Aigner and Ehrhard Behrends (Eds.). ACM SIGACT News, 47 (3):22–24, September 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nagaraj:2017:RAC

[Nag17a] S. V. Nagaraj. Review of Algebraic Coding Theory Revised Edition by Elwyn Berlekamp. ACM SIGACT News, 48(1):23–26, March 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nagaraj:2017:RHC

[Nag17b] S. V. Nagaraj. Review of Handbook of Computational Social Choice Edited by Felix Brandt, Vincent Conitzer, Ulle Endriss, Jérôme Lang, and Ariel D. Procaccia. ACM SIGACT News, 48(4):13– 17, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nagaraj:2018:RPU

[Nag18a] S. V. Nagaraj. Review of Power Up:: Unlocking the hidden mathematics in video games. ACM SIGACT News, 49(3):14–19, September 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nagaraj:2018:RTC

[Nag18b] S. V. Nagaraj. Review of Trends in Computational Social Choice Edited by Ulle Endriss. ACM SIGACT News, 49(2):14–17, June 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nagaraj:2019:RCC

[Nag19] S. V. Nagaraj. Review of Codes, Cryptology and Curves with Computer Algebra. ACM SIGACT News, 50(1):14–16, March 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nagaraj:2020:RMD

[Nag20] S. V. Nagaraj. Review of Market Design: a Linear Programming Approach to Auctions and Matching. ACM SIGACT News, 51 (2):11–14, June 2020. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3406678.3406683.

Nagaraj:2021:RNF

[Nag21a] S. V. Nagaraj. Review of Network Flow Algorithms, David P. Williamson. ACM SIGACT News, 52(1):12–15, March 2021. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

URL https://dl.acm.org/doi/10.1145/3457588.3457592.

Nagaraj:2021:RAA

[Nag21b] S. V. Nagaraj. Review of The Age of Algorithms by Serge Abiteboul and Gilles Dowek. ACM SIGACT News, 52(4):15–17, December 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3510382.
3510387.

Nagaraj:2023:RPP

[Nag23] S. V. Nagaraj. Review of Programming for the Puzzled: Learn to Program While Solving Puzzles, Srini Devadas. ACM SIGACT News, 54(1):18–21, March 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/

doi/10.1145/3586165.3586170.

Nahin:1998:ITS

[Nah98] Paul J. Nahin. An Imaginary Tale: the Story of \sqrt{−1} [the square root of minus one]. Princeton University Press, Princeton, NJ, USA, 1998. ISBN 0-691-02795-1. xvi + 257 pp. LCCN QA255 .N34 1998.

Nahin:2004:WLB

[Nah04] Paul J. Nahin. When Least Is Best: How Mathematicians Discovered Many Clever Ways to Make Things as Small (or as Large) as Possible. Princeton University Press, Princeton, NJ, USA, 2004. ISBN 0-691-07078-4 (hardcover). xviii + 370 pp. LCCN QA306.N34 2004.

Narendran:2000:BRB

[Nar00] Paliath Narendran. Book review: Term Rewriting and All That, by Franz Baader and Tobias Nipkow (Cambridge University Press, 313 pages). ACM SIGACT News, 31(3):24–26, September 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BN98].

Narayanan:2014:RLA

[Nar14] Krishnan Narayanan. Review of The LLL Algorithm, edited by Phong Q. Nguyen and Brigitte Vallée. ACM SIGACT News, 45 (4):24–31, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nasim:2013:RVL

[Nas13] Farhan Nasim. Review of Variable-length codes for data compression by David Salomon. ACM SIGACT News, 44(4):24–26, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Natarajan:2007:BRB

[Nat07] Rajesh Natarajan. Book review: The Political Mapping of Cyberspace, by Jeremy W. Crampton, The University of Chicago Press, 2003. ACM SIGACT News, 38(3):17–20, September 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Cra03a, Cra03b].

Natarajan:2010:BRB

[Nat10] Rajesh Natarajan. Book review: The Political Mapping of Cyberspace, by Jeremy W. Crampton, Publisher: The University of Chicago Press, 2003. ACM SIGACT News, 41(1):38–41, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Cra03a, Cra03b].

Navrat:2004:BRB

[Náv04] Pavol Návrat. Book review: Algorithm Design: Foundations, Analysis and Internet Examples, by Michael T. Goodrich and Roberto Tamassia. John Wiley & Sons, Inc. 2001. ACM SIGACT News, 35(2):14–16, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [GT02].

Nehaniv:1999:BRB

[Neh99] Chrystopher L. Nehaniv. Book review: Mathematical Support for Molecular Biology, edited by Martin Farach-Colton, Fred S. Roberts, Martin Vingron, and Michael Waterman. ACM SIGACT News, 30(4):4–7, December 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [FC^{+}99].

Neuburger:2010:BRB

[Neu10] Shoshana Neuburger. Book review: The Burrows–Wheeler Transform: Data Compression, Suffix Arrays, and Pattern Matching, by Donald Adjeroh, Timothy Bell and Amar Mukherjee Springer, 2008. ACM SIGACT News, 41(1):21–24, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [ABM08].

News:1976:RTR

[New76] SIGACT News. Recent technical reports. ACM SIGACT News, 8 (3):26–36, July 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ngo:2023:DTC

[Ngo23] Hung Q. Ngo. Database theory column. ACM SIGACT News, 54 (3):42–45, September 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3623800.3623806.

Nievergelt:2002:FLM

[Nie02] Yves Nievergelt. Foundations of Logic and Mathematics: Applications to Computer Science and Cryptography. Birkhäuser Verlag, Basel, Switzerland, 2002. ISBN 0-8176-4249-8 , 3-7643-4249-8. xvi + 415 pp. LCCN QA9 .N53 2002. URL http:

//www.loc.gov/catdir/enhancements/fy0812/2001052551-d.
html;   http://www.loc.gov/catdir/enhancements/fy0812/
2001052551-t.html.

Nirenberg:1986:PTM

[Nir86] Robert M. Nirenberg. A practical Turing machine representation. ACM SIGACT News, 17(3):35–44, Winter 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nisselbaum:1978:MOE

[Nis78] Y. Nisselbaum. On merging N ordered elements with three elements. ACM SIGACT News, 9(4):14–16, Winter 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nourani:1980:NCC

[Nou80] Farshid Nourani. A note on the constructors of the computable universe. ACM SIGACT News, 12(3):88–89, Fall 1980. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Nowakowski:2002:MGN

[Now02] Richard J. Nowakowski, editor. More Games of No Chance, volume 42 of Mathematical Sciences Research Institute publications. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-80832-4. xii + 535 pp. LCCN QA269 .M63 2002. URL http://

www.loc.gov/catdir/description/cam031/2002034798.html;
 http://www.loc.gov/catdir/toc/cam031/2002034798.html.

Nisan:2007:AGT

[NRTV07] Noam Nisan, Tim Roughgarden, Eva Tardos, and Vijay V. Vazirani, editors. Algorithmic Game Theory. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-87282-0 (hardcover). xxi + 754 pp. LCCN QA269.A43 2007. URL http:

//www.loc.gov/catdir/enhancements/fy0729/2007014231-b.
html;   http://www.loc.gov/catdir/enhancements/fy0729/
2007014231-d.html;   http://www.loc.gov/catdir/toc/ecip0715/
2007014231.html.

Nerode:1993:LA

[NS93] Anil Nerode and Richard A. Shore. Logic for Applications. Texts and monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-94129-0 (New York), 3-540-94129-0 (Berlin). xvii + 365 pp. LCCN QA76.63 .N45 1993.

Nerode:1997:LA

[NS97] Anil Nerode and Richard A. Shore. Logic for Applications. Graduate texts in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1997. ISBN 0-387-94893-7 (hardcover). xiii + 456 pp. LCCN

QA76.9.M35 N47 1997. URL http://www.loc.gov/catdir/
enhancements/fy0815/96043297-d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/96043297-t.html.

Olsen:1983:EMG

[OB83] Dan R. Olsen and Norman Badler. An expression model for graphical command languages. ACM SIGACT News, 15(2):76, Spring– Summer 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ogihara:2001:BRB

[Ogi01] Mitsunori Ogihara. Book review: DNA Based Computers V, by Eric Winfree and David K. Gifford. American Mathematics Society. ACM SIGACT News, 32(2):11–13, June 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [WG00].

OHearn:1998:POA

[O’H98] Peter W. O’Hearn. Polymorphism, objects and abstract types. ACM SIGACT News, 29(4):39–50, December 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Olderog:1991:NTF

[Old91] Ernst-Rüdiger Olderog. Nets, Terms, and Formulas: Three Views of Concurrent Processes and Their Relationship, volume 23 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-40044-9. x + 267 pp. LCCN QA76.58 .O43 1991. URL http://www.loc.gov/

catdir/description/cam025/92160256.html;   http://www.
loc.gov/catdir/toc/cam029/92160256.html.

Oliveira:2004:BRBa

[Oli04a] Carlos Oliveira. Book review: Essentials of Constraint Programming, by T. Fruhwirth and S. Abdennadher. Springer-Verlag. ACM SIGACT News, 35(3):17–20, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [FA03].

Oliveira:2004:BRBc

[Oli04b] Carlos A. S. Oliveira. Book review: Linear Optimization and Extensions — Problems and Solutions, by Dimitris Alevras and Manfred Padberg, Springer-Verlag, 450 pages, $54.95. ACM SIGACT News, 35(4):15–18, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [AP01, Pad10].

Oliveira:2004:BRBb

[Oli04c] Carlos A. S. Oliveira. Book review: Selected Papers on Discrete Mathematics, by D. Knuth, CSLI (Center for the Study of Language and Information Publication) paperback, $72.00. ACM SIGACT News, 35(4):11–14, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Knu03b].

Oliveira:2006:BRB

[Oli06] Carlos A. S. Oliveira. Book review: Computational Complexity: a Quantitative Perspective, by Marius Zimand, Elsevier, ISBN 0-444-82841-9. ACM SIGACT News, 37(1):21–25, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Zim04].

ORourke:1988:CGC

[O’R88] Joseph O’Rourke. Computational geometry column. ACM SIGACT News, 19(3–4):21–26, Fall 1988. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1989:CGCc

[O’R89a] J. O’Rourke. Computational geometry column 8. ACM SIGACT News, 20(4):30, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1989:CGCa

[O’R89b] Joseph O’Rourke. Computational geometry column. ACM SIGACT News, 20(2):10–11, March 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1989:CGCb

[O’R89c] Joseph O’Rourke. Computational geometry column. ACM SIGACT News, 20(3):25–26, July 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1990:CGC

[O’R90] Joseph O’Rourke. Computational geometry column 9. ACM SIGACT News, 21(1):18, January 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1992:BRB

[O’R92a] Joseph O’Rourke. Book review: Intersection and Decomposition Algorithms for Planar Arrangements, by Pankaj K. Agarwal. (Cambridge University Press, Cambridge, 1991. xvii + 277 pp. $39.50 cloth. ISBN 0-521-40446-0). ACM SIGACT News, 23(1): 35–36, Winter 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Aga91].

ORourke:1992:CG

[O’R92b] Joseph O’Rourke. Computational geometry. ACM SIGACT News, 23(2):26–28, Spring 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1993:CGC

[O’R93] Joseph O’Rourke. Computational geometry column 18. ACM SIGACT News, 24(1):20–25, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1994:CG

[O’R94a] Joseph O’Rourke. Computational geometry. ACM SIGACT News, 25(1):31–33, March 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1994:CGCa

[O’R94b] Joseph O’Rourke. Computational geometry column 23. ACM SIGACT News, 25(3):24–27, September 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1994:CGCb

[O’R94c] Joseph O’Rourke. Computational geometry column 24. ACM SIGACT News, 25(4):12–14, December 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1994:CGCc

[O’R94d] Joseph O’Rourke. Computational Geometry in C. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-44034-3, 0-521-44592-2 (paperback). xi + 346 pp. LCCN QA448.D38 O76 1994.

ORourke:1995:CG

[O’R95a] Joseph O’Rourke. Computational geometry. ACM SIGACT News, 26(1):14–16, March 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1995:CGCa

[O’R95b] Joseph O’Rourke. Computational geometry column 26. ACM SIGACT News, 26(2):15–17, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1995:CGCb

[O’R95c] Joseph O’Rourke. Computational geometry column 27. ACM SIGACT News, 26(4):19–21, December 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1996:CGCa

[O’R96a] Joseph O’Rourke. Computational geometry column 28. ACM SIGACT News, 27(2):18–19, June 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1996:CGCb

[O’R96b] Joseph O’Rourke. Computational Geometry Column 29. ACM SIGACT News, 27(3):55–59, September 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1997:CGCa

[O’R97a] Joseph O’Rourke. Computational Geometry Column 30. ACM SIGACT News, 28(1):7, March 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1997:CGCb

[O’R97b] Joseph O’Rourke. Computational geometry column 31. ACM SIGACT News, 28(2):20–23, June 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1997:CGCc

[O’R97c] Joseph O’Rourke. Computational geometry column 32. ACM SIGACT News, 28(3):12–16, September 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:1998:CGC

[O’R98] Joseph O’Rourke. Computational geometry column 33. ACM SIGACT News, 29(2):14–20, June 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ogihara:1999:BRB

[OR99a] Mitsunori Ogihara and Animesh Ray. Book review: DNA Based Computers II, edited by Laura F. Landweber and Eric K. Baum. ACM SIGACT News, 30(4):7–9, December 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [LB99].

ORourke:1999:CGC

[O’R99b] Joseph O’Rourke. Computational geometry column 36. ACM SIGACT News, 30(3):35–38, September 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2000:CGCa

[O’R00a] Joseph O’Rourke. Computational geometry column 38. ACM SIGACT News, 31(1):28–30, March 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2000:CGCb

[O’R00b] Joseph O’Rourke. Computational geometry column 39. ACM SIGACT News, 31(3):47–49, September 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2000:CGCc

[O’R00c] Joseph O’Rourke. Computational geometry column 40. ACM SIGACT News, 31(4):62–73, December 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2001:CGC

[O’R01] Joseph O’Rourke. Computational geometry column 41. ACM SIGACT News, 32(1):53–55, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2002:CGC

[O’R02] Joseph O’Rourke. Computational geometry column 43. ACM SIGACT News, 33(1):58–60, March 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2003:CGC

[O’R03] Joseph O’Rourke. Computational geometry column 44. ACM SIGACT News, 34(2):58–60, June 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2004:AMT

[O’R04a] Joseph O’Rourke. Approximating metrics by tree metrics. ACM SIGACT News, 35(2):60–70, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2004:CGCa

[O’R04b] Joseph O’Rourke. Computational geometry column 45. ACM SIGACT News, 35(2):71–73, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2004:CGCb

[O’R04c] Joseph O’Rourke. Computational geometry column 46. ACM SIGACT News, 35(3):42–45, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2004:GCE

[O’R04d] Joseph O’Rourke. Guest column: error-correcting codes and expander graphs. ACM SIGACT News, 35(3):25–41, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2006:CGCa

[O’R06a] Joseph O’Rourke. Computational geometry column 47. ACM SIGACT News, 37(2):47–49, June 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2006:CGCb

[O’R06b] Joseph O’Rourke. Computational geometry column 48. ACM SIGACT News, 37(3):55–57, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2006:SNC

[O’R06c] Joseph O’Rourke. SIGACT news complexity theory column 52. ACM SIGACT News, 37(3):36–54, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2007:CGC

[O’R07] Joseph O’Rourke. Computational geometry column 49. ACM SIGACT News, 38(2):51–55, June 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2008:CGCa

[O’R08a] Joseph O’Rourke. Computational geometry column 50. ACM SIGACT News, 39(1):73–76, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2008:CGCb

[O’R08b] Joseph O’Rourke. Computational geometry column 51. ACM SIGACT News, 39(3):58–62, September 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

ORourke:2012:CGC

[O’R12] Joseph O’Rourke. Computational geometry column 52. ACM SIGACT News, 43(1):82–85, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ornstein:1985:RR

[OS85] Severo M. Ornstein and Lucy A. Suchman. Reliability and responsibility. ACM SIGACT News, 16(4):53–55, April 1985. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Okhotin:2014:CID

[OS14] Alexander Okhotin and Kai Salomaa. Complexity of input-driven pushdown automata. ACM SIGACT News, 45(2):47–67, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Olariu:2009:VNT

[OW09] Stephan Olariu and Michele Aylene Clark Weigle, editors. Vehicular networks: from theory to practice. Chapman and Hall/CRC computer and information science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-8588-3 (hardcover), 1-4200-8589-1. xi + 473 pp. LCCN TE228.37 .V44 2009.

ODonnell:2017:GCP

[OW17] Ryan O’Donnell and John Wright. Guest column: a primer on the statistics of longest increasing subsequences and quantum states (shortened version). ACM SIGACT News, 48(3):37–59, September 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Olariu:2006:HBA

[OZ06] Stephan Olariu and Albert Y. Zomaya, editors. Handbook of Bioinspired Algorithms and Applications. Chapman and Hall/CRC computer and information science series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2006. ISBN 1-58488-475-4. xvii + 679 pp.

LCCN QA76.618 .H36 2006. URL http://www.loc.gov/catdir/
enhancements/fy0647/2005051859-d.html.

Ozsoyoglu:1997:RP

[Özs97] Z. Meral Özsoyoğlu. Report on PODS’97. ACM SIGACT News, 28(3):17–20, September 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pach:1993:EPG

[Pac93] Janos Pach. Extremal problems in graph drawings (abstract). ACM SIGACT News, 24(1):57, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Padberg:2010:LOE

[Pad10] Manfred W. Padberg. Linear Optimization and Extensions, volume 12 of Algorithms and combinatorics, 0937-5511. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2010. ISBN 3-642-08511-3 (paperback). xx + 501 pp. LCCN T57.74 .P3 2010.

Pager:1970:SOP

[Pag70] David Pager. A solution to an open problem by Knuth. ACM SIGACT News, 2(7):9–10, 1970. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Palopoli:1990:NPU

[Pal90] Luigi Palopoli. A new proof of undecidability of safety of logic queries. ACM SIGACT News, 21(4):69–72, Fall 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Panangaden:1993:BRB

[Pan93] Prakesh Panangaden. Book review: Programs, Recursion and Unbounded Choice. By Wim H. Hesselink. (Cambridge University Press, 1992. xii + 223 pages. ISBN 0-521-40436-3. $39.95). ACM SIGACT News, 24(3):34–37, October 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Hes92].

Papadimitriou:1995:DMA

[Pap95] Christos H. Papadimitriou. Database metatheory: asking the big queries. ACM SIGACT News, 26(3):13–30, September 1995. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Papanikolaou:2005:BRBb

[Pap05a] Nikolaos Papanikolaou. Book review: Classical and Quantum Computing with C++ and Java Simulations, by Yorick Hardy and Willi-Hans Steeb, Birkhäuser Verlag, 2001. ACM SIGACT News, 36(3):5–9, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [HS01].

Papanikolaou:2005:BRBa

[Pap05b] Nikolaos Papanikolaou. Book review: Data Privacy and Security, by David Salomon; Spring-Verlag, 2003, $51.48, Hardcover. ACM SIGACT News, 36(2):8–13, June 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sal03].

Papanikolaou:2010:BRB

[Pap10] Nick Papanikolaou. Book review: The Space and Motion of Communicating Agents, by Robin Milner, Cambridge University Press, 2009, ISBN 978-0-521-73833-0. ACM SIGACT News, 41(3):51– 55, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Mil09].

Papamichail:2011:BRB

[Pap11] Dimitris Papamichail. Book review: Applied Combinatorics, by Fred S. Roberts and Barry Tesman. ACM SIGACT News, 42 (3):29–32, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Rob84, RT05].

Papamichail:2012:BRB

[Pap12a] Dimitris Papamichail. Book review: Design Theory, by C. C. Lindner and C. A. Rodger. ACM SIGACT News, 43(1):35–37, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [LR97].

Papanikolaou:2012:RBT

[Pap12b] Nick Papanikolaou. Review of Algorithms and Theory of Computation Handbook by Mikhail J. Atallah and Marina Blanton. ACM SIGACT News, 43(2):29–32, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Papamichail:2013:RBB

[Pap13a] Dimitris Papamichail. Review of Bioinformatics for Biologists edited by Pavel Pevzner and Ron Shamir. ACM SIGACT News, 44(3):20–24, September 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Papamichail:2013:RIC

[Pap13b] Dimitris Papamichail. Review of Introduction to computational proteomics by Golan Yona. ACM SIGACT News, 44(4):38–41, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:1986:PSS

[Par86] I. Parberry. Parallel speedup of sequential machines: a defense of parallel computation thesis. ACM SIGACT News, 18(1):54– 67, Summer 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:1988:HPP

[Par88] Ian Parberry. How to present a paper in theoretical computer science: a speaker’s guide for students. ACM SIGACT News, 19 (2):42–47, ???? 1988. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:1989:GNR

[Par89a] I. Parberry. A guide for new referees in theoretical computer science. ACM SIGACT News, 20(4):92–99, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:1989:JRC

[Par89b] I. Parberry. The journal review column. ACM SIGACT News, 20 (4):31–37, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:1990:JRCa

[Par90a] Ian Parberry. The journal review column. ACM SIGACT News, 21 (1):21, January 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:1990:JRCb

[Par90b] Ian Parberry. The journal review column. ACM SIGACT News, 21(2):21, Spring 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parikh:1991:TFL

[Par91] Rohit Parikh. A test for fuzzy logic. ACM SIGACT News, 22(3): 49–50, Summer 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:1994:FRT

[Par94] Ian Parberry. A form for referees in theoretical computer science. ACM SIGACT News, 25(4):96–102, December 1994. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:1995:PA

[Par95a] Ian Parberry. Problems on Algorithms. ACM SIGACT News, 26 (2):50–56, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:1995:SW

[Par95b] Ian Parberry. Surfing the Web. ACM SIGACT News, 26(1):99– 101, March 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:1996:PE

[Par96] Ian Parberry. Paul Erdős (1913–1996). ACM SIGACT News, 27 (4):62–65, December 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:1998:EYW

[Par98a] Ian Parberry. Everything you wanted to know about the running time of Mergesort but were afraid to ask. ACM SIGACT News, 29 (2):50–57, June 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Paredaens:1998:DTC

[Par98b] Jan Paredaens. Database theory column: report on PODs ’98. ACM SIGACT News, 29(3):23–26, September 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:2000:EL

[Par00a] Ian Parberry. Editor’s Letter. ACM SIGACT News, 31(1):1, March 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:2000:HPP

[Par00b] Ian Parberry. How to present a paper in theoretical computer science: a speaker’s guide for students. ACM SIGACT News, 31(1): 77–86, March 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:2000:RIM

[Par00c] Ian Parberry. Report on the 6th international meeting on DNAbased computers. ACM SIGACT News, 31(4):118–120, December 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parberry:2002:TCS

[Par02a] Ian Parberry. Things a computer scientist rarely talks about. ACM SIGACT News, 33(4):14–15, December 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Parkes:2002:ILM

[Par02b] Alan Parkes. Introduction to Languages, Machines and Logic: Computable Languages, Abstract Machines and Formal Logic.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 1-85233-464-9 (paperback). xi + 351 pp. LCCN QA267.3 .P37 2002. URL http:

//www.loc.gov/catdir/enhancements/fy0816/2001054282-d.
html;   http://www.loc.gov/catdir/enhancements/fy0816/
2001054282-t.html.

Parkes:2008:CIL

[Par08] Alan Parkes. A Concise Introduction to Languages and Machines. Undergraduate topics in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-84800-120-7 (paperback). x, vi–vii, 343 pp. LCCN QA267.3 .P36 2008.

Parrish:2011:BRB

[Par11] Andy Parrish. Book review: Erdős on Graphs: His Legacy of Unsolved Problems, by Fan Chung and Ron Graham. ACM SIGACT News, 42(3):16–20, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [CEG98].

Pasanen:1999:PDT

[Pas99] Tomi Pasanen. Ph.D. Thesis: In-place algorithms for sorting problems. ACM SIGACT News, 30(2):61, June 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Paterson:1972:DPC

[Pat72] Michael S. Paterson. Decision problems in computational models. ACM SIGACT News, 4(14):74–82, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Paz:1970:EWR

[Paz70] A. Paz. Events which are not representable by a probabilistic automaton. ACM SIGACT News, 2(4):8–11, 1970. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Paz:2013:BWR

[Paz13] Ami Paz. Bremen workshop review. ACM SIGACT News, 44 (4):91–97, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2003:BRB

[PC03] Riccardo Pucella and Stephen Chong. Book review: Type-Logical Semantics. ACM SIGACT News, 34(1):6–17, March 2003. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Car98].

Pardalos:1998:NDC

[PD98] P. M. (Panos M.) Pardalos and Dingzhu Du, editors. Network Design: Connectivity and Facilities Location: DIMACS Workshop, April 28–30, 1997, volume 40 of DIMACS series in discrete mathematics and theoretical computer science. American Mathematical Society, Providence, RI, USA, 1998. ISBN 0-8218-0834-6. LCCN T57.85 .N4557 1998.

Peleg:2013:PID

[Pel13] David Peleg. Prize for innovation in distributed computing: awarded to Roger Wattenhofer. ACM SIGACT News, 44(2):89–91, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Petersen:1994:TWO

[Pet94] H. Petersen. Two-way one-counter automata accepting bounded languages. ACM SIGACT News, 25(3):102–105, September 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Peterson:2002:MTS

[Pet02] Ivars Peterson. Mathematical Treks: From Surreal Numbers to Magic Circles. Spectrum series; A science news book. Mathematical Association of America, Washington, DC, USA, 2002. ISBN 0-88385-537-2. x + 170 pp. LCCN QA99 .P47 2002. URL http://

www.loc.gov/catdir/description/cam022/2001098094.html;
 http://www.loc.gov/catdir/toc/cam027/2001098094.html.

Petride:2004:BRB

[Pet04] Sabina Petride. Book review: Concurrent and Real-Time Systems: the CSP Approach, by Steve Schneider. Wiley 1999. ACM SIGACT News, 35(2):4–12, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sch00].

Petzold:2008:ATG

[Pet08] Charles Petzold. The Annotated Turing: a Guided Tour Through Alan Turing’s Historic Paper on Computability and the Turing Machine. John Wiley, New York, NY, USA, 2008. ISBN 0-470-22905-5 (paperback). xii + 372 pp. LCCN QA267 .P48 2008.

Petkovic:2009:FPG

[Pet09] Miodrag Petković. Famous Puzzles of Great Mathematicians. American Mathematical Society, Providence, RI, USA, 2009. ISBN 0-8218-4814-3. xviii + 325 pp. LCCN QA95 .P4358 2009.

Papadimitriou:1996:FCC

[PGW^{+}96] Christos H. Papadimitriou, Oded Goldreich, Avi Wigderson, Alexander A. Razborov, and Michael Sipser. The future of computational complexity theory: part I. ACM SIGACT News, 27 (3):6–12, September 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pichler:2021:DTC

[Pic21] Reinhard Pichler. Database theory column report on PODS 2021. ACM SIGACT News, 52(3):69–72, September 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3494656.3494669.

Pierce:2002:TPL

[Pie02] Benjamin C. Pierce. Types and Programming Languages. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-16209-1 (hardcover). xxi + 623 pp. LCCN QA76.7 .P54 2002.

Pighizzini:2015:GCO

[Pig15] Giovanni Pighizzini. Guest column: One-tape Turing machine variants and language recognition. ACM SIGACT News, 46(3):37–55, September 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pincock:2010:BRB

[Pin10] Christopher Pincock. Book review: Proof and Other Dilemmas: Mathematics and Philosophy, edited by Bonnie Gold & Roger A. Simons, Spectrum Series, MAA, 2008 346 pages, hardcover. ACM SIGACT News, 41(2):28–33, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [GS08].

Pippenger:1997:TC

[Pip97] Nicholas Pippenger. Theories of Computability. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-55380-6 (hardcover). ix + 251 pp. LCCN QA267 .P553 1997. URL http:

//www.loc.gov/catdir/description/cam028/96044633.html;
 http://www.loc.gov/catdir/toc/cam023/96044633.html.

Pittenger:2000:IQC

[Pit00] Arthur O. Pittenger. An Introduction to Quantum Computing Algorithms, volume 19 of Progress in computer science and applied logic. Birkhäuser Verlag, Basel, Switzerland, 2000. ISBN 0-8176-4127-0 , 3-7643-4127-0. xii + 138 pp. LCCN QA76.889 .P58 2000.

Porto:1980:AS

[PM80] António G. Porto and Armando B. Matos. Ackermann and the superpowers. ACM SIGACT News, 12(3):90–95, Fall 1980. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pollack:2009:ABD

[Pol09] Paul Pollack. Not Always Buried Deep: a Second Course in Elementary Number Theory. American Mathematical Society, Providence, RI, USA, 2009. ISBN 0-8218-4880-1 (hardcover). xv + 303 pp. LCCN QA241 .P657 2009.

Pop:2012:RBA

[Pop12] Mihai Pop. Review of Parallel Algorithms by Henri Casanova, Arnaud Legrand, and Yves Robert. ACM SIGACT News, 43(2): 11–14, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Porreca:2012:RBA

[Por12] Antonio E. Porreca. Review of Algorithmic Adventures: From Knowledge to Magic, by Juraj Hromković. ACM SIGACT News, 43(3):22–24, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Postow:1998:BRB

[Pos98] Brian Postow. Book review: Basic Simple Type Theory, by J. Roger Hindley (Cambridge University Press 1997). ACM SIGACT News, 29(1):5–8, March 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Hin97].

Postow:1999:BRB

[Pos99] Brian Postow. Book review: A Theory of Objects, by Martin Abadi and Luca Cardelli (Springer-Verlag, 1996): Series–Monographs in Computer Science. ACM SIGACT News, 30(1):9–11, March 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [AC96].

Pritchard:2010:BRB

[Pri10] David Pritchard. Book review: Decisions and Elections: Explaining the Unexpected, by Donald G. Saari, Cambridge University Press, 2001, ISBN 0-521-80816-2. ACM SIGACT News, 41(3):30– 33, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Saa01].

Proskurowski:1990:RGG

[Pro90] Andrzej Proskurowski. Report on GRA-GRA: 4th International Workshop on Graph Grammars and Their Applications to Computer Science. ACM SIGACT News, 21(2):39, Spring 1990. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pruhs:1998:HDD

[Pru98] Kirk Pruhs. How to design dynamic programming algorithms sans recursion. ACM SIGACT News, 29(1):32–35, March 1998. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pruim:2000:BRB

[Pru00] Randall Pruim. Book review: Hilbert’s Tenth Problem, by Yttri Matiyasevich. ACM SIGACT News, 31(1):4, March 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Mat93].

Plotkin:2000:PLI

[PSTM00] G. (Gordon) Plotkin, Colin P. Stirling, Mads Tofte, and R. (Robin) Milner, editors. Proof, Language, and Interaction: Essays in Honour of Robin Milner. Foundations of computing. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-16188-5 (hardcover). xiv + 722 pp. LCCN QA76.P699 2000.

Parter:2014:DR

[PT14] Merav Parter and Edward Talmage. DISC 2014 review. ACM SIGACT News, 45(4):94–99, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2000:BRB

[Puc00] Riccardo Pucella. Book review: Communicating and Mobile Systems: the π-Calculus, by Robin Milner. ACM SIGACT News, 31 (4):24–26, December 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Mil99].

Pucella:2001:BRBb

[Puc01a] Riccardo Pucella. Book review: Dynamic Logic (Foundations of Computing), by D. Harel, D. Kozen and J. Tiuryn. ACM SIGACT News, 32(4):9–17, December 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [HKT00].

Pucella:2001:BRBa

[Puc01b] Riccardo Pucella. Book review: Proof, Language, and Interaction: Essays in Honour of Robin Milner, edited by Plotkin, Stirling and Tofte. ACM SIGACT News, 32(1):12–16, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [PSTM00].

Pucella:2003:JRB

[Puc03] Riccardo Pucella. Joint review of Foundations of Cryptography: Basic Tools, by O. Goldreich. Cambridge University Press, and Modelling and Analysis of Security Protocols, by P. Ryan and S. Schneider. Addison Wesley. ACM SIGACT News, 34(4):26–31, December 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Gol01].

Pucella:2004:CMC

[Puc04a] Riccardo Pucella. Coordination mechanisms for congestion games. ACM SIGACT News, 35(4):58–71, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2004:SC

[Puc04b] Riccardo Pucella. Specifying confidentiality. ACM SIGACT News, 35(4):72–83, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2005:ASN

[Puc05a] Riccardo Pucella. ACM SIGACT news distributed computing column 19. ACM SIGACT News, 36(3):41–50, September 2005. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2005:APP

[Puc05b] Riccardo Pucella. Algorithmic problems in power management. ACM SIGACT News, 36(2):63–76, June 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2005:BRB

[Puc05c] Riccardo Pucella. Book review: Logic for Computer Scientists, by Uwe Schöning, Birkhäuser, 1994. ACM SIGACT News, 36(3):17– 19, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sch08a, Sch89].

Pucella:2005:FIT

[Puc05d] Riccardo Pucella. The finite and the infinite in temporal logic. ACM SIGACT News, 36(1):86–99, March 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2005:LVE

[Puc05e] Riccardo Pucella. Logical verification and equational verification. ACM SIGACT News, 36(2):77–88, June 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2005:SNLa

[Puc05f] Riccardo Pucella. SIGACT news logic column 13. ACM SIGACT News, 36(3):51–66, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2005:SNLb

[Puc05g] Riccardo Pucella. SIGACT news logic column 14. ACM SIGACT News, 36(4):47–69, December 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2005:TDP

[Puc05h] Riccardo Pucella. Three dozen papers on online algorithms. ACM SIGACT News, 36(1):71–85, March 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2006:ASN

[Puc06a] Riccardo Pucella. ACM SIGACT news distributed computing column 22. ACM SIGACT News, 37(2):50–56, June 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2006:SCC

[Puc06b] Riccardo Pucella. Security and composition of cryptographic protocols: a tutorial (Part I). ACM SIGACT News, 37(3):67–92, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2006:SNL

[Puc06c] Riccardo Pucella. SIGACT news logic column 15. ACM SIGACT News, 37(2):57, June 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2007:ALB

[Puc07a] Riccardo Pucella. Alternative logics: a book review: SIGACT news logic column 18. ACM SIGACT News, 38(2):75–86, June 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2007:Ia

[Puc07b] Riccardo Pucella. Introduction. ACM SIGACT News, 38(3):76, September 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2007:Ib

[Puc07c] Riccardo Pucella. Introduction. ACM SIGACT News, 38(4):64, December 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pucella:2010:SNL

[Puc10] Riccardo Pucella. SIGACT news logic column 21. ACM SIGACT News, 41(1):74, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Pudlak:2006:GCA

[Pud06] Pavel Pudlák. Gödel and computations: a 100th anniversary retrospective. ACM SIGACT News, 37(4):13–21, December 2006. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Purushothaman:1993:BRB

[Pur93] S. Purushothaman. Book review: Nets, Terms and Formulas. By E.-R. Olderog. (Cambridge University Press, 1991. x + 267pp. ISBN 0-521-40044-9. $49.95). ACM SIGACT News, 24(1):26–27, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Old91].

Puxang:1997:BRB

[Pux97] K. Puxang. Book review: Metamathematics, Machines, and Gödels Proof, by N. Shankar. ACM SIGACT News, 28(4):16– 19, December 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sha94].

Pass:2021:GCA

[PV21] R. Pass and M. Venkitasubramaniam. Guest column: Average-case complexity through the lens of interactive puzzles. ACM SIGACT News, 52(1):47–69, March 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/

doi/10.1145/3457588.3457598.

Pottmann:2001:CLG

[PW01] Helmut Pottmann and Johannes Wallner. Computational Line Geometry. Mathematics and visualization. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42058-4. ix + 563 pp. LCCN QA608 .P68 2001. URL http:

//www.loc.gov/catdir/enhancements/fy0816/2001040054-d.
html;   http://www.loc.gov/catdir/enhancements/fy0816/
2001040054-t.html.

Pottmann:2009:CLG

[PW09] Helmut Pottmann and Johannes Wallner. Computational Line Geometry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-642-04017-9 (softcover). ix + 563 pp. LCCN QA608 .P68 2010.

Petkovsek:1996:AB

[PWZ96] Marko Petkovšek, Herbert S. Wilf, and Doron Zeilberger. A = B. A. K. Peters, Ltd., Wellesley, MA, USA, 1996. ISBN 1-56881-063-6. xii + 212 pp. LCCN QA164.9 .P48 1996.

Pyster:1976:LCD

[Pys76] Arthur Pyster. A language construct for “dovetailing”. ACM SIGACT News, 8(1):38–40, January 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Qiang:1986:PAS

[Qia87] Z. Qiang. An O(ln n) parallel algorithm for the subset sum problem. ACM SIGACT News, 18(2):57–63, September/March 1986– 1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Radeva:2014:RBW

[Rad14] Mira Radeva. Review of BDA Workshop 2014. ACM SIGACT News, 45(4):100–104, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ragde:1991:EPT

[Rag91] Prabhakar Ragde. Extracting poetry from technical papers. ACM SIGACT News, 22(2):50, Spring 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2000:PDC

[Raj00] Sergio Rajsbaum. Principles of distributed computing: an exciting challenge. ACM SIGACT News, 31(4):52–61, December 2000. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2001:ASNa

[Raj01a] Sergio Rajsbaum. ACM SIGACT news distributed computing column 2. ACM SIGACT News, 32(1):32, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2001:ASNb

[Raj01b] Sergio Rajsbaum. ACM SIGACT News distributed computing column 3. ACM SIGACT News, 32(2):44–45, June 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2001:ASNc

[Raj01c] Sergio Rajsbaum. ACM SIGACT news distributed computing column 5. ACM SIGACT News, 32(4):34–58, December 2001. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2001:DC

[Raj01d] Sergio Rajsbaum. Distributed Computing. ACM SIGACT News, 32(3):53–62, September 1, 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajaraman:2002:TCR

[Raj02a] Rajmohan Rajaraman. Topology control and routing in ad hoc networks: a survey. ACM SIGACT News, 33(2):60–73, June 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2002:ASNa

[Raj02b] Sergio Rajsbaum. ACM SIGACT news Distributed Computing Column 6. ACM SIGACT News, 33(1):46–53, March 2002. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2002:ASNb

[Raj02c] Sergio Rajsbaum. ACM SIGACT news distributed computing column 7. ACM SIGACT News, 33(2):48–51, June 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2002:ASNc

[Raj02d] Sergio Rajsbaum. ACM SIGACT news distributed computing column 8. ACM SIGACT News, 33(3):50, September 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2002:ASNd

[Raj02e] Sergio Rajsbaum. ACM SIGACT news distributed computing column 9. ACM SIGACT News, 33(4):37–54, December 2002. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2003:ASNa

[Raj03a] Sergio Rajsbaum. ACM SIGACT news distributed computing column 11. ACM SIGACT News, 34(2):42–57, June 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2003:ASNb

[Raj03b] Sergio Rajsbaum. ACM SIGACT News distributed computing column 12. ACM SIGACT News, 34(3):40–61, September 2003. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2003:ASNc

[Raj03c] Sergio Rajsbaum. ACM SIGACT news distributed computing column 13. ACM SIGACT News, 34(4):53–56, December 2003. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2003:PIC

[Raj03d] Sergio Rajsbaum. Partial information classes. ACM SIGACT News, 34(1):32–46, March 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2004:ASNa

[Raj04a] Sergio Rajsbaum. ACM SIGACT news distributed computing column 15. ACM SIGACT News, 35(3):46–57, September 2004. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2004:ASNb

[Raj04b] Sergio Rajsbaum. ACM SIGACT news distributed computing column 16. ACM SIGACT News, 35(4):38, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2004:CME

[Raj04c] Sergio Rajsbaum. The computation of market equilibria. ACM SIGACT News, 35(4):23–37, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2004:CGC

[Raj04d] Sergio Rajsbaum. Computational geometry column 46. ACM SIGACT News, 35(3):42–45, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2004:LS

[Raj04e] Sergio Rajsbaum. Larry Stockmeyer: 1948–2004. ACM SIGACT News, 35(4):39, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2005:ASNa

[Raj05a] Sergio Rajsbaum. ACM SIGACT news distributed computing column 19. ACM SIGACT News, 36(3):41–50, September 2005. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2005:ASNb

[Raj05b] Sergio Rajsbaum. ACM SIGACT news distributed computing column 20. ACM SIGACT News, 36(4):36–46, December 2005. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2005:FCS

[Raj05c] Sergio Rajsbaum. Four colors suffice! ACM SIGACT News, 36 (2):43–44, June 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2005:GCN

[Raj05d] Sergio Rajsbaum. Guest Column: NP-complete problems and physical reality. ACM SIGACT News, 36(1):30–52, March 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2005:RP

[Raj05e] Sergio Rajsbaum. Report on PODS 2005. ACM SIGACT News, 36(3):39–40, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2006:ASNa

[Raj06a] Sergio Rajsbaum. ACM SIGACT news distributed computing column 22. ACM SIGACT News, 37(2):50–56, June 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2006:ASNb

[Raj06b] Sergio Rajsbaum. ACM SIGACT news distributed computing column 24. ACM SIGACT News, 37(4):58–84, December 2006. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2006:CGCa

[Raj06c] Sergio Rajsbaum. Computational geometry column 47. ACM SIGACT News, 37(2):47–49, June 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2006:CGCb

[Raj06d] Sergio Rajsbaum. Computational geometry column 48. ACM SIGACT News, 37(3):55–57, September 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2006:OQT

[Raj06e] Sergio Rajsbaum. Open questions in the theory of semifeasible computation. ACM SIGACT News, 37(1):47–65, March 2006. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2007:I

[Raj07a] Sergio Rajsbaum. Introduction. ACM SIGACT News, 38(3):55, September 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rajsbaum:2007:IDR

[Raj07b] Sergio Rajsbaum. Invitation to data reduction and problem kernelization. ACM SIGACT News, 38(1):31–45, March 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ramshaw:1989:SON

[Ram89] L. Ramshaw. Suitening our nomenclature. ACM SIGACT News, 20(1):60–61, Winter 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Raman:1994:SAA

[Ram94] Rajeev Raman. A simpler analysis of algorithm 65 (find). ACM SIGACT News, 25(2):86–89, June 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Raman:1997:RRS

[Ram97] Rajeev Raman. Recent results on the single-source shortest paths problem. ACM SIGACT News, 28(2):81–87, June 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rauff:2008:BRB

[Rau08] James V. Rauff. Book review: A Beginner’s Guide to Discrete Mathematics, by W. D. Wallis, Birkhäuser, 2003. ACM SIGACT News, 39(1):16–18, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Wal03, Wal12].

Rauff:2018:RWG

[Rau18] James V. Rauff. Review of Words and Graphs by Sergey Kitaev and Vadim Lozin. ACM SIGACT News, 49(2):7–9, June 2018. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ravi:1990:JBR

[Rav90a] S. S. Ravi. Journal backlog report. ACM SIGACT News, 21(1): 26, January 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ravikumar:1990:SAT

[Rav90b] B. Ravikumar. Some applications of a technique of Sakoda and Sipser. ACM SIGACT News, 21(4):73–77, Fall 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Raynal:2005:SIF

[Ray05] M. Raynal. A short introduction to failure detectors for asynchronous distributed systems. ACM SIGACT News, 36(1):53–70, March 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Raynal:2021:DCF

[Ray21] Michel Raynal. Distributed computability: a few results masters students should know. ACM SIGACT News, 52(2):92–110, June 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3471469.
3471484.

Razborov:2007:EMM

[Raz07] Alexander (Sasha) A. Razborov. Eulogy: Michael (Misha) Alekhnovich 1978–2006. ACM SIGACT News, 38(1):70–71, March 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Razborov:2016:GCP

[Raz16] Alexander Razborov. Guest column: Proof complexity and beyond. ACM SIGACT News, 47(2):66–86, June 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Robinson:2011:IBO

[RB11] Peter N. (Peter Nicholas) Robinson and Sebastian Bauer. Introduction to bio-ontologies. Chapman and Hall/CRC mathematical and computational biology series. Taylor and Francis, Boca Raton, FL, USA, 2011. ISBN 1-4398-3665-5 (hardcover). xxvii + 488 pp. LCCN QH324.2 .R618 2011.

Reeker:1971:NBN

[Ree71] L. Reeker. A note on a Badly Nondeterministic automaton. ACM SIGACT News, 3(13):22–24, December 1971. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Reeker:1977:BRB

[Ree77] Larry H. Reeker. Book review: Programs, Machines, and Computation: an Introduction to the Theory of Computing, by K. L. Clark and D. F. Cowell. ACM SIGACT News, 9(3):20–22, Fall 1977. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [CC76].

Reeds:2001:BRB

[Ree01] Jim Reeds. Book review: The Code Book: the Evolution of Secrecy from Mary Queen Of Scots to Quantum Cryptography, by Simon Singh. Anchor Books. ACM SIGACT News, 32(2):6–11, June 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sin99].

Regan:1993:MML

[Reg93] Kenneth W. Regan. Machine models and linear time complexity. ACM SIGACT News, 24(3):5–15, October 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Reid:2006:ZIW

[Rei06] Constance Reid. From Zero to Infinity: What Makes Numbers Interesting. A. K. Peters, Ltd., Wellesley, MA, USA, fifth edition, 2006. ISBN 1-56881-273-6. xvii + 188 pp. LCCN QA93 .R42 2006.

Reynal:2005:SIF

[Rey05] Michel Reynal. A short introduction to failure detectors for asynchronous distributed systems. ACM SIGACT News, 36(1):53–70, March 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Reyzin:2011:BRB

[Rey11] Lev Reyzin. Book review: Famous Puzzles of Great Mathematicians, by Miodrag S. Petković. ACM SIGACT News, 42(3):36– 39, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Pet09].

Rosenstein:1997:DMS

[RFR97] Joseph G. Rosenstein, Deborah S. Franzblau, and Fred S. Roberts, editors. Discrete Mathematics in the Schools, volume 36 of DI-MACS series in discrete mathematics and theoretical computer science, 1052-1798. American Mathematical Society, National Council of Teachers of Mathematics, Providence, RI, USA, 1997. ISBN 0-8218-0448-0 (hardcover). LCCN QA11.A1 D57 1997.

Rosen:2006:CZK

[RG06] Alon Rosen and Oded Goldreich. Concurrent Zero-Knowledge. Information security and cryptography, 1619-7100. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-32938-2 (hardcover). xiii + 184 pp. LCCN QA76.9.A25 R657 2006.

Ravi:2010:TML

[RGL10] Srivatsan Ravi, Vincent Gramoli, and Victor Luchangco. Transactional memory, linking theory and practice. ACM SIGACT News, 41(4):109–115, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ricci:1973:SFC

[Ric73] G. Ricci. Some further comments about nothing. ACM SIGACT News, 5(3):29–30, July 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rice:1994:MEL

[Ric94] Rice University. In memoriam Eugene L. Lawler. ACM SIGACT News, 25(4):108–109, December 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Richards:2000:NRT

[Ric00] Dana Richards. NSF Report: Theory of Computing Program. ACM SIGACT News, 31(4):37–38, December 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Richards:2021:TN

[Ric21] Dana Richards. Teaching nondeterminism. ACM SIGACT News, 52(2):85–90, June 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3471469.3471482.

Riecke:1993:BRB

[Rie93] Jon G. Riecke. Book review: Action Semantics. By Peter D. Mosses. (Cambridge University Press, 1992. xx + 372pp. ISBN 0-521-40347-2. $49.95). ACM SIGACT News, 24(2):24–25, Spring 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Mos92].

Riecke:1997:APS

[Rie97] Jon G. Riecke. The analysis of programming structure. ACM SIGACT News, 28(2):24–31, June 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Riecke:1999:CGC

[Rie99a] Jon G. Riecke. Computational geometry column 35. ACM SIGACT News, 30(2):31–32, June 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Riecke:1999:RP

[Rie99b] Jon G. Riecke. Report on POPL 1999. ACM SIGACT News, 30(1): 28–29, March 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rieffel:2010:BRB

[Rie10] Eleanor Rieffel. Book review: Quantum Computer Science: an Introduction, by N. David Mermin, Cambridge Press, 2007, 236 pages. ACM SIGACT News, 41(3):39–44, September 2010. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Mer07].

Robertson:1970:CDL

[Rob70] Edward L. Robertson. A corrected definition of “local speed-up”. ACM SIGACT News, 2(6):15–16, 1970. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Roberts:1984:AC

[Rob84] Fred S. Roberts. Applied Combinatorics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-039313-4. xviii + 606 pp. LCCN QA164 .R6 1984. US$37.95.

Rodger:1996:RFI

[Rod96] Susan H. Rodger. Report on The First International Workshop on Implementing Automata 1996. ACM SIGACT News, 27(4):38– 45, December 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rodger:1999:TAT

[Rod99] Susan H. Rodger. Teaching automata theory with JFLAP. ACM SIGACT News, 30(4):53–56, December 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rodrigues:2009:BLT

[Rod09] Rodrigo Rodrigues. Barbara Liskov’s Turing award. ACM SIGACT News, 40(4):68–70, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rogers:2008:BRB

[Rog08] John D. Rogers. Book review: The Art of Computer Programming, Volume 4, Fascicles 2, 3, and 4, by Donald E. Knuth, Pearson Education (Addison-Wesley), 2005. ACM SIGACT News, 39(2): 32–35, June 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Knu11].

Rogers:2013:RBA

[Rog13] John D. Rogers. Review of The Art of Computer Programming: volume 4a by Donald E. Knuth. ACM SIGACT News, 44(2):36–39, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rogers:2018:RAC

[Rog18] John D. Rogers. Review of The Art of Computer Programming Fascicle 6 ‘Satisfiability’ by Donald E. Knuth. ACM SIGACT News, 49(1):12–15, March 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rojat:2012:BRB

[Roj12] Antoine Rojat. Book review: Cryptanalysis of RSA and its variants, by Jason Hinek. ACM SIGACT News, 43(1):16–18, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Hin10].

Rosenberg:1973:TRR

[Ros73a] Arnold L. Rosenberg. On the time required to recognize properties of graphs: a problem. ACM SIGACT News, 5(4):15–16, October 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rosenburg:1973:TRR

[Ros73b] A. L. Rosenburg. On the time required to recognize properties of graphs: a problem. ACM SIGACT News, 5(??):15–16, ???? 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rosen:1975:CRT

[Ros75] Barry K. Rosen. A Church–Rosser theorem for graph grammars. ACM SIGACT News, 7(3):26–31, Summer 1975. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rosen:1977:AGP

[Ros77] Barry K. Rosen. Arcs in graphs are NOT pairs of nodes. ACM SIGACT News, 9(3):25–27, Fall 1977. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1992:EFa

[Ros92a] Rocky Ross. Education forum. ACM SIGACT News, 23(1):18–19, Winter 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1992:EFb

[Ros92b] Rocky Ross. Education forum. ACM SIGACT News, 23(2):50–53, Spring 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1992:EFc

[Ros92c] Rocky Ross. Education forum. ACM SIGACT News, 23(3):33, Summer 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1993:CSL

[Ros93a] R. Ross. Computer science laboratories. ACM SIGACT News, 24 (4):45–48, December 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1993:EFI

[Ros93b] Rocky Ross. Education forum: An Introductory Computer Science Curriculum Incorporating Theory. ACM SIGACT News, 24(2): 27–29, Spring 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1993:EFL

[Ros93c] Rocky Ross. Education forum: Lecture and lab syllabus for a breadth-first introductory computer science course sequence following the data structures and algorithms paradigm. ACM SIGACT News, 24(3):38–43, October 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1993:EFN

[Ros93d] Rocky Ross. Education forum: New Courses on the Horizon. ACM SIGACT News, 24(1):30, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1994:EFP

[Ros94a] Rocky Ross. Education forum: Project Impact: NSF-funded science education projects. ACM SIGACT News, 25(3):49–52, September 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1994:EFD

[Ros94b] Rocky Ross. Education forum: The Dynalab Animation System. ACM SIGACT News, 25(4):49–54, December 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1995:EWW

[Ros95a] Rocky Ross. Education and the World Wide Web. ACM SIGACT News, 26(3):45–48, September 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1995:EFAa

[Ros95b] Rocky Ross. Education Forum: Animated Textbooks: A Current Example. ACM SIGACT News, 26(1):40–43, March 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1995:EFAb

[Ros95c] Rocky Ross. Education Forum: Animation Activities at SRC. ACM SIGACT News, 26(2):47–50, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1995:EFB

[Ros95d] Rocky Ross. Education Forum: Beware the Backlash: The Teaching vs. Research Conundrum Revisited. ACM SIGACT News, 26 (4):36–38, December 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1996:EFL

[Ros96a] Rockford J. Ross. Education forum: loops and induction proofs. ACM SIGACT News, 27(4):15–19, December 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1996:EFJ

[Ros96b] Rocky Ross. Education Forum: Java… Hot Java!: What is that brewing on the Web? ACM SIGACT News, 27(1):24–27, March 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1996:EFM

[Ros96c] Rocky Ross. Education Forum: Making Theory Come Alive. ACM SIGACT News, 27(2):58–64, June 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1996:EFR

[Ros96d] Rocky Ross. Education forum: retrospective. ACM SIGACT News, 27(3):69–70, September 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1997:EF

[Ros97a] Rocky Ross. Education forum. ACM SIGACT News, 28(1):13, March 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1997:EFC

[Ros97b] Rocky Ross. Education Forum: Computer Science, Engineering, or Home Economics? ACM SIGACT News, 28(4):37–40, December 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1997:EFM

[Ros97c] Rocky Ross. Education Forum: Matter of Salesmanship. ACM SIGACT News, 28(2):37–39, June 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1997:EFW

[Ros97d] Rocky Ross. Education forum: Where Have all the Women Gone? ACM SIGACT News, 28(3):41–49, September 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1998:EFR

[Ros98a] Rocky Ross. Education forum: Reflections on turning fifty. ACM SIGACT News, 29(1):21–31, March 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1998:EFW

[Ros98b] Rocky Ross. Education forum: Web Enhanced Textbooks. ACM SIGACT News, 29(4):51–57, December 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1998:GDW

[Ros98c] Rocky Ross. Graduate degrees on the Web. ACM SIGACT News, 29(3):56–57, September 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1998:MW

[Ros98d] Rocky Ross. Mathematics on the Web. ACM SIGACT News, 29 (2):33–41, June 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1999:EFa

[Ros99a] Ricky Ross. Education forum. ACM SIGACT News, 30(2):43–46, June 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1999:EFb

[Ros99b] Rocky Ross. Education Forum. ACM SIGACT News, 30(4):51– 52, December 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:1999:EFT

[Ros99c] Rocky Ross. Education forum: Taking the pulse: Assessing student perception of learning. ACM SIGACT News, 30(1):30–32, March 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2000:EFa

[Ros00a] Rocky Ross. Education Forum. ACM SIGACT News, 31(1):43–48, March 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2000:EFb

[Ros00b] Rocky Ross. Education forum. ACM SIGACT News, 31(2):34– 38, June 1, 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2000:GBI

[Ros00c] Rocky Ross. Going Backwards: Introductory Programming LAnguages. ACM SIGACT News, 31(4):65–73, December 2000. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2000:IEC

[Ros00d] Rocky Ross. International efforts in computer science education. ACM SIGACT News, 31(3):50–53, September 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2001:EFa

[Ros01a] Rocky Ross. Education forum. ACM SIGACT News, 32(1):56–59, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2001:EFb

[Ros01b] Rocky Ross. Education forum. ACM SIGACT News, 32(4):59– 65, December 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2001:MTL

[Ros01c] Rocky Ross. Molasses and theories of learning. ACM SIGACT News, 32(2):64–71, June 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2002:APT

[Ros02a] Rocky Ross. Accountability and the public trust. ACM SIGACT News, 33(1):61–66, March 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2002:SWO

[Ros02b] Rocky Ross. Space wars and other things. ACM SIGACT News, 33(3):71–74, September 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2003:SNO

[Ros03a] Ricky Ross. SIGACT news online algorithms column 1. ACM SIGACT News, 34(4):68–77, December 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2003:DP

[Ros03b] Rocky Ross. Deconstructing Paxos. ACM SIGACT News, 34(1): 47–67, March 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2003:EFa

[Ros03c] Rocky Ross. Education forum. ACM SIGACT News, 34(3):62– 67, September 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2003:EFb

[Ros03d] Rocky Ross. Education forum. ACM SIGACT News, 34(4):78– 83, December 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2003:TRS

[Ros03e] Rocky Ross. Teaching (and research) squeezed. ACM SIGACT News, 34(1):68–77, March 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2004:MM

[Ros04a] Rockey Ross. Mental models. ACM SIGACT News, 35(2):79–82, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2004:CB

[Ros04b] Rocky Ross. Chicago blues. ACM SIGACT News, 35(3):69–71, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2004:PSF

[Ros04c] Rocky Ross. A princess swimming in the fog looking for a monster cow. ACM SIGACT News, 35(2):74–78, June 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2004:SNOa

[Ros04d] Rocky Ross. SIGACT news online algorithms column 2. ACM SIGACT News, 35(1):38–48, March 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2004:SNOb

[Ros04e] Rocky Ross. SIGACT news online algorithms column 4. ACM SIGACT News, 35(3):58–66, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2004:ST

[Ros04f] Rocky Ross. Stability and theory. ACM SIGACT News, 35(1): 49–51, March 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2005:EF

[Ros05a] Rocky Ross. Education forum. ACM SIGACT News, 36(4):70– 71, December 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2005:EFT

[Ros05b] Rocky Ross. Education forum: trying again with MathML. ACM SIGACT News, 36(3):82–84, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2005:SNO

[Ros05c] Rocky Ross. SIGACT news online algorithms column 8. ACM SIGACT News, 36(3):67–81, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2005:T

[Ros05d] Rocky Ross. Textbooks. ACM SIGACT News, 36(3):84–93, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ross:2023:RDM

[Ros23] Peter Ross. Review of 1 The Discrete Mathematical Charms of Paul Erdős: A Simple Introduction by Vasek Chvatal. ACM SIGACT News, 54(1):15–17, March 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://

dl.acm.org/doi/10.1145/3586165.3586169.

Rothe:2005:CTC

[Rot05] Jörg Rothe. Complexity Theory and Cryptology: an Introduction to Cryptocomplexity. Texts in theoretical computer science: an EATCS series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-22147-6, 3-540-28520-2 (e-book). xi + 478 pp. LCCN QA76.9.A25 R672

2005. URL http://www.myilibrary.com?id=133022;   http:
//www.springerlink.com/openurl.asp?genre=book\%26isbn=
978-3-540-22147-0.

Rothe:2007:BRB

[Rot07] Jörg Rothe. Book review: Complexity and Cryptography: An Introduction, by John Talbot and Dominic Welsh, Cambridge University Press, 2006, 292 pages. ACM SIGACT News, 38(2):16–20, June 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [TW06].

Roughgarden:2005:IVT

[Rou05] Tim Roughgarden. An interview with Vladimir Trifonov 2005 Danny Lewin Best Student Paper Award winner. ACM SIGACT News, 36(4):111–114, December 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rota:1997:IT

[RP97] Gian-Carlo Rota and Fabrizio Palombi, editors. Indiscrete Thoughts. Birkhäuser Verlag, Basel, Switzerland, 1997. ISBN 0-8176-3866-0 , 3-7643-3866-0. xxii + 280 pp. LCCN QA7 .R65 1997.

Rajsbaum:2020:YMC

[RR20] Sergio Rajsbaum and Michel Raynal. 60 years of mastering concurrent computing through sequential thinking. ACM SIGACT News, 51(2):59–88, June 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3406678.3406690.

Reingold:1988:MNO

[RS88] Edward M. Reingold and Xiao-Jun Shen. More nearly optimal algorithms for unbounded searching. Report UIUCDCS-R-88-1471, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, 1988. 56 pp.

Reingold:1991:MNOb

[RS91a] Edward M. Reingold and Xiao Jun Shen. More nearly optimal algorithms for unbounded searching. II. the transfinite case. SIAM Journal on Computing, 20(1):184–208, February 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reingold:1991:MNOa

[RS91b] Edward M. Reingold and Xiao Jun Shen. More nearly optimal algorithms for unbounded searching, Part I: The finite case. SIAM Journal on Computing, 20(1):156–183, February 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rossman:2015:CTC

[RST15] Benjamin Rossman, Rocco A. Servedio, and Li-Yang Tan. Complexity theory column 89: The polynomial hierarchy, random oracles, and Boolean circuits. ACM SIGACT News, 46(4):50–68, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Romashchenko:2021:OPK

[RSZ21] Andrei Romashchenko, Alexander Shen, and Marius Zimand. 27 open problems in Kolmogorov complexity. ACM SIGACT News, 52(4):31–54, December 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3510382.3510389.

Roberts:2005:AC

[RT05] Fred S. Roberts and Barry Tesman. Applied Combinatorics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2005. ISBN 0-13-079603-4. xxii + 824 pp. LCCN QA164 .R6 2005.

Rubin:2013:RBR

[Rub13] Paul Rubin. Review of Information Retrieval by Buettcher, Clarke, Cormack. ACM SIGACT News, 44(1):29–33, March 2013. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rupert:1991:CRL

[Rup91] C. P. Rupert. Commutative regular languages. ACM SIGACT News, 22(4):48–49, Fall 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Robert:2010:IS

[RV10] Yves Robert and Frédéric Vivien, editors. Introduction to Scheduling. Chapman and Hall/CRC computational science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4200-7273-0 (hardcover). xx + 313 pp. LCCN QA76.9.D5 I673 2010.

Regan:1994:QIC

[RW94] Kenneth W. Regan and Jie Wang. The quasilinear isomorphism challenge. ACM SIGACT News, 25(3):106–113, September 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rubin:1999:DBC

[RW99] Harvey Rubin and David Harlan Wood, editors. DNA Based Computers III: DIMACS Workshop, June 23–25, 1997, volume 48 of DIMACS series in discrete mathematics and theoretical computer science. American Mathematical Society, Providence, RI, USA, 1999. ISBN 0-8218-0842-7. LCCN QA76.887 .D523 1999.

Rubinstein:2019:SVA

[RW19] Aviad Rubinstein and Virginia Vassilevska Williams. SETH vs approximation. ACM SIGACT News, 50(4):57–76, December 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3374857.
3374870.

Reed:2009:DSA

[RZ09] David M. Reed and John M. Zelle. Data Structures and Algorithms using Python and C++. Franklin, Beedle and Associates, Wilsonville, OR, USA, 2009. ISBN 1-59028-233-7. xiii + 568 pp. LCCN QA76.73.P98 R44 2009.

Saari:2001:DEE

[Saa01] D. (Donald) Saari. Decisions and Elections: Explaining the Unexpected. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-80816-2, 0-521-00404-7 (paperback). xiii + 240 pp. LCCN HB846.8 .S25 2001.

Sakarovitch:2011:EAT

[Sak11] Jacques Sakarovitch. Elements of Automata Theory. American Mathematical Society, Providence, RI, USA, 2011. ISBN 0-521-84425-8. LCCN ???? URL http://purl.utwente.nl/ publications/78134. Translation by Reuben Thomas from the French original.

Salomon:2003:DPS

[Sal03] David Salomon. Data Privacy and Security. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 0-387-00311-8. xiv + 465 pp. LCCN QA76.9.A25

S265 2003. US$59.95. URL http://www.booksbydavidsalomon.
com/.

Salomon:2005:CDC

[Sal05a] David Salomon. Coding for Data and Computer Communications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-21245-0. xv + 548 pp. LCCN

TK5102.94 .S35 2005. URL http://www.DavidSalomon.name/
Codes/Codes.html;   http://www.ecs.csun.edu/~dsalomon/.

Salomon:2005:FCS

[Sal05b] David Salomon. Foundations of Computer Security. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-84628-193-8, 1-84628-341-8. xxi + 368 pp. LCCN QA76.9.A25 S2656 2005. URL http://

www.springer.com/sgw/cda/frontpage/0,11855,4-40007-22-
65173048-0,00.html;   http://www.springer.com/sgw/cda/
pageitems/document/cda_downloaddocument/0,11855,0-0-45-
166687-p65173048,00.pdf;   http://www.springer.com/sgw/
cda/pageitems/document/cda_downloaddocument/0,11855,0-
0-45-166688-p65173048,00.pdf;   http://www.springer.com/
sgw/cda/pageitems/document/cda_downloaddocument/0,11855,
0-0-45-166689-p65173048,00.pdf.

Salomon:2007:VLC

[Sal07] David Salomon. Variable-length Codes for Data Compression. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-84628-958-0 (paperback), 1-84628-959-9 (ebook). xii + 191 pp. LCCN QA76.9.D33 S25 2007.

Santoro:1984:SDT

[San84] Nicola Santoro. Sense of direction, topological awareness and communication complexity. ACM SIGACT News, 16(2):50–56, Summer 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Santoro:1985:MP

[San85] Nicola Santoro. The un-merging problem. ACM SIGACT News, 17(1):5–6, June 1985. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sanford:2010:BRB

[San10] Michael Sanford. Book review: Computer Viruses and Malware, by John Aycock. ACM SIGACT News, 41(1):44–47, March 2010. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Ayc06].

Saridis:2003:BRB

[Sar03] Prodromos Saridis. Book review: Number Theory for Computing. ACM SIGACT News, 34(1):4–5, March 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Yan02].

Savage:1974:ACL

[Sav74] John E. Savage. An algorithm for the computation of linear forms. SIAM Journal on Computing, 3(2):150–158, ???? 1974. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Savage:1998:NAF

[Sav98] John E. Savage. A new approach to the first theory course. ACM SIGACT News, 29(4):58–62, December 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sharma:2014:TMM

[SB14] Gokarna Sharma and Costas Busch. Transactional memory: models and algorithms. ACM SIGACT News, 45(2):74–103, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Schneider:1975:ATW

[Sch75] Richard E. Schneider. Automata theory: what is it used for? ACM SIGACT News, 7(4):8–11, November 1975. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Schoning:1981:NCS

[Sch81] Uwe Schöning. A note on complete sets for the polynomial-time hierarchy. ACM SIGACT News, 13(1):30–34, Winter 1981. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Schoning:1982:NDS

[Sch82] Uwe Schöning. On NP-decomposable sets. ACM SIGACT News, 14(1):18–20, Winter 1982. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Schoning:1989:LCS

[Sch89] Uwe Schöning. Logic for Computer Scientists, volume 8 of Progress in computer science and applied logic. Birkhäuser Verlag, Basel, Switzerland, 1989. ISBN 0-8176-3453-3. 166 pp. LCCN QA9 .S363

Schorer:1997:SPS

[Sch97] Peter Schorer. Is a solution to the 3x + 1 problem in sight? ACM SIGACT News, 28(4):90–93, December 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Schneider:2000:CRT

[Sch00] S. A. (Steve A.) Schneider. Concurrent and Real Time Systems: the CSP Approach. Worldwide series in computer science. John Wiley, New York, NY, USA, 2000. ISBN 0-471-62373-3. xv + 510 pp. LCCN QA76.58 .S346 2000. URL http://www.loc.

gov/catdir/description/wiley032/99041795.html;   http:
//www.loc.gov/catdir/toc/onix03/99041795.html.

Schroder:2003:OSI

[Sch03] Bernd S. W. (Bernd Siegfried Walter) Schröder. Ordered Sets: an Introduction. Birkhäuser Verlag, Basel, Switzerland, 2003. ISBN 0-8176-4128-9, 3-7643-4128-9. xvii + 391 pp. LCCN QA171.48

.S47 2003. URL http://www.loc.gov/catdir/enhancements/
fy0812/2002018231-d.html;   http://www.loc.gov/catdir/
enhancements/fy0812/2002018231-t.html.

Schneider:2004:VRS

[Sch04] Klaus Schneider. Verification of Reactive Systems: Formal Methods and Algorithms. Texts in theoretical computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-00296-0. xiv + 600 pp. LCCN QA76.76.V47 S34 2004. URL http:

//www.loc.gov/catdir/enhancements/fy0817/2003066407-b.
html;   http://www.loc.gov/catdir/enhancements/fy0817/
2003066407-d.html;   http://www.loc.gov/catdir/enhancements/
fy0817/2003066407-t.html.

Schmeister:2006:BRB

[Sch06] William Schmeister. Book review: Algorithms: Sequential, Parallel, and Distributed, by Kenneth A. Berman and Jerome L. Paul, Thomson Course Technology, 2005. ACM SIGACT News, 37(2): 17–22, June 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BP05].

Schoning:2008:LCS

[Sch08a] Uwe Schöning. Logic for Computer Scientists. Modern Birkhäuser classics. Birkhäuser Verlag, Basel, Switzerland, 2008. ISBN 0-8176-4762-7, 0-8176-4763-5. ???? pp. LCCN QA9 .S363 2008.

Schroder:2008:BRB

[Sch08b] Bernd S. W. Schröder. Book review: Ordered Sets: An Introduction. ACM SIGACT News, 39(3):18–21, September 2008. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sch03].

Schae:2009:BRB

[Sch09] Elisa Schae. Book review: A Course on the Web Graph, by Anthony Bonato (American Mathematical Society, Providence, Rhode Island, USA). ACM SIGACT News, 40(2):35–37, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Bon08b].

Schwarzman:2016:PR

[Sch16] Gregory Schwarzman. PODC 2016 review. ACM SIGACT News, 47(4):35–38, December 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Schneider:2018:HCD

[Sch18] Fred Schneider. History and context for defining liveness: Winner 2018 Edsger W. Dijkstra Prize. ACM SIGACT News, 49(4):60– 63, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Scott:2015:TMT

[Sco15] Michael Scott. Transactional memory today. ACM SIGACT News, 46(2):96–104, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Simovici:2008:MTD

[SD08] Dan A. Simovici and Chabane Djeraba. Mathematical Tools for Data Mining: Set Theory, Partial Orders, Combinatorics. Advanced information and knowledge processing, 1610-3947. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-84800-200-9, 1-84800-201-7. xii + 616 pp. LCCN QA76.9.D343 S595 2008.

Seiferas:1974:NPR

[Sei74] Joel I. Seiferas. A note on prefixes of regular languages. ACM SIGACT News, 6(1):25–29, January 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiden:1996:TCS

[Sei96] Steve Seiden. Theoretical computer science cheat sheet. ACM SIGACT News, 27(4):52–61, December 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:1997:RCRa

[Sei97a] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 28(2):32–33, June 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:1997:RCRb

[Sei97b] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 28(3):30, September 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:1997:RCRc

[Sei97c] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 28(4):28–30, December 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:1998:RCRa

[Sei98a] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 29(1):13–14, March 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:1998:RCRb

[Sei98b] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 29(2):27–28, June 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:1998:RCRc

[Sei98c] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 29(3):53–54, September 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:1998:RCRd

[Sei98d] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 29(4):15–16, December 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:1999:RCRa

[Sei99a] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 30(1):12–13, March 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:1999:RCRb

[Sei99b] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 30(2):17–18, June 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:2000:RCRa

[Sei00a] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 31(2):12–13, June 1, 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:2000:RCRb

[Sei00b] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 31(3):27–28, September 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:2000:RCRc

[Sei00c] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 31(4):27, December 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiden:2001:CCP

[Sei01a] Steve Seiden. Can a computer proof be elegant? ACM SIGACT News, 32(1):111–114, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:2001:RCRa

[Sei01b] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 32(1):17–18, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:2001:RCRb

[Sei01c] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 32(2):20–21, June 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:2002:RCRa

[Sei02a] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 33(1):21–22, March 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:2002:RCRb

[Sei02b] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 33(2):21–23, June 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Seiferas:2002:RCRc

[Sei02c] Joel Seiferas. Reprints from Computing Reviews. ACM SIGACT News, 33(3):23–24, September 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sen:2012:RB

[Sen12] Siddhartha Sen. Review of PODC 2012. ACM SIGACT News, 43 (4):104–111, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sesiano:2009:IHA

[Ses09] Jacques Sesiano. An introduction to the history of algebra: solving equations from Mesopotamian times to the Renaissance, volume 27 of Mathematical world. American Mathematical Society, Providence, RI, USA, 2009. ISBN 0-8218-4473-3. viii + 174 pp. LCCN QA151 .S4513 2009. Translated by Anna Pierrehumbert.

Salenieks:1989:HAI

[SF89] Peter Salenieks and Michael G. Farringdon. How To Answer It. ACM SIGACT News, 20(3):60–63, July 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sgarbas:2009:BRB

[Sga09] Kyriakos N. Kyriakos N. Sgarbas. Book review: How to Think about Algorithms, by Je Edmonds (Cambridge University Press, 2008). ACM SIGACT News, 40(2):13–17, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Edm08].

Sgarbas:2010:BRB

[Sga10] Kyriakos N. Sgarbas. Book review: Automata Theory with Modern Applications, by James A. Anderson, Cambridge University Press, 2006, viii + 256 pages. ACM SIGACT News, 41(2):43–46, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [AH06].

Sgarbas:2012:RBT

[Sga12] Kyriakos N. Sgarbas. Review of Semantic Techniques in Quantum Computation, edited by Simon Gay and Ian Mackie. ACM SIGACT News, 43(4):44–48, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sgarbas:2013:RBC

[Sga13] Kyriakos N. Sgarbas. Review of Quantum Computing: a Gentle Introduction by Eleanor Rieffel and Wolfgang Polak. ACM SIGACT News, 44(2):31–35, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shallit:1992:RAP

[Sha92] J. Shallit. Randomized algorithms in “primitive cultures”. ACM SIGACT News, 23(4):77–80, Fall 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). Shallit, in a slightly tongue-in-cheek manner, traces back some of the concepts of randomized algorithms to the native American society of the Naskapi and the central African society of the Azande. Roots in the works of Pierre Laplace and Lord Kelvin are also pointed out.

Shankar:1994:MMG

[Sha94] N. Shankar. Metamathematics, Machines, and Gödel’s Proof, volume 38 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-42027-X (hardcover). xv + 202 pp. LCCN QA9.65 .S53 1994. URL http:

//www.loc.gov/catdir/description/cam032/94222366.html;
 http://www.loc.gov/catdir/samples/cam034/94222366.html;
 http://www.loc.gov/catdir/toc/cam032/94222366.html.

Shallit:2002:CCL

[Sha02] Jeffrey Shallit. The computational complexity of the local postage stamp problem. ACM SIGACT News, 33(1):90–94, March 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shackell:2004:SA

[Sha04] John R. Shackell. Symbolic Asymptotics, volume 12 of Algorithms and computation in mathematics, 1431-1550. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-21097-0 (hardcover). xi + 243 pp. LCCN QA297

.S435 2004. URL http://www.loc.gov/catdir/enhancements/
fy0829/2004045315-d.html;   http://www.loc.gov/catdir/
enhancements/fy0829/2004045315-t.html.

Shallit:2009:SCF

[Sha09] Jeffrey Outlaw Shallit. A Second Course in Formal Languages and Automata Theory. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-86572-7 (hardcover). xi + 240 pp. LCCN

QA267.3 .S53 2009. URL http://www.loc.gov/catdir/toc/
ecip0822/2008030065.html.

Shallit:2010:BRB

[Sha10a] Jeffrey Shallit. Book review: Cryptographic Applications of Analytic Number Theory: Lower Bounds and Pseudorandomness, by Igor Shparlinski, Birkäuser, 2003. ACM SIGACT News, 41(3):44– 45, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Shp03].

Shaltiel:2010:TCD

[Sha10b] Ronen Shaltiel. Typically-correct derandomization. ACM SIGACT News, 41(2):57–72, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shallit:2012:BRB

[Sha12a] Jeffrey Shallit. Book review: Finite Fields and Applications, by Gary L. Mullen and Carl Mummert. ACM SIGACT News, 43(1): 30–31, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [MM07].

Shallit:2012:RBN

[Sha12b] Jeffrey Shallit. Review of Biscuits of Number Theory by Arthur T. Benjamin and Ezra Brown. ACM SIGACT News, 43(2):21–24, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shallit:2013:RCC

[Sha13] Jeffrey Shallit. Review of Computability and complexity theory by Steven Homer and Alan L. Selman. ACM SIGACT News, 44 (4):41–42, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sharir:2017:CGC

[Sha17] Micha Sharir. Computational geometry column 65. ACM SIGACT News, 48(2):68–85, June 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shafei:2019:RKC

[Sha19] Hadi Shafei. Review of Kolmogorov Complexity and Algorithmic Randomness by A. Shen, V. A. Uspensky, and N. Vereshchagin. ACM SIGACT News, 50(4):9–13, December 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3374857.3374861.

Sherman:1991:SSFa

[She91a] Alan T. Sherman. On Superpolylogarithmic Subexponential Functions (Part I). ACM SIGACT News, 22(1):65, Winter 1991. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sherman:1991:SSFb

[She91b] Man T. Sherman. On Superpolylogarithmic Subexponential Functions (Part II). ACM SIGACT News, 22(2):51–56, Spring 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shen:1997:CEC

[She97] Zhizhang Shen. Correcting an error in a ’cheat sheet’. ACM SIGACT News, 28(1):49, March 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shen:2008:APP

[She08] A. (Alexander) Shen. Algorithms and Programming: Problems and Solutions. Modern Birkhäuser classics. Birkhäuser Verlag, Basel, Switzerland, 2008. ISBN 0-8176-4761-9, 0-8176-4760-0. viii + 217 pp. LCCN QA76.9.A43.

Shen:2010:APP

[She10] A. (Alexander) Shen. Algorithms and Programming: Problems and Solutions. Springer undergraduate texts in mathematics and technology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2010. ISBN 1-4419-1748-9, 1-4419-1747-0. ix + 272 pp. LCCN QA76.9.A43 S47 2010.

Shehab:2013:RBD

[She13] Omar Shehab. Review of The Dots and Boxes Game: Sophisticated Child’s Play by Elwyn Berlekamp. ACM SIGACT News, 44(1): 42–45, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shehab:2014:RBP

[She14] Omar Shehab. Review of In pursuit of the unknown: 17 equations that changed the world by Ian Stewart. ACM SIGACT News, 45(1): 11–15, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shehab:2015:RWN

[She15] Omar Shehab. Review of: A Wealth of Numbers: An Anthology of 500 Years of Popular Mathematics Writing by Benjamin Wardhaugh. ACM SIGACT News, 46(1):15–19, March 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shimi:2019:SMR

[Shi19] Adam Shimi. The splendors and miseries of rounds. ACM SIGACT News, 50(3):35–50, September 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shoup:2005:CIN

[Sho05] Victor Shoup. A Computational Introduction to Number Theory and Algebra. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-85154-8 (hardcover), 0-521-61725-1 (paperback). xvi + 517 pp. LCCN QA241 .S465 2005. URL http:

//www.loc.gov/catdir/enhancements/fy0664/2006530670-d.
html;   http://www.loc.gov/catdir/enhancements/fy0664/
2006530670-t.html;   http://www.loc.gov/catdir/enhancements/
fy0733/2006530670-b.html.

Shoup:2009:CIN

[Sho09] Victor Shoup. A Computational Introduction to Number Theory and Algebra. Cambridge University Press, Cambridge, UK, second edition, 2009. ISBN 0-521-51644-7 (hardcover). xvii + 580 pp. LCCN QA241 .S465 2009.

Shparlinski:2003:CAA

[Shp03] Igor E. Shparlinski. Cryptographic Applications of Analytic Number Theory: Complexity Lower Bounds and Pseudorandomness, volume 22 of Progress in computer science and applied logic. Birkhäuser Verlag, Basel, Switzerland, 2003. ISBN 3-7643-6654-0 , 0-8176-6654-0. viii + 411 pp. LCCN QA267.7 .S55 2003.

Shyamasundar:1978:NMM

[Shy78a] R. K. Shyamasundar. A note on the multiplication of 4 × 4 matrices. ACM SIGACT News, 10(1):31–32, Spring 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shyamasundar:1978:NTC

[Shy78b] R. K. Shyamasundar. A note on the transitive closure of a Boolean matrix. ACM SIGACT News, 9(4):18–21, Winter 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See notes [Yel78].

Staff:1980:TR

[SIG80] SIGACT News Staff. Technical reports. ACM SIGACT News, 12 (1):2–135, Winter–Spring 1980. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1981:TRa

[SIG81a] SIGACT News Staff. Technical reports. ACM SIGACT News, 13(1):39–87, Winter 1981. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1981:TRb

[SIG81b] SIGACT News Staff. Technical reports. ACM SIGACT News, 13 (2):16–65, Spring–Summer 1981. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1982:RTR

[SIG82a] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 14(1):44–65, Winter 1982. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1982:RAA

[SIG82b] SIGACT News Staff. Research advance abstracts. ACM SIGACT News, 14(4):72–83, Fall–Winter 1982. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1982:TR

[SIG82c] SIGACT News Staff. Technical reports. ACM SIGACT News, 14(2):20–41, Spring 1982. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1983:A

[SIG83a] SIGACT News Staff. Abstracts. ACM SIGACT News, 15(2):63– 74, Spring–Summer 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1983:AAT

[SIG83b] SIGACT News Staff. Abstracts and available technical reports. ACM SIGACT News, 15(3):72–95, Summer–Fall 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1983:ACR

[SIG83c] SIGACT News Staff. Abstracts computing research laboratory. ACM SIGACT News, 15(1):93–104, Winter–Spring 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1984:Aa

[SIG84a] SIGACT News Staff. Abstracts. ACM SIGACT News, 15(4): 48–56, Winter–Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1984:Ab

[SIG84b] SIGACT News Staff. Abstracts. ACM SIGACT News, 16(1): 48–56, Winter–Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1984:Ac

[SIG84c] SIGACT News Staff. Abstracts. ACM SIGACT News, 16(2):57– 62, Summer 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Silvestri:2012:RBT

[Sil12] Francesco Silvestri. Review of Graph Theory and Interconnection Networks, by Lih-Hsing Hsu and Cheng-Kuan Lin. ACM SIGACT News, 43(4):30–34, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Simon:1976:TRP

[Sim76] Istvan Simon. Two results on polynomial-time reducibilities. ACM SIGACT News, 8(1):33–37, January 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Simons:1984:RCSa

[Sim84a] Babra Simons. The reentry Computer Science Program at U.C. Berkeley. ACM SIGACT News, 15(4):36, Winter–Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Simons:1984:RCSb

[Sim84b] Babra Simons. The reentry Computer Science Program at U.C. Berkeley. ACM SIGACT News, 16(1):36, Winter–Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Simons:1985:STR

[Sim85] Barbara Simons. Some thoughts on running a conference. ACM SIGACT News, 16(4):48–52, April 1985. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Simmons:2000:DCT

[Sim00] Harold Simmons. Derivation and Computation: Taking the Curry– Howard Correspondence Seriously, volume 51 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-77173-0 (hardcover). xxv + 384 pp. LCCN QA9.54 .S55 2000.

Simovici:2004:BRB

[Sim04] Dan A. Simovici. Book review: The Classical Decision Problem, by Egon Börger, Erich Grädel and Yuri Gurevich. Springer-Verlag 1997. ACM SIGACT News, 35(1):4–7, March 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BGG97].

Simon:2022:RJ

[Sim22] Janos Simon. Remembering Juris. ACM SIGACT News, 53(4):41– 47, December 2022. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3577971.3577978.

Sintzoff:1972:CPP

[Sin72] Michel Sintzoff. Calculating properties of programs by valuations on specific models. ACM SIGACT News, 4(14):203–207, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sinha:1998:TRC

[Sin98] Rakesh K. Sinha. Technical report column. ACM SIGACT News, 29(2):30–32, June 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Singh:1999:CBE

[Sin99] Simon Singh. The Code Book: the Evolution of Secrecy from Mary, Queen of Scots, to Quantum Cryptography. Doubleday, New York, NY, USA, 1999. ISBN 0-385-49531-5. xiii + 402 pp. LCCN Z103 .S56 1999. US$24.95.

Sipser:1996:ITC

[Sip96] Michael Sipser. Introduction to the Theory of Computation. ACM SIGACT News, 27(1):27–29, March 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sistla:1993:BRB

[Sis93] A. P. Sistla. Book review: The Temporal Logic of Reactive and Concurrent Systems — Specification. By Zohar Manna and Amir Pnueli. (Springer-Verlag, 1991. xiv + 427pp. ISBN 0-387-97664-7. $49.95). ACM SIGACT News, 24(4):34–36, December 1993. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [MP92].

Skiena:1998:ADM

[Ski98] Steven S. Skiena. The Algorithm Design Manual. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-94860-0. xvi + 486 pp. LCCN QA76.9.A43S55 1997. US$54.95.

Skiena:1999:WIA

[Ski99] Steven S. Skiena. Who is interested in algorithms and why?: lessons from the Stony Brook algorithms repository. ACM SIGACT News, 30(3):65–74, September 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Skiena:2001:CBC

[Ski01] Steven S. Skiena. Calculated Bets: Computers, Gambling, and Mathematical Modeling to Win. Outlooks. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-80426-4, 0-521-00962-6 (paperback). xv + 232 pp. LCCN QA401 .S474

2001. URL http://www.loc.gov/catdir/description/cam021/
2001025483.html;   http://www.loc.gov/catdir/samples/cam031/
2001025483.html;   http://www.loc.gov/catdir/toc/cam024/
2001025483.html.

Skiena:2008:ADM

[Ski08] Steven S. Skiena. The Algorithm Design Manual. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2008. ISBN 1-84800-069-3 (hardcover), 1-84800-070-7 (ebook). xvi + 730 pp. LCCN QA76.9.A43 S55 2008.

Shoham:2009:MSA

[SLB09] Yoav Shoham and Kevin Leyton-Brown. Multiagent Systems: Algorithmic, Game-Theoretic, and Logical Foundations. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-89943-5 (hardcover). xx + 483 pp. LCCN QA76.76.I58 S75 2009.

Sloan:2002:NRTa

[Slo02a] Robert Sloan. NSF report: theory of computing program. ACM SIGACT News, 33(1):28, March 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sloan:2002:NRTb

[Slo02b] Robert Sloan. NSF report: theory of computing program. ACM SIGACT News, 33(2):27–28, June 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sloan:2007:JBN

[Slo07] Robert H. Sloan. The joys of being an NSF program director. ACM SIGACT News, 38(1):7–8, March 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Smith:1999:IR

[Smi99] Carl Smith. On the importance of refereeing. ACM SIGACT News, 30(3):75–76, September 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Smith:2000:BRB

[Smi00] Carl Smith. Book review: Systems that Learn (second edition) by Jain, Osherson, Royer, Sharma. ACM SIGACT News, 31(1): 12–13, March 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [JORS99].

Smith:2009:BRB

[Smi09] Brent Smith. Book review: How to Prove It: a Structured Approach, by Daniel J. Velleman (Cambridge University Press, 2006). ACM SIGACT News, 40(1):18–20, March 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Vel94].

Smolka:2021:KSC

[Smo21] Scott A. Smolka. Kanellakis–Smolka 1983: a convolution of circumstances. ACM SIGACT News, 52(4):78–79, December 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3510382.
3510395.

Snyder:1991:PTG

[Sny91] Wayne Snyder. A Proof Theory for General Unification, volume 11 of Progress in computer science and applied logic. Birkhäuser Verlag, Basel, Switzerland, 1991. ISBN 0-8176-3593-9 (hardcover), 3-7643-3593-9 (hardcover). vi + 175 pp. LCCN QA76.9.A96 S6 1991. US$28.00.

Sofo:2003:CTS

[Sof03] Anthony Sofo. Computational Techniques for the Summation of Series. Kluwer Academic/Plenum Publishers, New York, NY, USA, 2003. ISBN 0-306-47805-6. xv + 189 pp. LCCN QA295

.S6412 2003. URL http://www.loc.gov/catdir/enhancements/
fy0820/2003054692-d.html.

Soifer:2009:MCB

[Soi09] Alexander Soifer. The Mathematical Coloring Book: Mathematics of Coloring and the Colorful Life of its Creators. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-74640-4 (hardcover). xxx + 607 pp. LCCN QA166 .S65 2009.

Soltys:2012:RBP

[Sol12] Karolina Soltys. Review of Permutation Patterns, edited by Steve Linton, Nik Ruškuc, Vincent Vatter. ACM SIGACT News, 43 (3):54–55, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sottile:2010:BRB

[Sot10] Matthew J. Sottile. Book review: Curve and Surface Reconstruction: Algorithms with Mathematical Analysis, by Tamal K. Dey, Cambridge University Press. ACM SIGACT News, 41(1):24–27, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Dey07].

Schoning:1998:GTC

[SP98] Uwe Schöning and Randall J. Pruim. Gems of Theoretical Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-64425-3 (hardcover). x + 320 pp. LCCN QA76.9.A96 S42713 1998.

Sivilotti:2007:CKL

[SP07] Paolo A. G. Sivilotti and Scott M. Pike. A collection of kinesthetic learning activities for a course on distributed computing: ACM SIGACT news distributed computing column 26. ACM SIGACT News, 38(2):56–74, June 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Spencer:1993:CFL

[Spe93] Thomas H. Spencer. Context-free languages. ACM SIGACT News, 24(4):90–91, December 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Springer:2003:BRB

[Spr03] William M. Springer II. Book review: Cryptography: Theory and Practice, second edition by Douglas R. Stinson. CRC Press. ACM SIGACT News, 34(4):22–25, December 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sti95, Sti02, Sti06].

Springer:2006:BRB

[Spr06] William M. Springer II. Book review: Introduction to Coding Theory, by Juergen Bierbrauer, Chapman and Hall/CRC, 2005, ISBN 1-58488-421-5. ACM SIGACT News, 37(1):16–18, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Bie05].

Springer:2008:BRB

[Spr08] William M. Springer II. Book reviews: Introducing Game Theory and Its Applications, by Elliott Mendelson, CRC Press, and Game Theory and Strategy, by Philip D. Straffin, MAA Press. ACM SIGACT News, 39(1):24–27, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Str93, Men04].

Springer:2009:BRB

[Spr09] William M. Springer II. Book review: The Traveling Salesman Problem: a Computational Study, by Applegate, Bixby, Chvátal, and Cook (Princeton University Press). ACM SIGACT News, 40 (2):30–32, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [ABCC06].

Skiena:2003:PCP

[SR03] Steven S. Skiena and Miguel A. Revilla. Programming challenges: the programming contest training manual. ACM SIGACT News, 34(3):68–74, September 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Srinivasan:2006:BRB

[Sri06] Aravind Srinivasan. Book review: The Random Projection Method, by Santosh Vempala. ACM SIGACT News, 37(4):41–43, December 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Vem04].

Srinivasan:2010:BRB

[Sri10] Aravind Srinivasan. Book review: Concentration of Measure for the Analysis of Randomized Algorithms, by Devdatt P. Dubhashi and Alessandro Panconesi, Cambridge University Press, 2009. ACM SIGACT News, 41(1):28–30, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [DP09].

Srinivasan:2014:RVI

[Sri14] Aravind Srinivasan. Review of Visions of Infinity: The Great Mathematical Problems by Ian Stewart. ACM SIGACT News, 45 (4):41–45, December 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Srinivasan:2018:PC

[Sri18] Aravind Srinivasan. Probability and computing. ACM SIGACT News, 49(3):20–22, September 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Scholkopf:2002:LKS

[SS02] Bernhard Schölkopf and Alexander J. Smola. Learning with Kernels: Support Vector Machines, Regularization, Optimization, and Beyond. Adaptive computation and machine learning. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-19475-9. xviii + 626 pp. LCCN Q325.5 .S32 2002.

Shilova:2004:AET

[SS04a] S. O. Shilova and Nikolay V. Shilov. Addendum to Etude on theme of Dijkstra. ACM SIGACT News, 35(4):118, December 2004. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shilova:2004:ETD

[SS04b] S. O. Shilova and Nikolay V. Shilov. Etude on theme of Dijkstra. ACM SIGACT News, 35(3):102–108, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Swamy:2006:ACA

[SS06] Chaitanya Swamy and David B. Shmoys. Algorithms column: Approximation algorithms for 2-stage stochastic optimization problems. ACM SIGACT News, 37(1):33–46, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sally:2007:RRV

[SS07] Judith D. Sally and Paul Sally. Roots to Research: a Vertical Development of Mathematical Problems. American Mathematical Society, Providence, RI, USA, 2007. ISBN 0-8218-4403-2. xiii + 338 pp. LCCN QA11.2 .S25 2007. URL http://www.loc.gov/

catdir/toc/fy0804/2007060836.html.

Scheid:2014:IMM

[SSC14] Paul T. Scheid, Ari J. Spilo, and Ron K. Cytron. Inferring memory map instructions. ACM SIGACT News, 45(3):47–52, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1973:RTRa

[Sta73a] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 5(1):21–24, January 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1973:RTRb

[Sta73b] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 5(2):26–27, April 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1973:RTRc

[Sta73c] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 5(3):31–32, July 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1974:RTRb

[Sta74a] SIGACT News Staff. A recent technical report. ACM SIGACT News, 6(3):5, July 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1974:RTRa

[Sta74b] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 6(2):57–60, April 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1974:RTRc

[Sta74c] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 6(3):33–34, July 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1974:TR

[Sta74d] SIGACT News Staff. Technical reports. ACM SIGACT News, 6 (1):35, January 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1975:RTR

[Sta75a] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 7(3):35–39, Summer 1975. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1975:TR

[Sta75b] SIGACT News Staff. Technical reports. ACM SIGACT News, 7 (2):20–24, April 1975. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1976:ACS

[Sta76a] SIGACT News Staff. Abstracts from the computer science conference. ACM SIGACT News, 8(2):34–37, April/June 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1976:RTRa

[Sta76b] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 8(1):41–49, January 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1976:RTRb

[Sta76c] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 8(2):38–52, April/June 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1976:RTRc

[Sta76d] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 8(4):33–40, October 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1977:RTRa

[Sta77a] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 9(1):19–34, January 1977. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1977:RTRb

[Sta77b] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 9(2):37–44, Summer 1977. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1977:RTRc

[Sta77c] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 9(3):28–32, Fall 1977. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1978:IBM

[Sta78a] SIGACT News Staff. International Business Machines. ACM SIGACT News, 10(1):38–41, Spring 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1978:RTR

[Sta78b] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 9(4):24–28, Winter 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1979:RTRa

[Sta79a] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 11(1):44–80, Summer 1979. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staff:1979:RTRb

[Sta79b] SIGACT News Staff. Recent technical reports. ACM SIGACT News, 11(2):40–57, Fall 1979. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Stark:1984:CAa

[Sta84a] Richard Stark. Combinatory automaton. ACM SIGACT News, 15 (4):37, Winter–Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Stark:1984:CAb

[Sta84b] Richard Stark. Combinatory automaton. ACM SIGACT News, 16 (1):37, Winter–Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Staples:1987:THP

[Sta87] E. Staples. The tower of Hanoi problem with arbitrary start and end positions. ACM SIGACT News, 18(3):61–64, April 1987. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Stallman:1989:COC

[Sta89] M. Stallman. Course outline: course announcement (spring 1989) CSE/OR 691 I: Surviving Intractability. ACM SIGACT News, 20 (4):74–77, November 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Stasko:1992:AAX

[Sta92] John T. Stasko. Animating algorithms with XTANGO. ACM SIGACT News, 23(2):67–71, Spring 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Stansifer:1993:BRB

[Sta93] Ryan Stansifer. Book review: Semantics of Programming Languages: Structures and Techniques. By Carl A. Gunter. (MIT Press, 1992. xviii + 419pp. ISBN 0-262-07143-6 $37.50). ACM SIGACT News, 24(1):28–29, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Gun92].

Stamatiou:2007:BRB

[Sta07] Yannis C. Stamatiou. Book review: Probability and Computing: Randomized Algorithms and Probabilistic Analysis, by Michael Mitzenmacher and Eli Upfal, Cambridge University Press, 2005. ACM SIGACT News, 38(3):22–27, September 2007. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [MU05].

Stamatiou:2010:BRB

[Sta10] Yannis C. Stamatiou. Book review: Formal Correctness of Security Protocols, by Giampaolo Bella, Springer-Verlag. ACM SIGACT News, 41(1):47–50, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Bel07].

Stanhope:2011:BRB

[Sta11] Stephen Stanhope. Book review: Pólya Urn Models, by Hosam Mahmoud. ACM SIGACT News, 42(2):27–33, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Mah08].

Stewart:2009:PSC

[Ste09] Ian Stewart. Professor Stewart’s Cabinet of Mathematical Curiosities. Basic Books, New York, NY, USA, 2009. ISBN 0-465-01302-3 (paperback). x + 310 pp. LCCN QA95 .S726 2009.

Sterling:2011:BRBb

[Ste11a] Aaron Sterling. Book review: Handbook of Chemoinformatics Algorithms, by Faulon. ACM SIGACT News, 42(4):37–48, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [FB10].

Sterling:2011:BRBa

[Ste11b] Aaron Sterling. Book review: Handbook of Nature-Inspired and Innovative Computing, by Albert Y. Zomaya. ACM SIGACT News, 42(4):23–26, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Zom06].

Sterling:2012:RBB

[Ste12] Aaron Sterling. Review of Algorithmic Bioprocesses, edited by Condon, Harel, Kok, Salomaa, Winfree. ACM SIGACT News, 43 (4):17–24, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sterling:2013:RBM

[Ste13] Aaron Sterling. Review of The Mathematics of Life by Ian Stewart. ACM SIGACT News, 44(3):9–11, September 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Stichtenoth:1993:AFF

[Sti93] H. (Henning) Stichtenoth. Algebraic Function Fields and Codes. Universitext. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-56489-6. x + 260 pp. LCCN QA341 .S75 1993.

Stinson:1995:CTP

[Sti95] Douglas R. (Douglas Robert) Stinson. Cryptography: theory and practice. The CRC Press series on discrete mathematics and its applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1995. ISBN 0-8493-8521-0. 434 pp. LCCN QA268 .S75 1995.

Stirling:1999:DBE

[Sti99] Colin Stirling. Decidability of bisimulation equivalence for normed pushdown processes. ACM SIGACT News, 30(3):19–21, September 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Stirling:2001:MTP

[Sti01] Colin P. Stirling. Modal and Temporal Properties of Processes. Texts in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-98717-7. xii + 190 pp. LCCN QA76.9.L63

S75 2001. URL http://www.loc.gov/catdir/enhancements/
fy0818/00067924-d.html;   http://www.loc.gov/catdir/enhancements/
fy0818/00067924-t.html.

Stinson:2002:CTP

[Sti02] Douglas R. (Douglas Robert) Stinson. Cryptography: theory and practice. The CRC Press series on discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2002. ISBN 1-58488-206-9. 339 pp. LCCN QA268 .S75 2002.

Stinson:2006:CTP

[Sti06] Douglas R. (Douglas Robert) Stinson. Cryptography: theory and practice. The CRC Press series on discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, third edition, 2006. ISBN 1-58488-508-4 (hardcover). 593 pp. LCCN QA268 .S75 2006.

Stinson:2008:CDC

[Sti08] Douglas R. Stinson. Combinatorial designs: constructions and analysis. ACM SIGACT News, 39(4):17–21, December 2008. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Stichtenoth:2009:AFF

[Sti09] H. (Henning) Stichtenoth. Algebraic Function Fields and Codes, volume 254 of Graduate texts in mathematics, 0072-5285. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2009. ISBN 3-540-76877-7 (hardcover). xiii + 355 pp. LCCN QA341 .S75 2009.

Stockmeyer:1973:PCN

[Sto73a] L. Stockmeyer. Planar 3-colorability is NP-complete. ACM SIGACT News, 5(3):19–25, July 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). NCPY.

Stockmeyer:1973:PCP

[Sto73b] Larry Stockmeyer. Planar 3-colorability is polynomial complete. ACM SIGACT News, 5(3):19–25, July 1973. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Storer:2002:IDS

[Sto02] James A. (James Andrew) Storer. An Introduction to Data Structures and Algorithms. Birkhäuser Verlag, Basel, Switzer-

land, 2002.  ISBN 0-8176-4253-6 , 3-7643-4253-6.  xvii

+ 599 pp. LCCN QA76.9.D33 S73 2002. URL http:

//www.loc.gov/catdir/enhancements/fy0812/2001043503-d.
html;   http://www.loc.gov/catdir/enhancements/fy0812/
2001043503-t.html.

Straffin:1993:GTS

[Str93] Philip D. Straffin. Game Theory and Strategy, volume 36 of New mathematical library. Mathematical Association of America, Washington, DC, USA, 1993. ISBN 0-88385-637-9. x + 244 pp. LCCN QA269 .S77 1993. URL http://www.loc.gov/

catdir/description/cam025/92064176.html;   http://www.
loc.gov/catdir/toc/cam027/92064176.html.

Straubing:1994:FAF

[Str94] Howard Straubing. Finite Automata, Formal Logic, and Circuit Complexity. Progress in theoretical computer science. Birkhäuser Verlag, Basel, Switzerland, 1994. ISBN 3-7643-3719-2, 0-8176-3719-2. xii + 226 pp. LCCN QA76.9.M35 S77 1994.

Su:2009:DTC

[Su09] Jianwen Su. Database theory column: report on PODS 2009. ACM SIGACT News, 40(4):62–63, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Su:2016:DR

[Su16] Lili Su. DISC 2016 review. ACM SIGACT News, 47(4):39–43, December 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Subrahmanyam:1978:BRB

[Sub78] P. A. Subrahmanyam. Book review: Linguistic Structures Processing, by A. Zampolli. North Holland Publishing Company 1977. ACM SIGACT News, 10(2):39–41, Summer 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Zam77].

Subramanian:1995:TRA

[Sub95] Ashok Subramanian. Two recent algorithms for the global minimum cut problem. ACM SIGACT News, 26(2):78–87, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Suciu:1998:OSD

[Suc98] Dan Suciu. An overview of semistructured data. ACM SIGACT News, 29(4):28–38, December 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Suciu:2008:PD

[Suc08] Dan Suciu. Probabilistic databases. ACM SIGACT News, 39(2): 111–124, June 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sudan:2000:LDA

[Sud00] Madhu Sudan. List decoding: algorithms and applications. ACM SIGACT News, 31(1):16–27, March 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sudan:2011:GCT

[Sud11] Madhu Sudan. Guest column: Testing linear properties: Some general theme. ACM SIGACT News, 42(1):59–80, March 2011. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sugiyama:1993:DCD

[Sug93] Kozo Sugiyama. Drawing compound digraphs and its application to an idea organizer (abstract). ACM SIGACT News, 24(1):58, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Suomela:2020:URE

[Suo20] Jukka Suomela. Using round elimination to understand locality. ACM SIGACT News, 51(3):63–81, September 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3427361.3427374.

Smith:1974:MRA

[SvL74] Carl H. Smith and J. van Leeuwen. Microprogrammed random access stored program machines. ACM SIGACT News, 6(3):23–32, July 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Strong:1972:PPU

[SW72] H. R. Strong and S. A. Walker. Properties preserved under recursion removal. ACM SIGACT News, 4(14):97–103, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Stanat:1982:PTR

[SW82] Donald F. Stanat and Stephen F. Weiss. A pumping theorem for regular languages. ACM SIGACT News, 14(1):36–37, Winter 1982. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sangiorgi:2001:PCT

[SW01] Davide Sangiorgi and David Walker. The π-Calculus: a Theory of Mobile Processes. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-78177-9. xii + 580 pp. LCCN QA76.59 .S36 2001.

Swenson:1987:CPC

[Swe87] James M. Swenson. A constructive proof of the countability of Σ∗. ACM SIGACT News, 18(4):48–50, Summer 1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Simmons:1987:FFC

[SY87] B. Simmons and J. Yudken. Federal funding in computer science: a preliminary report. ACM SIGACT News, 19(1):54–63, Fall 1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Szarka:2012:BRB

[Sza12] Robert Szarka. Book review: Pearls of Discrete Mathematics, by Martin Erickson. ACM SIGACT News, 43(1):33–34, March 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Eri10].

Tan:1995:BRB

[TA95] Erkan Tan and Varol Akman. Book reviews: Logic for Applications, by Anil Nerode and Richard A. Shore. ACM SIGACT News, 26(1):20–22, March 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [NS93].

Talib:2009:BRB

[Tal09] Fatima Talib. Book review: Coding for Data and Computer Communications, by David Salomon. Springer, 2005. ACM SIGACT News, 40(4):36–41, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sal05a].

Tamassia:1993:AOD

[Tam93] Roberto Tamassia. Algorithms for orthogonal drawings (abstract). ACM SIGACT News, 24(1):59, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Tan:2016:DTC

[Tan16] Wang-Chiew Tan. Database theory column: Report on PODS 2016. ACM SIGACT News, 47(3):67–68, September 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Tao:2020:DTC

[Tao20] Yufei Tao. Database theory column. ACM SIGACT News, 51 (3):59–61, September 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3427361.3427372.

Tasic:1999:BRB

[Tas99] Vladimir Tasic. Book review: The Limits of Mathematics, by G. J. Chaitin (Springer Verlag 1998). ACM SIGACT News, 30(1):5, March 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Cha98].

Taubenfeld:2006:SAC

[Tau06] Gadi Taubenfeld. Synchronization Algorithms and Concurrent Programming. Pearson/Prentice Hall, Harlow, England, 2006. ISBN 0-13-197259-6 (paperback). xv + 423 pp. LCCN QA76.9.D5 .T38 2006.

Taylor:2002:MCA

[Tay02] R. Gregory Taylor. Modern computer algebra. ACM SIGACT News, 33(3):7–16, September 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Taylor:2004:BRB

[Tay04] R. Gregory Taylor. Book review: Boolean Functions and Computation Models, by Peter Clote and Evangelos Kranakis, Springer-Verlag, 2002. ACM SIGACT News, 35(4):5–11, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [CK02].

Tay:2005:WSC

[Tay05] Y. C. Tay. What should computer science students learn from mathematics? ACM SIGACT News, 36(2):131–143, June 2005.

CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Tate:1995:BR

[TBG95] Stephen R. Tate, David B. Benson, and Jonathan Goldstine. Book reviews. ACM SIGACT News, 26(2):26–32, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Tel:2000:IDA

[Tel00] Gerard Tel. Introduction to Distributed Algorithms. Cambridge University Press, Cambridge, UK, second edition, 2000. ISBN 0-521-79483-8. xii + 596 pp. LCCN QA76.9.D5 T44 2000. URL

http://www.loc.gov/catdir/description/cam021/00036292.
html;   http://www.loc.gov/catdir/toc/cam026/00036292.
html.

Teng:1987:CHE

[Ten87] Shang-Hua Teng. The construction of Huffman-equivalent prefix code in NC. ACM SIGACT News, 18(4):54–61, Summer 1987. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Teng:2018:KPA

[Ten18] Sheng-Hua Teng. 2018 Knuth Prize is awarded to Johan Håstad. ACM SIGACT News, 49(3):78–79, September 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Terai:2008:BRB

[Ter08] S. Terai. Book review: Cryptography in C and C++, by Michael Welschenbach, Apress, 2005. ACM SIGACT News, 39(1):12–16, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [WK01, Wel05].

Terai:2011:BRB

[Ter11] Saif Terai. Book review: Foundations of Logic and Mathematics Applications to Computer Science and Cryptography, by Yves Nievergelt. ACM SIGACT News, 42(4):17–21, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Nie02].

Teutsch:2013:RBR

[Teu13] Jason Teutsch. Review of Algorithmic Randomness and Complexity by Downey and Hirschfeldt. ACM SIGACT News, 44(1):25–28, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Tewari:2014:RAC

[Tew14] Raghunath Tewari. Review of Additive combinatorics by Terence Tao and Van H. Vu. ACM SIGACT News, 45(1):24–26, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Tsou:1982:DRS

[TF82] Don-Min Tsou and Patrick C. Fischer. Decomposition of a relation scheme into Boyce–Codd Normal Form. ACM SIGACT News, 14 (3):23–29, Summer 1982. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Thielecke:1999:CFJ

[Thi99] Hayo Thielecke. Continuations, functions and jumps. ACM SIGACT News, 30(2):33–42, June 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL http://www.cs.

bham.ac.uk/~hxt/research/Logiccolumn8.ps.

Thomborson:1993:WFF

[Tho93] Clark Thomborson. Why are fewer females obtaining bachelor’s degrees in computer science? ACM SIGACT News, 24(3):114– 116, October 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Thorup:2013:MPO

[Tho13] Mikkel Thorup. Mihai Pătrascu: obituary and open problems. ACM SIGACT News, 44(1):110–114, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Thurimella:1990:RWW

[Thu90] Ramki Thurimella. Report on WOPA: Workshop on Parallel Algorithms. ACM SIGACT News, 21(3):??, Summer 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). Published by: ACM Press.

Tielmann:2009:RD

[Tie09] Andreas Tielmann. A review of DISC 2009. ACM SIGACT News, 40(4):75–79, December 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Timnat:2013:DR

[Tim13] Shahar Timnat. DISC 2013 review. ACM SIGACT News, 44 (4):87–90, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Terese:2011:BRB

[TM11] Brittany Terese and David L. Millman. Book review: Geometric Algebra: an Algebraic System for Computer Games and Animation, by J. Vince. ACM SIGACT News, 42(1):46–48, March 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Vin09].

To:2006:BRB

[To06] Anthony Widjaja To. Book review: Algebraic Complexity Theory, by Peter Bürgisser, Michael Clausen and Amin Shokrollahi, Springer 1997. ACM SIGACT News, 37(2):22–27, June 2006. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BCS97].

Tollis:1993:VRP

[Tol93] Ioannis G. Tollis. Visibility representations of planar graphs (abstract). ACM SIGACT News, 24(1):57–58, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Tompa:1989:FM

[Tom89] M. Tompa. Figures of merit. ACM SIGACT News, 20(1):62–71, Winter 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Tompa:1990:FMS

[Tom90] Martin Tompa. Figures of merit: the sequel. ACM SIGACT News, 21(4):78–81, Fall 1990. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Toynbee:1993:OWC

[Toy93] Claire Toynbee. On the outer: women in computer science courses. ACM SIGACT News, 24(2):18–21, Spring 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Tran:2022:RAD

[Tra22] Nicholas Tran. Review of “The Algorithm Design Manual, 3rd ed. by Steven S. Skiena, Springer, 2020 793 pages, Hardcover, $99.99”. ACM SIGACT News, 53(3):21–23, September 2022. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

URL https://dl.acm.org/doi/10.1145/3561064.3561068.

Tran:2023:BRCa

[Tra23a] Nicholas Tran. The book review column. ACM SIGACT News, 54 (1):6–8, March 2023. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3586165.3586167.

Tran:2023:BRCb

[Tra23b] Nicholas Tran. The book review column. ACM SIGACT News, 54 (2):3–14, June 2023. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3604943.3604945.

Tran:2023:BRCc

[Tra23c] Nicholas Tran. The book review column. ACM SIGACT News, 54(3):3–21, September 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3623800.3623802.

Tran:2023:BRCd

[Tra23d] Nicholas Tran. The book review column. ACM SIGACT News, 54(4):7–26, December 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3639528.3639530.

Tran:2024:BRC

[Tra24] Nicholas Tran. The book review column. ACM SIGACT News, 55 (1):6–19, March 2024. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3654780.3654782.

Trevisan:2009:GCA

[Tre09] Luca Trevisan. Guest column: additive combinatorics and theoretical computer science. ACM SIGACT News, 40(2):50–66, June 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Trickey:1988:UL

[Tri88] Howard Trickey. Using LaTEX. ACM SIGACT News, 19(2):55– 57, Summer 1988. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Troelstra:1996:BPT

[TS96] A. S. (Anne Sjerp) Troelstra and Helmut Schwichtenberg. Basic Proof Theory, volume 43 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-57223-1 (hardcover). xi + 343 pp. LCCN QA9.54 .T76

1996. URL http://www.loc.gov/catdir/description/cam027/
96028596.html;   http://www.loc.gov/catdir/toc/cam027/
96028596.html.

Ta-Shma:1999:CVQ

[TS99] Amnon Ta-Shma. Classical versus quantum communication complexity. ACM SIGACT News, 30(3):25–34, September 1999. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Troelstra:2000:BPT

[TS00] A. S. (Anne Sjerp) Troelstra and Helmut Schwichtenberg. Basic Proof Theory, volume 43 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, second edition, 2000. ISBN 0-521-77911-1 (paperback). xii + 417 pp. LCCN QA9.54 .T76 2000. URL http://www.loc.gov/

catdir/description/cam021/00702855.html;   http://www.
loc.gov/catdir/samples/cam033/00702855.html;   http://
www.loc.gov/catdir/toc/cam027/00702855.html.

Tse:1991:UFS

[Tse91] T. H. Tse. A Unifying Framework for Structured Analysis and Design Models: an Approach Using Initial Algebra Semantics and Category Theory, volume 11 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-39196-2. xi + 179 pp. LCCN QA76.9.S84 T74

1991. URL http://www.loc.gov/catdir/description/cam025/
91204091.html;   http://www.loc.gov/catdir/toc/cam029/
91204091.html.

Tseng:2015:PR

[Tse15] Lewis Tseng. PODC 2015 review. ACM SIGACT News, 46(4):94– 102, December 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Tseng:2017:SR

[Tse17] Lewis Tseng. SIROCCO 2016 review. ACM SIGACT News, 48(1): 91–100, March 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Tarakanov:2003:IPA

[TSS03] A. O. Tarakanov, Victor A. Skormin, and S. P. (Svetlana P.) Sokolova. Immunocomputing: Principles and Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 0-387-95533-X. xi + 193 pp. LCCN QA76.875 .T37 2003. URL http://www.loc.gov/

catdir/enhancements/fy0817/2002044508-d.html;   http:/
/www.loc.gov/catdir/enhancements/fy0817/2002044508-t.html.

Tamassia:1995:RGD

[TT95] Roberto Tamassia and Ioannis G. Tollis. Report on Graph Drawing ’94. ACM SIGACT News, 26(1):87–91, March 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Tuchman:1966:ZT

[Tuc66] Barbara W. Tuchman. The Zimmermann telegram. Macmillan Publishing Company, New York, NY, USA, 1966. xii + 244 pp. LCCN D511 .T77 1966. Reprint of original 1958 edition. Kahn [Kah96] describes this book as “recount[ing] the political effects of the most important cryptogram solution in history”.

Turner:1984:GGEa

[Tur84a] Jonathan S. Turner. On the general graph embedding problem with applications to circuit layout. ACM SIGACT News, 15(4):59, Winter–Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Turner:1984:GGEb

[Tur84b] Jonathan S. Turner. On the general graph embedding problem with applications to circuit layout. ACM SIGACT News, 16(1):59, Winter–Spring 1984. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Tseng:2016:NFT

[TV16] Lewis Tseng and Nitin H. Vaidya. A note on fault-tolerant consensus in directed networks. ACM SIGACT News, 47(3):70–91, September 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Talbot:2006:CCI

[TW06] John Talbot and D. J. A. Welsh. Complexity and Cryptography: an Introduction. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-85231-5 (hardcover). LCCN Z103 .T35 2006.

Teutsch:2016:BSD

[TZ16] Jason Teutsch and Marius Zimand. A brief on short descriptions. ACM SIGACT News, 47(1):42–67, March 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Upadhyay:2010:BRB

[Upa10] Sarvagya Upadhyay. Book review: Concurrent Zero-Knowledge, by Alon Rosen, Springer-Verlag, 2006. ACM SIGACT News, 41 (4):38–42, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [RG06].

Upadhyay:2016:RDC

[Upa16] Jalaj Upadhyay. Review of: Distributed Computing Through Combinatorial Topology by Maurice Herlihy and Dmitry Kozlov and Sergio Rajsbaum. ACM SIGACT News, 47(2):18–20, June 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Upadhyay:2020:RIP

[Upa20] Sarvagya Upadhyay. Review of Introduction To Property Testing by Oded Goldreich. ACM SIGACT News, 51(4):6–10, December 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827

(electronic). URL https://dl.acm.org/doi/10.1145/3444815.
3444818.

Upadhyay:2021:RAA

[Upa21] Sarvagya Upadhyay. Review of Algorithmic Aspects of Machine Learning By Ankur Moitra. ACM SIGACT News, 52(1):9– 11, March 2021. CODEN SIGNDM. ISSN 0163-5700 (print),

1943-5827 (electronic). URL https://dl.acm.org/doi/10.1145/
3457588.3457591.

Vadhan:2007:UTP

[Vad07] Salil P. Vadhan. The unified theory of pseudorandomness: guest column. ACM SIGACT News, 38(3):39–54, September 2007. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Vakilian:2013:JRB

[Vak13] Mohsen Vakilian. Joint review of How to solve it: a new aspect of mathematical method by George Polya and Street-fighting mathematics by Sanjoy Mahajan. ACM SIGACT News, 44(2):10–12, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Valiente:2009:CPM

[Val09] Gabriel Valiente. Combinatorial Pattern Matching Algorithms in Computational Biology Using Perl and R. Chapman and Hall/CRC mathematical and computational biology series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-6973-X (hardcover). 352 pp. LCCN QH324.2 .V35 2009.

vanBenthemJutting:1979:CLG

[van79] L. S. van Benthem Jutting. Checking Landau’s “Grundlagen” in the Automath System. Mathematical Centre tracts; 83. Mathematisch Centrum, Amsterdam, The Netherlands, 1979. ISBN 90-6196-147-5 (paperback). 120 pp. LCCN QA241 .B46.

VandenBussche:2006:DTC

[Van06] Jan Van den Bussche. Database theory column: report on PODS 2006. ACM SIGACT News, 37(4):56–57, December 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

VanGucht:2010:DTC

[Van10] Dirk Van Gucht. Database theory column report on PODS 2010. ACM SIGACT News, 41(3):86–87, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vandenBroek:2004:CTA

[vdBN04] Pim van den Broek and Joost Noppen. Comparison of two approaches to dynamic programming. ACM SIGACT News, 35(4): 111–116, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Velu:1978:TPU

[Vél78] Jacques Vélu. Tests for primality under the Riemann hypothesis. ACM SIGACT News, 10(2):58–59, Summer 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Velleman:1994:HPI

[Vel94] Daniel J. Velleman. How to Prove It: a Structured Approach. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-44116-1 (hardcover), 0-521-44663-5 (paperback). ix + 309 pp. LCCN QA9 .V38 1994.

Vempala:2004:RPM

[Vem04] Santosh S. (Santosh Srinivas) Vempala. The Random Projection Method, volume 65 of DIMACS series in discrete mathematics and theoretical computer science. American Mathematical Society, Providence, RI, USA, 2004. ISBN 0-8218-2018-4. ix + 105 pp. LCCN QA501 .V45 2004.

Venkatasubramanian:1999:TRW

[Ven99] Suresh Venkatasubramanian. A theory repository on the Web: a proposal. ACM SIGACT News, 30(4):91–95, December 1999. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Venkatasubramanian:2002:DMP

[Ven02] Suresh Venkatasubramanian. Discrete mathematical problems with medical applications DIMACS volume 55. ACM SIGACT News, 33(4):9–11, December 2002. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Venkatasubramanian:2013:CGC

[Ven13a] Suresh Venkatasubramanian. Computational geometry column 55: new developments in nonnegative matrix factorization. ACM SIGACT News, 44(1):70–78, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Venkatasubramanian:2013:MHE

[Ven13b] Suresh Venkatasubramanian. Moving heaven and earth: distances between distributions. ACM SIGACT News, 44(3):56–68, September 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Verissimo:2006:TTW

[Ver06] Paulo Veríssimo. Travelling through wormholes: a new look at distributed systems models. ACM SIGACT News, 37(1):66–81, March 2006. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Vesely:2021:PSP

[Ves21] Pavel Veselý. Packet scheduling: Plans, monotonicity, and the golden ratio. ACM SIGACT News, 52(2):72–84, June 2021. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

URL https://dl.acm.org/doi/10.1145/3471469.3471481.

Vovk:2005:ALR

[VGS05] Vladimir Vovk, A. (Alexander) Gammerman, and Glenn Shafer. Algorithmic Learning in a Random World. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-00152-2, 0-387-25061-1 (e-book). xiv + 324 pp. LCCN QA279.2 .V68 2005. URL http://www.loc.gov/

catdir/enhancements/fy0662/2005042556-d.html;   http:/
/www.loc.gov/catdir/enhancements/fy0817/2005042556-t.html.

Viann:2005:SNC

[Via05] Victor Viann. SIGACT news complexity theory column 48. ACM SIGACT News, 36(3):24–38, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Vianu:2008:DTC

[Via08a] Victor Vianu. Database theory column. ACM SIGACT News, 39(1):77–79, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Vianu:2008:MFC

[Via08b] Victor Vianu. Modularity: a first class concept to address distributed systems. ACM SIGACT News, 39(2):91–110, June 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Vince:2009:GAA

[Vin09] John Vince. Geometric Algebra: An Algebraic System for Computer Games and Animation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-84882-379-7. 300 pp. LCCN QA199 .V56 2009.

Viola:2004:BRB

[Vio04] Lorenza Viola. Book review: Quantum Computing: Mika Hirvensalo, (2001) Springer-Verlag, Berlin ISBN 3-540-66783-0. Computer Physics Communications, 163(1):65–66, October 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504004114. See [Hir01] and re-

view [dW07].

Viola:2009:GCC

[Vio09] Emanuele Viola. Guest Column: correlation bounds for polynomials over {0, 1}. ACM SIGACT News, 40(1):27–44, March 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Viola:2017:SCC

[Vio17] Emanuele Viola. Selected challenges in computational lower bounds. ACM SIGACT News, 48(1):39–45, March 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Viola:2019:GCN

[Vio19] Emanuele Viola. Guest column: Non-abelian combinatorics and communication complexity. ACM SIGACT News, 50(3):52–74, September 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Vishkin:1991:CPA

[Vis91] Uzi Vishkin. Can parallel algorithms enhance serial implementation. ACM SIGACT News, 22(4):63, Fall 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Vitter:2000:ASA

[Vit00] Jeffery S. Vitter. ACM SIGACT 1999–2000 annual report. ACM SIGACT News, 31(3):2–6, September 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanLeeuwen:1974:FCB

[vL74] Jan van Leeuwen. A forgotten connection between tag-systems and parallel-rewriting. ACM SIGACT News, 6(4):19–20, October 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanLeeuwen:1976:RCP

[vL76] Jan van Leeuwen. A regularity condition for parallel rewriting systems. ACM SIGACT News, 8(4):24–27, October 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanLeeuwen:1978:EPR

[vL78] Jan van Leeuwen. Evaluating a polynomial and its reverse. ACM SIGACT News, 10(1):18–21, Spring 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Vlach:1994:BRB

[Vla94] Frank Vlach. Book review: The Deductive Foundations of Computer Programming. By Zohar Manna and Richard Waldinger. (Addison-Wesley, 1993. xiv + 717pp. ISBN 0-201-54886-0. $46.25). ACM SIGACT News, 25(2):26–27, June 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [MW93].

vanLeunen:1976:HYA

[vLL76] Mary-Claire van Leunen and Richard Lipton. How to have your abstract rejected. ACM SIGACT News, 8(3):21–24, July 1976. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Vitanyi:1985:BOV

[VM85] Paul M. B. Vitányi and Lambert Meertens. Big omega versus the wild functions. ACM SIGACT News, 16(4):56–59, April 1985. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Knu76a, Bra85].

Vollmer:1999:UCC

[Vol99] Heribert Vollmer. Uniform characterizations of complexity classes. ACM SIGACT News, 30(1):17–27, March 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Volk:2021:GCA

[Vol21] Ben Lee Volk. Guest column: Algebraic natural proofs Ben Lee Volk. ACM SIGACT News, 52(4):56–73, December 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL

https://dl.acm.org/doi/10.1145/3510382.3510392.

Volk:2023:SNCa

[Vol23a] Ben Lee Volk. SIGACT news complexity theory column 118. ACM SIGACT News, 54(2):43, June 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.

org/doi/10.1145/3604943.3604949.

Volk:2023:SNCb

[Vol23b] Ben Lee Volk. SIGACT news complexity theory column 119. ACM SIGACT News, 54(4):53, December 2023. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://

dl.acm.org/doi/10.1145/3639528.3639534.

Volk:2024:SNC

[Vol24] Ben Lee Volk. SIGACT news complexity theory column 120. ACM SIGACT News, 55(1):66, March 2024. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.

org/doi/10.1145/3654780.3654787.

Vega-Redondo:2007:CSN

[VR07] Fernando Vega-Redondo. Complex Social Networks, volume 44 of Econometric Society monographs. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-85740-6 (hardcover), 0-521-67409-3 (paperback). xiv + 294 pp. LCCN HM741

.V44 2007. URL http://www.loc.gov/catdir/enhancements/
fy0702/2006030242-d.html;   http://www.loc.gov/catdir/
enhancements/fy0702/2006030242-t.html;   http://www.loc.
gov/catdir/enhancements/fy0729/2006030242-b.html.

Volpano:1998:CPP

[VS98] Dennis Volpano and Geoffrey Smith. Confinement properties for programming languages. ACM SIGACT News, 29(3):33–42, September 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanStee:2012:SNOa

[vS12a] Rob van Stee. SIGACT news online algorithms column 20: the power of harmony. ACM SIGACT News, 43(2):127–136, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanStee:2012:SNOb

[vS12b] Rob van Stee. SIGACT news online algorithms column 21: AP-PROX and ALGO. ACM SIGACT News, 43(4):123–129, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanStee:2013:SNO

[vS13] Rob van Stee. SIGACT News online algorithms column 22. ACM SIGACT News, 44(2):125, June 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanStee:2014:SNOa

[vS14a] Rob van Stee. SIGACT news online algorithms column 23. ACM SIGACT News, 45(1):90, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanStee:2014:SNOb

[vS14b] Rob van Stee. SIGACT news online algorithms column 24: 2014 so far. ACM SIGACT News, 45(3):105–111, September 2014. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanStee:2015:SNO

[vS15] Rob van Stee. SIGACT News online algorithms column 26: Bin packing in multiple dimensions. ACM SIGACT News, 46(2):105– 112, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanStee:2016:SNO

[vS16a] Rob van Stee. SIGACT News online algorithms column 29. ACM SIGACT News, 47(3):92, September 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanStee:2016:SNOa

[vS16b] Rob van Stee. SIGACT news online algorithms column 27: Online matching on the line, Part 1. ACM SIGACT News, 47(1):99–110, March 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanStee:2016:SNOb

[vS16c] Rob van Stee. SIGACT news online algorithms column 28: Online matching on the line, Part 2. ACM SIGACT News, 47(2):40–51, June 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanStee:2017:SNOa

[vS17a] Rob van Stee. SIGACT News online algorithms column 30: 2016 in review. ACM SIGACT News, 48(1):46–53, March 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanStee:2017:SNOb

[vS17b] Rob van Stee. SIGACT News online algorithms column 32: 2017 in review. ACM SIGACT News, 48(4):100–109, December 2017. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanStee:2018:SNOa

[vS18a] Rob van Stee. SIGACT news online algorithms column 33. ACM SIGACT News, 49(1):66, March 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanStee:2018:SNOb

[vS18b] Rob van Stee. SIGACT news online algorithms column 34: 2018 in review. ACM SIGACT News, 49(4):36–45, December 2018. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

vanStee:2021:SNO

[vS21] Rob van Stee. SIGACT News online algorithms column 37. ACM SIGACT News, 52(2):71, June 2021. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.

org/doi/10.1145/3471469.3471480.

vanStee:2022:SNO

[vS22] Rob van Stee. SIGACT news online algorithms column 39. ACM SIGACT News, 53(2):83, June 2022. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.

org/doi/10.1145/3544979.3544992.

Vukolic:2010:BEI

[Vuk10] Marko Vukolić. The Byzantine Empire in the intercloud. ACM SIGACT News, 41(3):105–111, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

VanDerMeulen:1978:TPS

[VV78] S. G. Van Der Meulen and M. Veldhorst. TORIX: a Programming System for Operations on Vectors and Matrices over Arbitrary Fields and of Variable Size, volume 1. Mathematisch Centrum, Amsterdam, The Netherlands, 1978. ISBN ???? ???? pp. LCCN ????

Wagstaff:2003:CNT

[Wag03] Samuel S. Wagstaff. Cryptanalysis of Number Theoretic Ciphers. Computational mathematics. Chapman and Hall/CRC, Boca Raton, FL, USA, 2003. ISBN 1-58488-153-4. xv + 318 pp. LCCN QA76.9.A25 W33 2003. URL http://www.loc.gov/

catdir/enhancements/fy0646/2002034919-d.html.

Wall:1963:CM

[Wal63] H. S. (Hubert Stanley) Wall. Creative Mathematics. University of Texas Press, Austin, TX, USA, 1963. 189 pp. LCCN QA37 .W29.

Walsh:1982:BBO

[Wal82] T. R. S. Walsh. The busy beaver on a one-way infinite tape. ACM SIGACT News, 14(1):38–43, Winter 1982. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Wallis:2003:BGD

[Wal03] W. D. Wallis. A Beginner’s Guide to Discrete Mathematics. Birkhäuser Verlag, Basel, Switzerland, 2003. ISBN 0-8176-4269-2 , 3-7643-4269-2. xi + 367 pp. LCCN QA39.3 .W35 2003. URL http:

//www.loc.gov/catdir/enhancements/fy0812/2002026043-d.
html;   http://www.loc.gov/catdir/enhancements/fy0812/
2002026043-t.html.

Wall:2009:CM

[Wal09] H. S. (Hubert Stanley) Wall. Creative Mathematics. Classroom resource materials. Mathematical Association of America, Washington, DC, USA, revised edition, 2009. ISBN 0-88385-750-2. xx + 195 pp. LCCN QA37 .W29 2009. URL http:

//www.loc.gov/catdir/enhancements/fy1114/2008938215-b.
html;   http://www.loc.gov/catdir/enhancements/fy1114/
2008938215-d.html;   http://www.loc.gov/catdir/enhancements/
fy1114/2008938215-t.html.

Wallis:2012:BGD

[Wal12] W. D. Wallis. A Beginner’s Guide to Discrete Mathematics. Birkhäuser Verlag, Basel, Switzerland, second edition, 2012. ISBN 0-8176-8285-6 (hardcover), 0-8176-8286-4 (e-book). xiii + 427 pp. LCCN QA39.3 .W35 2012.

Wang:1997:RTA

[Wan97] Jie Wang. Report on the Third Annual International Computing and Combinatorics Conference (COCOON’97): Shanghai, China, August 20–23, 1997. ACM SIGACT News, 28(4):69–76, December 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Wan:2005:BRB

[Wan05] Raymond Wan. Book review: Genomic Perl From Bioinformatics Basics to Working Code, by Rex A. Dwyer, Cambridge University Press 2003, 0-521-80177-X. ACM SIGACT News, 36(4):9–12, December 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Dwy03].

Wang:2012:RBN

[Wan12] Yu Wang. Review of Vehicular Networks, from Theory to Practice, edited by Stephan Olariu and Michele C. Weigle. ACM SIGACT News, 43(4):25–29, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Wapner:2005:PSM

[Wap05] Leonard M. Wapner. The Pea and the Sun: a Mathematical Paradox. A. K. Peters, Ltd., Wellesley, MA, USA, 2005. ISBN 1-56881-213-2. xiv + 218 pp. LCCN QA248 .W29 2005.

Washington:2008:BRB

[Was08a] Lawrence C. Washington. Book review: Handbook of Elliptic and Hyperelliptic Curve Cryptography, by H. Cohen and G. Frey, Chapman & Hall/CRC, 2006, 1-58488-518-1. ACM SIGACT News, 39(1):19–22, March 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [CFA^{+}06].

Washington:2008:ECN

[Was08b] Lawrence C. Washington. Elliptic Curves: Number Theory and Cryptography. Discrete mathematics and its applications. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2008. ISBN 1-4200-7146-7 (hardcover). xviii + 513 pp. LCCN

QA567.2.E44 W37 2008. URL http://www.loc.gov/catdir/
toc/ecip0810/2008006296.html.

Watrous:2011:GCI

[Wat11] John Watrous. Guest column: An introduction to quantum information and quantum circuits 1. ACM SIGACT News, 42(2):52–67, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Wegner:1972:OSP

[Weg72] Peter Wegner. Operational semantics of programming languages. ACM SIGACT News, 4(14):128–141, January 1972. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Wegner:1974:MAU

[Weg74] Peter Wegner. Modification of Aho and Ullman’s correctness proof of Warshall’s algorithm. ACM SIGACT News, 6(1):32–35, January 1974. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Wegman:1986:WLP

[Weg86] M. N. Wegman. What it’s like to be a POPL referee. ACM SIGACT News, 17(4):50–51, Spring 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Wegener:2000:BPB

[Weg00] Ingo Wegener. Branching Programs and Binary Decision Diagrams: Theory and Applications. SIAM monographs on discrete mathematics and applications. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2000. ISBN 0-89871-458-3. x + 408 pp. LCCN QA274.76 .W44 2000.

Weiss:1993:TRC

[Wei93] Mark A. Weiss. Technical report column. ACM SIGACT News, 24(1):10–16, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Weiss:1995:JBRa

[Wei95a] Mark A. Weiss. Journal Backlog Report. ACM SIGACT News, 26 (2):39–44, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Weiss:1995:JBRb

[Wei95b] Mark A. Weiss. Journal Backlog Report. ACM SIGACT News, 26(4):24–31, December 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Weiss:1995:TRCa

[Wei95c] Mark A. Weiss. Technical report column. ACM SIGACT News, 26(1):34–39, March 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Weiss:1995:TRCb

[Wei95d] Mark A. Weiss. Technical report column. ACM SIGACT News, 26 (2):45–46, June 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Weiss:1995:APS

[Wei95e] Mark Allen Weiss. Algorithms, Problem Solving, and Data Structures with C + + . ACM SIGACT News, 26(4):39–50, December 1995. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Weiss:1996:JBR

[Wei96a] Mark A. Weiss. Journal Backlog Report. ACM SIGACT News, 27 (2):38–44, June 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Weiss:1996:TRC

[Wei96b] Mark A. Weiss. Technical report column. ACM SIGACT News, 27(1):16–23, March 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Weiss:1998:DSP

[Wei98] Mark Allen Weiss. Data structures and problem solving using Java. ACM SIGACT News, 29(2):42–49, June 1998. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Weissman:2003:BRB

[Wei03] Vicky Weissman. Book review: Modal and Temporal Properties of Processes, by Colin Stirling. Springer Verlag 2001. ACM SIGACT News, 34(2):8–15, June 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sti01].

Weinstein:2015:ICQ

[Wei15] O. Weinstein. Information complexity and the quest for interactive compression. ACM SIGACT News, 46(2):41–64, June 2015. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welschenbach:2005:CCC

[Wel05] Michael Welschenbach. Cryptography in C and C++. Apress, Berkeley, CA, USA, second edition, 2005. ISBN 1-59059-502-5. xxv + 478 pp. LCCN QA76.9.A25 W4313 2005. URL http:

//www.loc.gov/catdir/enhancements/fy0663/2005002553-d.
html;   http://www.loc.gov/catdir/toc/ecip057/2005002553.
html.

Welch:2013:DCC

[Wel13] Jennifer L. Welch. Distributed computing column 52: annual review 2013. ACM SIGACT News, 44(4):79–80, December 2013. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2014:DCCa

[Wel14a] Jennifer L. Welch. Distributed computing column 53: Dagstuhl seminar review: consistency in distributed systems. ACM SIGACT News, 45(1):66, March 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2014:DCCb

[Wel14b] Jennifer L. Welch. Distributed computing column 54: Transactional memory: models and algorithms. ACM SIGACT News, 45 (2):73, June 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2014:DCCc

[Wel14c] Jennifer L. Welch. Distributed computing column 55: WTTM 2013 review, and lower bounds for distributed quantum computing. ACM SIGACT News, 45(3):71, September 2014. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2014:DCCd

[Wel14d] Jennifer L. Welch. Distributed computing column 56: Annual review 2014. ACM SIGACT News, 45(4):86–88, December 2014. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2015:DCCa

[Wel15a] Jennifer L. Welch. Distributed computing column 57: Distributed algorithms as combinatorial structures. ACM SIGACT News, 46 (1):62, March 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2015:DCCb

[Wel15b] Jennifer L. Welch. Distributed computing column 58: Maurice Herlihy’s 60th birthday celebration. ACM SIGACT News, 46(2):

78, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2015:DCCc

[Wel15c] Jennifer L. Welch. Distributed computing column 59: Resourcecompetitive algorithms. ACM SIGACT News, 46(3):56, September 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2015:DCCd

[Wel15d] Jennifer L. Welch. Distributed computing column 60: Annual review 2015. ACM SIGACT News, 46(4):82–83, December 2015. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2016:DCCa

[Wel16a] Jennifer L. Welch. Distributed computing column 61: Distributed algorithmic foundations of dynamic networks. ACM SIGACT News, 47(1):68, March 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2016:DCCb

[Wel16b] Jennifer L. Welch. Distributed computing column 62: Decidability in parameterized verification. ACM SIGACT News, 47(2):52, June 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2016:DCCc

[Wel16c] Jennifer L. Welch. Distributed computing column 63: a note on fault-tolerant consensus in directed networks. ACM SIGACT News, 47(3):69, September 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2016:DCCd

[Wel16d] Jennifer L. Welch. Distributed computing column 64: Annual review 2016. ACM SIGACT News, 47(4):33–34, December 2016. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2017:DCCa

[Wel17a] Jennifer L. Welch. Distributed computing column 65: Automatic synthesis of distributed protocols and SIROCCO 2016 review. ACM SIGACT News, 48(1):54, March 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2017:DCCb

[Wel17b] Jennifer L. Welch. Distributed computing column 66: Algorithmic foundations of programmable matter Dagstuhl Seminar 16271. ACM SIGACT News, 48(2):86, June 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2017:DCCc

[Wel17c] Jennifer L. Welch. Distributed computing column 67: Review of 2016 BIRS CMO Workshop on Complexity and Analysis of Distributed Algorithms. ACM SIGACT News, 48(3):60–67, September 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2017:DCCd

[Wel17d] Jennifer L. Welch. Distributed computing column 68. ACM SIGACT News, 48(4):78–79, December 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2018:DCCa

[Wel18a] Jennifer L. Welch. Distributed computing column 69 Proving PACELC and Concurrent Computing Summer School. ACM SIGACT News, 49(1):72, March 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2018:DCCb

[Wel18b] Jennifer L. Welch. Distributed computing column 70: Formalizing and implementing distributed ledger objects. ACM SIGACT News, 49(2):57, June 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2018:DCCc

[Wel18c] Jennifer L. Welch. Distributed computing column 71: Recent algorithmic advances in population protocols. ACM SIGACT News, 49(3):62, September 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2018:DCCd

[Wel18d] Jennifer L. Welch. Distributed computing column 72: Annual review 2018. ACM SIGACT News, 49(4):58–59, December 2018. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2019:DCCa

[Wel19a] Jennifer L. Welch. Distributed computing column 73 SPAA 2018 review. ACM SIGACT News, 50(1):60, March 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2019:DCCb

[Wel19b] Jennifer L. Welch. Distributed computing column 74 survey of reconfigurable data center networks. ACM SIGACT News, 50(2): 61, June 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welch:2019:DCCc

[Wel19c] Jennifer L. Welch. Distributed computing column 75 the splendors and miseries of rounds. ACM SIGACT News, 50(3):34, September 2019. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Werden:2010:BRB

[Wer10] David A. Werden. Book review: Coding for Data and Computer Communications, by David Salomon, Springer, 2005. ACM SIGACT News, 41(4):32–34, December 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sal05a].

Werden:2013:RBA

[Wer13] Dave Werden. Review of Spyware and Adware by John Aycock. ACM SIGACT News, 44(1):17–19, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Winfree:2000:DBC

[WG00] Erik Winfree and David K. Gifford, editors. DNA Based Computers V: DIMACS Workshop DNA Based Computers V, June 14–15, 1999, Massachusetts Institute of Technology, volume 54 of DIMACS series in discrete mathematics and theoretical computer science. American Mathematical Society, Providence, RI, USA, 2000. ISBN 0-8218-2053-2. LCCN QA76.887. D5 1999.

Widjaja:2003:BRB

[Wid03] Anthony Widjaja. Book review: Algorithms Sequential & Parallel: a Unified Approach, by R. Miller & L. Boxer. Prentice Hall 2000. ACM SIGACT News, 34(2):3–5, June 2003. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [MB05].

Wiesner:1983:CC

[Wie83] Stephen Wiesner. Conjugate coding. ACM SIGACT News, 15(1): 78–88, Winter–Spring 1983. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Wiedemann:1986:QC

[Wie87] D. Wiedemann. Quantum cryptography. ACM SIGACT News, 18 (2):48–51, September/March 1986–1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Wilson:1986:HPT

[Wil86] D. M. Wilson. The halting problem for Turning machines. ACM SIGACT News, 18(1):52, Summer 1986. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Wille:2004:BRB

[Wil04] Luc T. Wille. Book review: Learning Kernel Classifiers: Theory and Algorithms, by Ralf Herbrich. MIT Press, Cambridge, Mass., 2002. ISBN 0-262-08306-X, 384 pages; and Learning with Kernels: Support Vector Machines, Regularization Optimization and Beyond by Bernhard Scholkopf and Alexander J. Smola. MIT Press, Cambridge, Mass., 2002, ISBN 0-262-19475-9, 644 pages. ACM SIGACT News, 35(3):13–17, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Her02a, SS02].

Willinsky:2006:APC

[Wil06] John Willinsky. The Access Principle: the Case for Open Access to Research and Scholarship. Digital libraries and electronic publishing. MIT Press, Cambridge, MA, USA, 2006. ISBN 0-262-23242-1. xv + 287 pp. LCCN Z286.O63 W55 2006. URL http://edrev.

asu.edu/reviews/rev478.htm;   http://mitpress.mit.edu/
catalog/item/default.asp?tid=10611\%26ttype=2.

Williams:2008:APT

[Wil08] Ryan Williams. Applying practice to theory. ACM SIGACT News, 39(4):37–52, December 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Williams:2010:BRB

[Wil10a] Mike Williams. Book review: Introduction to Languages and Machines, by Alan P. Parkes, Springer, 2008. ACM SIGACT News, 41 (2):37–40, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Par08].

Wilson:2010:BRBb

[Wil10b] Kevin A. Wilson. Book review: A Second Course in Formal Languages and Automata Theory, by Jeffrey Shallit, publisher: Cambridge University Press, 2008. ACM SIGACT News, 41(2):40–43, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Sha09].

Wilson:2010:BRBa

[Wil10c] Kevin A. Wilson. Book review: The Annotated Turing, by Charles Petzold, Publisher Wiley, 2008. ACM SIGACT News, 41(2):22–26, June 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Pet08].

Wilson:2010:BRBc

[Wil10d] Mark C. Wilson. Book review: The Mathematics of Voting and Elections: a Hands-On Approach, by Jonathan K. Hodge and Richard E. Klima, American Mathematical Society (Mathematical World Series, volume 22) 226 + xiv pages. ACM SIGACT News, 41(3):34–36, September 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [HK05].

Williams:2011:GCC

[Wil11] Ryan Williams. Guest column: a casual tour around a circuit complexity bound. ACM SIGACT News, 42(3):54–76, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Wilson:2013:RBC

[Wil13] Mark C. Wilson. Review of Models of Conflict and Cooperation by Rick Gillman and David Housman. ACM SIGACT News, 44(1): 34–35, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Williams:2018:SOP

[Wil18] Virginia Vassilevska Williams. Some open problems in fine-grained complexity. ACM SIGACT News, 49(4):29–35, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Welschenbach:2001:CCC

[WK01] Michael Welschenbach and David Kramer. Cryptography in C and C++. Apress, Berkeley, CA, USA, second edition, 2001. ISBN 1-893115-95-X. xix + 432 pp. LCCN QA76.9.A25 W4313 2001. URL

http://www.loc.gov/catdir/toc/fy0702/2005298150.html.

Williams:1987:RPS

[WM87] K. L. Williams and M. R. Meybodi. Representing problems as string transformations. ACM SIGACT News, 18(3):29–30, April 1987. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Wolfram:1993:CTT

[Wol93] D. A. Wolfram. The Clausal Theory of Types, volume 21 of Cambridge tracts in theoretical computer science. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-39538-0 (hardcover). viii + 124 pp. LCCN QA76.63 .W64 1993. $124.95. URL http://

www.loc.gov/catdir/description/cam026/93246461.html;
http://www.loc.gov/catdir/toc/cam029/93246461.html.

Wolsey:1998:IP

[Wol98] Laurence A. Wolsey. Integer programming. Wiley-Interscience series in discrete mathematics and optimization. John Wiley, New York, NY, USA, 1998. ISBN 0-471-28366-5. xviii + 264 pp. LCCN

T57.74 .W67 1998. URL http://www.loc.gov/catdir/bios/
wiley047/98007296.html;   http://www.loc.gov/catdir/description/
wiley032/98007296.html;   http://www.loc.gov/catdir/toc/
onix02/98007296.html.

Wolfram:2002:NKS

[Wol02] Stephen Wolfram. A New Kind of Science. Wolfram Media, Champaign, IL, USA, 2002. ISBN 1-57955-008-8. xiv + 1197 pp. LCCN QA267.5.C45 W67 2002. URL http://www.loc.gov/

catdir/toc/fy031/2001046603.html.

Wood:1978:OSH

[Woo78] Derick Wood. One-sided height-balanced search trees. ACM SIGACT News, 10(2):60–62, Summer 1978. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Williamson:2011:DAA

[WS11] David P. Williamson and David Bernard Shmoys. The design of approximation algorithms. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-19527-6 (hardcover). xi + 504 pp.

LCCN QA221 .W55 2011. URL http://assets.cambridge.org/
97805211/95270/cover/9780521195270.jpg.

Xie:2012:RBA

[Xie12] Yulai Xie. Review of Applied Algebra: Codes, Ciphers and Discrete Algorithms, by Darel W. Hardy, Fred Richman, and Carol L. Walker. ACM SIGACT News, 43(3):25–27, September 2012. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Xiao-Long:1993:RPB

[XL93] Hu Xiao-Long. The representation of a program in the Blum– Shub–Smale theory of computation over an arbitrary ring. ACM SIGACT News, 24(3):117–119, October 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Yannakakis:1996:PDT

[Yan96] Mihalis Yannakakis. Perspectives on database theory. ACM SIGACT News, 27(3):25–49, September 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Yan:2002:NTC

[Yan02] Song Y. Yan. Number Theory for Computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2002. ISBN 3-540-43072-5. xxii + 435 pp. LCCN QA241 .Y27 2002. URL http:

//www.loc.gov/catdir/enhancements/fy0812/2002510398-d.
html;   http://www.loc.gov/catdir/enhancements/fy0812/

2002510398-t.html. Foreword by Martin E. Hellman.

Yan:2007:CAR

[Yan07] Song Y. Yan. Cryptanalytic attacks on RSA. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-48741-7. xx + 254 pp. LCCN QA76.9.A25 Y34 2007.

Yan:2011:BRBa

[Yan11a] Song Yan. Book review: A Guide to Elementary Number Theory, by Underwood Dudley. ACM SIGACT News, 42(2):41–42, June 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Dud09].

Yan:2011:BRBb

[Yan11b] Song Yan. Book review: Dynamic Fuzzy Logic and Its Applications, by Fanzhang Li. ACM SIGACT News, 42(4):48–49, December 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Li08].

Yan:2012:RBC

[Yan12] Song Yan. Review of Modern Computer Arithmetic, by Richard Brent and Paul Zimmermann. ACM SIGACT News, 43(4):49– 51, December 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Yan:2020:RAN

[Yan20] Song Y. Yan. Review of Applied Number Theory. ACM SIGACT News, 51(2):9–10, June 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/doi/

10.1145/3406678.3406682.

Cai:2008:HAG

[yC08] Jin yi Cai. Holographic algorithms: guest column. ACM SIGACT News, 39(2):51–81, June 2008. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Yang:1997:SBR

[YDX97] Ruiguang Yang, Longyun Ding, and Shurun Xu. Some better results estimating the shift function in terms of busy beaver function. ACM SIGACT News, 28(1):43–48, March 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

years:2000:BRC

[yea00a] Publication years. The book review column. ACM SIGACT News, 31(1):101, March 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

years:2000:DM

[yea00b] Publication years. Death of a monster. ACM SIGACT News, 31 (4):130–133, December 2000. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Yehudayoff:2012:PET

[Yeh12] Amir Yehudayoff. Proving expansion in three steps. ACM SIGACT News, 43(3):67–84, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Yelowitz:1978:NNT

[Yel78] Lawrence Yelowitz. Notes on “A note on the transitive closure of a Boolean matrix”. ACM SIGACT News, 10(1):30, Spring 1978. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Shy78b].

Ylikoski:2004:SSI

[Yli04] Antti Ylikoski. Some simple but interesting results concerning the P ?= NP problem. ACM SIGACT News, 35(3):94–97, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Ylikoski:2005:HPF

[Yli05] Antti Ylikoski. The halting problem on finite and infinite computers. ACM SIGACT News, 36(1):132–138, March 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Yanofsky:2008:QCC

[YM08] Noson S. Yanofsky and Mirco A. Mannucci. Quantum Computing for Computer Scientists. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-87996-5 (hardcover). xvi + 384 pp. LCCN

QA76.889 .Y35 2008. URL http://www.loc.gov/catdir/toc/
ecip0818/2008020507.html.

Young:1992:HRS

[You92] Paul Young. How reductions to sparse sets collapse the polynomialtime hierarchy: a primer; part I: polynomial-time Turing reductions. ACM SIGACT News, 23(3):107–117, Summer 1992. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Young:1999:BRB

[You99] Neal E. Young. Book review: Online Computation and Competitive Analysis, by Allan Borodin and Ran El-Yaniv. ACM SIGACT News, 30(4):13–17, December 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [BEY05].

Yu:2011:SDS

[Yu11] Haifeng Yu. Sybil defenses via social networks: a tutorial and survey. ACM SIGACT News, 42(3):80–101, September 2011. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Yuval:1976:GMD

[Yuv76] G. Yuval. The geometric mean distance. ACM SIGACT News, 8 (3):24–25, July 1976. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Yu:1991:SCI

[YZ91] Sheng Yu and Qingyu Zhuang. On the state complexity of intersection of regular languages. ACM SIGACT News, 22(3):52–54, Summer 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Zalcstein:1997:NRTa

[Zal97a] Zeke Zalcstein. NSF Report: Theory of Computing Program. ACM SIGACT News, 28(2):34, June 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Zalcstein:1997:NRTb

[Zal97b] Zeke Zalcstein. NSF report: theory of computing program. ACM SIGACT News, 28(3):32, September 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Zalcstein:1997:NRTc

[Zal97c] Zeke Zalcstein. NSF Report: Theory of Computing Program. ACM SIGACT News, 28(4):31, December 1997. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Zalcstein:1999:NRT

[Zal99] Zeke Zalcstein. NSF report: theory of computing program. ACM SIGACT News, 30(4):20–21, December 1999. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Zampolli:1977:LSP

[Zam77] Antonio Zampolli, editor. Linguistic structures processing, volume 5 of Fundamental studies in computer science. North-Holland Publishing Co., Amsterdam, The Netherlands, 1977. ISBN 0-444-85017-1. xii + 585 pp. LCCN P98 .I575 1974.

Zeugmann:1989:BRB

[Zeu89] T. Zeugmann. Book review: Baase, Computer Algorithms: Introduction to Design and Analysis (2nd ed.) (1988). ACM SIGACT News, 20(??), ???? 1989. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Baa78, Baa88, BV00].

Zeugmann:1992:BRR

[Zeu92] T. Zeugmann. Book review: Reingold and Shen, More Nearly Optimal Algorithms for Unbounded Searching (1991). ACM SIGACT News, 23(??):??, ???? 1992. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [RS88, RS91b, RS91a].

Zeugmann:1994:RC

[Zeu94] Thomas Zeugmann. Report on COLT 1994. ACM SIGACT News, 25(4):88–95, December 1994. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Zhao:2004:BRB

[Zha04] Wenzhong Zhao. Book review: Reasoning about Uncertainty, by Joseph Y. Halpern. The MIT Press, 2003. ACM SIGACT News, 35(3):8–12, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Hal03].

Zhao:2005:BRB

[Zha05] Wenzhong Zhao. Book review: Immunocomputing: Principles and Applications, by Alexander O. Tarakanov, Victor A. Skormin, Svetlana P. Sokolova, Springer-Verlag New York, Inc. 2003. ACM SIGACT News, 36(4):14–17, December 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [TSS03].

Zhou:2009:BRL

[Zho09] Lidong Zhou. Building reliable large-scale distributed systems: when theory meets practice. ACM SIGACT News, 40(3):78–85, September 2009. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Zhu:2020:CGC

[Zhu20] Binhai Zhu. Computational geometry column 70: Processing persistence diagrams as purely geometric objects. ACM SIGACT News, 51(1):105–117, March 2020. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL https://dl.acm.org/

doi/10.1145/3388392.3388404.

Zimand:2004:CCQ

[Zim04] Marius Zimand. Computational Complexity: a Quantitative Perspective, volume 196 of North-Holland mathematics studies, 0304-0208. Elsevier, Amsterdam, The Netherlands, 2004. ISBN 0-444-82841-9. xii + 340 pp. LCCN QA267.7 .Z55 2004. URL http:

//www.loc.gov/catdir/enhancements/fy0632/2006273471-d.
html.

Zimand:2013:RBE

[Zim13] Marius Zimand. Review of Deterministic Extraction from Weak Random Sources by Ariel Gabizon. ACM SIGACT News, 44(1): 36–37, March 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Zomaya:2006:HNI

[Zom06] Albert Y. Zomaya, editor. Handbook of Nature-Inspired and Innovative Computing: Integrating Classical Models with Emerging Technologies. Springer Science+Business Media, New York, NY, USA, 2006. ISBN 0-387-40532-1. xiv + 736 pp. LCCN MLCM 2006/40164 (Q); QA76.618 .H363x 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0663/2005933256-d.
html;   http://www.loc.gov/catdir/enhancements/fy0818/
2005933256-b.html;   http://www.loc.gov/catdir/enhancements/
fy0818/2005933256-t.html.

Zwarico:1993:BRB

[Zwa93] Amy Zwarico. Book review: Action Refinement in Process Algebras. By Luca Aceto. (Cambridge University Press, 1992. ix + 273pp. ISBN 0-521-43111-5. $49.95). ACM SIGACT News, 24(2): 25–26, Spring 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [Ace92].